A Bibliography of IEEE Transactions on Computers (1968–1969) and its predecessors: IEEE Transactions on Electronic Computers, IRE Transactions on Electronic Computers, and Transactions of the I.R.E. Professional

Group on Electronic Computers

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

01 October 2024 Version 1.14

Title word cross-reference

(d, k) [Ake65c, Fri66e, Kor67c]. (d, k, μ) [Qua69]. 1 [Wei69b]. 15 [Sei64]. 2 [EKP67, Fia66, Fre67a, Roy68b, SA64, Sch69b, Wei69b]. 21/2 [Gil66d]. 2 × 1 [Set66a]. 3 [AS66, Fre67a, Kei64, SA64, Wol69]. 4 [Fre67a]. 48 [Eps68]. [n, k] [Lec64]. _{2} [MS66]. 3 [MS66]. a [Mul59d, Mul60]. cos x [Spe65]. cot^{−1} x [Spe65]. exp x [Spe65]. F [Sed68, Hsi68]. γ [MS66]. k [Lem69, Liu63b, Liu64a]. k = 4 [Bra66]. ln x [Spe65]. m [Fro65]. N [Ber58, Ble61b, Dav65, Hig61b, MN61, BB62, Col69a, Fro65, Lee62a, Ull69b]. P^{n} [Mow66]. π [GH59]. r [HS69c]. Σ(k) [Bra66]. sin x [Spe65]. t∗ [Tay65]. tan^{−1} x [Spe65]. θ [Bro60b]. X [Ort59]. x^{y} [Kri64]. Z [Tou60, Fuk64].

-Channel [Eps68, Hsi68]. -Classification [Wol69]. -Dimensional [Col69a]. -Fe [MS66]. -Input [Wei69b]. -Multiprogramming [Wel68]. -N [GH59]. -Number [Mul59d, Mul60]. -Output [Wei69b]. -Stable [Lem69]. -Summability [Roy68b]. -Tuple [BB62, Ull69b, Ble61b, Hig61b]. -Valued [Kei64, Ber58].

1 [Ben69b, BJST64, Les64, Min69, MT59]. 10 [Dre67, Fif68a, Wak69]. 104 [Mel63]. 105 [Sha63, Wei52]. 105-Bit [Mei65]. 106 [Bas63b]. 107 [Woo63a]. 11 [Bek65, Den67, Gil69, Lau68]. 11-Cryotron [Dav63b, Squ62]. 12 [Ard65, Don67, Hor69, Nie68, Wai66]. 1299/000/01 [Mur65c]. 13 [Hor68, Nie69b, Paq67, Sal65]. 14 [Ano65s, Arm65, Cur68a, Har69a, Wat67]. 15 [Ano66p, Ano66-31, Don65b, Mar68, Sch69a, WS67a]. 16 [Ano67a, Bre67, Gre68a, Kar67, Led63, Tor69]. 160 [GJN59]. 160-MC [GJN59]. 1620 [Gas65]. 17 [Cro68, Gre64b, Leh63b, Nie69a]. 18 [Ano69b, Gai68, Gra64b, Gre69b, Sem63]. 19 [Bed64, Ull69a, Who63, Wie68]. 1961 [Ano60e]. 1963 [Ano62-43].

2 [Ano68-43, Hau69]. 2.18-Microsecond [ABC61]. 20 [Lar69, Smi68c, Sto63]. 20-ns [Sei67]. 21 [Ben69a, FK65, Kat68, Sal63]. 22 [Els69, Gar65, Mur68]. 23 [Gre69c, Kuc68b, Sam65]. 24 [Hop67a, Muk69a, Wai68b, Wir65]. 25 [Bob65, Fre68, Kop69, Mcc67a, New66a]. 26 [Bar68, Bro69a, Can67a, Her65, Ros66b]. 27 [Aho68, Cof69b, Mer65, Mur65c, Rid66, Rub67]. 28 [Den69c, Fog67, McK68]. 29 [Dar68, Fir65, Kor67a].

3 [Kli69]. 30 [Alo64, Bek67, Col69b, Lew65b, Lit68, McK68]. 300 [Car64, Coo64b]. 300-Mc [Car64, Coo64b]. 31 [Cro64, Fis63, Gle68, Nil69, Vic67a]. 32 [Gar69, Kar63a, Nol64, Ser68a]. 33 [Den69a, Kir64b, Lew63a, Ser68b]. 34 [Gil63b, Kle64, Sil68a]. 35 [Cop68, Els64, Lev63b, Met65a]. 36 [Fra64b, Mcc68b, Tru63]. 360 [Gea65, Nie68, Wie68]. 360/67 [Wie68]. 37 [Cra64, Flo68, Pre63]. 38 [Bin68, Gre63, Les64]. 39 [Hei63, Rog66, Rub68, Wag67].

4 [Kar69, Ulb63]. 4-79 [ROK58]. 40

[Bol67, Car68, Fit63]. 41

[Kor65, Leo63b, Luk63, Mas67, Sal68]. 4102 [Don65b]. 4102-S [Don65b]. 42 [CW68, Rub65, Sze67]. 43 [Len65, Str68a, Wit67]. 44 [Gas65]. 440 [Fra64b]. 45 [Far68, Mei65]. 46 [Ear68]. 47 [Fif68b]. 48 [Hen64b, New68c]. 49 [Boo64, McG68].

5 [Epl68, Kle69]. 50

[Ama64b, Jac68, Sto66a]. 5000 [Lew65a]. 51 [Han68, Sho64, Sto66b]. 52 [Lec64, Rap66, Wel68]. 53 [Ols65, Sut68]. 54 [Ham65a]. 55 [Ard63, Mcg65]. 56 [Coa63, Gea65]. 57 [Con67a, Hum63, Sch65b]. 58 [Hig65, Pau67]. 59 [Kuc66, Tam67].

6 [Sas69, Str68b]. 60 [Gas66, Vid67]. 61 [Ble64, How67, Sag66]. 62 [Con67b, New66b, Sto64]. 63 [Gra64a, KF66, Kor67b]. 64 [Ble66, Mac64]. 65 [Kud64]. 66 [Mei66, New64]. 67 [Kov66, Wie68]. 68 [Str66, Wit63]. 68-44 [de 68b]. 69 [Dun65, Isb63, Lar66a].

7 [Bie68, O’N69]. 70 [Fir63b, Nad65]. 704 [Dom57]. 705 [Mer56]. 71 [Pet63]. 72 [Ihm63]. 73 [Gin63]. 74 [Els63b]. 75 [Man63b]. 76 [Nel64]. 77 [Fis64]. 78 [Has66, Sri64a]. 79 [Kru64, Sci66].

8 [Gre69a, Kei68, Law67, VB69, VBW69]. 80 [Gab64b, Gre66a, Knu65a]. 81 [Car64, Lew65a, Skr66]. 82 [Cly65, Miu66]. 825 [Bas63b]. 83 [Lew66a, Mar65a]. 84 [Gli66b]. 85 [Van66]. 86 [Fis66]. 87 [Bit66]. 88 [Bro63b, Can66]. 89 [Fel63].

9 [Bow67, Har69b, Tho68b]. 90 [Ste63a]. 91 [Hig63b]. 92 [Ano63-40]. 93 [And63]. 94 [Kel63].

A. [Nol62, Shu62]. Abbreviated [CM65, CD66b]. Abort [Mar68]. Absolute [Abh59]. Absorption [Dow58]. Abstract [Gin62, Wak69]. Abstracts [Ano59b, Ano59c, Ano59d, Ano59e, Ano59a, Ano60a, Ano60b, Ano60c, Ano60d, Ano60-31, Ano61a, Ano61b, Ano61c, Ano61d, Ano62c, Ano62a, Ano62b, Ano62d, Ano62e, Ano62f, Ano63a, Ano63b, Ano63c, Ano63d, Ano63e, Ano63f, Ano63g, Ano64a, Ano64b, Ano64c, Ano64d, Ano64e, Ano64f, Ano65a, Ano65b, Ano65c, Ano65d, Ano65e, Ano65f, Ano66a, Ano66b, Ano66c, Ano66d, Ano66e, Ano66f, Ano67b, Ano67c, Ano67d, Ano67e, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano69c, Ano69d, Ano69e, Ano69f, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k, Ano69n, Ano69l, Ano69m, Brz63, Bur63, Cho66, Coa64, Fir64a, Fre67b, Gan64, Gau60, Gil66b, Har63b, Har67b, Her67, Hoa64]. Abstracts [Kau64, Kau66b, Kaz65, Kim67, Nas67, Nic65a, Pre64a, Rob66, Sha65, She63b, Sho66, Top67, Ung63a, Wil65a]. Ac [ZP66, KL58, Pic66b]. Access [AH62, CM68, Hoa62, Loe64, Til60, Wij68]. Accumulation [WHB56]. Accumulative [Bri67]. Accumulator [BC65, Loo66, Luc69a, QK67]. Accuracy [Bar68, Cah53, Fuc57, Gre66b, HS64a, Hen60, Hol63c, KS67, MC62, Nat57, PD62, Sch58, Ver64]. Accurate [Bin68, FS59, KS61]. Achieve [NP64]. Achieving [Gil64b]. Acoustic [LBB66, RG65, Rub67]. Acquisition [Eps68]. Action [Wak69]. Active [Fit63]. Activities [Goo55]. Activity [Die68, Svo63b]. Acyclic [Epl68, ME67a]. ADAC [Vid67]. ADAC-A [Vid67]. Adaptation [May64]. Adaptive [Ama67, Bar66, CP68, Coo64a, Coo66, Gab67, Glu66, IT67, IS65, KS63, KT65, Lew65b, Lew66b, May64, Rub65, RV68, SH68, SW65, Sto63, WF68, Wee68, Whi69c]. Adaptive-Element [Whi69c]. Added [Bon66]. Adder [Ale67, Che58, Dav63b, HS63b, Lin66, MW67, Men61a, Men61b, RR62a, RR62b, Sal60, Squ62, Svo69, WS56, Bed62]. Adders [Amo64, Edw61, LB60, Maj67, Skl60b, Skl63b, Wel69]. Addition [Bri67, Fen69, Hen60, KLP68, Rei60, Sas67, Skl60a, YY66b]. Additive [Sas68]. Address [Chu65b, Els64]. Addressable [Cho67, Han66]. Addressed [CS66, Rux69, SH66c, SH67b]. Addressing [Hol61a, Len65]. Adjustable [May65b]. Aeroplan [Lan63b]. Aerospace [Jos67, MMB66b, Pas67a]. Affecting [Jon69b]. Affiliate [Bak57]. Affine [SJ69]. AFIPS [Ano62g, Den63, Lo63, Rab63, Ung63b]. Ahead [SG69]. Aid [ALS68, Bow67, GHR67, WB62]. Aided [Che68b, Hay69b, Mck69, Sch64c, Yen69]. Aids [McG69, Ros61, Ung65]. Air [AW56, Hea56, Kir64b, Nie56, PVW59]. Aircraft [Kor65]. Aizerman [Auf62]. Aktiebolaget [Lan63b]. ALERT [FY69a, FY69b]. Algebra [Hol61b, Lin60, Mul54]. Algebraic [Ake61a, AH63, Bas64, Bou68, GM65, Han63c, Hel66b, KF66, Kei64, Liu66a, ML61, Mul59a, Rao58, Nel54]. Algirdas [Met65a]. ALGOL [Her65, Wai68b, WS67a, Woo64c]. Algorithm [Ake67, And68, Arm62b, Cha65b, DM68, Dav69a, DS65, Ein61, Ein62, FL67, Gab65, Har57, Has69, HK65, Hoc62, IT67, Kri64, Lam64, Law67, Lee61, Liu64b, Mac62, Mac69, MS69, Min62, Nec68b, Nec68a, Pra69, Ris67, SD68, SE66, SC68, Sto66a, Sus61, Svo63a, Tor68, Tun68, Wan66, WL61, Win68b]. Algorithms [MFF63, May65b, Mor67, Ros65b, RBS67, SSG69, Spe65]. All-Digital [JSS66]. All-Magnetic [BCE61, Cra60b, CV61, Cra61]. All-Pass [HF66]. Allocation [Chu69a, Den69c, Den67, HM67a, Kop69, Ros67b, SSG69, Wil65b]. Alloy [Tam67]. Almost [Fie64]. Alpha [Win62]. Alpha-State [Win62]. Alterable [Far68, Gra60b, Til60]. Among [Lec63b, dL69]. Amplification [BW63, Pic66b, Tsu62]. Amplifier [Agg64, Agg65, CKLP60, Gil69, HL57, Han62, KH63, Kin66a, Kon60, Kur64, Nis60, Sei64, Tsu62, Wad64, Wad68]. Amplifiers [CM62b, CM63, GK62, Gra58, Hel55, How56, Kor62a, Kor63a, KL58, MC62, Pfe59, RR66]. Amplitude [CKLP60, DH69, Gil66g, Goo64, Kae62, Ort59, Roy61, ZP66]. Amplitude-Distribution [CKLP60]. Analog [AM67a, AM67b, And62, And63, Ano62-32, Ano62-38, Ano65-32, Ano68o, Azg68, Bar68, BBP57, BW61, BW62, Bek59, Bek65, Bek67, Ber60b, Bie68, BH65b, Bin68, Bir67, Bla62, Boh60, Bri63, Bri66, Bru57, BO59, Cah53, CK65b, Che61a, Cly59, Con62, CM61, Con67b, Cox61, Cro68, Cun54, Del62, Dia56, DW67, Eck62, Fel65, Fog67, FP54, Fuc56, Gas65, Gau64, Gel62, GM65, Gon65, GRT55, GDD62, Gre68b, HKM63, Hau65a, Hau65c, Hen62, Hig61a, Hol63c, How67, HG57, Kae61, Kah68, Kal57, Kas62, Kel62, Kel63, Ker56, KS61, KII62, Knu65a, Kon68, Kor60, Kor63b, KC54, KC60, Kov66, KM61, Lar69, LC59, Lip55, MT68, Mar56b, Mar67a, MMR60, Mar66b, Mas67, MN58, MR59, Mye57, Nad63, Nag63, Nat57, NM57, Nel63, Nel64, NW64, O’G67, O’N67b]. Analog [Ott64, Pic65, Pic66b, Pre55, Ran64, Rat67, Raw59, Rid66, RM65, Rob67, Rob55b, Rub62a, Rub62b, RG65, San64, Sas57, Sav59, Sca53, Sch63a, Sco57, Ses63, She64a, She63a, Shu62, Smi65a, SH68, Ste63c, SD57, SS68, Swi63, SOS56, Tal64, Tav65, Ter62, Tho60, Tho67, Tho62, Tom57, Tru62, Tru63, Van57, Ver64, Vid67, VS62a, VS62b, WW55, Wai66, Wat59, Wei57, Whi65, Whi66, Wie65a, Wil63, Wit63, Wit67, dM66, Wal53]. Analog-Computer [Eck62, Hen62, Kas62, Kor60, Ses63]. Analog-Digital [Con62, Cox61, HKM63, LC59, Nad63, Nel64, Shu62, Tru62, VS62a, VS62b]. Analog-Hybrid [Ano62-38, Tav65]. Analog-To-Digital [KII62, Che61a, Fel65, Kae61, Lip55, Sca53, Ver64]. Analog-Type [Sch63a]. Analog/Hybrid [How67, Kor63b]. Analogs [ACGN60, Bal62a]. Analogue [Gli66b, Kar69, Tam67]. Analogy [Bla56b]. Analysis [Ama64b, Ara57, AH57, Auf58, BH67, Ber66a, BB61, Bie68, BW63, Cam57, Cho63a, Cla56, Dow57, Dow58, ENB61, FDBW63, Gab65, Gam68, GMB61, Gil64a, GA58, Gil66e, GR67, Hol63a, Jac68, Kin66b, Kle64, Kno67b, Kos62, Lan68, Lar66a, Lyn69a, Mar56b, Mil58, Miu66, Mow69, Nar60, Nis60, Ott58, Pau67, Sam69a, Sam69b, Sch64c, Sch69c, Sed68, Ses63, SS66, SG69, Slo62, Str66, Str68b, Ulz63, Van59, Vic67b, VB65, Wee69, Wie65b, WK68, Wom65, Woo63b, Yao63]. Analytical [MS58]. Analyzer [Far68, Gar69, Gil59b, Gre56, HH53, KS68, Lam64, MPP59, Pet61, Skr66, TP60]. Analyzers [Dow57, Dow58, Eck65, Fuc57, Gil59c, HP64b, KS67, Knu65b, Nat57, Nat58, Woo65]. Analyzing [Dor69, SHB68, SHB69]. And-Invert [Hel63]. Anderson [Ano62-43, Low62]. Angle [Bro60b, Gra60a, WP62]. Anneals [ROK58]. Annotated [Bre65a, SS61, SS62, Tod63]. Announcement [Ano68m]. Annual [Ano65s, Ano66p, Ano66-33, Ash66, BH65a, Cho65a, Har66a, Mei67b, Pre66a, Sac65, She66, Win67b]. Answering [Col69b]. Antiparallel [GS65]. APACHE [GDD62, Wit63]. APACHE-A [GDD62]. APE [How67]. Apictorial [Fis64, FG64].

Apollo [Mar68, PMRC67, Sil68a]. Appearing [MP64, MP65]. Application [Cha59b, Els64, GS58, Gra64a, Gra64c, JO68a, JO68b, Kel63, Knu65a, KK68, Kuc68a, LK60, Mar69b, Mul54, Nar60, RR66, Sta65, Tho67, Tis67]. Applications [And53, BBP57, Brz62, Cha59a, Cov65, Den69a, Dia56, DS67b, FF63a, Gaf68, GH59, HK65, Kam60a, KB64, Law67, Lee61, OKL59, PC56, SSPR60, Smi68c, SP63, Tom58, Wei52, YI68a, YI68b]. Applied [Lar69, Mii61, Raw59]. Apply [GR68]. Applying [Dei65]. Appreciation [Tan65]. Approach [Ake65b, CP68, Dav67, Der64, Dor69, Fil67, FCC67, HS69c, MF68a, SM69, SP62, Tal64, Tor66, Wee68, Who63, Yoe65]. Approaches [Ott58]. Approximants [HF66]. Approximate [Nec68b]. Approximately [Her66]. Approximation [Coo64a, Gil59b, KW65, McC68a, Vic69a, YS68, Zac61]. Approximations [Elf62, Gon65, Kor62b, McA68, Rid66, Win69]. Arabic [Kam61]. Arbeitsweise [Ulb63]. Arbitrarily [Gim65a]. Arbitrary [AM67b, BH65b, Cop68, Cro68, Fre61b, Miy63, Miy64a, Sut69, Urb65a, Wei69b, Woo65, YI68a]. Arbitrary-Wall [Sut69]. Architecture [Kat68]. Argument [Win65a, Win67a]. Arithmetic [AS66, Avi61, Bro60c, Cro61, DK63, FP62, Gar66, Gre66b, Ham65a, HS69b, LB61, Leh63a, Luc59, Man67a, McK67, MA58, MA63b, Rao68, RM65, Roz62, SA64, Svo69, Tun68, Met65a]. Arithmetic-Type [Rao68]. Arm [FS65]. Armstrong [Nic63]. Arnold [Ano60e]. Arrangement [Cro60]. Array [Ale67, Kau67c, PGB64]. Arrays [CP66, Col69a, KLW68, Kau69b, Wee63, Wee69, Wei69b]. Art [Hob66, Tau62]. Aside [Lee69]. Aspects [ABSH63, Bau58, Mil67, Sam65, Vic67b, Wel64]. Assemblies [Law62a]. Assertion [Lev63c]. Assessment [Kau66c]. Assign [Bob65]. Assigning [Arm62b]. Assignment [AYT68, Arm62a, Arm62c, Bow69, Dav67, ET63, Har61b, Hum63, JO68a, JO68b, Kar64a, Kle64, Koh64b, Nic63, PMC67, SH61, Ung68]. Assignments [BF61, FY66, FGU69, Har68e, MU59, NB65, Rus65, Tor68, Tra66, WS67b]. Assimilations [Leh63a]. Associate [Ano64u, Ano65-32, Pra66a]. Associations [Pet63]. Associative [Ahr65a, Chu65a, Cra68, Fos68a, Han66, HH68, Hil66, Mii64, NT69, Ols65, RC64, Ros66b, Wol69, YY66a, YY66b, YY66c]. Assumed [Ben69a, Gre69a]. Assuming [AFM69]. Assumptions [FT68, Lan68]. Astrac [Nel64]. Astronomical [Lew63b]. Asymmetric [Gre68a, McC68a, Wol62]. Asymmetry [Mur65a]. Asynchronous [Ano69-53, AFM68, AFM69, BS68b, Cat66, Fri66c, FM68, FGU69, Gra62b, Haz65, Lan68, Lan69c, Ler65b, MTS69, MK69, PW67, Ree65, Rob56, Sau67, SF62, Sin69, Stu68a, TMF69, Tra66, Wai64b]. Atmospheric [Ses63]. Attempt [Har68d]. Attenuationless [Ano69o]. Augmented [Lin60]. Author [And66, Ano59f, Ano59g, Ano59h, Ano60f, Ano60g, Ano60h, Ano61e, Ano61f, Ano61g, Ano62h, Ano62i, Ano62j, Ano62k, Ano64m, Ano65g, Ano68n, Bra64, Brz67, Cho64, Cho65a, Con68, Cow63, DM69, Deo69, Der66, Ein62, Eve66, Gea63, Gli66a, Hay69a, Kau69a, Knu67, Koh65a, Kru63, Kuo69, Len63c, Mai63, Mar67b, Mat65, Mei67b, Mik65a, Mil65, MK68, O’D64, Pre66a, Ren69a, Ren69b, Ren64, RH67c, Spa66, Sto67, Tab67, Wai64a, Wal67]. Authors [Ano59q, Ano60p, Ano61q, Ano61r, Ano63r, Ano63s, Ano63t, Ano63u, Ano63v, Ano63w, Ano64n, Ano64o, Ano64p, Ano64q, Ano64r, Ano64s, Ano65z, Ano65-27, Ano65-28, Ano65-29, Ano65-30, Ano65-31, Ano66w, Ano66x, Ano66y, Ano66z, Ano66-27, Ano66-28, Ano67g, Ano67h, Ano67i, Ano67j, Ano67k, Ano67l, Ano68-29, Ano68-30, Ano68-31, Ano68-32, Ano68-33, Ano68-34, Ano68-35, Ano68-36, Ano68-37, Ano68-38, Ano68-39, Ano69-41, Ano69-42, Ano69-43, Ano69-44, Ano69-45, Ano69-46, Ano69-47, Ano69-48, Ano69-49, Ano69-50, Ano69-51, Ano69-52]. Autocorrelation [Dia56]. Autocorrelations [EG61]. Automata [Aho68, Ano66-30, Car68, Coh64, Cur68b, Cut68, Eve65b, Ghi62, Gil60a, Gin66, Gre69b, Har69b, Kli67, MY60, Mez61, Nie69a, PRS63, Sta65, SO68]. Automated [HK69, MA63a]. Automatic [Bas63b, Bla62, Bla56d, Ble64, Bob65, Col69b, ET63, FF63a, FY69a, FY69b, Han68, Hig63b, Hig65, Kel63, Kle53, LS60, Lee62b, LBB66, Nec68b, PMRC67, Per63, Rub67, SD63, Sch65a, She64a, Ste63c, SS68, Sus61, Tav65, WHB56]. Automation [DS69, Ger68, Wat66]. Automaton [Auf62, Liu63a, Liu63b, Liu64a, MW64, Win62, Ric67]. Autonomous [BD64, CN68, Fuk64, Har62, Kin66b, Pre64b, Pre65, YIK68, Yau65]. Availability [Lei67]. Available [Ano68m]. Avalanche [Ham60]. Average [Hel66a, MP64, MP65, Sob62]. Avisienis [Met65a]. Awards [Ano56l]. Axiom [Ano68-45]. Axiomatic [CL61, Lar61]. Axon [Fit63].

B [Mac62, Mad62, Nol62]. B69 [Min69]. B69-1 [Min69]. Background [Cof69a]. Balanced [BW63, Den69b, Lan63a, NP64]. Balanced-Pair [BW63]. Band [Lar66b, Ort59, Sze67, Tam67]. Band-Pass [Lar66b, Sze67]. Bandwidth [Dow57]. Base [AS66, CVV65, Has66, SA64, Son63, Wad57, Wad61]. Based [Ano66q, Ano68-46, Che61b, Sha66a, Sin62, VBW69, Win69]. Basic [Ata64, Ble61a, Fos68b, FF63b, Kli67, MA63b, MK67]. Basis [Sas68]. Batch [Gra69]. Batch-Fabricated [Gra69]. Bayes [Coo63a, FK69]. Bayesian [SSG69]. be [Uhr61]. Beam [Cly59, Gre62, Kle53, TBS64]. Beam-Vibration [Gre62]. Beat [Hor69].

Beck [Alo62]. Becker [Rub62b]. Behavior [Ano56r, Bro69a, VBW69]. Bell [Rap66]. Beneath [FR58]. Bessel [BH65b, Bin68]. Best [Gir66, Gur65, Sho64]. Between [Cof69a, Lec67, Mow66, RBS67]. Bias [CM62a]. Bias-Controlled [CM62a]. Bibliographical [BCE61]. Bibliographies [Ano58r]. Bibliography [Bre65a, Hol61b, Mor59, Net58, Net59a, Rot61, SS61, SS62, Tod63]. BIDEC [Cou58]. Bidirectional [FS59, Gol57]. Bilateral [AE61, Mil58]. Bilogic [BE69]. Binary [Bri61, Bro55b, Bro60c, CP66, Col67, CVV65, Cou58, Cro61, Dag59, Dou68, DR69, Fen69, Fia66, Fre67a, Gam68, GM63a, Ger60, Gil60a, Gol57, HM65a, Has66, Hen60, Hop64, Kau62, KLP68, Kri63, KO66, LB61, Leh63a, Lin66, Luc59, Mac62, MW67, Mar63a, McK67, Met65a, Met62, Mit62b, Mur55, New67, New68b, New68a, O’B57, Oli65, Per68, Pre66b, Pre68, RR62a, RY66, Rei60, RR62b, RM65, Rob55a, Rob56, Rot59, Roz62, Sal61, Sch68, Skl60b, Tom56, Van59, Van60a, Wan66, Wel69, WL61, Yoe63b]. Binary-Compatible [Met65a]. Binary-Decimal [O’B57, Tom56]. Binary-to-Decimal [Cou58, Cro61]. Biological [Del62, McC57, Qua57]. Biomedical [Ble66]. Bistable [Dav57a, Eck62]. Bistables [Brc65]. Bit [Amo66, Cho67, Fra69, Gab64b, Gil64b, Lip64, Mei65, Mil55, NP56, Pas67b, SS59]. Bit-Organized [Pas67b]. Bit-Reversed [Fra69]. Bit-Steering [Cho67]. Bits [BBB59, Fre67a]. Bledsoe [HK60, Ste62]. Block [Cam69, Cly65, DKP69, Kli67, May60, Sin62, Tav65]. Block-Diagram [DKP69]. Block-Operators [Tav65]. Block-Oriented [Cam69]. Blocking [GL67]. Blocks [CW69, Har66c, Whi69c]. Bloom [Hoa62]. Bluffing [Fin61]. Board [Koh64a]. Body [She64a]. BOMARC [LW59]. Book [Str62]. Books [Ahr65b, Ake65a, Ake66, And67a, Ano59v, Ano59w, Ano60x, Ano60y, Ano60z, Ano60-27, Ano61-28, Ano61-29, Ano61-30, Ano61-31, Ano62-33, Ano62-34, Ano62-35, Ano62-36, Ano62-37, Ano68-42, Ash65, Avi63, Axe63, Can67b, Car65, Cha67, Epl66, Fir64b, Gil65a, Gol63, Hop67b, Jon64, Lev63a, Man67b, Mil66, Min63, Ros64, Sci64, Ses65a, Son64, Sri64b, Tai63, Tra67, Urb66, War66, Yau66]. Boolean [CD66a, Ein62, Min62, SHB69, Abh58, Abh59, AH63, Bro63a, Cam57, Dan66, DS67a, DD69, Dvo68, Ein61, EG61, Els60, ED67, EM69, Gab62, GBC69, Gim65a, Har63c, Hru69, JA67, Kri66, Led59a, Led59b, Lin60, Lor64, MP64, MP65, MN61, MC64, Mul54, Nec67, Nec68b, Pre66b, Rou58, SHB68, Set66b, She65a, She65d, She65e, Svo63a, Tis67, UM56, Urb63, Van60a, Wan67]. Bordered [Set67]. Both [Sal68]. Bound [ES64, Gil65b, Lew66b, MT66, Vai69, YI65]. Boundaries [Cad68, Gre68a]. Boundary [Coo66, May65a, Ter62]. Bounded [Ard65, Wak69]. Bounds [BE69, Cof68, Die68, Mur65b, Smi66, Win63]. Box [Eve63, Sin66]. Boxcar [HS64a]. Braid [AA66]. Brain [Ano56r]. Breakthrough [GDD62, Wit63]. Brigham [Kau62]. Browning [HK60, Ste62]. Buchholz [Str62]. Buffer [Dor67]. Build [Ham60]. Build-Up [Ham60]. Building [CW69, Har66c, Whi69c]. Built [Cow64]. Built-In [Cow64]. Bulk [CG65, Wie68]. Bundle [Rub68]. Burroughs [VB69]. Bylaws [Ano65r].

C [Kau62, Low62, Mad62, Mit62a, Gil69, Ano69b]. C-18 [Ano69b]. C. [Mad62, Rub62b]. Calculation [Fre67a, FK69, Kah68, RW67, Sto66a, Whi69a]. Calculations [Ahr65a]. Calculus [Val67]. Call [Ano63h, Ano63i, Ano63j, Ano65h]. Canonical [Bro69b, Coh62, Lec63b, Mor67, SKS65, SK65]. Capabilities [KL67, WP64]. Capacitive [MY61, SH68]. Capacitor [Ano68o, Dow58]. Capacitor-IGFET [Ano68o]. Capacitors [PM58]. Capacity [BHRS61, FS65, Lar69]. Captive [Bra63]. Card [BHRS61, BP59, Lar69, LBG65, Hoa62]. Card-Changeable [BHRS61]. Card-to-Tape [BP59]. Cardinality [Men65]. Cards [YR65b]. Cares [SH67a]. Carlo [Han67, Kor67a, Lit68, Van66]. Carriers [Wei57]. Carry [AS66, Bed62, Fer68, GPW55, Hen64b, HS63b, LB60, LB61, Leh63a, Maj67, MW67, Rei60, Wel69]. Carry-Dependent [HS63b]. Carry-Propagation [AS66, Fer68, LB61]. Carry-select [Bed62]. Carry-Storage [Leh63a]. Cascade [Cut68, Dvo68, Nat65, NM68, Sho64, Wei69a, Yoe61, Yoe63a]. Cascade-Parallel [Yoe63a]. Cascaded [Cad68, Gil61a, Mai62, Mar63a]. Cascades [Fan68, SK65, Yoe65, Yoe68]. Case [ABSH63, BdT65, Lyn69a, PLS61, Wra62]. Casting [Ger60]. Catalog [Hel63, YR65b]. Categories [Fro65]. Category [Coo63a, Smi69]. Cathode [DG61]. Caution [Kor62b]. Cell [GL65a]. Cells [Kil62, Mai62]. Cellular [FG69, Kau67c, KLW68, Kau69b, Min64, Muk69b, SK65, YY66a, Yoe68]. Cent [Dei65, Sch58]. Central [Cra64]. Certain [Dow57, Dow58, Har68b, Mau67, Mul59d, Mul60]. Chains [GO63, Kar63a, LE57, Mar62, SK63]. Chairman [Ano62-43, Ano58a, Ano59i, Ano59j, Ano59k, Ano60e, Ano64t, End59, Unc64]. Change [Bla56c]. Changeable [BHRS61]. Changes [FM68]. Changing [She64a]. Channel [Eps68, Hsi68, SC67, SH66b, Wol62]. Character [Cho57, Cho59, Cho63b, FR58, Flo58, FG60, Flo62, Fre61a, HK60, Hig61a, Hig63a, Liu64b, Nad63, Per63, Ros62]. Character-Recognition [Fre61a]. Characteristic [Kno69]. Characteristics [BS63, Bro66, Dav57a, Dun65, DR69, Eas58, Gel62, Kau67b, LM62, Ols65, OKL59, ROK58, TNM64, TF64]. Characterization [Gre68b, Men65, Wei69a]. Characterizing [Gil60a, Lev68, Mez61, Yen68a]. Characters [Bro64, CN66, GHR67, Kno69, Mer65, MW68]. Charge [DS67b, HM67b, Kra67, MWR68]. Charge-Storage [MWR68]. Chart [CB62]. Charts [Muk63]. Chebyshev [HF66, KW65]. Checking [AW56, BC66, Gar58, Kau58, Kim66b, Lee62b, Rao68, Roy68b]. Checkout [Bed64, PMRC67]. Checkout-Status [Bed64]. Chess [Fis66, Kuc68b]. Chief [Ano61s, Ano65-33]. Chinese [CN66, GHR67]. Chip [Luc69a]. Chow [Win69]. Ci [Lew65a]. Ci-5000 [Lew65a]. Circle [Con67a]. Circuit [ALS68, Ano66-33, AS66, Bas64, BG56b, BB61, BW63, BM63, Brz65b, Cha59a, Cra64, CV61, Cur68b, EM66, FF63b, GMB61, Gin63, Gol60, Gon63, Hal69, Har66a, HS64a, Kur64, Led59a, Lo61, MN61, Nar60, Nie56, PLS61, Sed68, Sil68b, Tam67, Tho67, Tsu62, Wei63a, Wel69, Whi69a, WP62, EW53, Mul54]. Circuitry [Cho60, CS67, Har63a, Har58, LK60, PLS61, Wan55, WS56, Yao65]. Circuits [Agg65, AFM68, AFM69, BGR66, Bal62a, Ber60a, Ber58, BF61, BS59, BdT65, BB56b, BM64, Brz64b, Brz65a, BS68b, BvB65, CM62a, Car64, Cha60, CS64, Cho63a, Chu69b, Col61, Coo64b, Cos69, Cra60b, DM64, Dom57, Dwy69, Eas58, Eps58, FP62, Fri66c, Fri66d, Fri67a, FM68, FK66, GJN59, GM65, Gil64a, Gil66c, Gol64, GR68, GM62, GMN^{+}60, Gra64b, Hay61, Haz65, Hel63, Hen64b, Hol63b, Hol61b, HS64b, Jar63, Kau68, Kim66b, Lan68, Lan69c, LS60, Law61, Law62b, Ler65b, MTS69, Man67a, Man64b, MS58, Mas60, MS68, MRS60, Mau67, MU59, McG69, MK69, MWR68, Mor59, Muk69a, Mul56, Mul59a, NM68, NBC60, NP64, Pie63a, Pie63c, Por64, Pra65, Pra66b, PC56, RC57, Ree65, RY66, Ros58, RBS67, San64, SI59, SM69, Sch63b, Seb57]. Circuits [Sei64, Sho64, Sin69, Smi65b, SD69, Sus61, TMF69, Tho55, Tor68, Uzu62, WR63, Wan65, War58a, Woo64b, Wra62, Yen68a, Yen69, YR65a, YBS61, HS69a]. Circular [Pie63b]. Circulating [Lei57]. CIRRUS [APP^{+}63]. Clamped [MRS60]. Class [Bir67, CB63, Ito69, Mar66b, MFF63, Met62, Rob58, SP69, Spe65, Tal64, Whi66, Wit67, YO68]. Classes [Ano69-54, Cla67, Har68b, Har63c, Lec64, Lec67, Luc69b, PS66, Pre65]. Classical [WD68]. Classification [Ano69p, DF66, GHR67, HF69, Kno67a, SW65, WF68, Wee68, Wol69]. Classifier [Glu65, Smi68b, Smi69]. Classifiers [Ama67, Smi69, YS68]. Classifying [Fro65]. Clipping [Tsu62]. Clock [GJN59]. Clocking [Gab67]. Clocks [Har62, Yau65]. Closed [DBC69, Gra66, Set66b, Tei55]. Closed-Loop [Tei55]. Cluster [Ihm63, PF69]. Clustering [LLT69]. Cobalt [MS66]. Cobalt-Substituted [MS66]. COBLOC [JSS66]. Cockpit [GNV59]. Code [Bey63, Bey64, Bro65a, Che68c, Chu65b, CE69b, Fis57, HW69, Lip55, Luc59, O’B57, Skr66, Wee60]. Coded [LC53]. Codes [AM67b, Arm62a, Arm62b, Bro60c, Chi59, Cla67, Cro68, Dav67, Eve63, Fia66, Gar59b, Gar66, Gol64, Hsi62, Hum63, Kau58, KT65, Lec67, Lip64, Neu62a, Oli65, SH66b, Sin66, Tom56]. Coding [DM64, Els63b, Els64, Gab67, LE57, Smi65a, Wol62, Zah66]. Coefficient [McC68a]. Coefficients [HH53]. Cognitive [SS61, SS62]. Cohen [Ano60e]. Cohn [Miy64b]. Coincidence [Bla56a, SC60]. Coincident [BB56a, BAL61, Lan64a, LM62, Wom65]. Coincident-Current [BB56a, BAL61, LM62, Wom65]. Cold [DG61]. Cold-Cathode [DG61]. Column [Ano58a, Ano59i, Ano59j, Ano59k, End59]. Columns [Luc66b]. Combination [Kno68, Mas57]. Combinational [Ake61b, AC67, Arm66, Cos63, DM68, Dwy69, Ell65, Has69, HS69a, Kau68, Kil59, Kil62, McC58, McG69, Mul59c, NB65, SD68, Smi68a]. Combinations [Fis66, Mar56a]. Combinatorial [Sam65, Wel64]. Combined [Shu62]. Comcor [Lew65a]. Command [Ano66-35, Sam69b]. Comment [Ano69p, Azg68, Bar63, Bir67, Bro69b, CD66a, Dav69b, Dou68, FF67, Fri67b, Gir69, HH62, Jon69a, Mck69, Mis66, Miy64a, Ogb69, Ott64, Pow65, SH67b, Tae69, Tho69, Vic69b]. Comments [Bri63, Bri66, Cho59, DM68, Dwy69, Eve65a, Fos68b, Has69, HK60, Hig61b, Kno67a, Lew65b, Mac69, Min62, Nat64, Nic63, Rib64, Ros68, Roy68a, SBM63, Sur68, VS62a, VS62b, Whi69b, Ein62]. Common [Sch65b]. Communication [AJ60, Ade62, Ros65a, SC67, Ade62]. Communications [Ano66q, Fra64a, Mcg65, Sha66a, WP64]. Communications-Based [Ano66q, Sha66a]. Commutation [Tom56]. Comparative [MRS60]. Comparator [Far69, Goo64, HS64a, Kae62, Kon68, MC65b, Roy61]. Comparators [Amo64, Mii61]. Compare [Wei63a]. Comparison [Gol60, Gre62, HP64b, Man67a, SW65, Whi65]. Compatibility [Ano69-54, Gra64b, Luc69b, Pie63a]. Compatible [Ano68-43, FK66, Met65a]. Compensating [Con67a]. Compensation [Bri67, Gle68, SB61]. Compilation [Kel63]. Compiling [Ear68]. Complement [Fia66, Gar59b, Rob55a]. Complementary [Els60, MF68b, Sho62]. Complementary-Output [Sho62]. Complementation [Lan69a, Lan69b]. Complete [Ano68-45, KO66, LE57, LW65, PS68, Set67]. Completely [BS68a, Bor69, YI68a, YI68b]. Completeness [Loo65]. Completion [MW67, Wai64b]. Complex [Gra62b, Hau65a, Kov66, Sto66a].

Complexes [Gre66a]. Complexity [Fra58, Mul56, Qua57]. Component [BW63, Hea56, KC60, MR68]. Components [Bru57, LC63b, Lo61, SS59]. Composed [Kil62, Van60b, VO65]. Composition [BHNW57]. Compositions [Kei64]. Compound [She65a]. Compressed [Gre66a]. Computable [Yam62b]. Computation [Bar68, Bic60, Bla56b, Boo64, Col69a, Col67, CVV65, Con62, CM61, Con67b, DW67, Dou68, Els69, Gle68, Gre63, Has66, Hau65c, Kor65, Mar68, O’N67b, Rat67, Raw59, Sch63a, SM69, Sha69, Ter62, Tom57, VS62a, VS62b, Wag67, Whi65, Yam62b]. Computational [Epl68, Gre66b, HP64b, Kar63a, Mar62, ME67a, Nie69b]. Computations [AK69, BE69, ET63, Hen66, Hol63c, KS68, ME67b, ME69, Sto66b]. Compute [RBS67]. Computer [Abe69, Ahr65b, Ake65a, Ake66, APP^{+}63, ABSH63, Alr55, And53, And62, And63, And64, And67a, And68, Ano57k, Ano57p, Ano59b, Ano59c, Ano59d, Ano59e, Ano59a, Ano59v, Ano59w, Ano60a, Ano60b, Ano60c, Ano60d, Ano60x, Ano60y, Ano60z, Ano60-27, Ano60-31, Ano61a, Ano61b, Ano61c, Ano61d, Ano61-28, Ano61-29, Ano61-30, Ano61-31, Ano62c, Ano62a, Ano62b, Ano62d, Ano62e, Ano62f, Ano62-33, Ano62-34, Ano62-35, Ano62-36, Ano62-37, Ano62-38, Ano63a, Ano63b, Ano63c, Ano63d, Ano63e, Ano63f, Ano63g, Ano64a, Ano64b, Ano64c, Ano64d, Ano64e, Ano64f, Ano64g, Ano64h, Ano64i, Ano64j, Ano64t, Ano64-27, Ano65a, Ano65b, Ano65c, Ano65d, Ano65e, Ano65f, Ano65i, Ano65j, Ano65k, Ano65l, Ano65m, Ano65n, Ano65r, Ano66a, Ano66b, Ano66c]. Computer [Ano66d, Ano66e, Ano66f, Ano66g, Ano66h, Ano66i, Ano66j, Ano66o, Ano66-34, Ano67b, Ano67c, Ano67d, Ano67e, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano68-42, Ano69c, Ano69d, Ano69e, Ano69f, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k, Ano69n, Ano69l, Ano69m, Ano69a, Ash65, Avi63, Axe63, BFL^{+}66b, BBK^{+}68, Bar61, BC65, BBP57, Bau64, BW61, BW62, Bek59, Bek67, Ber66b, Ber60b, BH65b, Bin68, Bit66, Bla62, Bob65, Boh60, BB56b, Bri63, Bro55a, BTW61, BJST64, Bro64, Bro65b, Bru57, Brz63, Buc62, Bur63, Cah53, Cam69, Can66, Can67b, CET62, Car65, CS58, CK65b, Cha67, Che68b, Che68a, Cho66, Chu66a, Chu67a, Chu69a, Coa64, CL64, Cof69a, Cra64, Cun54, Dag65, DS61]. Computer [Del62, Den69a, Den69b, Den65, Deo66, Der65, Dia56, DS65, Dor69, Dru69, DEL55, Eck62, Eng61a, Eng61b, Epl66, ET63, EBTB63, Far69, Fei61, Fif68b, Fin61, Fir64a, Fir64b, Fir65, Fis64, Fog57, FK65, Fra57, Fra58, Fra59, Fra64b, Fra65, Fre61a, FG64, Fre67b, Gan64, Gar65, Gau64, Gel61, Gil69, Gil65a, Gil66b, Gli66b, Gol63, GM62, Gon63, GHI^{+}63, Gre61, Gre66a, Gre69a, GM63b, Gre66b, Gre64b, GHR67, Guf62, Ham65a, Han67, HS67a, Har63b, Har67b, Hau62a, Hau62b, Hau65a, HM63, Hay69b, Hei63, Hen62, Hen67, Her67, Hig66, HH57, Hoa64, Hob66, Hol63b, Hop67b, Hor68, How67, HS69b, Isb63, JSS66, Jon64, Kam60a, Kam60b, Kar63a, Kar64b, Kar55, Kas62, Kau64, Kau66b, Kaz65, Kei68, Kel62, Kil59, Kim66a, Kim67, Kir64a, Kir64b]. Computer [Kle64, Knu65a, KW69b, Kon68, Kor60, Kor67a, Kov66, Kri63, Kuc66, Lam59, LC59, LEN63a, Les64, Lev63a, Lew64, Lew67b, LBB66, Lit68, Liu66c, Luk63, Mac69, Mac64, Mad62, MA63a, Man63b, Man66b, Man66c, Man67b, Mar62, Mar68, MT68, Mar65b, MR68, McK67, McL62, McQ62, Mck69, Men69b, Mer64b, MA58, MW61, Mil66, Mil67, Min63, Min69, Mit62b, MN58, Miu66, MTI67, ML61, MR59, MT59, MS63, Nad65, Nas56, Nas67, NM57, New64, Nic65a, O’G67, PAB68, PF69, PMRC67, PGB64, Poh64a, Pre64a, Pre55, PC56, QK67, Ran64, Rob66, RW53, Rog66, Ros67a, Ros64, Ros65b, Ros68, RG65, Rub65, Rub67, RV68, Sal60, Sal65, SW63, Sch64c, Sch69c, Sci64, Ses63, Ses65a, Sha63, Sha65, She64a, She63a, SS66, She63b, Sho66, SBM63, Slo63, Sob62]. Computer [Son64, Sri64b, Sto64, Str62, Stu69, Stu68a, Stu68b, SD69, SS68, Sut69, Swi63, Tai63, Tal64, Tam67, Tar65, Tay62, Tei55, Teo62, Tho67, Tho68b, Tho62, Top67, Tra67, Tru62, Tru63, Tsu61, TR58, Ung63a, Urb66, VB69, Van66, Vic67b, Vic69a, Vid67, WW55, Wag66, Wak69, War65, War60, War66, Was67, Wat66, Wee63, Whi66, Whi67, Wil65a, WB62, Win66a, Yau66, Yen69, Nel54, Flo62, HS61, McC63a, MMB66a, Rub62b, She65b]. Computer-Aided [Hay69b, Mck69, Sch64c, Yen69]. Computer-Oriented [DS65]. Computer-Stored [Gre64b]. Computers [Ano56q, Ano62-32, Ano65-32, Bek65, Bla56d, Bla58, Ble61a, Bow69, Bra58, Bre67, Bri66, BO59, Cor64, Deo68, Fuc56, GPW55, Gir69, Gon65, Goo55, Gun58, HK69, HMSG59, HP64b, HG57, Jos67, Kuc66, LWCL61, Luc59, Mar56b, Mar67a, McC57, Mea59, Nel63, Ott64, Pas67a, Pau67, Qua57, Rib64, Rid66, Rob55a, Rub62a, Sco57, Ste63c, SD57, Tav65, Wai66, Wan55, Wie65a, Wil61, Yao63, Yao65, ZA59, dM66, Wal53, Ano55a, Ano56g, Ano57f, Ano58l, Ano59p, Ano60o, Ano61p, Ano62t, Ano63p, Ano63x, Ano63y, Ano63z, Ano65s, Ano66p, Ano67a, Ano69b, Ano69-32, Ano69-33, Ano69-34, Ano69-35, Ano69-36, Ano69-37, Ano69-38, Ano69-39, Ano69-40, Bar64, Dav63a, Fir63a, Har63d, Kar63b, Kar67]. Computing [AW67, AW56, BMRS69, CM65, CD66b, Con67a, CV61, Cra61, Ger63, GR67, GDD62, HM65a, Hea56, HH57, HS69b, Ker56, Kle64, Kon60, Kor62a, Kor67b, KL58, Lew65a, MN58, Nat58, Nie56, O’G67, Rob55b, Ros61, Sal61, Sam65, Sch63a, Ser68a, Ser68b, Vol59, Wel64, Wit63, Yov59, Mar54]. Computing-Load [Kle64]. Computor [Man68, Vic67a]. Concatenative [PP68]. Concentrator [BMRS69]. Concept [Gre64b, Loe64, Ser68b, Sto63]. Concerning [Gau60]. Conclusion [Uhr61]. Concurrent [FS65]. Conditional [NM57, Skl60a]. Conditional-Sum [Skl60a]. Conditions [Law61, Law62b, MW61, NP64]. Conductance [RR66]. Conference [Ano69a]. Configuration [Den69b, Kum62, Kur64, Por64]. Configurations [Fre61b]. Confluence [Gil64a]. Confocal [CM60, Elf62]. Conic [Rob67]. Conjectured [Bra66]. Connected [BS64]. Connectedness [Fle67]. Connecting [Wag66]. Connection [Ake67, PMC67]. Connections [GR68, Lee61]. Connectivity [Ram65]. Consensus [Bro68, Mot60, Tis67]. Consideration [MN58]. Considerations [BdT65, Lot69, McC58, MF68c, Pas67b, Ram65]. Considering [GR68]. Consisting [Gos65]. Consoles [Ano68-40]. Constant [Dor67, Her66, Kau60, She63a, Tho62]. Constant-Weight [Kau60]. Constitution [Ano55a, Ano65r]. Constrained [Lan69c]. Constraint [Mul59d, Mul60]. Constraints [Dav69a, Gre56]. Constructing [Wat59]. Construction [Ake65c, BP63, Fos68b, GG68, Haz64, MK67]. Contact [ACGN60, Gil60b, Gou58]. Contained [MO65]. Containing [Har63c, Tei55]. Content [Cho67, CS66, Han66, Rux69, SH66c, SH67b]. Content-Addressable [Cho67, Han66]. Content-Addressed [CS66, Rux69, SH66c, SH67b]. Contents [Ano56a, Kov63]. Context [Ard65]. Continuity [HDW66]. Continuous [Ano68-43, Bra58, Kor64, Rub62a, Str66, Str68a, Wai67, Wai68a]. Contribution [WS67a]. Contributions [Ade62]. Contributors [Ano53a, Ano53b, Ano55b, Ano55c, Ano55d, Ano55e, Ano56b, Ano56c, Ano56d, Ano56e, Ano57a, Ano57b, Ano57c, Ano57d, Ano58b, Ano58c, Ano58d, Ano58e, Ano59l, Ano59m, Ano59n, Ano59o, Ano60i, Ano60j, Ano60k, Ano60l, Ano61h, Ano61i, Ano61j, Ano61k, Ano62l, Ano62m, Ano62n, Ano62o, Ano62p, Ano63k, Ano63l, Ano64k, Ano64l, Ano65o, Ano65p, Ano66k, Ano66l, Ano67f, Ano68p, Ano68q, Ano68r, Ano68s, Ano68t, Ano68u, Ano68v, Ano68w, Ano68x, Ano68y, Ano69q, Ano69r, Ano69s, Ano69t, Ano69u, Ano69v, Ano69w, Ano69x, Ano69y, Ano69z, Bar67, Bas63a, Boo67, Dol64, Gia64, Ham65b, Kat66, Kli66, Ler65a, Lo64, Mcc67b, Mcc63b, Mei63, Mer67a, Rap67, She65c, Ung63c, Wee66, Win66b, Win64, Win65d]. Control [Ale67, APP^{+}63, ABSH63, Ano66-35, Bed64, BR69, Bra58, Cur68a, Eck65, Fog57, Ger63, GS65, Gra62b, Gra62a, Gun58, Haw67, HH68, Kar55, Kir64b, Leo63b, MV55, Mel63, MTI67, Pau67, Ros68, RV68, Sch69c, SS68, Sus58, Tei55, Tru63, VB69, WP64, Zuc65]. Controlled [CM62a, CM62b, CM63, Kam60b, MC62, New66a]. Controller [Hor69]. Controlling [Cra64]. Conventional [SAP65, Wad61]. Conventions [Fei63b]. Convergence [Lew66b, May64, Set67, Urb64, Urb65b]. Conversational [Ble64, SD63]. Conversion [AM67b, Che61a, Cro68, Cro61, Die63, Goe64, Lip55, Roz62, Sch68, Wad61, Wan66, Dav62]. Conversions [Dag59]. Converter [Ano68o, BP59, Cou58, Cox61, KII62, Pic65, Pic66b, Sav59, Sca53, ZP66]. Converters [AM67a, Fel65, Kae61, Rob67, Ver64]. Convexity [DF67]. Convolution [SC68]. Convolutional [Pre66b, Pre68]. Cooley [SC68]. Coordinate [DS61, Mad62]. Coordinate-Rotation [DS61, Mad62]. Coping [GR64]. Copper [SB61]. Copper-Mandrel [SB61]. CORDIC [Dag59, Vol59]. Core [ABC61, AH62, BB61, BP59, Bro66, Car60, Che58, Chr61, Gol64, Hsi62, KI66, Lan64a, LM62, Lin63, MS59, NP56, Nit64, Ros58, SAP65, SC60, Sin62, Wee69, Wid54, Wie68, You57]. Cores [Bal62a, Ein61, Ein62, Lam59, Min62, Pie63b, ROK58, Til60]. Correcting [Bro60c, Chi59, FY66, Gol64, Kil59, Lev63b]. Correction [Ano55f, Ano56f, Ano57e, Ano58f, Ano58g, Ano58h, Ano60m, Ano60n, Ano61l, Ano61m, Ano62q, Ano63m, Ano63n, Ano63o, Ano65q, BFL^{+}66a, BW62, BK62a, BV65, CK65a, Chu66b, Eng61a, Fer69, FY69a, Gil66f, Gra69, Hau62a, Hsi68, Lan69a, Law62b, Ler65b, Mel63, Men61a, Mul60, SHB69, Wai68a, Win67a]. Corrections [Ano58i, Ano58j, Fri66a, Lew66b]. Correlation [BW61, BW62, Gon65, Lar66a, Len65, O’G67, Rid66, Rub62b, Wie65b, WK68]. Correlator [Che61b, Lar69, Sas57]. Correspondence [Lec67]. Corresponding [MO65]. Cost [Chr61, Kel68, Kli69]. Costs [Cof69a, RH67b]. Counter [CE69b, Dav69b, Fis57, GM63a, Hal69, Har56, Jon69a, Lev63c, MC65a, Ran63, Ren68, Tae69]. Countermeasures [Fif68b]. Counters [Ano68-46, Bei68, Brc65, Bri61, Bro55b, Gol57, Kai67, Kau60, Kau62, MM55, Mar63a, Oli65, Rob56, Rus65]. Counting [BHK62, Cha59b, Mis66, Pri65, Yoe63b]. Coupled [Eas58, Har58, Hsi67, Jos66, KH63, Lyn69b, MS58, RW67, WM59]. Couplers [KI63]. Coupling [Con66, MY61, McC68a]. Covariance [CM61]. Covering [GHL^{+}66, RH67b]. CPSS [Sch65b]. CRAM [Hoa62]. CRC [Wei52]. CRC-105 [Wei52]. Criteria [Urb65a]. Criterion [Dei65, Glu66, Nat57]. Critical [Fri66b, SW65]. Critique [Gas66]. Cross [Con66, Pie63b, RW67, Sas57]. Cross-Coupled [RW67]. Crossed [ENB61]. Crossed-Film [ENB61]. Crosstalk [Cat67]. Crustacean [Rub68]. Cryogenic [Sus61]. Cryosar [Joh61, Mit62a]. Cryoton [ENB61]. Cryotron [Dav63b, Har63a, Por64, Squ62]. CSX [BJST64]. CSX-1 [BJST64]. Cube [Dav65, Fro65]. Cumulative [Ano60-31, BH65a, Cho65a, Mei67b, Pre66a, She66, Win67b]. Current [Alo62, Ano55k, Ano55l, Ano55m, Ano55n, Ano56m, Ano56n, Ano56o, Ano56p, Ano57l, Ano57m, Ano57n, Ano57o, Ano58s, Ano59b, Ano59c, Ano59d, Ano59e, Ano59a, Ano60a, Ano60b, Ano60c, Ano60d, Ano60-31, Ano61a, Ano61b, Ano61c, Ano61d, Ano62c, Ano62a, Ano62b, Ano62d, Ano62e, Ano62f, Ano63a, Ano63b, Ano63c, Ano63d, Ano63e, Ano63f, Ano63g, Ano64a, Ano64b, Ano64c, Ano64d, Ano64e, Ano64f, Ano65a, Ano65b, Ano65c, Ano65d, Ano65e, Ano65f, Ano66a, Ano66b, Ano66c, Ano66d, Ano66e, Ano66f, Ano67b, Ano67c, Ano67d, Ano67e, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano69c, Ano69d, Ano69e, Ano69f, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k]. Current [Ano69n, Ano69l, Ano69m, Ano69o, BB56a, Brz63, Bue60, BAL61, Bur63, Cho66, Coa64, Fir64a, Fre67b, Gan64, Gil66b, Ham60, Har63b, Har67b, Her67, Hoa64, JMW60, Kau64, Kau66b, Kaz65, Ker56, Kim67, Kle53, Lan64a, LM62, MRS60, Mur68, Nas67, Nic65a, Pre64a, RC57, Rob66, Sha65, She63a, She63b, Sho66, Tho62, Top67, Ung63a, Wil65a, Wom65]. Current-Analog [Ker56]. Current-Clamped [MRS60]. Currently [Dav57a]. Curve [Wat59]. Curved [Swi63]. Cutpoint [Min64, YY66a]. Cybernetics [Fei61]. Cycle [Lav65a, Sei64, Sei67]. Cycles [Fen69, Mow66]. Cyclic [Epl68, Gol64, ME67a, Neu62a, SH66b]. Cyclical [Auf62]. Cyclo [ASC68]. Cyclo-Decompositions [ASC68]. Cylinders [Elf62]. Cylindrical [TF64].

D [Alo62, Uhr62, Gil69, Bas63b, Gil66d]. d-c [Gil69]. D. [Bit62, Flo62, Tau62]. D21 [Lan63b]. Damping [Pow59]. Dangers [Gil63b]. Danish [FK65, GHI^{+}63]. Data [AM62, BBB59, Bro69a, BMRS69, Che68a, Eps68, FH63, FK66, Hig63a, Lan63b, Led63, Lew63b, Man63a, Mul59b, MDH68, NW64, Rib64, Sam69a, WHB56, Mur65c]. Data-Accumulation [WHB56]. Data-Acquisition [Eps68]. Dataless [Mcc68b]. Datum [Ric67]. Dc [ZP66, HS64a, Kon60, Lyn69b, Nis60]. Dc-Ac [ZP66]. DDA [LA68, Wei52]. DDL [DD68, DD69]. Dead [Hor69, McA68]. Dead-Time [McA68]. Debugging [Kno68]. DEC [VB69]. Decimal [BC66, Brc65, Cou58, Cro60, Cro61, Dag59, LC53, Lip55, O’B57, Roz62, Sch68, Svo69, Tom56, Wei52]. Decimal-Coded [LC53]. Decimal-to-Binary [Cou58, Roz62]. Decision [Cad68, Cho57, CL61, Coo63a, Coo66, Das69, Flo58, Gaf68, Lan64b, Lar61, Sch65a, Toh64]. Decision-Making [Gaf68]. Decision-Table [Sch65a]. Decode [Sei67]. Decoder [Hsi62]. Decoding [Cha64, CB63, Kau60, Neu62a]. Decomposable [KMM69]. Decomposition [Cut68, Dav69a, Muk63, TMF69, Yoe61]. Decompositions [AYT68, ASC68, PP68, Rud65, Toh64, Yoe63a]. Deductive [Nil69]. Deering [Mad62]. Defense [AW56, Hea56, Nie56]. Definite [BS68b, Gin66, PRS63]. Definition [KM64b, Luc69b]. Definitions [RW53, Wai68b]. Degeneration [Gra64b, Pie63a]. Delay [AJ60, And53, Ard63, AFM68, Ben69b, Bri66, Bud66, Che68c, CS66, Cun54, GG69, Gle68, Gli62, Gon65, Har61a, HP64a, HF66, Her66, Lan68, Lan69c, LLT69, LTW63, MWR68, Rid66, Rot61, Rux69, SJ68, SD57, Tho68a, Tho55, Ung68, Wai66, Wie65a, Zac61]. Delay-Free [Lan69c, Ung68]. Delay-Line [CS66, Har61a, HP64a, LTW63]. Delayed [Loo65]. Delayed-Logic [Loo65]. Delays [AFM69, Hep65, Kor62b, Lan69c, MO63, SM69]. Demagnetizing [Elf62]. Demand [Cof69b]. Demonstrating [Pro64]. Demonstration [Bit62, NM57, Rog66]. Densities [Amo66, Gab64b]. Density [Gil64b, HB60, Hop64, Mat68, MS66, She62, Won55]. Dependence [Cho62]. Dependencies [WS67c]. Dependency [Koh65b]. Dependent [HS63b]. Deriving [Lin60]. DES-1 [Les64]. Descent [Kor64, Sas69]. Describing [Sch64a]. Description [Alr55, Bre65a, Rap66]. Descriptors [Bob65]. Design [AM67a, ABSH63, ALS68, Ano56r, Ano66-32, Ano66-33, Ano69p, AFM69, Bar61, BC65, BSK67, BBP57, Bow67, Brz65b, Cha66, Che68b, CE69a, Col61, CV61, Cra61, CS66, Dav69b, Dav69d, Dei65, DS65, DS69, DD68, Ear64, Fel65, Fif68a, Fra64a, Fra57, Fra59, Fri66e, FY69a, FY69b, FI65, Ger68, Gil59a, GM62, GK62, GMPS59, Gra62a, Gre66b, GRU58, Han68, HS67a, Har66a, Har68d, Hay69b, Hor69, Joh61, Jon69a, Kam60b, Kau66c, Kei68, Kin59, KW69a, KW69b, KL67, KK68, Kor62a, KL58, LS60, Led59a, LWCL61, Lot69, Luk63, MTS69, Man63b, Man66c, MS58, MG60, Mas60, Mau67, McC58, McQ62, Mck69, Mei65, ML61, Nat57, NK66, Net58, Net59a, Neu62b, New66b, Nie56, Pas67b, Pre63, Pro64, Ren68, Sam65, SW63, Sch64c, Seb57, SSG69, She62]. Design [Sho62, Smi68b, Smi69, Ste60, Sus61, Tae69, Teo62, Uhr62, VBW69, Wat66, Wel64, Wra62, YS68, YR65a, LNSW54, Mul54, Nel54, Mit62a, Ros62]. Designing [BdT65, CW68, Ger68, Liu64b, Wai67, Wai68a]. Designs [Sin62]. Destructive [Chu65a, Kru64, Ros66b]. Destructive-Readout [Chu65a, Ros66b]. Detect [RBS67]. Detecting [Bro60c, Kil59]. Detection [Arm66, BB69, BFK63, CD66a, Coo64a, Cra60b, Fri67a, KL67, KK68, Leh63b, MW67, Mar56a, McG69, Mel63, Mer67b, Muk63, Nis60, Sem63, She65d, Sza62, Wei63b, KCT62, Mul54]. Detector [Bri63, Hen62, Hil66, Lyn69b]. Detectors [Che61a, Per68]. Deterioration [Cor64]. Determination [Ano69-54, FF63a, Fos68a, Koh63, Liu63a, Maj67, Meo65, Mor67, Mot60, Nec67, PS68, Rei60, UM56]. Determined [Auf62, DF66, Kno67a]. Determining [Har57, Har67a, MO65, Sho64]. Deterministic [Har69b, ID63]. Developed [Mar57]. Developing [Ano57p]. Development [Hea56, Ros62, Ros61, Sal68, WS67a]. Device [AP66, Alo64, AD63, DS67b, Gaf68, Kas63, Mat66, Roy68b, SC68, TF64, Wie55, Woo63b]. Devices [Ata64, Don67, Hay61, Hor65a, Loo65, MC65b, Nag63, RR66, SSPR60, Tay65, Van60b, WM59]. Diagnosability [Deo66]. Diagnosable [PMC67]. Diagnosing [Chu67b]. Diagnosis [AC67, BFK63, Dwy69, HS69a, Kau68, Man66b, Man66c, SF62, Ses65b]. Diagnostic [Cha65b, Had67, Ihm63, Man64a, Pow69]. Diagram [DKP69, May60]. Diagrammatic [Ake65b]. Diagrams [BM63, Cly65, Col69b, Gin63, Kli67]. Dielectric [Dow58]. Difference [Gil66e, Gre62, SHB68, SHB69]. Different [Lan68, ROK58]. Differential [Ben69a, Bir67, Bro69b, Den69a, Dow57, Dow58, Eck65, Far68, Fuc57, Gil59b, Gil59c, Gre56, Gre69a, Han67, Hau62a, Hau62b, HP64b, HH53, Kar64b, KS67, KS68, Knu65b, Koh63, Kor67a, Lam64, Lit68, MPP59, Mar66b, May65b, Nat57, Nat58, Pet61, Ros65b, Skr66, TP60, Vic69a, Whi66, Wit67, Woo65]. Differentiation [Tho60]. Differentiator [Kin66a]. Differentiators [KP67]. Diffusion [Cha69]. Digit [Avi61, Lyn69b, MA58, Met65a, Svo69, Tun68]. Digital [AM67a, AM67b, Alr55, And53, Ano68o, Ano69-27, AK69, BH67, Bla56d, Ble61a, BB56b, Boo68, Bra58, Cam57, CET62, Cat67, Cha64, Che61a, Che61b, Cho60, Chu66a, Con62, Con66, Con67b, Cor64, Cox61, Cra64, Cro68, Deo68, DW67, Don67, DD68, DD69, DEL55, Dun65, Dwy69, EGP56, Far69, Far68, Fel65, Gab64b, Gar65, Gas65, Gas66, Gel61, Ger68, GPW55, Gil59c, Gir69, Gle68, GNV59, Ham58, HKM63, Hay69b, HB60, Hol63a, HP64b, Hor69, Hug67, JSS66, JC69, Kae61, Kar55, Kau68, KS67, KS68, Kel63, Kil59, KII62, Kno67b, Knu65b, KLP68, Kuc66, Lam64, LLT69, Led59a, LC59, Leh57, LWCL61, Leo63b, Les64, Lip55, LE57, Lo61, Luc59, Lun63, Mck69, Men69b, Mer64b, Min69, Mor59, MT59, Nad63, Nel64, O’h69, OKL59, Pic65, Pic66b, Pie63c, PC56, RR66]. Digital [Rob67, Rob55a, Rob58, RW67, Sag66, Sal60, SAL66, Sav59, Sca53, Sch63a, SW63, Sch64c, SC67, SC68, Shi68, Sho61, Shu62, Sie65, Ste63c, Sto63, Stu68a, Stu68b, Sus58, Tav65, Tay62, Tei55, TBS64, TNM64, Tou60, Tru62, Tsu61, Ver64, VS62a, VS62b, Wai67, Wai68a, Wan55, Whi67, Whi69c, Woo64a, Woo65, Yao63, Yao65, MGN54, Nel54, Dav62, Ric67]. Digital-Analog [Rob67]. Digital-to-Analog [AM67a, AM67b, Ano68o, Cro68]. Digital-to-Voice [Dav62]. Digitalrechenanlage [Ulb63]. Digitizer [LC53]. Dimension [PAB68]. Dimensional [Atr65, Bla56a, CP66, Col69a, FL67, Gre68a]. Dimensions [MMR60]. Diode [Alo62, BS64, Ber60a, BW63, Car64, CS64, Cho63a, Cho60, Coo64b, Dei65, Ear64, Goo64, GMN^{+}60, GRT55, HP64a, Kae62, KH63, KII62, Kra67, Kru64, Man64b, Mas60, MS59, Men61a, Men61b, MAN57, NBC60, Oli65, RR62b, Sal64, SP62, Ulz63, Uzu62, Wil63, YBS61]. Diode-Charge [Kra67]. Diode-Steered [MS59]. Diode-Transistor [Ear64, Mas60]. Diodeless [Pry58]. Diodes [DK63, FS59, HMSG59, Kae61, Pfe59, SI59, Uzu62]. Dioes [Amo64]. Dip [Sas57]. Direct [Chu65b, Eas58, Fle67, Har58, KH63, Ros65a, Sas69, Vid67, WM59]. Direct-Coupled [Eas58, Har58, KH63, WM59]. Directed [Gar69, Kei68]. Directional [KI63]. Discipline [Cof69a]. Discontinuities [Hau69]. discovery [Mar54]. Discrete [Kar69, Kle69, Ram65, She63c, War58a]. Discrete-Event [Kle69]. Discriminant [DF66, Kno67a, Spe67]. Discrimination [Smi65a]. Discriminator [Ulz63]. Discriminators [KH63]. Discussion [Ano58q, WD68]. Disjunctively [Coa63, KMM69, Yam62a]. Disk [FS65, Man63a, VB69]. Displacements [Col67, Dou68]. Display [BFL^{+}66b, Der67, HW69, MW68, PAB68, Rob67]. Displays [Far69, Kon68]. Distance [Kau58, O’B57, Rus65, Tom56]. Distinct [WS67b, dL69]. Distinguish [RBS67]. Distinguishability [Gil61b]. Distinguishing [KK68]. Distortions [Chu66a]. Distributed [CG65, Pow59]. Distribution [Bro63b, CKLP60, DH69, Fie64, Gil66g, Gre66a, Har64]. Distributions [Coo63b, FK69, Maj67]. Distributive [Wil64]. Divider [KS61, KB64, MGN54]. Division [Fer67, Fer69, Hau62a, Hau62b, KCT62, Kri63, Lam64, Leh63b, Lil56, Mit62b, Rob58, Sal61, Sch58, Tun68, Whi69b, WL61, Atk68, Mac62]. Divisions [Met62]. Divison [Nit64]. Divisor [Whi69b, Atk68]. Docking [Sil68a]. Document [Pet63]. Documents [Bob65]. Does [GR68]. Dolotta [Cha68]. Domain [Bit62, Cro64, Hau69, Rea63, Smi61, Wil64]. Domain-Wall [Smi61]. Don’t [SH67a]. DONUT [CL64]. Double [Val67]. Down [Gar69, LTW63, Kai67]. DPSS [Mur65c]. Drawn [Mer65, Woo64a]. Drive [Chr61, Sei67]. Driven [Kir64b, You57]. Driver [Sei64]. Driving [Thi63, Thi64]. Drum [AD69, Hol61a, Jac64b, MV55]. Drum-Like [AD69]. DTL [FK66, Slo62].

Dual [Kin66a, Kin59, Liu65, MF68b, Mul59d, Mul60, Mur65a, Tod62]. Dual-Monotonicity [Liu65]. Dual-Polarity [Kin59]. Duel [PVW59]. Duplicative [CS67]. During [Ano57k, Bro55a, Nas56]. Dwyer [Ano69-55]. Dynamic [Cro64, Den69c, EW53, FCC67, Hol63c, Kas62, Mar67a, Mei66, Nat57, Rea63, SB61, TPM65, Vic67b, Wil65b]. DYSEAC [EW53, LNSW54, LA54].

E. [Alo62, Cha65a, Dav62]. E.Stearns [Arm62c]. Easily [Bar66]. EC [Ano65s, Ano66p, Ano67a, Kar67]. EC-14 [Ano65s]. EC-15 [Ano66p]. EC-16 [Ano67a, Kar67]. Economical [APP^{+}63, GM65, Ros65a]. Edge [Sma68]. Edge-Weighted [Sma68]. Edited [Str62]. Editor [Fre68, Hau65b, And67b, Ano61s, Ano64u, Ano65-32, Ano65-33, Ano66n, Ano68z, Ano68-27, Ano68-28, Ano69-28, Ano69-29, Mar66a, MMB66a, O’N67a, Pra66a]. Editor-in-Chief [Ano61s, Ano65-33]. Editorial [Ano53c, Ano61n, Ano61o, Ano64-27, Ano66m]. Editors [Pra66a]. Editorship [Bla56c]. EDP [Ano66q, Sha66a, Zuc65]. EDPM [Mer56]. Education [Ano58t, Ano58u, Ano58v, Goo55]. Effect [Agg65, BW63, Chu67a, Gle68, Hol63b, Jos66, Tho68b, WR63, KM61]. Effectiveness [Ano63-40, Lei67]. Effects [Jar63, May64, Smi68c, Ulz66]. Efficiency [Jon69b, Man64a]. Efficient [Arm62a, Hum63, Sed68]. Elected [Ano60e]. Electrical [And53, Chu67a, Law62a, Ott58, She65b, Tho68b, Wee63]. Electrically [Far68, Gra60b, Til60]. Electro [Bra59, Mil55, SH66c, SH67b]. Electro-Optical [Bra59, SH66c, SH67b, Mil55]. ElectroData [Alr55]. Electrodeposited [SS59].

Electron [TBS64]. Electronic [Ano55a, Ano56g, Ano57f, Ano57k, Ano58l, Ano59p, Ano60o, Ano61p, Ano62t, Ano63p, Ano63x, Ano63y, Ano63z, Ano65s, Ano66p, Ano67a, Bar64, Bar63, Bro55a, BTW61, Dav63a, Dia56, DS62, Dow57, Dow58, Eps58, Fir63a, Gre56, Har63d, Hor65a, HH53, Kal57, Kar63b, Kar67, Kor63a, Lew67b, Mar56b, Mar67a, MN58, MB52, Mul56, Mul59a, Nas56, Nat57, Rob55b, RG65, San64, Sas57, Sch69a, Sil68b, Tho67, Tom58, WW55, Wei57, Wal53]. Electronic-Analog [Nat57, RG65]. Electronic-Computer [BTW61]. Electronics [She65b]. Electrostatic [GMPS59]. Element [Dav69c, Der64, Eck62, Gab62, Gab64a, Gab65, GT68, Gos65, HS69b, MF69, Nec68a, PGB64, Sal64, SAP65, Ser68a, Whi69c]. Elementary [Kam60b, MS63]. Elements [AE61, Ard63, AFM68, CP68, Die68, Ell65, GG69, GBC69, HC69, Har66d, HD67, HO67, Har68b, HG57, Hru69, IS65, Kas66, Kau61, Kro59, May64, May65b, Mei68a, Mei68b, Miy63, Miy64a, MF68a, MF68c, Ris67, Roy67, Sch63a, Ser68b, She65e, Sob62, Toh64, Tor69, VO65, Wei69b, YIK68, Low62]. Eliminating [Hau62a, Hau62b]. Elimination [Bro63a, Ull69a]. Embedding [BW64, de 68a]. Emission [Nar60]. Emitter [Jos66]. Employing [Bal62a, KH63, KP67, KMM69, RR66]. Encoder [CG64]. Encoding [Fre61b, Haz65, MC65b, Neu62a, Sau67]. Endmarkers [Ull69a]. Endres [Ano59x]. Enforcing [Kor64]. Engineering [Alr55, Ano69-27, Bak57, Bek65, Gea65, Kli67, LC63a, Ste66]. Engineers [Gas64, She65b]. English [Col69b, Kir64a]. English-Like [Col69b]. Enhance [WP64]. Entering [MW64]. Entry [FH62, Kas62, Shu62]. Enumeration [Win65a, Win67a]. Environment [Bro69a, Kir64b, PVW59, Sch69a]. Equation [Cha69, Koh63, MMR60, Rao58, Vic69a]. Equations [Ben69a, Bir67, Bro69b, Cow64, DS61, Den69a, DD69, EM69, GM65, Gra60a, GD67, Gre69a, Guf62, Han67, Hau62b, Hau65c, HH53, Kar64b, Kor62a, Kor67a, Kuo68, Led59a, Led59b, Lit68, MPP59, MTS69, Mar66b, May65b, Ogb69, Ros65b, Rou58, Sil68b, Sur68, Svo63a, Tay69, Tho69, TP60, Vic69b, Whi66, Wit67, Mad62]. Equipment [Den69b, Lew63b, Sch65a]. Equitable [Bro63b]. Equivalence [DS67a, Har63c, Har68e, Har63e, Iba67, Lec67, PS66, dL69]. Equivalent [CS64, GMB61, Kur64, Tam67]. Erasure [CS67]. Error [Bro60c, Bru57, Chi59, Con67a, Dei65, DR69, FY66, FK69, Gar66, Gol64, Gra69, Hel55, Hol63c, Hsi68, Kau58, Kil59, Mar56b, McK67, Mel63, Miu66, Nat58, Nel63, Ott64, Rao68, Rus65, SB61, Vic67b, WP64, Woo63b, Mul54]. Error-Checking [Kau58, Rao68]. Error-Correcting [Chi59, FY66, Kil59]. Error-Detecting [Kil59]. Error-Tolerant [Rus65]. Errors [Bri67, Dow57, Dow58, HS63a, Mar67a, MN58, Nis60, SHB68, SHB69, VB65, Wol62]. Esaki [GMN^{+}60, KII62, NBC60, YBS61]. ESIAC [ML61]. Essential [Mar63b, Ung68]. Estimates [Gir66, Gur65, Whi69b, Atk68]. Estimation [HM67a]. Euclidean [Kel62]. EULER [McK68, Gra60a]. Europe [Bla56d]. Evaluating [And62, Bru57, HL57, Kri64]. Evaluation [AW67, BS63, CET62, Con67a, DH69, Ihm63, Skl60b]. Event [Kle69]. Events [Eve64, PP68]. Evolutionary [MG66]. Exact [Kor62a]. Examination [Luk67]. Examining [Pet61]. Example [Lev63c]. Examples [Gin62]. Excess [Lip64]. Exchanger [Her62]. Excitation [Pow59]. Execution [Chu65b, Tsu61]. Executive [Sal68]. Expansions [BP69]. Experience [Gre63]. Experiment [BHNW57, FF67, Kno67a, Pro64, RH66]. Experimental [BS63, Bri66, Bro65b, Bro69a, Cho63b, IT67, KI66, LS66, Man66b, MFF63, Nie68, PF69, Sei67, Wai66, Wie65a]. Experiments [Bob65, Fir63b, Gil60a, Gil66g, Kim66b, Kno69, KK68, ME67b, Mez61, MDH68, NF64, Pet63, Ull69b, Whi67]. Exploitation [NP64]. Explosion [GR64]. Exponential [CET62]. Expression [Bic60, EM66]. Expressions [Abh58, Abh59, Ano68-45, Brz62, BP63, Brz64b, Brz65a, DC65, Ghi62, Hel66b, KF66, MY60]. Extended [Ano68-45, RH67b]. Extending [Glu66, Luc69b, Mye57]. Extension [BW64, CLS64, Coo63b, Miy64b]. Extraction [Amo66]. Extrapolating [Hag56].

F [Ano69-55, Cha65a]. Fabricated [Gra69]. Fabrication [Chi63]. Facilitate [CLS64]. Facilities [Ros65a]. Facility [BW64, Lew66a]. Factor [Sci66]. Factoring [DS65]. Factors [Elf62, Jon69b, Ros62]. Fail [Fos68b, MK67]. Fail-Safe [Fos68b, MK67]. Failure [Bei68, CS67]. Failure-Erasure [CS67]. Failure-Masking [CS67]. Failure-Tolerant [Bei68]. Failures [Chu67b, Hol63a, HS69a, RBS67]. Fall [Thi63, Thi64]. Families [SJ69]. Fan [DS69, Hol63b]. Fan-In [DS69, Hol63b]. Fan-Out [Hol63b]. Fano [Gil62]. Fast [AS66, Avi61, BS59, Brc65, Eck65, Fer68, GPW55, HS64a, Hen60, Jac64a, Lan64a, Lei57, Mat66, Pra69, SW63, Sha69, Ver64, Wag67, Wal64, Whi65]. Fast-Serial-Parallel [Ver64]. Fast-Time [Wag67]. Faster [Por64]. Fault [Arm66, Dwy69, Fri67a, HS67a, Kau68, KL67, KK68, Lee62b]. Fault-Detection [KL67, KK68]. Fault-Locating [Lee62b]. Fe [MS66]. Feature [TR58]. Features [MV55]. Feed [Hug67]. Feed-Forward [Hug67]. Feedback [Ano68-46, Bri61, Bro60b, BHK62, BK62a, BK62b, Dab69, Fri66c, Fri66d, GG69, Gra58, HS63a, Har66e, Hen64b, HS64b, KP67, Lem69, Mar63a, Mil55, Mow66, Mow69, Pau67, Sin69, Yoe63b, MGN54, Kau62]. Feeding [Gos65]. Ferranti [Sha63]. Ferranti-Packard [Sha63]. Ferrite [BB61, Bro66, KI65, KI66, Mei65, SAP65, Til60, Wee69]. Ferroelectric [And56, PM58]. Ferroelectrics [Tay65]. Ferromagnetic [Kau67a]. Ferroresonant [Bie68]. Field [Ahr65b, Ake65a, Ake66, And67a, Ano59v, Ano59w, Ano60x, Ano60y, Ano60z, Ano60-27, Ano61-28, Ano61-29, Ano61-30, Ano61-31, Ano62-33, Ano62-34, Ano62-35, Ano62-36, Ano62-37, Ano68-42, Ash65, Avi63, Axe63, Can67b, Car65, Cha67, Epl66, Fir64b, FH63, Gil65a, Gol63, Hop67b, Jon64, Lev63a, Lot69, Man67b, Mil66, Min63, Ros64, Sci64, Ses65a, Son64, Sri64b, Tai63, Tra67, Urb66, WR63, War66, Yau66]. Field-Effect [WR63]. Field-Length [FH63]. Fields [Boo64, CM59, CM60, Els69, Gra64c, Gre68a, Mul59c]. Figure [Hor65a, Mul59b, Sci66]. File [CW68, Chu69a, Els64, FG61, Man63a, Smi68c]. Film [Bit59, Bro60a, Cro64, ENB61, FK66, Gra64c, Hig66, PM60, Poh66, Rea63, SSPR60, Sei64, Sei67, Sob62, TF64, Ulz66, Yao66]. Films [CM59, Mat66, RC64]. Filter [JC69, Sie65]. Filtering [SC68]. Filters [Lar66b, Sze67]. Final [Liu63a]. Finding [Arm66, Cra68, Dab69, Had67, MPP59, Tor68]. Finite [Agg65, Auf62, Boo64, CE69a, Col69a, Cur68b, Eve65b, Ghi62, Gil60a, Gil61a, Gil66e, GO63, Har68a, Iba67, Kli67, Lew63a, Liu63b, Liu64a, Mas66, Mez61, MW64, New67, New68b, Nie69a, Sta65, SO68, Vai69, Win62, dL69]. Finite-Memory [Gil60a, Mas66, New67, New68b]. Finite-State [Col69a, Gil61a, Har68a, Iba67, dL69]. First [Bro66, Mar65b, Nad65]. First-and [Bro66]. Fixed [AM67b, AK69, Cro68, DF68a, FH63, Gre66b, LBG65, Wai67, Wai68a].

Fixed-Point [AK69]. Fixed-Word-Length [Gre66b]. Flaws [WD68]. Flexible [Mai62, Sam69b, Tsu61]. Flight [DEL55, FH62, Fog57, Gau64, GNV59, Rog66, Shu62]. Flip [Ano69-53, Ara57, BG56a, Cur69, GD67, Mar69a, RR62a, Sur68, Uzu62]. Flip-Flop [Ara57, Cur69, GD67, Mar69a, RR62a, Sur68]. Flip-Flops [Ano69-53, BG56a, Uzu62]. Floating [Chu65b]. Flop [Ara57, Cur69, GD67, Mar69a, RR62a, Sur68]. Flops [Ano69-53, BG56a, Uzu62]. Flow [BM63, BM64, DBC69, Gin63, Gra66, Luc66b, Ros66a, Ung65, Ung68]. Flowrate [Ber60b]. Flux [Bal62b, NP64, Nit64]. Fluxlok [Til60]. Folding [Wee58]. Font [LS66]. Foreground [Cof69a]. Foreground-Background [Cof69a]. Foreword [Ano69-30, Ano69-31, Pas67a]. Form [Ano68m, Har63a, SK65, Wag66]. FORMAC [SB64, Wir65]. Formal [Hop67a, KM64b, Mil58, Nil69, Sch64a, Ste63a]. Format [Sch65a]. Formation [Ihm63]. Forms [CM65, CD66b, Coh62, Hoc62, Lec63b, Mor67, Mot60, UM56, Wan67]. Formula [Kel68, Kli69]. Formulas [Boo62, CR62, GG68]. FORTRAN [Ano68-43, BSK67, CLS64, BH64]. FORTRAN-IV [Ano68-43]. Forward [Hug67]. foundations [Mar54]. Four [Gou58, JA67, Kor62a, Pfe59, Sch58, Yen68a, Yen69]. Four-Phase [Yen68a, Yen69]. Four-Quadrant [Pfe59, Sch58]. Four-Terminal [Kor62a]. Four-Variable [JA67]. Fourier [Mac69, And68, Cla56, Con67b, DW67, Rat67, Sha69, Sto66a]. FP6000 [Sha63]. Frame [Boo62]. Free [Kae62, Lan69c, LS60, MF69, Seb57, Ung68, Wil64]. Freeman [Flo62]. Frequencies [HL57]. Frequency [Ber60b, Bic60, Bri63, Hen62, MRS60, VB65, Wid54]. Frequency-to-Period-to-Analog [Ber60b]. Full [Dav63b, RR62a, RR62b, Squ62]. Function [AK69, CET62, DG61, Gil59a, Gim65a, GRU58, Har56, Har57, Hru69, JA67, Kah68, Kri64, Kri66, Lec63b, Lev68, Luk67, MO65, Meo65, MP64, MAN57, Mot60, MC64, Muk63, Nec67, O’N69, PD62, Par62, Pic66a, Pol57, Rub68, Sch62, Ste69, SD57, Ter62, Tom58, UM56, Urb68, WH62]. Functional [Ama64b, Den69a, Klo61, MC65a, Vic69a]. Functions [Abh58, Abh59, AP66, AH63, Art61, BH65b, Bin68, BS68a, Bor69, CM65, CB62, CD66a, CD66b, Cho57, Cho65b, Chu61, CKL62, Coh62, Col61, Cop68, DC65, Dav57b, DS67a, DF66, Dvo68, Ein61, Ein62, EG61, Els60, ES64, Eps58, EKP67, FT68, Flo58, FI65, Gab62, Gab64a, GBC69, Gil59b, Gos65, Gou58, Har63c, Haz64, Hen64a, Her66, Hoa62, How56, HS69c, Ito69, KW65, Kas66, Kau61, KMM69, Kno67a, KC58, KO66, Lec67, LC63b, Lis63, Liu65, Liu66a, Lor64, MS69, Mar56a, MG60, May65a, McN61, Mei69, Men65, Mer66, Mer67b, MN61, Min62, Miy63, Miy64a, MF68b, Muk66, Mul59d, Mul60, Mur65a, Mur65b, MT66, MK61, Nam64, Nar61, Nat64, Nec68b, O’G67, PU59, PW67, PS68, PS66, Pre66b, Pug65, RC66, Roy67, Roy68a, Rud65, Set66b, SP69]. Functions [She64b, She65a, She65d, She65e, SH66a, SH67a, Smi66, Spe67, SJ69, Tis67, Tod62, Toh64, Tor66, Uhr62, Van59, Van60a, Vic67a, War58b, Wee58, Wei69a, Wei69b, Wil63, Win65a, Win65b, Win67a, YI65, YI68a, YI68b, Yam62b, YO68, Yen68b]. Fundamental [Nel63, Ott64]. Fundamentals [Eng61a, Eng61b, WB62]. Further [Ble61b, Bro69a, Hig61b, NF64, Roy68a]. Future [Gil66d, Hob66]. Fuzzy [Mar69b].

Gain [Agg65, Got64, NP64]. Galois [Els69]. Gambling [Fin61]. Games [Eng61a, Eng61b, WB62]. Gapless [WL67]. Gate [ACW64, AFM69, BC68, CL63, CL64, CW69, FS59, Gra64b, HS69a, Lew66b, MW67, Mau67, Meo66, Pie63a, Slo62, SH68, SL65, Win63, Win69]. Gates [Cad68, Mii61]. Gating [Cat66]. Gaussian [DH69, FK69, Mat68, O’N69]. General [BC65, Fog67, Fra57, Fra58, GMB61, GBC69, HP64b, Hru69, Kam60b, KM64a, McC58, QK67, RG65, She65a, Skl63a, Stu68a, Stu68b, Who63, Wig63]. General-Purpose [Fog67, Kam60b, QK67, RG65, Stu68a, Stu68b]. Generalization [Chu61, Lei67, Tis67]. Generalizations [Man66c]. Generalized [AP66, Bek59, Cha59a, Cha60, CW68, Deo68, Gar58, Gir69, MMB66b, Nol64, PS66, Sin66, Ste63b, Wee68]. Generate [Shi68, Tal64]. Generated [FR58, Gil66g, Ste69]. Generating [Har56, MN61, Nat64, Pol57, Wil63]. Generation [And68, AK69, BH65b, Bin68, Bla62, Cam64, Cop68, Gil59a, HKM63, Hen67, Her62, LB60, LEN63a, Mac69, MTS69, Mat68, MF68b, Mur65a, Pet63, SP69, Spe67, Stu69, Yen69]. Generator [Bar63, BG56b, DS62, DG61, Far62, FY69a, FY69b, Ham58, MAN57, Pic66a, Rob67, Sch62, Tom58, WH62]. Generators [CE69a, Col67, Dou68, Luk67, NK66, PD62, RH67a, Ste63a, SW61, Whi67]. Geometric [Art61, Fre61b, Hor65b, MW62]. Geometrical [Cov65]. Geometry [CM59, Nie69b]. GIER [FK65, GHI^{+}63]. Gimpel [RH67b]. Given [EM66, Gin59c, MO65, Svo63b]. Giving [Hoc62]. Glass [Rux69]. Global [HS69b]. Glow [Cha59b]. Gradual [GS58]. Grammars [Hop67a, Lan64b]. Graph [BE69, BM63, BM64, Epl68, Gin63, LA68, ME67a, ME69]. Graphic [Ano68-40, Hor68]. Graphical [Der67, Ros67a, She65e, Sut68]. Graphical-Interface [Ros67a]. Graphics [Can66, MW68, PF69]. Graphics-Ten [Can66]. Graphs [AYT68, Ake65c, Fri66e, Kor67c, MY60, Qua69, Ram65, Sma68, Ste63a, Wan67].

Gray [Fis57, CE69b, Wan66]. Gray-to-Binary [Wan66]. Green [Ter62, Uhr62]. Greenblatt [Kuc68b]. Grey [Stu69]. Group [And64, Ano55a, Ano63x, Ano63y, Ano63z, Ano64g, Ano64h, Ano64i, Ano64j, Ano65i, Ano65j, Ano65k, Ano65l, Ano65m, Ano65n, Ano66g, Ano66h, Ano66i, Ano66j, Ano69a, Bau64, Ber66b, Dag65, Dav63a, Den65, Der65, Fir63a, Fra65, Har63d, Kar63b, Kim66a, Lav65b, Lec67, Lew64, Liu66c, MMB66a, Poh64a, Tar65, War65, Win66a, Yoe65, Ano64t, Ano65r, She65b]. Group-Theoretical [Yoe65]. Groups [Har63c, Sto64]. Guaranteed [Uhr61]. Guidance [Hol63a, LW59, MR68, Men69b, PMRC67]. Guide [Dor67, Gas66]. GUS [Kud64, MA63c]. Gusev [Auf62].

H [Ano58w, Low62]. H. [SBM63]. Hadamard [Pra69]. Hall [KM61]. Hall-Effect [KM61]. Hand [Mer65, Woo64a]. Hand-Drawn [Mer65, Woo64a]. Handprinted [Bro64, GHR67, Kno69]. Handwritten [Hig65, Kam61, Mar65b, Nad65]. Hardening [Mau67]. Hardware [AP66, Kno68, Sal68]. Hardware-Software [Kno68]. Harmonic [Gre68b, Kos62]. Hartmanis [Arm62c]. Having [Gim65a, Kil59, Mat68]. Hazard [Ler65b]. Hazards [HC68, Mar63b, McG69, MK69, Ung68, Woo64b]. Head [FR58, She62]. Heat [Her62]. Heater [GS58]. Heijn [SBM63]. Helicopter [Paq67]. Helium [Chi63]. Help [DBC69]. Hexadecimal [Goe64]. Hexagonal [Gol69]. Hidden [FL67, Wal53]. Hidden-Line [FL67]. Hierarchies [AW67]. High [Amo64, Amo66, And68, BG56a, BG56b, BP59, Cho63a, Con66, DK63, DF68a, Edw61, EGP56, Gam68, GT68, Gil69, Gil66d, Gil64b, GMN^{+}60, Hen60, HB60, Hop64, Jar63, Kae61, KH63, Kau66a, LM62, LC59, Leh57, LB60, LB61, Lin66, Lyn69b, Mac69, MR65, MRS60, McQ62, Mei65, MS66, NP56, O’G67, PC56, PLS61, Sal60, Sav59, Sch68, She62, SW61, Til60, Waa67, Waa68, Wei63a, Wel69, Wie55, Wij68, Won55, Yao63]. High-Accuracy [Hen60]. High-Density [Hop64, MS66, She62]. High-Frequency [MRS60]. High-Order [SW61]. High-Speed [Amo64, And68, BG56a, BG56b, BP59, Cho63a, DK63, DF68a, Edw61, EGP56, Gam68, GT68, GMN^{+}60, Jar63, Kae61, KH63, Kau66a, LM62, LC59, Leh57, LB60, LB61, Lyn69b, Mac69, MR65, McQ62, Mei65, NP56, O’G67, PC56, PLS61, Sal60, Sav59, Til60, Waa67, Waa68, Wei63a, Wel69, Wie55, Wij68]. Higher [Atk68, Gre62, SH66a, Whi69b]. Higher-Order [Gre62]. Higher-Radix [Whi69b, Atk68]. Highest [Bra66]. Highleyman [MDH68]. Highly [HS69b, KW69b]. Hill [Str62]. History [Mea59, PLS61]. Hit [Van57]. Hollow [CM60, Elf62]. Homogeneous [Set66a]. Horizons [MR68]. Hospital [Led63]. Howard [Ano58k]. Human [Ano56r, Fog57, Sto64]. Hybrid [Ano62-38, Ano65-32, Ben69a, Cha69, Cly65, Con67a, Den69a, Gle68, GR67, Gre63, Gre69a, HK69, HKM63, Han67, Hen67, How67, JSS66, Kar64b, Kel63, Kor63b, Kor65, Kor67a, Lew65a, Lew66a, Lit68, LK60, Mar68, MMB66b, Miu66, MTI67, Nel64, O’N69, Paq67, Pau67, Rib64, Rog66, Ros68, Rub65, Sch63a, Ser68a, Ser68b, Sil68a, Skr66, Tav65, Van66, Whi65, dM66]. Hybrid-Code [Skr66]. Hydro [Dar68]. Hydrodynamic [CK65b]. Hyperplanes [San66]. Hypersphere [Coo66]. Hysteresis [Bie68, Kae62, Swi63]. Hysteresis-Free [Kae62].

IADIC [Ser68a]. IBM [Dom57, Gas65, Gea65, McK68, Mer56].

Idealized [Kil59]. Ideally [Brc65]. Identification [CE65, Cos63, DS67a, JA67, KMM69, Mar56a, Rub65, Str66, Str68b]. Identifiers [MS69]. IEEE [Ano63p, Ano64t, Ano65r, Ano65s, Ano66o, Ano66p, Ano67a, Ano69a, Ano69b, Ano69-32, Ano69-33, Ano69-34, Ano69-35, Ano69-36, Ano69-37, Ano69-38, Ano69-39, Ano69-40, Bar64, Kar67]. IFIP [Ano62r, Ano62s, Ano63q, Seb63]. IGFET [Ano68o]. II [WM53, Arm62c, Auf58, Bey63, Bre65a, Cra61, Dow58, How67, JO68b, Man66c, Nel64, Net59a, SS62, SH61]. III [Bey64, McC63a, PMRC67]. ILLIAC [BBK^{+}68, Bre65a, Dav69c, Kuc68a, McC63a]. Illinois [McC63a]. Image [Rus68]. Imaginary [Die63]. Immunity [Yao65]. Impact [Ano56q, Dau66, Jos67]. Impedance [GM63a, Hel55, Thi63, Thi64]. Implementation [Ano66q, APVS66, Bei68, Bre65b, McK68, Ols65, Sas68, Sha66a, SNS69]. Implementing [HH68]. Implicant [Ano66-31, Gim65a, Gim65b, Hoc62, Mor67, PM62]. Implicants [DC65, Fri67b, Luc66a, Meo65, Mot60, Nec67, Svo67]. Improved [Bud66, Cro60, GL65a, Gra56, Har68d, Hep65, MS69, NM68, Ses65b]. Improvement [Bar68, BTW61, Cha65a, Glu66]. Improvements [Bue60, PD62]. Improving [FY66, Gar65, Har68b, MT68, Mer64b, Tsu62]. Impulsive [Ses63]. Inc. [She65b]. Including [Bie68]. Incomplete [Bou68, Luc69b, MG66]. Incompletely [DBC69, GL65b, Gra66, Mei67a, Nar61, PU59, Ros66a]. Inconsistent [Coh62]. Increase [Mei66, TPM65, Ver64]. Increased [Gab64b]. Increasing [RA59, Tay62]. Incremental [DS61, HW69, Mad62, SH66a]. Incrementing [Fra69]. Independence [Cho65b]. Independent [Pol57]. Index [And66, Ano56g, Ano57f, Ano58l, Ano59f, Ano59g, Ano59h, Ano59a, Ano59p, Ano59-28, Ano59-29, Ano59-30, Ano60f, Ano60g, Ano60h, Ano60o, Ano60-28, Ano60-29, Ano60-30, Ano60-31, Ano61e, Ano61f, Ano61g, Ano61p, Ano61-32, Ano61-33, Ano61-34, Ano62h, Ano62i, Ano62j, Ano62k, Ano62t, Ano62-39, Ano62-40, Ano62-41, Ano62-42, Ano64m, Ano65g, Ano65s, Ano66p, Ano67a, Ano68n, Ano68-44, Ano69b, Ash66, BH65a, Bra64, Brz67, Cho64, Cho65a, Coh63, Cow63, Den67, Der66, Eve66, Fle63b, Gea63, Gli66a, Har65a, Har66b, Hop66, Kar67, Knu67, Koe63, Koh65a, Kon64, Kru63, Lec63a, Len63c, Lew67a, Liu66b, Loc67, Mai63, Man66a, Mat65, MR66, Mei67b, Met65b, Mik65a, Mik65b, Mil65, Mun67, O’D64, Ove63, Pag67, Pie65, Poh64b, Pre66a, Ren64, Sac65, Sch64b, She66, Spa66, Sto67, Tab67]. Index [Wai64a, Wal67, Wid64, Wid67, Win65c, Win67b, Wyn64]. Indexing [Bob65]. Inductor [Cro64, Rea63]. Industry [Gun58]. Inelastic [Kor60]. Inequalities [Cov65, HK65, Law67]. Inessentiality [CR62]. Inexpensive [Tsu61]. Inferiority [Uhr61]. Infinite [WH62]. Infinite-Resolution [WH62]. Information [Ano59q, Ano60p, Ano61q, Ano61r, Ano63r, Ano63s, Ano63t, Ano63u, Ano63v, Ano63w, Ano64n, Ano64o, Ano64p, Ano64q, Ano64r, Ano64s, Ano65t, Ano65u, Ano65v, Ano65w, Ano65x, Ano65y, Ano65z, Ano65-27, Ano65-28, Ano65-29, Ano65-30, Ano65-31, Ano66r, Ano66s, Ano66t, Ano66u, Ano66v, Ano66w, Ano66x, Ano66y, Ano66z, Ano66-27, Ano66-28, Ano66q, Ano67g, Ano67h, Ano67i, Ano67j, Ano67k, Ano67l, Ano68-29, Ano68-30, Ano68-31, Ano68-32, Ano68-33, Ano68-34, Ano68-35, Ano68-36, Ano68-37, Ano68-38, Ano68-39, Ano69-41, Ano69-42, Ano69-43, Ano69-44, Ano69-45, Ano69-46, Ano69-47, Ano69-48, Ano69-49, Ano69-50, Ano69-51, Ano69-52, Arb66, AM62, Ban65, BL66, CW68, Dav66, Don65a, Ear64, Eve65b, Fel63, Gaf68, Gil66a, GR64, Lan63a, LR65, Sha66a, Sha66b, Ste65, Sto65, Stu69]. Information [Sus58, TBS64, Wai65, Woo64a, Gil62, Tau62]. Information-Processing [Gaf68]. Informational [YR65b]. Initial [Liu63a, MW61, NP64, Vic69a]. Ink [FR58]. Inner [Win68b]. Input [Auf62, Bro60b, Dav67, Dor67, FM68, GD67, Hel55, Hor65b, Isb63, Kau61, Lew67b, Mai62, Min61, MF68c, MK61, Neg64, New67, New68b, New68a, Ris67, Sch65a, Sur68, Tru63, TR58, Wei69b, Wil64, WP62, Woo64a]. Input-Binary [New67, New68b, New68a]. Input-Output [Isb63]. Inputs [Men61a, Men61b, SS68]. Inserted [AFM68]. Installations [WW55]. Institute [She65b]. Instruction [How67, Law68, Mil67]. Instructions [WP64]. Integer [Bre65b, Cam64, Cro61, Fra69]. Integral [Bru57, Con67a, Dei65, TP60]. Integral-Error-Squared [Bru57]. Integrals [And62, And63, Con67b, DW67, Rat67]. Integrated [ALS68, DF68a, GR68, WM59, Whi69a]. Integrating [Ter62]. Integration [Bek59, GG68, Gre69a, Jos67, Kah68, Kat68, LLMS67, Mur68, SNS69, Tho60, Whi69c]. Integrator [Lun63]. Integrators [Hau67, Mye57, Sco57]. Interaction [Sto64]. Interactive [Sut68]. Intercommunication [Flo68, Lau68]. Interconnected [Die68]. Interconnection [KLW68, Yao65]. Interconnections [Jar63, Law62a, Woo65]. Interface [Ros67a]. Intergraphic [Ros67a]. Interlaced [Waa68]. Internal [Arm62a, Arm62b, BF61, DM64, Gel62, GL65b, Hum63, MU59, Mei67a, Tra66]. International [Bla56b]. Interpolation [AK69]. Interpolators [Nat63]. Interpretation [Chu65b, Kir64a, She65e]. Interrogation [AM62, SH66c, SH67b].

Interrupt [TR58]. Intersection [Hoc62]. Interval [Wol69]. Interwoven [Arm65, Sri64a]. Introduction [Dre67, Hor68, Lew63a, Mea59, Nie69b, SB64, Wir65]. Invariance [BS67]. Inverse [Wee68]. Inverses [MS68, RY66]. Inversion [Ake68, Tou60]. Invert [Hel63]. Inverters [Ake68, RR66]. Invertible [Lor64]. Investigation [Gre68a, IT67, LS66, MFF63, PT66]. Investigations [Gra58]. Involve [Ard63]. IRE [Ano55a, Ano56g, Ano57f, Ano58l, Ano59p, Ano60o, Ano61p, Ano62t, Ano62-32, Bak57]. IRETEC [Ano61s]. Irreducible [AYT68]. Irredundant [CM65, CD66b, Haz64, Hoc62, Mot60]. Isograph [Rao58]. Isomorphic [Wan65]. Issue [Ano62-38, Ano64-27, Ano66o, Ano66-34, Flo64a, Kau66a, Pas67a, Slo63]. Iterated [Bro68, Mot60, Urb64, Urb65a]. Iterative [Atr65, Col69a, Eck65, Fer68, Gon63, Hen64b, Kil62, Kin66b, McC58, O’G67, O’N67b, Ros65b, Wai64b]. Iteratively [DF66, Kno67a, Stu68a, Stu68b]. IV [Ano68-43, BBK^{+}68, Dav69c, Kuc68a].

J [Arm62c, Ros62, Rub62b, SBM63]. J. [Mac62, Ros62]. Jackson [Lev63c]. Jigsaw [Fis64, FG64]. Johnston [Mit62a]. Jump [BHK62]. Junction [Cho63a, GMB61, Gel62, PC56, Tam67]. Junction-Transistor [GMB61].

Karnaugh [KMM69, MS69]. Keating [Hoa62]. Keith [Ano64t]. Keyboard [Lew67b]. Kinds [SW65].

L [Ano62-43, Auf62, Dav62, Rub62b]. L. [Hoa62]. Laboratories [Kor65, Rap66]. Laboratory [How67]. Ladder [Ahr65a]. Ladders [AM67b, Cro68]. Laddics [Bal62b]. Lagrange [Hau65c, Pie63c]. Laguerre [BP69]. Language [Ano68-43, Ano68-45, BW64, BR69, DD68, Gas64, Kei68, KM64a, Law68, Mul64, Pro64, Sal65, Sam65, Sam69b, Sch65a, Sch64a, Sch64c, Wel64, Rap66]. Languages [Aho68, Ano64-27, Ano66-35, Cur68a, Flo64b, Gas66, Gre61, Haw67, Ste63a]. LARC [PLS61]. Large [AC67, BHRS61, Bin68, Bro60b, BMRS69, Cro64, DF68a, GMB61, GMPS59, Jos67, Kat68, LLMS67, Mur68, Rea63, Sag66, SNS69, Tal64, Whi69c, Yao66]. Large-Scale [DF68a, LLMS67, SNS69, Whi69c, Yao66]. Large-Signal [GMB61]. Latin [MES63]. Lattice [Wil64]. Leads [Hen64b]. Learning [Fir63b, IS65, Lew65b, SP63]. Least [McA68]. Least-Square [McA68]. Ledley [Mac62]. Lee [Ake67]. Leimer [Ros62]. Length [Dor67, FH63, Gre66b, HW69, KK68, ME69, Rei60]. Less [GG69]. Letdown [Men69b]. Letter [Boo62, Bra63, Hau65b]. Level [BW61, BW62, KELS62, Lav65a, Rap66, SD69, WD68, Nol62, Rub62b]. Lexical [Gal62, Sal63]. Lifetime [Bol67]. Light [Ano68-40, Kon68, RR66, Ros65a]. Light-Pen [Ros65a]. Like [AD69, Col69b, Die68]. Limit [Set66a]. Limitation [Ble61a]. Limitations [Dow57, Hol63b]. Limited [DS69]. Lincoln [Can67a]. Lindaman [Miy64b]. Line [Ano69o, Ben69b, Bro64, Bro65b, Che68b, CS66, Der67, DKP69, Fit63, Fre68, FL67, Gre69c, GHR67, Har61a, HP64a, Hen66, Lan69c, LTW63, Lun63, McC68a, MWR68, Rux69, Tho68a, WK68]. Linear [Ano69p, Ben69a, Bre65b, Brz65a, Coa63, Coh64, CE65, Col67, Cov65, Dou68, DF66, Gam68, Gas63, Gil59b, Gil64a, Gil66c, Glu66, GM62, Gos65, Guf62, Har68c, Hig61c, HK65, HH53, HS64b, ID63, Kah68, Kau61, Kno67a, Lav65a, Lav65b, Law67, Man67a, MT68, MS68, Min61, MK61, Nat57, Nat58, Nat63, Par62, Pet61, Pug65, San66, Sch62, Sie65, Smi68b, Tod65, Wai67, Wai68a, Wan67, Wan69, Yam62a]. Linear-Feedback [HS64b]. Linear-Input [Kau61, Min61, MK61]. Linear-Passive [Gam68]. Linear-Segment [Sch62]. Linearity [BD64, DB66, DF68b, Har65b, YW66]. Linearly [CKL62, KT65, Tor66]. Linearly-Separable [CKL62, Tor66]. Lines [AJ60, And53, Che68c, Gli62, Gra64c, Kau67b, Rot61, Was67, Was68]. Lines-Model [Was67]. Link [Isb63]. Linkage [Abe69]. Linked [Rap66]. Liquid [Chi63]. List [BW64, Gel61, Hoc62, LTW63, Mor67, MS63, Rap66, Wig63]. List-Processing [MS63]. Literature [Ano55k, Ano55l, Ano55m, Ano55n, Ano56m, Ano56n, Ano56o, Ano56p, Ano57l, Ano57m, Ano57n, Ano57o, Ano58s, Ano59b, Ano59c, Ano59d, Ano59e, Ano59a, Ano60a, Ano60b, Ano60c, Ano60d, Ano60-31, Ano61a, Ano61b, Ano61c, Ano61d, Ano62c, Ano62a, Ano62b, Ano62d, Ano62e, Ano62f, Ano63a, Ano63b, Ano63c, Ano63d, Ano63e, Ano63f, Ano63g, Ano64a, Ano64b, Ano64c, Ano64d, Ano64e, Ano64f, Ano65a, Ano65b, Ano65c, Ano65d, Ano65e, Ano65f, Ano66a, Ano66b, Ano66c, Ano66d, Ano66e, Ano66f, Ano67b, Ano67c, Ano67d, Ano67e, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano69c, Ano69d, Ano69e, Ano69f, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k, Ano69n]. Literature [Ano69l, Ano69m, Brz63, Bur63, Cho66, Coa64, Fir64a, Fre67b, Gan64, Gil66b, Har63b, Har67b, Her67, Hoa64, Kau64, Kau66b, Kaz65, Kim67, Nas67, Nic65a, Pre64a, Rob66, Sha65, She63b, Sho66, Top67, Ung63a, Wil65a]. LLLLLL [Rap66]. Load [Chi59, Kle64, Neu62b, Sin62]. Load-Sharing [Chi59, Neu62b, Sin62]. Loading [Chu67a, Tho68b, Was68]. Loads [Ham60]. Locating [Lee62b]. Location [HM65a]. Loci [Kno69]. Locked [OKL59].

LOCS [Zuc65]. Logarithm

[CVV65, Has66, Ste69]. Logarithmic [CET62, Far62, GB55]. Logarithms [Kri63, Mit62b]. Logic [Ake61a, Ake61b, Ake65b, AG68, Ano64u, ACGN60, Arm65, Arm66, BCE61, Ber60a, BS59, BP59, BB56b, Bow67, BS67, CM62a, Car64, Cat66, Cha59a, Cho63a, Chu69b, Coa63, CL61, Coo64b, Cos69, Cos63, Cow64, CG65, DM68, DF67, DS65, DS69, Dun65, Eas58, Ell65, Fra59, FY69a, FY69b, FF63b, FK66, GBC69, GPW55, Gir66, Gol57, GS65, Gra64b, Gur65, Ham65a, Han63c, Han63b, Har57, Har58, Has69, HM67b, Hay61, Hay69b, HM65b, Hor63, Hor65b, HC68, Hru69, Hug67, Jar63, Jos66, Kau69b, KMM69, Kin59, KW69a, KI63, Lar61, Lin60, Loo65, LW65, LK60, Man64a, Man64b, Mar69b, Mar54, Mau67, Mck69, Men69a, Mer64a, Mer67b, MW62, Min61, Min64, Mis66, MWR68, Muk69a, Muk69b, NBC60, Neg64, New66a, Ort59, Pie63a, Pre64b, Pre63, Pri65, Pro64]. Logic [Rao68, RBS67, Roy67, Roy68a, SI59, Sch64a, SD68, SM69, Sci66, She65a, She65e, Skl60a, Smi61, SP62, Smi65a, Smi65b, SH68, Sri64a, Sta65, Ste60, Tod63, TNM64, Tor69, WM59, Wil63, Win68a, Woo63a, YIK68, YI68a, YI68b, Yam62a, YO68, Yen68a, YBS61, Zuc65, de 68b, Bit62]. Logic-in-Memory [Kau69b]. Logical [ABSH63, Ano66-33, Ara57, Ard63, Bar61, BdT65, Bro55b, Col61, Cra61, Dav69d, Els63b, Fel65, Fis63, Fos68b, Fra57, Fra58, Fra64b, GHL^{+}66, Gos65, GMN^{+}60, Got64, GRU58, Har66a, Har63a, Hel63, Hoa55, Hol61b, Kau61, Kau66c, Kel68, Kli69, LC63b, Luk63, MS58, Mas57, MK67, Miy63, Miy64a, Net58, Net59a, Neu62b, PS68, Rud65, SSPR60, Sus61, Svo63b, Tod62, Toh64, YR65a]. Logically [Bla58]. Logics [Liu64b]. Logs [Sas57]. London [Str62]. Longest [Dav65]. Look [Lee69, SG69]. Look-Ahead [SG69]. Look-Aside [Lee69]. Lookup/Interpolation [AK69]. Loop [Bie68, CM59, Gle68, Hay61, MF69, Ott58, Swi63, Tei55]. Loop- [Ott58]. Loop-Free [MF69]. Loops [Dav65, Hen64b, Wal53]. Loss [Cat66, Chu67a, Tho68b]. Losses [Ano69o]. Lossless [Eve65b]. LOTIS [Sch64a]. Low [Chi63, Chr61, HL57, Rap66]. Low-Cost [Chr61]. Lower [Lew66b, Mur65b, MT66, Vai69, YI65]. LSI [Yen69]. Lukasiewicz [Val67].

M [Auf62, Gil62, Uhr62]. M. [Auf62]. Machine [BFK63, BSK67, Bou68, CN68, Cla56, Cra64, Cro61, Fir63b, FY66, Gil65b, Gin59c, Goe64, HM65a, Har67a, Har63e, Hen66, Hig63b, Iba67, Luc69b, Mul64, Net58, Net59a, Ros65a, Ros61, Sch64a, Sto66a, Sus58, Tod65, Wig63, Zuc65, Ros62]. Machine-Tool [Sus58]. machinery [Mar54]. Machines [Ano56r, Ano68-41, Ano69-53, Ard63, Arm62a, Arm62b, Arm62c, AH57, Auf58, Bra66, Bro65a, BP63, BD64, Car68, CE65, Col69a, Cur69, Dab69, DB66, Dav68, Eve65a, Gai68, GG69, Gil61a, Gil63b, GA58, Gin59a, Gin59b, Gin62, Gra64a, Gre69c, HC69, HS67b, Har68a, Har61b, Har62, HS63a, Har65b, Har66e, Hum63, Kam61, Kar64a, Kin66b, Koh64b, Koh65b, KL67, Lav65a, Lav65b, Lew63a, Lew65b, Mas66, Net59b, New67, New68b, New68a, New68c, Nic65b, Pug65, Rib67, RA59, SH61, SY68, Tra66, Vai69, Wak69, Wan69, WS67b, WS67c, WD69, Yau65, YW66, Yoe61, Yoe63a, Zah66, de 68a, de 69]. Magnet [BHRS61]. Magnetic [AH62, AD63, ACGN60, Bal62a, BB56a, Ben60, BCE61, Bla56a, BP59, Bro60a, Car60, CM59, CM60, Che58, Chr61, Chu66a, Chu67a, Cra60b, CV61, Cra61, Cro64, Ein61, Ein62, FR58, Gab64b, Gau64, Gil64b, Gra58, Gre68b, Hay61, Hoa55, HB60, Hsi62, KH63, Kno67b, Kos62, Kum62, Lan64a, Lin63, Lot69, MV55, Mar65a, Mat66, MS59, MN61, Min62, MH59, Mor59, NP64, Pet65, PM60, Poh66, Pry58, RC64, RC57, Rea63, Ros58, SSPR60, Sei67, She62, She63c, Sho61, Sie65, Smi61, Tho68b, Wid54, Wij68, Yao66, You57, Hay52]. Magnetic-Core [Chr61, Hsi62, Lin63, MS59, Wid54, You57]. Magnetic-Drum [MV55]. Magnetic-Field [Lot69]. Magnetically [Mar65a, New66a, Pet65]. Magneto [FDBW63, SH66c, SH67b]. Magneto- [SH66c]. Magneto-Optic [FDBW63]. Magneto-or [SH67b]. Magnetostrictive [AJ60, Gli62, Har61a, HH62, Rot61]. Main [SG69]. Maintenance [AW56, LWCL61, MA63a]. Maitra [Fan68]. Majority [Ake61a, ACW64, CL61, GR68, Hor63, Hor65b, Lar61, Lin60, Meo66, Mii61, MW62, Mis66, Miy63, Miy64a, Neg64, Pri65, Ris67, Toh64, VO65]. Majority-Decision [CL61, Lar61]. Majority-Logic [Hor63, Lin60, MW62, Mis66, Pri65]. Making [Das69, Gaf68]. Malfunctions [BFK63]. Man [Bro65b, Mil57, Ros65a]. Man-Computer [Bro65b]. Man-Machine [Ros65a]. Management [CW68]. Manchester [Che68c]. Mandrel [SB61]. Manipulate [Ghi62]. Manipulation [Ake61a, FH63, Gel61, Gre61, GO63]. Manual [Cro61]. Manuscript [Ano55o]. Map [Kel63, KMM69, MS69]. Mapping [AP66, Art61, PAB68, PF69, Sam69a]. Marcus [CB63]. Margin [Han62]. Marginal [AW56]. Marill [Uhr62]. Mark [Bla62]. Markov [CM68, GO63, SK63]. Masking [CS67]. Matched [Sie65, Tsu62]. Materials [Mor59]. Mathematical [Chu67b, Dun65, Gre68a, Mei69, Mei66, SO68, TNM64, TPM65, Wee69, Yen68a, de 68b]. Mathematics [LC63a, Ste66]. Mating [Fis66]. Matric [Vic67a]. Matrices [BB62, Bro63a, Cam57, Chi59, CM61, DF68a, Ear68, GO63, Hsi62, Nat58, SP63, Urb63]. Matrix [Car60, Chi59, CB63, Chr61, Hil66, Led59a, Leh61, Lyn69a, May65a, McQ62, Neu62b, Pra69, Sed68, Urb65a, Wij68]. Maximal [Har62, Meo65, New68b].

Maximum [Ake68, Ano69-54, Kor64, Loo66, MTI67, Mur65b, Ros68]. Maxium [BE69]. Maxterm [DC65]. Mayne [Hoa62]. Mazes [Sut69]. MC [GJN59, Car64, Coo64b]. McCluskey [Cha68]. McGraw [Str62]. McGraw-Hill [Str62]. Mean [Fre67a]. Means [BHK62, Ein61, Ein62, Gre56, How56, HG57, Min62, TR58]. Measure [ES64, Man64a, PVW59]. Measurement [Ber60b, Han63a, Jac68]. Measurements [Gon65, Rid66, Wat67]. Measures [Sci66]. Measuring [Han62]. Mechanisms [Sto63]. Mechanization [Ein61, Ein62, Min62]. Mechanizations [HM63]. Mechanized [Bar66, CB62]. Medical [Led63]. Medium [FK65, GHI^{+}63, Kos62, MH59]. Mediums [Kar69]. Meeting [Bla56b]. Megabit [ABC61, McQ62]. Megacycle [WS56]. Megapulse [GM63a]. Members [Koh64a]. Membership [Ano63x, Ano63y, Ano63z, Dav63a, Fir63a, Har63d, Kar63b, MO57, Unc59, Unc61]. Memories [Ahr65a, Bit59, Bla56a, Cho67, Dan66, DF68a, ED67, Fos68a, Gra69, Hil66, Hob66, Kau66a, Kle53, MH58, PT66, PM60, Poh66, SC60, Sei64, Wid54, Wil65b, Wol69, Yao66]. Memory [AA66, Alo64, AW67, BFL^{+}66b, BHRS61, BB56a, BG56b, BB62, Bro66, BAL61, Che68b, Chr61, Chu65a, Cof68, CG65, Cra68, CS66, Cur69, Fil67, Flo68, FH63, GL65a, GT68, Gil60a, Gil60c, Gil65b, Gil66d, Gol64, GK62, GMPS59, Gra60b, Gra64c, Gre68b, Han66, HP64a, HH68, Hig66, Hoa62, Hsi67, Joh61, Jon69b, KH63, Kau67a, Kau69b, KI65, Lan64a, Lau68, LM62, Lee69, Lei57, Len65, LBG65, Lot69, Lyn69a, MMT66, Mas66, Mat66, McG68, Mei65, MS59, Mer56, Mez61, Mii64, Mil57, Mit62a, MS63, Nag63, NT69, New67, New68b, Ols65, Pas67b, PM58, RC64, RR67, Ros67b, Ros66b, Rux69, Sal64, Sei67, SG69, SH66c, SH67b, Smi68c, Tay65, Til60, TF64, Vai69, Waa67, Waa68, Wee63, Wee69, WD69, Wom65, YY66a, YY66c, You57, Hay52]. Merit [Hor65a, Mul59b, Sci66]. Merits [Sch63b]. Message [Unc64]. Metallic [Lam59]. Metallic-Tape [Lam59]. Method [Ake61b, And62, ALS68, Ble61b, BB62, Boo62, Bou68, Bri66, Bro60b, Bro68, Bru57, Cho63a, Cho62, Chu69b, Con67a, Cow64, Dab69, Das69, Fer67, Fer69, FR58, Fos68b, FG61, Fri67b, Ger68, Gim65a, Gra56, GL65b, HL57, HO67, HK60, Hig61a, Hig61b, Hor65b, Hor69, JA67, Kar55, Koh63, Lar66a, Lee62b, LE57, Luc66a, MO65, MK67, Miy64b, Nat64, O’G67, O’h69, Per63, Pet61, Pol57, Ree65, She64b, Smi68a, Ste60, Sut69, Tay62, Tom57, Tsu61, Uhr61, Ull69b, UM56, Van57, Ver64, Wai66, Wat59, Wie65a, Wie65b, Wil63]. Methodical [Dor69]. Methodology [Fif68a]. Methods [Auf62, Bek65, Bek67, Bri67, Cla56, Fif68b, FY69a, FY69b, Gre62, Han63c, Hep65, Hru69, Kor67a, MW62, Mul59a, Nis60, Rob58, Str66, Vic69a, Wai67, Wai68a]. Micro [Bla58, Gra62a]. Micro-Programmed [Bla58, Gra62a]. Microfiche [Ano68m]. Micrologic [Low62]. Microprogram [APP^{+}63, Kam60b]. Microprogram-Controlled [Kam60b]. Microprogrammable [Fra64b]. Microprogrammed [Cur68a, Ger63, Haw67, McK68, Ros67a]. Microsecond [ABC61, BP64, WS56]. Microstrip [Kau67b]. Microwave [BS59, Fra59, HMSG59, KI63, KI65, KI66, Mea59, SI59, Sch69a, Yov59]. Midcourse [Mar68]. Miller [Nol62]. Millimicrosecond [BG56a]. Minimal [Abh58, Abh59, Arm66, BC68, Boo62, CR62, Cam64, EKP67, Gin59a, Gin59b, Gin59c, Gra66, Had67, Har57, Har66e, HM65b, Met65c, Mez61, Net59b, Nic65b, Smi65b, UM56]. Minimal-State [Gin59a, Gin59c]. Minimally [Met62]. Minimization [AG68, CE65, DF67, Eve65a, Gil60b, Gim67, Law61, Law62b, Leh63a, Luc66b, Mar57, Mei67a, Mer66, MP64, MP65, MC64, Nec68b, PW67, Sch63b, Tis67, WD68]. Minimize [LLT69]. Minimizing [Bou68, GL65b, Nar61, PU59]. Minimum [Ake68, Fra58, Law62a, Oli65, Roy61, Rus65]. Minimum-Amplitude [Roy61]. Minuend [Lan69a, Lan69b]. Misalignment [Kon60]. Misleading [Uhr61]. Mission [HM67a]. Mixed [LS66, WD69]. Mixed-Font [LS66]. Mod [Sch69b]. Mod- [Sch69b]. Mode [Ben69a, Eck65, Kor63a]. Mode-Control [Eck65]. Model [BE69, Chu67b, CS68, Cof68, Cro64, Dar68, Fin61, Gar59b, Gre68a, Koh63, McK68, Nie68, Rea63, Sch69c, Ste62, SO68, Was67, Yen68a, WM53]. Modelling [Str68a]. Models [Che68b, Dun65, Epl68, Har63e, Iba67, ME67b, ME67a, ME69, Mei66, TNM64, TPM65, War58a, dL69]. Modem [How67]. Modern [Led63]. Modifiable [Gra62a]. Modification [Ake67]. Modified [Hal69, Luc59, RH67a]. Modular [Cof68, Meo68, Muk69a, Wag66]. Modulation [Boh60, Kor63b, Ort59]. Modulator [Kam60a, SOS56]. Modulator-Switches [Kam60a]. Module [Hel66a, LLT69, SG69]. Modules [Wag66, Wan67]. Modulo [EKP67, Men61a, Men61b, Rot59]. Monitor [New68c]. Monitoring [Tsu61]. Monolithic [FK66, SL65]. Monotonic [YI68a, YI68b]. Monotonicity [Liu65]. Monte [Han67, Kor67a, Lit68, Van66]. Moore [Cha65a, Gil61b, New68a]. MOS [Pas67b, PT66, Whi69c, Yen68a, Yen69]. Motion [CM68]. Multi [Har66c, Har66d, Liu66a, Pic66a, SG69, Nol62]. Multi-Level [Nol62]. Multi-Module [SG69]. Multi-Purpose [Pic66a]. Multi-Threshold [Har66c, Har66d, Liu66a]. Multiaccess [NT69]. Multicategory [Smi69]. Multicellular [Cly59]. Multichannel [Roy61, Wol62]. Multiclass [WF68, Wee68]. Multicomputer [Kud64, MA63c]. Multics [Dre67, McG68].

Multidimensional [Hay52, PAB68].

Multifont [Liu64b]. Multifunction [FK66]. Multigate [GBC69]. Multigrid [SOS56]. Multilayer [Gon63]. Multilevel [Ake65b, BS64, Haz64, Man64b]. Multipath [Bal62a, LM62, Nit64]. Multiple [Abe69, AG68, And63, Bar61, Bla56a, Chu69a, Coo63a, DM68, FG61, FM68, GHL^{+}66, GNV59, Has69, Hel66a, Hil66, Isb63, Mii64, MP65, New66b, Pra66b, Ros65a, SC60, SD68, SH66b, SD69, Van60b, Was67, Was68, WD68, Wei63b, Wol62]. Multiple-Category [Coo63a]. Multiple-Cockpit [GNV59]. Multiple-Coincidence [Bla56a]. Multiple-Input [FM68]. Multiple-Module [Hel66a]. Multiple-Output [Bar61, DM68, Has69, SD68, SD69, Van60b, WD68]. Multiple-Valued [AG68]. Multiplex [Cos69]. Multiplexed [BC68]. Multiplexer [GRT55]. Multiplexing [Raw59]. Multiplication [EGP56, Fen69, Gar59b, Har68d, Kri63, Leh57, MR65, Mit62b, Rob55a, Win62]. Multiplier [Atr65, Azg68, Cro60, DR69, Fer67, Fer69, FP54, Fre67a, HG57, Jac64a, Kal57, KS61, KC54, KM61, KB64, Lil56, Mas67, MB52, Nat65, NM68, Pfe59, Pie63c, Sch58, Wal64, Wei57]. Multipliers [SOS56]. Multiplying [Gil59a, Rob67]. Multiprocessing [New66b, Nol64, RR67]. MultiProcessor [Kle64]. Multiprocessors [Ros67b]. Multiprogram [APP^{+}63]. Multiprogrammed [Sto66b]. Multiprogramming [Jac68, New68c, Nol64, Wel68]. Multipurpose [Dia56]. Multistep [GG68]. Multitape [Gre69b]. Multithreshold [HD67, HO67, MF68a, MF68c, MG66, Nec68a, SP69, Yen68b]. Multivariate [Coo63b, FK69]. Multivibrators [Dav57a]. MUSASINO [MT59]. MUSASINO-1 [MT59]. Muscle [Rub68]. Musical [BHNW57]. My [Ano55o].

N [Flo62, GH59, Shu62, GH59]. Names [Mar65b, Nad65]. NAND [BvB65, Dav69a, DS69, Ell65, Klo61, Pow65, Smi65b, Tod63]. Nanosecond [Ort59, SP62, SL65]. Nanoseconds [Sei64]. Narrowband [Man68]. Natural [Sal65]. Nature [Auf62]. NCR [Hig66]. NDRO [MMT66]. Near [FF67, Kno67a, RH66, Sil68b]. Near-Optimal [Sil68b]. Near-Optimum [FF67, Kno67a, RH66]. Nearly [Arm66]. Needed [Loe64]. Negation [Har63c]. Negative [Die63, Lip64, RR66, Son63, Wad57, Wad61]. Negative-Base [Son63, Wad61]. Negative-Bit-Weight [Lip64]. Neighbor [Cho62]. Nerve [Fit63]. Nets [Arm66, Bre65b, Cam64, Coa63, Fuk64, HSA57, Hug67, Mei68a, Set66b, Set66a, Set67, Urb63, Urb64, Urb65a, Urb65b, Urb68, YIK68, Yam62a]. Network [Ahr65a, Bat65, Bow69, BS67, Dav69a, Ger68, Gir66, Gur65, Hel55, IS65, Kel68, Kli69, Kor62b, LC63b, MO65, She65a]. Networks [Ale67, AC67, ACW64, ACGN60, Bar61, CR62, CB63, CL63, Cun54, Die68, DS69, Dru69, Fer68, FG69, Gil60b, Gim67, GL67, GHL^{+}66, Gos65, Gou58, GL65b, Hen64b, HM65b, HF57, Hor65b, HC68, KT65, Kil62, KW69a, Klo61, Kor62a, LLT69, Lec64, Lev63b, LWM64, Lin60, Mai62, McC58, Mcg65, Mei67a, Men69a, Meo66, Mil58, Mis66, MG66, Neg64, NB65, Ott58, PS66, Pre64b, Pre65, Pri65, Roy68a, Sau67, Sho62, Skl63a, SKS65, SW65, Van60b, VO65, VB65, Wai64b, WD68, YO68]. Neural [Urb63, Urb64, Urb65a]. Neuristor [Cra60a]. Neuron [Die68, NW64]. Neuron-Like [Die68]. Neutron [Gre66a]. News [Ama64a, And64, Ano55g, Ano55h, Ano55i, Ano55j, Ano56h, Ano56i, Ano56j, Ano56k, Ano57g, Ano57h, Ano57i, Ano57j, Ano58m, Ano58n, Ano58o, Ano58p, Ano59r, Ano59s, Ano59t, Ano59u, Ano60v, Ano60t, Ano60u, Ano61x, Ano61y, Ano61z, Ano61-27, Ano62g, Ano62r, Ano62s, Ano62-27, Ano62-28, Ano62-29, Ano62-30, Ano62-31, Ano63q, Ano63-34, Ano63-35, Ano63-33, Ano63-36, Ano63-37, Ano63-38, Ano64g, Ano64h, Ano64i, Ano64j, Ano64y, Ano64z, Ano65i, Ano65j, Ano65k, Ano65l, Ano65m, Ano65n, Ano66g, Ano66h, Ano66i, Ano66j, Ban63, Bau64, Ber66b, Brz64a, Dag65, Den63, Den65, Der65, Epl63, Fle63a, Fra65, Gil63a, Kim66a, Lew64, Liu66c, Lo63, MMB66a, Poh64a, Rab63, Sag63, Seb63, Sur63, Tar65, Ung63b, War65, Win66a]. Newsletter [Ano58t, Ano58u, Ano58v]. NLI [NK66]. Node [Bro63a, Ott58]. Node-Analysis [Ott58]. Node-Elimination [Bro63a]. Noise [Cat67, DH69, Gil66g, HKM63, Hen67, Mat68, Pow59, RW67, Ses63, Tsu62, Whi67, Yao65]. Noise-Matched [Tsu62]. Noiseless [Neu62b]. Noiselike [EG61]. Non [And63, AS66, Ben69a, Kru64, LW59, SE66]. Non-Destructive [Kru64]. Non-Linear [Ben69a]. Non-Parametric [SE66]. Non-Real-Time [LW59]. Non-Repetitive [And63]. Nonbulk [YY66b]. Noncatastrophic [Hol63a]. Nonconstant [Oli65]. Noncontact [She62]. Noncontrollable [Bir67, Mar66b, Wit67]. Nondestructive [Ear64, Gra60b, KI66, Lam59, Mat66, NP64, Sho61, Til60, Wid54]. Nonlinear [Bro69b, BK62a, BK62b, Cad68, Fuk64, How56, Kar64b, Kas63, Koh63, KC58, KC60, Kuo68, Leo63b, Mow69, MG66, Nat63, Ogb69, RY66, Rub65, Sam69a, SS68, Tay69, Tho69, Vic69b, Whi66, Yoe63b, de 68a]. Nonlinear-Summation [Kas63]. Nonobservable [Bir67, Mar66b, Wit67]. Nonparametric [HF69]. Nonquiescent [Wad68]. Nonredundant [Sem63, Sza62]. Nonreturn [Hoa55]. Nonreturn-to-Zero [Hoa55]. Nonsaturated [Gol60].

Nonsaturation [Hop64]. Nonstable [Lav65a]. Nonstandard [Cha69]. Nonstationary [Lar66a, Wie65b, WK68]. Nonsupervised [IT67]. Nonsymmetric [AE61]. Nonwriting [Gre69b]. NOR [Bat65, DS65, Ell65, Kin66b, Klo61, MW67, Mas60, Mau67, Pow65, Smi65b, Tod63]. NOR-Gate [MW67]. Normal [CM65, CD66b, Hoc62, Mot60]. Norman [Low62]. NORNAND [Fan68]. Notation [Klo61, LS60, Wan65]. Note [AM67a, Ben60, CM63, Coo66, EGP56, Fro65, Fuc57, Gel61, Gil61b, Got64, Gou58, HM65a, Hig61c, Hig61d, Ito69, KS67, Kli67, Lee62a, LB60, LWM64, Mas66, MU59, Mei67a, Mil57, Nam64, New67, PW67, RC66, Tay69, War58b, Woo63b, de 68a]. Notes [Bro55b, Hen64a]. Notice [And67b, Ano53d, Ano66n, Ano68z, Ano68-27, Ano68-28, Ano69-28, Ano69-29, Mar66a, O’N67a]. Notices [Ano59s, Ano59t, Ano59u, Ano60q, Ano60r, Ano60s, Ano60v, Ano60w, Ano61t, Ano61u, Ano61v, Ano61w, Ano62u, Ano62v, Ano62w, Ano62x, Ano62y, Ano63-27, Ano63-28, Ano63-29, Ano63-30, Ano63-31, Ano63-32, Ano64v, Ano64w, Ano64x, Ano66-29, Bow66, Byr65, Cho63c, Dal65, Emb66, Gli65, Gre64a, Joh63, Kov64, Len63b, Leo63a, Mel66, Mul63, Pet64, Raf64, Ral64, Tsy63, You64]. Notions [dL69]. Novel [Rao58]. NRZ [BS63]. ns [Sei67]. Number [Avi61, BB69, Bat65, Bou68, Che61b, Flo56, Gar59a, Gar65, GL65b, Guf62, Har63c, Har67a, Has66, Law62a, Lec64, Leh63b, Lew66b, Luc66b, MO65, MU59, Mer64b, MW64, Mul59d, Mul60, Mur65b, MT66, PU59, PS66, RH67a, Sas67, Sas68, Smi66, Son63, Tay62, Tod62, Wad57, Wad61, WS67b, Wil64, YI65, KCT62]. Number-Representation [Son63]. Numbers [BC66, Bar63, CVV65, DS62, Eve64, Ger60, Kam61, KLP68, McK67, Rot59]. Numerical [MC64, Nec67].

O [Ano59x, Bit62, MS66]. Obituary [Ano62z]. Oblate [CM59]. Observatory [Lew63b]. Obtain [PVW59]. Obtained [DH69]. Obtaining [Boo62, Bro60b, Gel62, LE57]. OCR [Lar69]. Octal [Roz62]. Odd [Rob56]. Off [Gre69c, WK68]. Off-Line [Gre69c, WK68]. Office [MV55]. Officers [Koh64a]. Ohmic [Ano69o]. On-Line [Bro64, Bro65b, Der67, DKP69, Fre68, GHR67, Hen66, Lun63]. One [Agg65, Atr65, Gre69b, Gre69c, KELS62, NP56, PAB68, Uhr61, WS56, Wol62]. One-Amplifier [Agg65]. One-Dimensional [Atr65]. One-Level [KELS62]. One-Megacycle [WS56]. One-Microsecond [WS56]. One-Tape [Gre69c]. One-Way [Gre69b]. Only [Gam68, Kil59, Waa68, Hsi67]. ONR [Yov59]. Operating [Bas63b]. Operation [Ara57, Eng61a, Eng61b, Luk67, Mei65, SG69, Stu68a, Thi63, Thi64, Waa68, WB62]. Operational [Agg64, CM62b, CM63, Cly59, DEL55, Fra64a, Gli66b, GNV59, Han62, Hel55, How56, KP67, Kor62a, Kor63a, Kur64, MC62, Mye57, Pfe59, Sch63a, Wad64, Wad68, MGN54]. Operations [Bro60c, Els63b, Gar66, Gra62b, Luc59, MA63b, Rao68, Sas68]. Operator [Har69b]. Operators [Gil66e, Tav65]. Optic [FDBW63]. Optical [Bra59, Ros62, SH66c, SH67b, Mil55]. Optimal [Chu69a, Cop68, Gil60c, Maj67, Man63a, MTI67, MC69, NK66, Ros68, Sil68b, Wei69b]. Optimization [Bek67, Dar68, Den69a, Dru69, FG60, MR59, Pie63c, PLS61, Sas69, Sch69c, SS68, Wag67, WS67c]. Optimizer [Nel64]. Optimizing [AD69, Leo63b]. Optimum [Cha65b, Cho57, Cho59, CP68, Dei65, FF67, Flo58, Hig61d, Kno67a, RH66, SC67]. Or-Invert [Hel63]. Orbital [FH62].

Orbiting [Fog67, Lew63b]. Order [Agg64, BH65b, Bro66, Eve65b, Gre62, HS69c, Kri66, Liu63b, Liu64a, Mei66, Pra69, SW61, TPM65, Wad64, Wad68]. Ordered [Wei63b]. Ordering [SH66a, Sho64, Sil68b, Svo67]. Orders [Tay62]. Ordinary [Hau62a, Hau62b, Whi66]. Organization [Ano66-30, Fra64b, Gil60c, GNV59, HM63, Hol61a, Kim66b, KW69b, Kop69, MA63a, MS63, Wig63, LA54]. Organized [Pas67b]. Organizing [FI65, Jac64b]. Oriented [Cam69, Cur68a, DS65, Gas64, Haw67, Law68, Mcc67a, Mul64]. Origin [Bro68]. Orthogonal [Chi59, Col61]. Orthonormal [Gil59b]. Oscillation [Urb65b]. Oscillator [OKL59]. Oscillators [MWR68]. Oscillatory [Die68]. Oscillographic [Luk67]. Other [Car60, Fia66, Ros65b]. Outline [Luk63]. Output [Bar61, Bro60b, DM68, Dor67, GHL^{+}66, Has69, Isb63, Koh65b, MP65, New67, New68b, New68a, PAB68, Pra66b, SD68, Sho62, SD69, Tru63, TR58, Van60b, WD68, Wei69b, WP62]. Outputs [Wan69]. Over-All [Kil59]. Overflow [Leh63b, KCT62]. Overlapped [Fen69]. Overview [Dre67].

P [GH59]. P-N- [GH59]. Packages [Ros58]. Packard [Sha63]. Packing [Gil64b]. PACTOLUS [Gas65]. Pad [Poh66]. Padé [Zac61]. Paged [SS66]. Paging [Bro69a, Cof69b, SSG69]. Pair [BW63, CM62a, Sma68]. Pair-Graphs [Sma68]. Pairs [Fri66b, Zah66]. Paper [Mas57]. Papers [Ahr65b, Ake65a, Ake66, And67a, Ano56l, Ano59v, Ano59w, Ano60x, Ano60y, Ano60z, Ano60-27, Ano61-28, Ano61-29, Ano61-30, Ano61-31, Ano62-33, Ano62-34, Ano62-35, Ano62-36, Ano62-37, Ano63h, Ano63i, Ano63j, Ano65h, Ano68-42, Ash65, Avi63, Axe63, Can67b, Car65, Cha67, Epl66, Fir64b, Gil65a, Gol63, Hop67b, Jon64, Lev63a, Man67b, Mil66, Min63, Ros64, Sci64, Ses65a, Son64, Sri64b, Tai63, Tra67, Urb66, War66, Yau66]. Parallel [Ale67, Avi61, Ber66a, Che58, Deo68, EBTB63, Fer67, Fer69, FG61, Gil66c, Gir69, Gol69, HM63, HM67b, Hel66b, Hen60, HS64b, HS69b, KF66, KW69b, Lin66, NF64, New64, Pas67b, QK67, Rob55a, SH66b, Ste69, Ver64, Yoe63a]. Parallelism [BE69]. Parameter [Bek67, Kro59, Nel64, Pow59, Str66, Str68b, SS68]. Parameters [Lev68, May64, Mei66, TPM65, Win69]. Parametric [Hau62a, Hau62b, HMSG59, HH57, MWR68, OKL59, SE66]. Parametron [Ran64, MT59]. Parametrons [SSPR60, Ulz66]. Parapropagation [Glu65]. Pardo [Hoa62]. Parenthesis [LS60]. Parenthesis-Free [LS60]. Parity [Gar58]. Part [Kil59, CV61, Cra61, Cur69, JO68a, JO68b, Man66b, Man66c]. Partial [Ben69a, Den69a, Gre69a, Han67, Kar64b, Kor67a, Lit68, Muk63, Ros65b, Tay65, Vic69a, Whi69b, Atk68]. Partial-Differential [Kor67a]. Partially [AH63, Mar57]. Partially-Developed [Mar57]. Partition [Zah66]. Partitioned [Mor67]. Partitioning [HF69, Smi68a]. Partitions [Dab69, Gra66, Sma68]. PASCAL [HS61]. Pass [HF66, Lar66b, Mul59b, Sze67]. Passing [FR58]. Passive [Gam68, Sie65]. Past [Gil66d]. Patching [Han68]. Path [Ake67, Cra68, Lee61, Lee62a, ME69]. Paths [Dav65, GG69]. Pattem [Fis64]. Patteme [Who63]. Pattern [Ama67, Ano69p, Bar66, Ble61b, BB62, Ble66, Bon66, CF68, Coo64a, Cov65, Das69, DF66, FF67, Fis63, FG64, FCC67, Glu65, Gol69, HF69, Hig61b, Hig61d, Hug67, KS63, Kno67a, Lew65b, MG60, MFF63, McC63a, Mei69, MG66, RH66, SE66, Smi68b, Smi69, Spe67, Ste60, Ste62, Uhr61, Uhr62, Ull69b, WF68, Wee68, YY66c, YS68, FF63a, Kno67a]. pattern-Recognition [FF63a]. Patterns [Kir64a]. Paull [Ree65]. PB [Fra64b]. PCM [AJ60, Che68a, Eps68]. PDP [VB69, VBW69]. PDP-8 [VB69, VBW69]. Pedestal [Seb57]. Pedestal-Free [Seb57]. Pen [Ano68-40, Kon68, Ros65a]. People [Ano63-39, Els63a]. Perceptrons [Nie69b]. Perform [SC68]. Performance [AD69, AW67, Bro66, Cof68, Gar65, Gil69, Jac68, Kor63a, MRS60, Mer64b, RV68, Sci66, Tsu62]. Period [Ber60b, Die68]. Periodic [EG61, Van59, Van60a]. Permalloy [ROK58]. Permanent [Alo64, BHRS61, Fil67, Len65, Wie55]. Permanent-Magnet-Twistor [BHRS61]. Permitting [Eck62]. Permutation [Neu62a]. Perspective [Cof69b]. PERT [Bow67]. Perturbation [BO59]. PGEC [Ano60e, Ano55g, Ano55h, Ano55i, Ano55j, Ano56h, Ano56i, Ano56j, Ano56k, Ano56l, Ano57g, Ano57h, Ano57i, Ano57j, Ano58m, Ano58n, Ano58o, Ano58p, Ano59r, Ano59s, Ano59t, Ano59u, Ano60v, Ano60w, Ano60t, Ano60u, Ano61x, Ano61y, Ano61z, Ano61-27, Ano62-27, Ano62-28, Ano62-29, Ano62-30, Ano62-31, Ano63-33, Goo55, MO57, Sag63, Unc59, Unc61]. Phase [Bla62, Col67, Dou68, HL57, Han62, Han63a, OKL59, Yen68a, Yen69]. Phase-Locked [OKL59]. Phase-Plane [Bla62]. PHASEPLOT [Der67]. PHENO [Ser68b]. Phenomena [MH58]. Philips [HS61]. Philosophy [Lew65a]. Photoelectric [LC53]. Photoresistive [Azg68, Mas67]. Phrase [Lan64b]. Phrase-Structure [Lan64b]. Physiological [Led63]. Picard [Tom57]. Picture [Kir64a]. Pictures [NF64]. Pieces [BBB59]. Piecewise [DF66, Kno67a]. Placement [Gir66, Gur65]. Plan [Bak57]. Planarity [EM66]. Plane [Bla62]. Planning [Buc62, Str62]. Plated [Cho67, Lot69, Waa67, Waa68]. Plated-Wire [Lot69, Waa67, Waa68]. Plot [Wom65]. Plots [O’h69]. Plots-Digital [O’h69]. PNPN [Uzu62]. Point [AK69, HM65a, HM67a, Kli67, Ter62].

Poisson [BP69, Dor67, MMR60]. Polarity [Kin59]. Polyfunctional [Set66b, Urb65b, Urb68]. Polylinear [Cur68b]. Polynomial [MPP59, Ran63, Spe67]. Pontryagin [Kor64]. Portable [Lew67b]. Position [Boh60, Gil59a, Val67]. Positioning [FS65]. Positive [AM67b, Cro68, Die63, KP67, Kri64]. Possibly [Uhr61]. Potential [Mei69, PT66]. Potentiometer [PD62, SB61]. Powered [BGR66]. pp [Str62]. Practical [Neg64, Sed68]. Pre [Tsu62]. Pre-Amplification [Tsu62]. Precision [CKLP60]. Predicted [Con67a]. Predicting [Dun65, Gra64b, Pie63a, TNM64]. Predictor [BBP57]. Preemption [Cof69a]. Preemptive [MC69]. Preface [Ano65r]. Prefix [Cut68]. Preparation [CLS64]. Prescribed [Gim65a]. Presence [CM65, CD66b]. Present [Gil66d, Hob66]. Pressure [Kas62]. Prewired [But64]. Primary [Lew63b]. Prime [Ano66-31, DC65, DBC69, Fri67b, Gim65a, Gim65b, Hoc62, Luc66a, Luc69b, Mor67, Mot60, Nec67, PM62]. Primitives [LW65]. Principle [Kor64, MTI67, Ros68]. Principles [Ade62, Cra64, Man66c, Min69]. Print [LS66]. Printed [CN66, FR58]. Printing [Dau66]. Priority [Bow69, Fos68a]. Probability [Cof68, SW61, Van57, War58a]. Problem [Arm62c, CS68, Cur68a, Els64, Fel63, Fis64, FG64, FL67, Fro65, Gas64, GHL^{+}66, Har61b, Haw67, Lee62a, Nil69, PMC67, RH67b, Set66b, She64a, SH61, Vic67b]. Problem-Oriented [Cur68a, Gas64, Haw67]. Problem-Solving [Nil69]. Problems [Bek65, Brz65b, Can66, ED67, Gli66b, Gre62, Hau65a, Kov66, Lan64b, McQ62, MTI67, Rob55b, Ros68, Ter62, Van57, Vic69a]. Procedure [CKL62, CL63, Coo63a, HD67, HF69, Lew66b, Mau67, Mer66, Mul64, Nec67, Pow69, WF68]. Procedure-Oriented [Mul64]. Procedures [Cur69, HM67a, MC64, Mow69, Smi69]. Proceedings [Ano66o]. Process [BR69, Glu66, Gun58, Leo63b, Rub65, RV68, Sas69, Tru63, Wat67]. Process-Control [BR69]. Processes [McG68, SS61, SS62]. Processing [AM62, Ber66a, BW64, BBB59, Che68a, CG65, Dav69c, EBTB63, FK66, Gab64b, Gaf68, Gal62, Gre64b, Hau69, Hel66b, KF66, Lan63b, Lun63, Mur65c, MS63, NF64, New64, Ric67, Sal63, Sam69b, Sto63, Sus58]. Processor [HH68, Lew63b, Man66b, MC69, Rao68, Rus68, SG69, Shi68, WP64, Wig63]. Processors [Flo68, Lau68, YY66b]. Producing [Gim65a]. Product [Fle67, Win68b]. Production [Ano57p, Wai64b]. Products [Abh58, EKP67, Sch69b]. Professional [Ano55a, Ano63x, Ano63y, Ano63z, Dav63a, Fir63a, Har63d, Kar63b]. Program [Bas63b, Bol67, Cam64, Don65b, Fir65, Fis66, Fis63, Fog67, Gas65, Gra62a, HK53, JSS66, Kuc68b, Leo63b, Mcg65, Ses65b, Tsu61, dM66]. Program-Modifiable [Gra62a]. Programmable [BMRS69, Mow69]. Programmed [Arm62b, Bla58, Gra62a, Har56, Liu64b, RBS67, Vid67, WP64]. Programmer [Rap66]. Programming [Aho68, Ano66-35, Ano69p, AW56, BFK63, Bit66, Bre65b, Bro69b, Flo64b, FCC67, GM62, Knu65a, Kuc66, Kuc68a, Law68, MA63a, Mcc68b, Nol62, Sch65a, Sch65b, Smi68b, Ste63c, Sto66b, Sut68, Tal64, Tou60]. Programming-Language-Oriented [Law68]. Programs [Ber66a, Bro69a, Chu65b, Kar63a, Mar62, PMRC67, Sed68, Str68a]. Progress [Ano57k, Bro55a, CS58, Kau66c, Kor65, May60, Nas56]. Project [Ros61, Str62, Buc62]. Prolate [CM60, Elf62]. Propagating [Kro59]. Propagation [Ano69o, AS66, Fer68, LB61, Rei60, Yao66]. Properties [AH63, Boo68, Bri61, Cov65, Fos68b, Gab64a, Gin66, Hop67a, Kau62, Kei64, Kro59, Lav65b, Liu66a, Mer64a, MK67, NW64, Pie63b, Pre66b, Rou58, Sob62, Tay65, Wee63, Wei69a, Win65b, Yen68b]. Property [YS68]. Proportional [Bro60b, WP62]. Proposal [CLS64, RC64, Smi61, Wai68b, Woo64a, Bit62]. Proposals [Wei63b]. Proposed [Ano62-32, Gau64]. Proposition [Boo62]. Proposition-Letter [Boo62]. Propositional [Val67]. Prospective [Ano53d]. Protecting [Gol64]. Protection [RR67]. Provides [Sch63a]. ps [Meo65]. Pseudo [Gil66g, Kon68, Whi67]. Pseudo-Light [Kon68]. Pseudo-Random [Gil66g, Whi67]. Pseudorandom [DH69, Mat68, RH67a]. Psychiatric [Ihm63]. PTGEC [Ama64a, Ano63-34, Ano63-35, Ano63-36, Ano63-37, Ano63-38, Ano63-39, Ano64y, Ano64z, Ban63, Brz64a, Els63a, Epl63, Fei63a, Fle63a, Gil63a, Sur63]. Publication [YR65b]. Pulse [BGR66, BG56b, Boh60, Dun65, Far62, Fit63, Gai68, GJN59, Ham58, KW69a, Kor63b, MRS60, Nar60, Pie63c, Rib67, Ros58, Ste63b, TNM64, Was67, Was68, Woo64b]. Pulse-Signal [Was68]. Pulse-Switching [Ros58]. Pulse-Type [KW69a]. Pulsed [Rib64]. Pulsed-Data [Rib64]. Pulses [Cat66]. Punched [Mas57]. Purpose [BC65, Fog67, Fra57, Fra58, HP64b, Kam60b, KM64a, Pic66a, QK67, Ros65b, RG65, Stu68a, Stu68b, Wig63]. Purposes [Car60, GG68]. Push [LTW63]. Push-Down [LTW63]. Pushdown [Gra64a]. Pushdown-Store [Gra64a]. Puzzles [Fis64, FG64].

Quadrant [Pfe59, Sch58]. Quadrature [Lar66b, Sze67]. Quantising [CG64].

Quantities [MP64, MP65]. Quantizer [GB55]. Quantum [Sci66]. Quasi [DH69, Kar69]. Quasi-Analogue [Kar69]. Quasi-Gaussian [DH69]. Question [Col69b]. Question-Answering [Col69b]. Questions [Ano58q, Col69b, SNS69]. Queue [CM68, CS68]. Queueing [Sch69c]. Quine [Chu61]. Quotient [Har69b]. Quotients [Met62].

R [Alo62, Low62, Mac62, Mar69a, Tau62, de 68b]. R. [Arm62c, Gil62, Kau62, Low62, Mit62a]. R63 [And63, Ano63-40, Ard63, Bas63b, Bro63b, Coa63, Els63b, Fel63, Fir63b, Fis63, Fit63, Gil63b, Gin63, Gre63, Hei63, Hig63b, Hum63, Ihm63, Isb63, Kar63a, Kel63, Led63, Leh63b, Leo63b, Lev63b, Lew63a, Man63b, Mel63, Pet63, Pre63, Sal63, Sem63, Sha63, Ste63a, Sto63, Tru63, Who63, Wit63, Woo63a]. R63-104 [Mel63]. R63-105 [Sha63]. R63-106 [Bas63b]. R63-107 [Woo63a]. R63-16 [Led63]. R63-18 [Sem63]. R63-19 [Who63]. R63-20 [Sto63]. R63-21 [Sal63]. R63-31 [Fis63]. R63-32 [Kar63a]. R63-33 [Lew63a]. R63-34 [Gil63b]. R63-35 [Lev63b]. R63-36 [Tru63]. R63-37 [Pre63]. R63-38 [Gre63]. R63-39 [Hei63]. R63-40 [Fit63]. R63-41 [Leo63b]. R63-55 [Ard63]. R63-56 [Coa63]. R63-57 [Hum63]. R63-68 [Wit63]. R63-69 [Isb63]. R63-70 [Fir63b]. R63-71 [Pet63]. R63-72 [Ihm63]. R63-73 [Gin63]. R63-74 [Els63b]. R63-75 [Man63b]. R63-88 [Bro63b]. R63-89 [Fel63]. R63-90 [Ste63a]. R63-91 [Hig63b]. R63-92 [Ano63-40]. R63-93 [And63]. R63-94 [Kel63]. R64 [Alo64, Ama64b, Bed64, Ble64, Boo64, Car64, Cra64, Cro64, Els64, Fis64, Fra64b, Gab64b, Gra64a, Gra64b, Gre64b, Hen64b, Kir64b, Kle64, Kru64, Kud64, Lec64, Les64, Mac64, Nel64, New64, Nol64, Sho64, Sri64a, Sto64]. R64-17 [Gre64b]. R64-18 [Gra64b].

R64-19 [Bed64]. R64-30 [Alo64]. R64-31 [Cro64]. R64-32 [Nol64]. R64-33 [Kir64b]. R64-34 [Kle64]. R64-35 [Els64]. R64-36 [Fra64b]. R64-37 [Cra64]. R64-38 [Les64]. R64-48 [Hen64b]. R64-49 [Boo64]. R64-50 [Ama64b]. R64-51 [Sho64]. R64-52 [Lec64]. R64-61 [Ble64]. R64-62 [Sto64]. R64-63 [Gra64a]. R64-64 [Mac64]. R64-65 [Kud64]. R64-66 [New64]. R64-76 [Nel64]. R64-77 [Fis64]. R64-78 [Sri64a]. R64-79 [Kru64]. R64-80 [Gab64b]. R64-81 [Car64]. R65 [Ard65, Arm65, Bek65, Bob65, Cly65, Don65b, Dun65, Fir65, FK65, Gar65, Gas65, Gea65, Ham65a, Her65, Hig65, Knu65a, Kor65, Len65, Lew65b, Lew65a, Mar65a, Mcg65, Mei65, Mer65, Met65a, Mur65c, Nad65, Ols65, Rub65, Sal65, Sam65, Sch65b, Wir65]. R65-11 [Bek65]. R65-12 [Ard65]. R65-13 [Sal65]. R65-14 [Arm65]. R65-15 [Don65b]. R65-21 [FK65]. R65-22 [Gar65]. R65-23 [Sam65]. R65-24 [Wir65]. R65-25 [Bob65]. R65-26 [Her65]. R65-27 [Mer65, Mur65c]. R65-29 [Fir65]. R65-30 [Lew65b]. R65-35 [Met65a]. R65-41 [Kor65]. R65-42 [Rub65]. R65-43 [Len65]. R65-44 [Gas65]. R65-45 [Mei65]. R65-53 [Ols65]. R65-54 [Ham65a]. R65-55 [Mcg65]. R65-56 [Gea65]. R65-57 [Sch65b]. R65-58 [Hig65]. R65-69 [Dun65]. R65-70 [Nad65]. R65-80 [Knu65a]. R65-81 [Lew65a]. R65-82 [Cly65]. R65-83 [Mar65a]. R66 [Ano66-31, Bit66, Ble66, Can66, Fis66, Gas66, Gli66b, Gre66a, Has66, KF66, Kov66, Kuc66, Lar66a, Lew66a, Mei66, Miu66, New66a, New66b, Rap66, Rid66, Rog66, Ros66b, Sag66, Sci66, Skr66, Sto66a, Sto66b, Str66, Van66, Wai66]. R66-12 [Wai66]. R66-15 [Ano66-31]. R66-25 [New66a]. R66-26 [Ros66b]. R66-27 [Rid66]. R66-39 [Rog66]. R66-50 [Sto66a]. R66-51 [Sto66b]. R66-52 [Rap66]. R66-59 [Kuc66]. R66-60 [Gas66]. R66-61 [Sag66]. R66-62 [New66b]. R66-63 [KF66]. R66-64 [Ble66]. R66-66 [Mei66]. R66-67 [Kov66]. R66-68 [Str66].

R66-69 [Lar66a]. R66-78 [Has66]. R66-79 [Sci66]. R66-80 [Gre66a]. R66-81 [Skr66]. R66-82 [Miu66]. R66-83 [Lew66a]. R66-84 [Gli66b]. R66-85 [Van66]. R66-86 [Fis66]. R66-87 [Bit66]. R66-88 [Can66]. R67 [Bek67, Bol67, Bow67, Bre67, Can67a, Con67a, Con67b, Den67, Don67, Dre67, Fog67, Hop67a, How67, Kor67a, Kor67b, Law67, Mas67, Mcc67a, Paq67, Pau67, Rub67, Sze67, Tam67, Vic67a, Vid67, Wag67, WS67a, Wat67, Wit67]. R67-10 [Dre67]. R67-11 [Den67]. R67-12 [Don67]. R67-13 [Paq67]. R67-14 [Wat67]. R67-15 [WS67a]. R67-16 [Bre67]. R67-24 [Hop67a]. R67-25 [Mcc67a]. R67-26 [Can67a]. R67-27 [Rub67]. R67-28 [Fog67]. R67-29 [Kor67a]. R67-30 [Bek67]. R67-31 [Vic67a]. R67-39 [Wag67]. R67-40 [Bol67]. R67-41 [Mas67]. R67-42 [Sze67]. R67-43 [Wit67]. R67-57 [Con67a]. R67-58 [Pau67]. R67-59 [Tam67]. R67-60 [Vid67]. R67-61 [How67]. R67-62 [Con67b]. R67-63 [Kor67b]. R67-8 [Law67]. R67-9 [Bow67]. R68 [Aho68, Bar68, Bie68, Bin68, Car68, CW68, Cop68, Cro68, Cur68a, Dar68, Ear68, Epl68, Far68, Fif68b, Fif68a, Flo68, Fre68, Gai68, Gle68, Gre68a, Han68, Hor68, Jac68, Kat68, Kei68, Kuc68b, Lau68, Lit68, Mar68, McG68, McK68, Mcc68b, Mur68, New68c, Nie68, Rub68, Sal68, Ser68a, Ser68b, Sil68a, Smi68c, Str68b, Str68a, Sut68, Tho68b, Wai68b, Wel68, Wie68]. R68-10 [Fif68a]. R68-11 [Lau68]. R68-12 [Nie68]. R68-13 [Hor68]. R68-14 [Cur68a]. R68-15 [Mar68]. R68-16 [Gre68a]. R68-17 [Cro68]. R68-18 [Gai68]. R68-19 [Wie68]. R68-20 [Smi68c]. R68-21 [Kat68]. R68-22 [Mur68]. R68-23 [Kuc68b]. R68-24 [Wai68b]. R68-25 [Fre68]. R68-26 [Bar68]. R68-27 [Aho68]. R68-28 [McK68]. R68-29 [Dar68]. R68-30 [Lit68]. R68-31 [Gle68]. R68-32 [Ser68a]. R68-33 [Ser68b]. R68-34 [Sil68a]. R68-35 [Cop68]. R68-36 [Mcc68b]. R68-37 [Flo68]. R68-38 [Bin68]. R68-39 [Rub68]. R68-40 [Car68]. R68-41 [Sal68]. R68-42 [CW68]. R68-43 [Str68a]. R68-45 [Far68]. R68-46 [Ear68]. R68-47 [Fif68b]. R68-48 [New68c]. R68-49 [McG68]. R68-5 [Epl68]. R68-50 [Jac68]. R68-51 [Han68]. R68-52 [Wel68]. R68-53 [Sut68]. R68-6 [Str68b]. R68-7 [Bie68]. R68-8 [Kei68]. R68-9 [Tho68b]. R69 [Ben69a, Ben69b, Bro69a, Cof69b, Col69b, Den69a, Den69c, Els69, Gar69, Gil69, Gre69a, Gre69b, Gre69c, Har69b, Har69a, Hau69, Hor69, Kar69, Kle69, Kli69, Kop69, Lar69, Muk69a, Nie69a, Nie69b, Nil69, O’N69, Sas69, Sch69a, Tor69, Ull69a, Wak69]. R69-1 [Ben69b]. R69-10 [Wak69]. R69-11 [Gil69]. R69-12 [Hor69]. R69-13 [Nie69b]. R69-14 [Har69a]. R69-15 [Sch69a]. R69-16 [Tor69]. R69-17 [Nie69a]. R69-18 [Gre69b]. R69-19 [Ull69a]. R69-2 [Hau69]. R69-20 [Lar69]. R69-21 [Ben69a]. R69-22 [Els69]. R69-23 [Gre69c]. R69-24 [Muk69a]. R69-25 [Kop69]. R69-26 [Bro69a]. R69-27 [Cof69b]. R69-28 [Den69c]. R69-3 [Kli69]. R69-30 [Col69b]. R69-31 [Nil69]. R69-32 [Gar69]. R69-33 [Den69a]. R69-4 [Kar69]. R69-5 [Kle69]. R69-6 [Sas69]. R69-7 [O’N69]. R69-8 [Gre69a]. R69-9 [Har69b]. Radar [O’h69, Sch69a]. Radiation [Che68a, Mau67]. Radio [Ses63, Wid54]. Radio-Frequency [Wid54]. Radio-Noise [Ses63]. Radix [Ano68-46, Cro61, Die63, Goe64, Roz62, Whi69b, Atk68]. Rado [Bra66]. Ragland [Dav62]. Rail [Dvo68, Yoe65]. Random [Bar63, Boo68, Bri67, CM68, CS68, DS62, Dor67, Gai68, Gil66g, HKM63, Hen67, Hoa62, MR59, Rib67, SO68, Til60, Wat67, Whi67]. Random-Access [Til60]. Random-Noise [HKM63]. Random-Process [Wat67]. Random-Pulse [Rib67]. Random-Walk [CS68]. Randomly [Die68, TR58]. Ranges [Bin68]. Rapid [Cro61, JA67, Mac62, WL61]. Ratchet [San64]. Rate [DR69, Har56, Loo66]. Rates [Dor67, GJN59]. Rational [Eve64]. Ray [LBB66, RG65, Rub67]. RC [VB65, Was68]. Re [FH62, Kas62, Shu62]. Re-Entry [FH62, Kas62, Shu62]. Read [Cha59b, Ear64, Gam68, Kam61, Kru64, Mat66, Sei64, Sho61, Hsi67, Sei67, Waa68]. READ-Cycle [Sei67]. Read-Only [Gam68, Hsi67, Waa68]. Read-Out [Cha59b, Ear64, Kru64, Sho61]. Reader [YR65b]. Readily [Mow69]. Reading [CN68, Che68c, FR58, Hig63b, Hoa55, WL67]. Readout [Chu65a, FDBW63, Hil66, KI65, KI66, Lam59, NP64, Ros66b, Wid54]. Real [Atr65, Bau58, BH67, Col69a, Con62, Gre66a, LW59, Les64, McK67, Mer65, Nol62, Rog66, Sch69c, Shi68, Tru62, Tru63, VS62a, VS62b, Yam62b]. Real-Time [Atr65, Bau58, BH67, Col69a, Con62, Gre66a, Les64, Nol62, Rog66, Sch69c, Tru62, Tru63, VS62a, VS62b, Yam62b]. Realizability [HDW66, Pre65, Win63]. Realizable [PS66, Wei69a, Yen68b]. Realization [Ano68-41, AFM68, CKL62, CL63, FP62, HC69, HD67, HS69c, JC69, Kau61, Kel62, LC63b, Mer67b, Miy63, Miy64a, MF68a, Ris67, RC66, She64b, She65a, She65e, SH66a, SH67a, Tor66, TR58, YI68a, YO68]. Realizations [BC68, Cur68b, Dav68, HS67b, Har68a, Har66e, Muk69a, Nic65b, SY68, Ung68, WD69]. Realizing [Auf62, Cur69]. Receptors [Ano63-40]. Recognition [Ano63-40, Ble61b, BB62, Ble64, Ble66, Bon66, Bra63, BP64, Bro64, CF68, CN66, Cho57, Cho59, Cho62, Cho63b, Coo64a, Cov65, Das69, Dru69, FF63a, FF67, Fis63, Fis64, Flo58, FG60, Flo62, Fre61a, FG64, FCC67, FI65, FK69, HK60, Hig61a, Hig61b, Hig61d, Hig63a, Hig65, Hug67, Ito69, KS63, Kno67a, Kno69, Lew65b, Liu64b, LS66, MG60, Mar65b, MFF63, McC63a, Mei69, Mer65, MG66, Nad63, Nad65, Par62, Per63, RH66, Ros62, SD63, SE66, Spe67, Ste60, Ste62, Uhr61, Uhr62, Ull69b, Who63, YY66c]. Recognize [Fir63b]. Recognizer [Bar66].

Recognizers [MG60, Uhr62]. Recording [BS63, Chu67a, Gab64b, Gab67, Hoa55, HB60, Hop64, Kno67b, Kos62, Kum62, MH59, MS66, Mul59b, She62, She63c, Sie65, Ste63b, Tho68b, TBS64]. Recordings [Chu66a]. Records [CW68]. Recovery [CS64, Pre68]. Rectangular [Cly59]. Rectifiers [DS67b]. Recursive [SC68, Yam62b]. Reduced [WS67c]. Reducing [Sal61]. Reduction [Ano66-31, Gil63b, Gim65b, Gin59c, Gre69c, Hen64b, Koh65b, Luc66b, Luc69b, PM62, Ree65, RH67b, SD69, War58b, de 69]. Redundancy [BTW61, Deo68, DS67a, Gir69, PM62, WR63]. Redundant [Arm65, Fri67a, HM67a, Man63b, RA59, Sco57, Sri64a, Teo62, AS66]. Reentry [Fog67]. Reference [FG60]. Refill [MH58]. Reflected [Flo56, Luc59]. Reflections [Was67, Was68]. Reflectivity [O’h69]. Regardless [Thi63, Thi64]. regenerative [Wal53]. Region [BG56a]. Regions [Wil64]. Register [And56, BS64, Bra59, Bro60a, CE69a, Dav68, Den67, ENB61, Gil66g, Har61a, Har68a, HH62, HS64b, Kra67, Nic65b, Sch64c, SY68]. Registers [Ano68-46, Ben60, BHK62, BK62a, BK62b, Dav69d, JO68a, JO68b, Lem69, Mow66, Mow69, NP56, Pry58, Yoe63b]. Regression [Rub62a, Str68b]. Regular [Ano68-45, Brz62, BP63, Brz64b, Brz65a, Eve64, Ghi62, MY60, PP68, Wei69b]. Related [Gin66, Van57]. Relation [Van60a]. Relations [Mow66, Pre64b]. Relationship [BP69, Bro66, May65b]. Relationships [Pro64, dL69]. Relative [Sch63b]. Relay [ACGN60, Brz65b, Mar63b, Mul59c]. Reliability [Ano66-32, AW56, BTW61, Cha66, FY66, Gil60b, HM67a, Har68b, Hea56, Law61, Law62b, Mil67, Nie56, RA59, Set66b, Urb64, Urb65a, Urb65b]. Reliable [Ano66-30]. Remainders [Whi69b, Atk68]. Remark [Har64, Har67a]. Remarkable [Mil57]. Remarks [Has69, Ree65]. Remote [Ano68-40]. Renaming [Fei63a]. Repairable [Man63b, Teo62]. Repairing [Dor69, Wil61]. Repetitive [And63, Fuc56, MPP59, Pet61, TP60, Wag67]. Reply [Con68, DM69, Deo69, Ein62, Hay69a, Kau69a, Kuo69, Mar67b, MK68, Ren69a, Ren69b, RH67c]. Report [Ano66-35, Bla56b, Unc59, Unc61]. Reports [Brz66, Fif66]. Representation [Ben69b, Gre56, How56, Kor60, MMR60, McK67, O’N67b, O’N69, Son63, Tho68a, Wad61]. Representations [Avi61, Bey63, Bey64, Har57, Wee60]. Represented [CM60, Elf62, Met62]. Representing [Ram65]. Reproduction [MH59]. Require [Kor62b]. Required [Dor67, Fra58, Lew66b, MK61]. Requirements [Bon66, Gel61, Tor69]. Research [Gli66b, Kor65]. Residue [BB69, Che61b, Gar59a, Gar65, Guf62, Leh63b, Mer64b, Sas67, Sas68, Sem63, Sza62, KCT62]. Residues [Rot59]. Resistance [Ham60]. Resistances [Chi63]. Resistive [AM67b, Cro68]. Resistor [Cha59a, DF68a, LBG65, Lyn69a, MS58, SP62]. Resistor-Card [LBG65]. Resistor-Coupled [MS58]. Resistor-Transistor [Cha59a]. Resistor-Transistor-Tunnel-Diode [SP62]. Resistors [Pfe59, Ste69]. Resolution [HG57, WH62]. Resolving [FG61, Mii64]. Resource [Kop69]. Response [Cof69a, Hil66, ID63]. Responses [FG61, Mii64, Wei63b]. Restricted [She64a]. Restructurable [EBTB63, New64]. Result [Cho63b, Lew65a]. Results [Ble61b, FF67, Gil66g, Kno67a, RH66, Urb65b]. Resynchronizing [Pre66b]. Retrieval [Fel63, Lan63a, Led63, Tau62, YR65b]. Reversal [Bal62b]. Reversals [Gre69c]. Reverse [Gin66]. Reverse-Definite [Gin66]. Reversed [Fra69]. Reversibility [KS68]. Review [Ade62, Alo62, Ano53e, Ano53f, Ano57k, Arm62c, Auf62, Bit62, Bro55a, CS58, CM63, Dav62, Flo62, Gil62, Hoa62, Hus54, Kau62, Low62, Mac62, Mad62, Met65a, Mit62a, Nas56, Nol62, Ros62, Rub62b, Shu62, SBM63, Tau62, Uhr62]. Reviews [Ahr65b, Ake65a, Ake66, And67a, Ano55k, Ano55l, Ano55m, Ano55n, Ano56m, Ano56n, Ano56o, Ano56p, Ano57l, Ano57m, Ano57n, Ano57o, Ano58r, Ano58s, Ano59v, Ano59w, Ano60x, Ano60y, Ano60z, Ano60-27, Ano61-28, Ano61-29, Ano61-30, Ano61-31, Ano62-33, Ano62-34, Ano62-35, Ano62-36, Ano62-37, Ano68-42, Ash65, Avi63, Axe63, Can67b, Car65, Cha67, Epl66, Fir64b, Gil65a, Gol63, Hop67b, Jon64, Lev63a, Man67b, Mil66, Min63, Ros64, Sci64, Ses65a, Son64, Sri64b, Str62, Tai63, Tra67, Urb66, War66, Yau66, Pra66a]. Richard [Ano59x]. Rigorous [MO63]. Ring [Gar59b, Hal69]. Rings [Wan67]. Rise [Thi63, Thi64]. RMC [Hig66]. Robot [Hag56]. Rod [Hig66]. Role [Ano66-32, Bed64, Cha66, Fra64a]. Rolling [GR67]. Root [Cow64, Dei65]. Rooting [Met65c]. Roots [MPP59]. Rope [Alo64]. Rose [Nol62]. Rosenblith [Ade62]. Rotation [DS61, Mad62]. Routines [Sal68]. Routing [JMW60]. Row [Ung68]. Rows [Kam61]. Rozonoer [Auf62]. RST [Sur68, GD67]. RTL [Cha60]. Rule [Roy68b]. Rules [Ghi62]. Run [ES64]. Rung [Bal62b]. Russian [ZA59]. RZ [BS63].

S [Mac62, Mad62, Don65b, Mar69a]. S-R-T [Mar69a]. Sabrac [LEN63a]. Safe [Fos68b, MK67]. SAGE [LW59]. SAHYB [Ano68-43, dM66]. SAHYB-2 [Ano68-43]. Sampled [NW64]. Sampled-Data [NW64]. Sampling [HH57]. Satellite [Bol67]. Satisfaction [Tor69]. Saturable [Nit64]. Saturated [Gol60]. Saturating [MRS60]. Saturation [Kos62, MH59, She63c].

Saturation-Type [MH59]. Saturn [HS67a]. Savitt [Alo62]. Scale [DF68a, Jos67, Kat68, LLMS67, Mur68, Sag66, She64a, SNS69, Whi69c, Yao66]. Scale-Changing [She64a]. Scaler [Mur55]. Scaling [Gil59c, HK69, Knu65b, LA68]. Scan [Bra63]. Scanned [Mar65a, Pet65]. Scanners [PM58]. Scheduling [Bas63b, Fif68b, Man63a, MC69, Smi68c]. Scheme [Bar66, Hsi62, Ste62]. Schemes [BCE61, Bir67, Mar66b, Wit67]. Schmoo [Wom65]. Science [Ano56q, Ano58t, Ano58u, Ano58v]. Sciences [Ble66, Fei61]. Scientific [Fei63b]. Scoring [Bra66]. Scratch [Poh66]. Scratch-Pad [Poh66]. SEAC [EW53, LNSW54]. Search [FG61, MR59]. Second [Bro66, GM63a, Gre68b, Uhr61]. Second-Harmonic [Gre68b]. Second-Order [Bro66]. Secondary [JO68a, JO68b, Koh64b, Nar60, Tor68]. Secondary-Emission [Nar60]. Section [Ano53e, Ano53f, Hus54]. Sections [Pie63b]. SEER [Hag56]. Segment [Sch62]. Segmentation [SSG69]. Segmented [SS66]. select [Bed62]. Selected [Net58, Net59a]. Selecting [Cha65b, Pow69]. Selection [Fri67b, Luc66a, Hay52]. Selector [DG61, NM68]. Self [Deo66, Dor69, Els60, FI65, Gab67, Jac64b, Lee62b, Lev63b, Man66b, Man66c, MF68b, Mul59d, Mul60, Mur65a, Tod62, Wil61]. Self-Checking [Lee62b]. Self-Clocking [Gab67]. Self-Complementary [Els60, MF68b]. Self-Correcting [Lev63b]. Self-Diagnosability [Deo66]. Self-Diagnosis [Man66b, Man66c]. Self-Dual [MF68b, Mul59d, Mul60, Mur65a, Tod62]. Self-Organizing [FI65, Jac64b]. Self-Repairing [Dor69, Wil61]. Semantics [Sto66b]. Semi [Fia66, Len65]. Semi-Permanent [Len65]. Semi-Systematic [Fia66]. Semiconductor [FS59, HMSG59]. Semipermanent [MY61]. Send [Ano55o]. SENEWS [Ano58t, Ano58u, Ano58v, Ano59y, Ano59z, Ano59-27]. Sense [GK62, KH63, Tsu62, Yao66]. Sense-Amplifier [Tsu62]. Sensing [Alo62]. Sensitivity [Kah68, LC63b, Mei66, O’G67, TPM65, Tor69]. Sensitivity-Function [Kah68]. Sensor [Che68a]. Sensory [Ade62, Ade62]. Separability [Gas63]. Separable [CKL62, KT65, Tor66]. Separated [Dav65]. Separation [ALS68, Glu66, Gre69a, Hig61c]. Separator [San66]. Sequence [Col67, Cra60b, Dou68, HDW66, Per68, Ste63a, Hag56]. Sequences [Boo68, CE69a, EG61, KK68, Man67a, Pre66b, Pre68, Ros66a]. Sequencing [Gra62b, Sch64a, Tav65]. Sequential [Ano68-41, Ara57, Ard63, Arm62a, Arm62b, Arm62c, AFM68, AH57, Auf58, Bas64, BF61, Bou68, Bro65a, BP63, BM63, BM64, BD64, Brz64b, Brz65a, BS68b, CF68, Car68, Cur68b, Cur69, Dab69, DB66, Dav68, DM64, EM69, FG69, FY66, Fri66c, Fri66d, FM68, FCC67, Fuk64, Gaf68, Ger68, GG69, Gil63b, Gil64a, Gil65b, Gil66c, GA58, Gin59b, Gin63, GL65b, HC69, HS67b, Har68a, Har67a, Har61b, Har62, Har63e, HS63a, Har65b, Har66e, Haz65, Hen64b, Hum63, Iba67, Kar64a, Kim66b, Kin66b, KW69a, Koh64b, Koh65b, KL67, Lav65a, Lav65b, Ler65b, Lev63b, Lin63, Luc69b, MTS69, Man64a, Man67a, Mas66, MS68, MU59, Mei67a, MK69, Nar61, Net59b, New67, New68b, New68a, Nic65b, PU59, PW67, Pre64b, Pre65, Pug65, Ram65, RY66]. Sequential [Sau67, SF62, Sin69, Sta65, SH61, SO68, SY68, TMF69, Tor68, Tra66, Wan69, WS67b, WS67c, WD69, Wei63b, Woo64b, Yau65, YW66, Yoe61, Yoe63a, Zah66]. Serial [Cha69, Den69a, Gil60c, HS64b, Leh61, LEN63a, Ver64]. Serial-to-Parallel [HS64b]. Series [BS64, Con67b, DW67, Gil66c, Rat67, Sto66a, Van59, Van60a].

Series-to-Parallel [Gil66c]. Service [Bak57, Cof69a, Hig63b, YR65b]. Servos [Gil59a]. Set [Arm66, Cha65b, Fri66b, Had67, Tod65]. Sets [DBC69, Gil64a, KO66, Lav65a, Loo65, LW65, PS68, Set66b, Set66a]. Setup [Eck62]. Seven [Win65a, Win67a]. Seven-Argument [Win65a, Win67a]. Several [KLP68, Nat64, Skl60b, Wil63]. Shafritz [Nol62]. Shaft [LC53]. Shannon [Fie64]. Shape [BS67]. Shared [Ano68-40, BFL^{+}66b, HH68, JC69, Kah68, Lew66a]. Sharing [Chi59, FP54, McG68, Neu62b, Sha63, SC67, Sin62, VBW69, Wie68]. Shelman [Mad62]. Shift [And56, Ano68-46, Ben60, Bra59, Bro60a, BHK62, BK62a, BK62b, CE69a, Dav68, Dav69d, ENB61, Fre67a, Gil66g, HL57, Han63a, Har61a, Har68a, HS64b, JO68a, JO68b, Kra67, Lem69, Mow66, Mow69, NP56, Nic65b, Pry58, SY68, Yoe63b]. Shift-Register [Dav68, Har68a, HS64b, Nic65b, SY68]. Shift-Registers [Ano68-46]. Shiftregisters [Els69]. Shiftrix [MR65]. Short [Bre65a]. Shortest [Lee62a]. Should [Ano55o]. Sided [Bic60]. Sign [BB69, Sem63, Sza62]. Signal [Bic60, Bro60b, BM63, BM64, Cat66, Coo64a, Cor64, Cro64, Gab64b, GMB61, Gin63, GL67, Gra64b, Lun63, Pie63a, Rea63, Was67, Was68, Yao63]. Signal-Processing [Gab64b]. Signals [DH69, FG60, Hau69, Lar66a, MH59, Mul59a, O’N67b, Wai64b, Wie65b, WK68, Yao66]. Signed [AS66, Avi61, Met65a, Svo69, Tun68]. Signed-Digit [Avi61, Met65a, Tun68]. Significance [Ihm63]. Significant [MA58]. Simple [CS64, CB63, Cof68, Col69b, Fra57, Gas63, Koh63, Kri64, MW68, WP64, Wij68]. Simplex [Ein61, Ein62, Min62]. Simplification [CB62, DBC69, GHL^{+}66, Kar63a, MS69, Mar62, TMF69, Ung65, Van60b].

Simplification-Some [Ung65]. Simplified [CKL62, SG69]. Simplifying [Chu61]. Simulate [Ano56r]. Simulating [ALS68, Fit63, Hep65, Kar69]. Simulation [Agg64, Ano68-43, Bau58, Ben69b, Bie68, Bri63, Bud66, Chu66a, Chu67a, Cly59, Del62, DKP69, Dia56, Dom57, FH62, Fra64a, Gas66, Gel62, Gil59b, GG68, HS67a, HF66, Hay69b, Hen62, Her62, JSS66, Kam61, Kas62, Kir64b, Kle69, KM64a, LW59, LC59, Les64, Lew66a, Man68, Mar68, McL62, Mcg65, Mck69, MW61, MMB66b, O’N69, Ott58, Paq67, PVW59, Rog66, Rub68, Sag66, SJ68, She64a, Sil68a, SS61, SS62, Sto64, Stu69, Swi63, Tam67, Tav65, Tho67, Tho68b, Tho68a, Vic67b, Wad64, Wad68, Wat67, Wee63, Yen68a, dM66, Shu62]. Simulations [Wai67, Wai68a, Whi67]. Simulator [Gas65, Mur65c, Sch69a, Zuc65]. Simulators [Gau64]. Simultaneous [Guf62, Hil66, KLP68, LB60, Mii61, SG69, Wei63b]. Simultaneous-Carry-Generation [LB60]. Sine [Har56, NK66]. Single [Agg64, Cro64, Dav68, Der64, FGU69, Gab62, Gab64a, HS69a, Hsi68, Lew66b, Men61a, Men61b, MN61, Mul59b, Rea63, Wad64, Wad68]. Single-Channel [Hsi68]. Single-Domain [Cro64, Rea63]. Single-Gate [HS69a]. Single-Pass [Mul59b]. Single-Threshold-Element [Der64]. Singular [Lav65b]. Singularities [Hau62a, Hau62b]. SIOUX [Cly65]. Sixth [Ano66-33, Har66a]. Size [FK65, GHI^{+}63, MK61]. Sketch [BCE61]. Skip [LB61]. Skips [Maj67]. Slave [Wil65b]. Slide [Roy68b]. Slow [Mat66, MN58]. Small [BB56a, New68c, SW63, Sto64]. Smirnova [Auf62]. Smith [Bit62]. Smoothing [Kar55]. Snake [Eve63, Sin66]. Snake-in-the-Box [Sin66]. Society [Ano56q, Bak57, Ano65t, Ano65u, Ano65v, Ano65w, Ano65x, Ano65y, Ano66r, Ano66s, Ano66t, Ano66u, Ano66v, Arb66, Ban65, BL66, Dav66, Don65a, Gil66a, LR65, Sha66b, Ste65, Sto65, Wai65]. SODAS [Fif68a]. Software [Kno68, Kuc66, Kuc68a, Sal68, Wat66]. SOL [KM64b, KM64a]. Solder [GR68]. Solenoid [Hsi67, PGB64]. Solenoid-Coupled [Hsi67]. Solid [KC60]. Solid-State [KC60]. SOLOMON [SBM63, GM63b, Hei63, Mac64]. Solution [Ben69a, Bir67, Cha69, Den69a, Fel63, Fis64, FG64, FL67, GM65, Gra60a, Gre62, Han67, Hau65c, HH53, Kuo68, Led59b, Lit68, Mar66b, MTI67, Ogb69, Pre55, Rob55b, Ros65a, Ros65b, Ros68, Tay69, Tho69, Van57, Vic69a, Vic69b, Was67, Wit67]. Solutions [Gre66a, Hau62a, Hau62b, Kor67a, MT68]. Solver [Rao58]. Solving [Bek65, DS61, Guf62, Mad62, Nil69, PM62, Sut69, Svo63a, TP60, Whi66]. Some [ABSH63, Bar68, Bla56d, Bri61, Bro55b, Brz65b, Cha59a, Cra64, Gil63b, Gil66g, Gin66, Gli66b, Hen64a, Kar64a, Lav65b, Lev63b, Liu66a, Lo61, McG69, Mer64a, NF64, Nis60, Pet63, Ree65, Rou58, SSPR60, Ung65, Urb65b, WD68, Yen68b, Kau62]. Sound [Ble64, SD63]. Source [She63a, Thi63, Thi64, Tho62]. Soviet [Ano57p, Fei61, Kau66c, War60]. SP [Mur65c]. SP-1299 [Mur65c]. SP-1299/000/01 [Mur65c]. Space [Don65b, PAB68, Wai67, Wai68a]. Spacing [Far62]. Spatial [Jos66]. Speakers [Ano53d]. Special [Ano62-38, Ano64-27, Ano66-34, Bir67, Kau66a, Mar66b, Pas67a, Pre65, Ros65b, Wit67]. Specific [SS68]. Specification [DD69]. Specifications [Gil60b]. Specified [DBC69, GL65b, Gra66, LC63b, Mei67a, Nar61, PU59, Ros66a]. Spectra [Shi68]. Spectral [BH67, Gil66e, Mat68]. Spectrum [Bic60, Kno67b]. Speech [Ble64, O’N69, SD63]. Speed [Ahr65a, Amo64, And68, BG56a, BG56b, Ble61a, BP59, Cho63a, Con66, DK63, DF68a, Edw61, EGP56, Gam68, GT68, Gil66d, GMN^{+}60, Hel66a, Jar63, Kae61, KH63, Kau66a, LM62, LC59, Leh57, LB60, LB61, Lin66, Lyn69b, Mac69, MR65, McQ62, Mei65, NP56, O’G67, PC56, PLS61, Sal60, Sav59, Sch68, Skl63b, Til60, Waa67, Waa68, Wei63a, Wel69, Wie55, Wij68, Yao63]. Speeds [HP64b]. Spheroidal [CM59]. Spheroids [CM60, Elf62]. Square [CM59, Cow64, Dei65, Hay61, McA68, Met65c]. Square-Loop [CM59, Hay61]. Squared [Bru57, Dei65]. Squarer [Dei65]. Squarers [NK66, SOS56]. Squares [MES63]. Squaring [Cox61]. Stability [CK65b, Gil66e, Kar55, NM57, Pet61, Urb63]. Stabilization [Kle53]. Stabilized [MB52]. Stable [Bat65, FF63b, Lem69, Mow66]. Stack [Har69b, KII62]. Standard [Ros58, Til60]. Standards [Ano62-32, May60]. Standpoint [Gre66b]. State [Arm62c, BM63, Col69a, FY66, FGU69, Gil61a, Gil63b, Gin59a, Gin59b, Gin59c, Gin63, GO63, Har68a, Har68e, Har61b, Hob66, Iba67, JO68a, JO68b, Kar64a, Knu65a, Koh64b, KC60, Lew63a, Liu63a, MY60, Mei67a, Nic63, NB65, PW67, Pre64b, Ree65, Rus65, SH61, Tau62, Tra66, Vai69, Wai67, Wai68a, WS67b, Win62, dL69, de 69, Auf62]. State-Logic [Pre64b]. State-of-the-Art [Hob66]. State-Space [Wai67, Wai68a]. Statement [Bas64]. States [Ard63, Bat65, Bou68, DM64, GL65b, Har67a, Luc69b, MW64, PU59, Zah66]. Static [Sho61]. Statistical [Boo68, Cho65b, Cov65, Ito69, MG60, SM69, SS66, Slo62, Van57, Uhr62]. Statistics [Kle69]. Status [Bed64, Mur68]. Steam [Her62]. Steel [GR67]. Steepest [Kor64]. Steered [MS59]. Steering [Cho67, GM63a, RC57]. Steinbuch [San66]. Stenotypy [Sal63, Gal62]. Stenowriter [Gal62, Sal63]. Step [CS64]. Stieltjes [And62]. Stochastic [Coo64a, Eve65a, ID63, Kor67b, Ste62, YS68, dL69, de 69]. Stops [Kor60]. Storage [AD69, ABC61, Alo62, Alo64, Bit62, But64, CM68, CS68, Den69c, Dor67, FS65, Gra56, HM67b, Hel66a, HH68, KELS62, Kor63b, Kra67, KI65, Kru64, Led63, Leh61, Leh63a, Lew63b, Lin63, MY61, McQ62, Mil55, MWR68, Ros65a, San64, SAP65, SSG69, Sho61, Smi61, Wie55, Wil65b, Won55]. Store [Ear64, Gra64a]. Stored [Gre64b, Woo65]. Stores [Car60, Gam68]. Straightforward [MN61]. Streams [Law68]. Stress [Ulz66]. Stretch [Buc62, Str62]. Stricter [Glu66]. Strip [Gra64c, McC68a]. Strip-Line [McC68a]. Strip-Lines [Gra64c]. Stripline [KI66]. Strobe [Hol61a]. Strobing [Tsu62]. Stroke [GRU58]. Strong [Fle67]. Structural [Pow59, TMF69]. Structure [Bak57, CET62, Cly59, Cos63, ET63, FG69, Kam60b, Lan64b, Nie69a, Sam69a, Sch65a, Urb68, Wak69, Wel68]. Structured [Stu68a, Stu68b]. Structures [Gel61, Kro59, Meo68, MF69, SJ69]. Struktur [Ulb63]. Student [Goo55]. Studies [Ble66, Hig63a, Kas62, Kir64b]. Study [ABSH63, Bri66, Bro65b, CK65b, Gar59b, HS63a, Lee69, Lev63b, Man66b, MH58, Ran64, Wai66, Wie65a]. Stylus [Kum62]. Subclass [FF63a]. Subcommittee [Ano58t, Ano58u, Ano58v]. Subharmonic [MWR68]. Subject [Ano59-28, Ano59-29, Ano59-30, Ano60-28, Ano60-29, Ano60-30, Ano61-32, Ano61-33, Ano61-34, Ano62-39, Ano62-40, Ano62-41, Ano62-42, Ano68-44, BH65a, Coh63, Fle63b, Gil60b, Har65a, Har66b, Hop66, Koe63, Kon64, Law61, Law62b, Lec63a, Lew67a, Liu66b, Loc67, Man66a, MR66, Met65b, Mik65b, Mun67, Ove63, Pag67, Pie65, Poh64b, Sch64b, She66, Wid64, Wid67, Win65c, Win67b, Wyn64]. Submicrosecond [SC60]. Suboptimal [CF68]. Substituted [MS66]. Subtraction [Lan69a, Lan69b, Sas67]. Successive [SH66a]. Suggestion [Jac64a, Wal64]. Sum [Abh58, Gre69a, HS63b, Skl60a]. Summability [BC66, Roy68b]. Summand [Skl60b]. Summary [Ano66-35, Flo64a]. Summation [Gos65, Kas63]. Sums [Abh58, EKP67, Sch69b]. Superconducting [Sob62]. Superconductive [Ahr65a, BAL61]. Superconductors [CM62b, CM63, MC62]. Superior [Uhr61]. Supplementary [Cha60]. Support [Sch65b]. Survey [Brz62, Dav57a, Flo64b, Han66, Hay61, Kau66c, MO57, Nag63, PM60, Unc59, WW55]. Svenska [Lan63b]. Swanson [Tau62]. Sweden [Lan63b]. Switch [Car60, Chr61, Dia56, Eck65, Wij68]. Switches [AH62, Chi59, CB63, Kam60a, Neu62b, Sin62, Val67]. Switching [Ano64u, Ano66-33, AE61, Art61, Ber58, BF61, BS68a, Bor69, Bue60, CS64, CB62, CKL62, Coh62, Col61, Cos63, DC65, Dav57b, Dom57, Ein61, Ein62, ES64, EKP67, FT68, Fri66d, GMB61, GL67, Gol60, Gou58, GH59, Har66a, Haz64, Hol61b, Hor65a, Hug67, JMW60, Kas66, Kau61, Kau66c, Kil62, Kor63a, LS60, Law61, Law62b, Lec63b, Lec64, Lec67, LWM64, LC63a, Mai62, MS69, Mar56a, Mar69b, McC58, MU59, McG69, Meo65, Mer66, Mer67b, Mil58, Min62, Muk63, Muk66, Mul56, Mul59a, Mul59c, Mul59d, Mul60, MK61, Nar61, PU59, PC56, Ros58, ROK58, Roy67, Roy68a, Sch63b, Seb57, SF62, Sho64, Skl63a, Ste66, SJ69, SD69, Tay65, Thi63, Thi64, Tor66, Van60b, Wan65, War58b, War58a, WD68, Wei69a, Wei69b, YO68, YR65a, Mul54]. Symbol [Gre61]. Symbolic [KM64a]. Symmetric [AH63, BS68a, Bor69, CD66a, Eps58, FT68, Kau61, Mar56a, Mer67b, MES63, Muk66, Mul59d, Mul60, Roy67, Roy68a, She65d, She65e, Wee58]. Symmetrical [Eck62]. Symmetry [Ama64b, DS67a, Els60, Muk63, Win68a]. Symposium [Ade62, Ano56r, Ano56q, Ano66-33, Har66a, Yov59]. Symptom [Ihm63]. Synchro [Bro60b, WP62]. Synchronizing [Ros66a]. Synchronous [Cla67, Cur69, Fri66d, HS67b, Kar64a, New68a, Sal61, SA64, Tor68, WS67b, WS67c]. Syntax [Flo64b, Gar69, Mcc67a]. Syntax-Directed [Gar69]. Syntax-Oriented [Mcc67a]. Synthesis [AM67b, Ake61b, Ake65b, Ber58, Cad68, Cam57, Chu69b, Cro68, DM68, Der64, Dvo68, Ell65, Eps58, FT68, Flo62, Fre61a, FM68, Gab62, GBC69, Gin59a, Gin59b, GL67, Gos65, Han63c, HO67, Has69, Hau69, HM65b, HF57, Hor63, Hor65b, Hru69, Kai67, Kas66, MFF63, Mil58, MW62, Nec68a, Neg64, New68a, Oli65, Per68, Pow65, Roy67, Roy68a, Rus65, Sam69b, SAL66, SD68, She65a, Skl63a, Smi68a, Sta65, SK65, Svo63b, Wan69, Wei69b]. Synthesized [de 69]. Synthesizing [Dor69, FR58]. System [AJ60, Ano68-43, AW56, BFL^{+}66b, BFK63, BSK67, Bie68, Ble64, BP64, Bro65b, Bro66, Buc62, BMRS69, Che61b, Cho57, Chr61, Chu67b, Chu69a, Cly65, CV61, Cra61, Dar68, Del62, Dre67, DD68, DD69, Eps68, ET63, EBTB63, FDBW63, Fif68a, Flo58, Fra64a, Gal62, Gar59a, Gea65, Gel61, Ger68, Gil59b, GNV59, Guf62, Han68, HM67a, Har68d, HW69, Hea56, Hel66a, Hig65, Hoa55, Hsi68, Jac68, Kat68, Kel63, KELS62, Kno68, Kop69, KLP68, Kru64, Kud64, LB60, Lei67, LNSW54, LA54, LLMS67, Lew65a, Lin63, LS66, LK60, MMT66, MA63a, McK68, Mei65, Mer56, MA63b, MA63c, Min69, Mur65c, Nad63, New64, New66b, Nie56, Nil69, Ols65, Rux69, SD63, Sal63, Sas67, Sas68, Sch63a, SW63, Sch64c, Sch69c, Sch65b, Sei67, Sha63, Shu62, Slo63, Str62, Str68a, Sut68]. System [Tru63, VBW69, Waa67, Waa68, WHB56, Wel68, Wol62, Woo64a, Lan63b, Nie68, Nol62, Wie68]. System-Design [Mei65]. System/360 [McK68, Nie68]. Systematic [Cur69, Fia66, Gil59c, Hsi62, Tal64]. Systems [Agg64, AG68, AW67, Ano63-40, Ano66q, Ano66-32, Ano68-45, APVS66, BB69, Bed64, Bra58, Cam57, Cat67, Cha66, CW68, Cho59, Cof68, Con66, CS67, Cov65, Den69b, Den69c, DKP69, Dun65, Epl68, FS65, Flo56, FG60, Fos68b, Fre61a, Fri66b, FI65, Ger63, Gil66d, Ham58, Han66, Har69a, Hig61d, Hol63a, ID63, Isb63, Jos66, Kah68, KS63, KM64a, Koh63, Kon60, Kuc66, Leh61, Leh63b, Leo63b, Man68, Mar69b, MT68, ME67b, ME67a, Mei66, MK67, Miu66, Mul59b, MC69, Nol64, OKL59, Pau67, PMRC67, Pet61, PMC67, PVW59, Pro64, Ram65, RR67, RW67, RV68, Sag66, SAL66, Sch65a, Sem63, SF62, Sha66a, SS66, SH66b, SNS69, Son63, Stu69, SS68, Svo63b, Sza62, Tei55, TNM64, TPM65, Vic67b, Wad57, Wad64, Wad68, Wai67, Wai68a, Wom65]. Systems [dL69, Hay52, KCT62, Flo62, Yov59]. Systems-Cyclic [Epl68, ME67a].

T [Mar69a]. T. [Ano69-55, Uhr62]. Table [Ake61b, AK69, Gim65a, Luc66b, Sch65a, Ung65]. Table-Lookup [AK69]. Table-Lookup/Interpolation [AK69]. Tables [Ano66-31, DBC69, Gim65b, Gra66, Mar56a, PM62, Ros66a, Ung68]. Tabular [Ara57, HO67, Mer66]. Tal’ [Auf62]. Tangents [Pre55]. TANT [Gim67]. Tape [BP59, Eps68, Gil64b, Gre69c, Lam59, Mar65a, MS66, Pet65, WL67]. Tapes [Mas57]. Tapped [PD62]. Tapped-Potentiometer [PD62]. Taylor [San66]. Taylor-Steinbuch [San66]. Teach [Eng61a, Eng61b, WB62]. Teaching [WB62]. Technical [Ano63x, Ano63y, Ano63z, Dav63a, Fei63b, Fir63a, Har63d, Kar63b]. Technique [AG68, ALS68, Ano66-31, BHK62, BO59, Cha68, Cha59b, Cha64, Cho67, CS67, Cro61, Dav69b, DS61, Der67, Gil64b, Gim65b, Gin59b, Gin59c, Gos65, Han62, Har68b, Jon69a, Kar64b, Kas62, Mad62, New68a, Pie63c, Ren68, Rob55b, RH67b, Sam69b, Seb57, Ses63, Tae69, Tho60, Til60, Vol59, Whi66, YY66b]. Techniques [AYT68, Bar68, BM63, BM64, Cha69, Cra64, CS66, Den69a, Gin63, Gol60, GM62, GR64, Hau62a, Hau62b, Hau65a, HM67b, HH57, HB60, Hop64, JMW60, Kar64a, Kel63, Knu65a, Kov66, LB61, Lo61, Mea59, Nic63, RM65, Rub62a, Sch63b, She64a, Van66, YS68, Yov59, EW53]. Technology [MR68, Mur68, War60]. Telefunken [Ulb63]. Telefunken-Digitalrechenanlage [Ulb63]. Telegraph [Hig63b]. Telephone [Lew67b, MV55, Rap66]. Temperature [Jos66]. Temperatures [Chi63]. Ten [Can66, McL62]. Terminal [Kor62a, Pie63b]. Terminals [Hor68]. Ternary [AD63, APVS66, ASC68, Bor69, Han63b, MM55, Mer64a, Mer66, Mer67b, Muk66, Sal64, SAP65, SAL66, VO65, Woo63a, YR65a, Yoe68]. Test [Con67a, EM66, Gas63, HM67a, Hru69, Lis63, Sch65a, Yen69]. Testing [Dwy69, Kau68, RC66, She64b, SH66a, SH67a, Low62]. Tests [Arm66, Cha65b, Had67, Har65b, Man64a, Pow69, RBS67]. Text [Kir64a]. th [Lee62a, Liu63b, Liu64a]. th-Order [Liu63b, Liu64a]. Their [Bro66, Brz62, DC65, Fif68b, KK68, Mer67b, Muk69a, Pre66b, RY66, SP63, Van60a]. Theorem [Ata64, Bro63a, Cha65a, Chu61, Coh64, DF68b, Lin60, de 69, Gil61b]. Theoretic [Lav65b]. Theoretical [BS63, Cho63a, MN58, Yen68b, Yoe65]. Theory [Ama67, Ano64u, Ano66-33, Car68, Els64, Fuk64, FI65, Gar65, Har66a, HSA57, Kau66c, Lar66a, Lew63a, LC63a, Mer64b, Nel63, Ott64, PRS63, Ric67, Ste66, Tis67, Wak69, War58a, WD68, Wie65b, YI68b, Nel54]. Thermal [Sob62]. Thermionic [GS58]. Thermistors [GS58]. Thesaurus [Gre64b]. Thick [Mat66]. Thin [Bit59, Bro60a, CM59, Cro64, FK66, Hig66, Rea63, Sei64, Sob62, Ulz66, Yao66]. Thin-Film [Bit59, FK66, Hig66, Sob62, Ulz66]. Third [Agg64, Wad64, Wad68]. Third-Order [Agg64, Wad68]. Thought [Bre67]. Thoughts [Lo61]. Three [Bal62b, Bon66, Dav57b, Gre68a, Hel63, Hor65b, Kam61, KO66, Men61a, Men61b, Neg64, Pra66b, Ris67, Rot59, Rus65, She64a, Smi65b, Val67]. Three-Body [She64a]. Three-Dimensional [Gre68a]. Three-Input [Hor65b, Neg64, Ris67]. Three-Position [Val67]. Three-Rung [Bal62b]. Three-Valued [Val67]. Three-Variable [Hel63, Smi65b]. Threes [Ger60]. Threshold [APVS66, Ata64, BC68, Bre65b, BS67, Cad68, Cam64, Che61a, Cho65b, CL63, CL64, CW69, CP68, Der64, DF67, FP62, FT68, Gab62, Gab64a, Gab65, GT68, GBC69, HC69, Han63c, Han63b, Har66c, Har66d, HD67, HO67, Har68b, HM65b, HC68, Hru69, Hug67, HS69c, IS65, KW65, Kas66, Kas63, Kau67c, KT65, KW69a, Kri66, Lev68, LC63b, Lew66b, Liu65, Liu66a, May64, May65a, May65b, Mei68a, Mei68b, Mer64a, Mer67b, MF68a, MF68c, MF69, Mur65a, Mur65b, MT66, Nec68a, RC66, Roy67, Roy68a, SAL66, She64b, She65a, She65e, SH66a, SH67a, Smi65a, Smi66, SH68, Tor69, Ulz63, Win63, Win65a, Win65b, Win67a, Win68a, Win69, Woo63b, Woo63a, Wra62, YI65, YIK68, YI68a, YI68b, YO68]. Threshold-Logic [Han63c, She65a, YO68]. Thyrite [KC54, KC58]. Time [Ano68-40, Ara57, Atr65, Bau58, Ben69b, BH67, Bla62, Bri66, Bro69b, Bud66, Cat66, Col69a, Col67, Con62, Cun54, Dou68, FP54, Fre67a, FGU69, Fuc56, Gre66a, HF66, Hau69, HH68, Hep65, Her66, ID63, Kah68, Kor62b, LW59, Les64, Lew66a, Lil56, Loe64, MO63, May64, McA68, Mer65, Mye57, Nie69a, Nol62, Raw59, Rog66, Sal61, Sch58, Sch69c, Sei64, Sei67, Sha63, SC67, Shi68, Sob62, Tho68a, Tho55, Tru62, Tru63, VBW69, Van59, Van60a, VS62a, VS62b, Wag67, Wai66, Wie68, Wie65a, Yam62b, Zac61]. Time-Delay [Ben69b, Cun54, HF66, Her66, Tho68a, Tho55]. Time-Division [Lil56, Sch58]. Time-Mark [Bla62]. Time-Phase [Col67, Dou68]. Time-Sequential [Ara57]. Time-Shared [Ano68-40, HH68, Kah68, Lew66a]. Time-Sharing [FP54, VBW69]. Time-Variant [Nie69a]. Time-Varying [Bro69b]. Timed [JC69, TR58]. Timed-Shared [JC69]. Times [Thi63, Thi64]. Timing [Amo66, Sch64a]. Titan [PMRC67]. Together [Wag66]. Tohma [Rud65]. Tolerance [CP68, May64, MF68c, Ulz63]. Tolerances [BW63]. Tolerant [Bei68, Rus65]. Tompkins [Ano58k]. Tones [Stu69]. Tool [Kin59, Sus58]. Tools [Cra64, Kel62]. Top [Gar69]. Top-Down [Gar69]. Topological [UM56, War58a]. Toroid [BB61]. Toroidal [Bal62a, Pie63b]. Total [Ano66-32, Cha66, Muk63]. Totally [CD66a, She65d]. TR [Ulb63]. Tracing [LBB66, RG65, Rub67]. Track [Don65b, Tom56]. Tracking [LW59]. Tracks [She63c]. Tradeoff [Cof69a]. Traffic [Kir64b]. Trainer [DEL55, GNV59]. Training [Bob65, Don67]. Trajectories [Bla62, Tal64]. Trajectory [Fog67, Mar68, Sas69, Wag67]. Transactions [Ano56g, Ano57f, Ano58l, Ano59p, Ano60o, Ano61p, Ano62t, Ano63p, Ano65s, Ano66p, Ano67a, Ano68m, Ano69b, Ano69-32, Ano69-33, Ano69-34, Ano69-35, Ano69-36, Ano69-37, Ano69-38, Ano69-39, Ano69-40, Bar64, Kar67]. Transducer [Mar65a, Pet65]. Transfer [Her66, KC58, Mar57, Rub68, Sch64c, Vic67a]. Transfers [Gil60c, Man63a]. Transfluxors [Pry58]. Transform [BP69, Men69a, Pra69, Sha69]. Transformation [AYT68, BS68a, Bor69, Bro65a, Gil66c, Gra60a, HS64b, Pre68]. Transformations [Epl68, Gab64a, Gol69, HDW66, Lec63b, ME67a, Pre66b]. Transformer [AA66, Har68c]. Transforms [And68, Fuk64, Mac69, Tou60]. Transient [GL67, Gre66a]. Transients [SK63]. Transistor [BB56b, Cha59a, CS64, Dom57, Ear64, Eas58, GMB61, Gel62, GJN59, Ham58, Har58, JMW60, Kur64, MS58, Mas60, PT66, PC56, RR62a, SP62, Tam67, WM59, WR63, Wan55, You57]. Transistor-Driven [You57]. Transistorized [Edw61, KB64, Pic65, Pic66b, Sal60, Sch58, Yao63, ZP66]. Transistors [Amo64, ALS68, Ham60, Ker56, KL58, Thi63, Thi64]. Transition [Chu69b, Ear68, FGU69, GO63, Mul59a]. Transitions [MW64]. Translation [Ard65, Bas64, DD69, GHR67]. Translational [Fog67]. Translator [Har69a, Mcc67a, Pre63, Pro64]. Translators [O’B57]. Transmission [Ben69b, Che68b, FS65, Fit63, FS59, Kau67b, MO65, Ric67, Tho68a, Was67, Was68, Yao63, Gil62]. Transport [Ben69b, Tho68a]. Treating [Hau62a, Hau62b, Kar64b]. Treatment [Mul59a]. Treatments [MO63]. Tree [Har68c, Lan63a, Mar57, Meo68, Pra65, Pra66b, Tod65]. Trees [Kau60]. Triangular [Pfe59]. Tributary [LWM64, PS66, Skl63a, SKS65]. Triggering [Eck62]. Trigonometric [HG57, Rob55b, Vol59]. Triode [GH59]. Triplicated [Gir66, GR68, Gur65]. TRL [Wra62]. Truncation [Bri67, VB65]. Truth [Ake61b, Chu61, Lis63, McN61, Mot60]. TSAM [FP54]. TSS [VBW69]. TSS/8 [VBW69]. TTL [SL65]. Tube [Cha59b, HK53, Kle53, MH58]. Tubes [DG61, GS58, Ros65a]. Tukey [SC68]. Tunnel [Amo64, BS64, Ber60a, BW63, CM62a, Car64, Cho60, Coo64b, DK63, Ear64, Goo64, HP64a, Kae61, Kae62, KH63, Kra67, Kru64, Men61a, Men61b, RR62b, Sal64, SP62, Ulz63, Uzu62, Alo62].

Tunnel-Diode [BS64, BW63, Car64, Coo64b, Goo64, Kae62, KH63, RR62b, Ulz63]. Tunnel-Pair [CM62a]. Tunnels [WHB56]. Tuple [BB62, Ble61b, Hig61b, Ull69b]. Turing [Gre69c, Hen66]. Twelve [Pra69]. Twisted [Hal69]. Twisted-Ring [Hal69]. Twistor [BHRS61, Gra60b, SS59]. Twistors [CM60]. Two [Ano60-31, Ano68-45, BW61, BW62, Bic60, Bla56a, CP66, CVV65, Dvo68, ED67, FL67, FK69, Har65b, Has66, Hru69, Kam60a, KO66, Mai62, MMR60, Men61a, Men61b, MC69, Pol57, Rob55a, Skl60b, Smi69, SW65, Str68a, SD69, Ter62, Tom56, WD68, Yoe65, Rub62b]. Two-Category [Smi69]. Two-Dimensional [Bla56a, CP66, FL67]. Two-Input [Mai62]. Two-Level [BW61, BW62, SD69, WD68, Rub62b]. Two-Point [Ter62]. Two-Processor [MC69]. Two-Rail [Dvo68, Yoe65]. Two-Sided [Bic60]. Two-Summand [Skl60b]. Two-Track [Tom56]. Two-Year [Ano60-31]. Type [And56, DC65, JA67, KW69a, MN58, MH59, Pic65, Rao58, Rao68, Sch63a, WD69]. Types [Els60, Tod62, Win68a].

U.S. [ZA59]. Ultimate [Skl63b, Urb64]. Ultimate-Speed [Skl63b]. Ultra [Yao63]. Ultrasonic [AJ60]. Unate [Lis63, McN61, Muk69b]. Unateness [Hru69]. Unbounded [AFM69, CM68]. Uncapher [Ano64t]. ’Understand [Fir65]. Unger [Ree65]. Unified [Wol69]. Uniform [Fie64, Har64, Loe64]. Uniform-Access [Loe64]. Uniformly [Wag66]. Unilateral [Van60b]. Union [Kau66c]. Unipolar [WM59, WR63]. Unique [MS69]. Uniqueness [Bey63, Bey64, Wee60]. UNISIM [Mcg65]. Unit [ABC61, But64, DK63, Gre66b, HS69b, Kau58, O’B57, Tom56, VB69]. Unit-Distance [Kau58, O’B57, Tom56].

Unitary [SY68]. United [Kor65]. Units [Gra62b, Gra62a, LB61, Leh63a, RM65, SA64]. Unityper [WM53]. UNIVAC [HK53, PLS61]. Univariate [Coo63b]. Universal [FGU69, Gar69, Har68a, Men65, Muk69a]. Unknown [Liu63a]. Unnormalized [MA63b]. Unsolved [Can66]. Unspecified [Wan69]. Updating [YS68]. Upon [Man64a]. Use [BTW61, Che68c, DEL55, Ear68, Ein61, Ein62, Far69, Gau64, GMB61, Gel62, Gli66b, GR67, Gre69a, Gre64b, HS67a, HH68, LS60, MPP59, Mar56a, McQ62, Min62, Muk63, Pie63c, RM65, RA59, Sco57, Sed68, SC68, She63a, Tho62, Vic69a, Zah66, Nel54, Hoa62, Shu62]. Used [Dav57a, FY69a, FY69b, May65b, EW53]. Useful [Ung65]. User [New66b]. Using [Alo62, Amo64, AE61, Bek65, Bra63, Bri67, CK65b, Cha69, Che68b, CM62b, CM63, Cho57, Cho62, Coo64a, Cra60b, Cro61, Cur69, DK63, DF68a, Dav69d, DG61, Els69, Fer67, Fer69, Flo58, Fre67a, Fuk64, FS59, Gab62, Gar65, GM62, Gon65, GRU58, Guf62, Han67, Hay61, Hop64, Hru69, Hsi62, IS65, Kah68, KC54, KL58, Kri63, KI66, LWCL61, Lit68, Luc59, MS69, Mat66, Men61a, Men61b, Mer64b, MN61, Mit62b, MH59, Miy63, Miy64a, NP56, Pet61, Pfe59, RC64, Rid66, Ris67, Rob67, Roy67, Roy68a, Roz62, Rub62a, Rus65, Sal64, SAP65, SI59, Sch65a, SC60, Sch64c, SH66c, SH67b, SH68, SOS56, Til60, Toh64, Uzu62, Wag67, WM59, Wei57, WS56, Whi69b, Wil63, YY66c, YS68, Atk68]. Utilization [Lan63a, LLMS67, Tay65]. Utilizing [Kae61, KII62, Len65, MC62, MWR68, O’N69, Pry58, Sob62].

V [Rub62b]. Value

[Bra66, Hol63b, Ter62, Vic69a]. Valued [AG68, Ber58, Kei64, Val67]. Values [Kri64, MP64, MP65]. Variable [Ano68-46, BF61, CET62, Dav69b, ET63, FH63, Har56, HW69, Hel63, HH53, ID63, JA67, Jon69a, Knu65a, KK68, Kor62b, Kro59, MO63, MU59, Mei68a, Mei68b, Mil67, Mur55, New66a, Pra66b, Ren68, SJ68, Smi65b, SD57, Tae69, Wra62, Zac61]. Variable-Length [HW69, KK68]. Variable-Radix [Ano68-46]. Variable-Rate [Har56]. Variable-Threshold [Mei68a, Mei68b, Wra62]. Variables [Dav57b, Gou58, Gre69a, Hau65a, Kov66, KO66, MN61, Nat64, Pol57, Sho64, Smi65a, Wil63]. Variant [Nie69a]. Variation [Auf62, Gau64]. Variational [Vic67b]. Variations [Jos66]. Varying [Bro69b]. Vector [Hau67, HF57, Par62]. Vehicle [MMB66b, Shu62]. Vehicles [FH62, Fog67]. Velocity [Gil59a]. Venture [Bak57]. Versatile [Gas64, MC65b]. Vertex [Boo62]. Vertex-Frame [Boo62]. Vertical [Men69b]. Vertices [Fro65, MO65]. Via [HS64a, Mil55]. Vibration [Cly59, Gre62, Pow59]. View [Kli67, Ros65a]. Virtual [Jon69b, McG68]. Visit [ZA59]. Visual [Rus68, Stu69]. Voice [Dav62]. Vol [Ano67a, Ano69b, Kar67]. Voltage [Ano69o, GS58, GB55, Jos66, NM68]. Voltages [GRT55]. Volume [Ano65s, Ano66p]. Voters [Gir66, Gur65]. Voting [GR68]. vs [SA64].

W [Ano64t, Hoa62]. W. [Ade62]. Wait [Rub62b]. Walk [CS68]. Wall [Bit62, Smi61, Sut69]. Walsh [Hen64a, Nam64, Sha69]. Walter [Ano62-43]. Wand [Can67a]. Ware [Ano58w]. Wave [O’N69]. Waveform [Chu66a, FR58]. Waveforms [Kno67b, O’N69]. Waves [Pfe59]. Way [Gre69b, MN61]. Ways [Ano57p]. Weapon [Bed64]. Weight [Kau60, Lip64, Mur65b, Oli65]. Weighted [Bey63, Bey64, Sma68, Wee60]. Weights [AM67b, Cro68, MK61, SH66a, SH68].

Werner [Str62]. Western [Bla56d]. Where [Ano55o]. Which [Fir65, Kam61, Mei66, TPM65, WP64, Uhr61]. Whiteside [Alo62]. Whose [Ard63, Auf62, Liu63a]. Wide [Tam67]. Williams [Gra56, Kle53, MH58, Won55]. Willis [Ano58w]. Wilson [Mac62, Shu62]. Wind [WHB56]. Wire [Cho67, Kau67a, Lot69, MMT66, SS59, Sho61, Waa67, Waa68]. Wiring [Bla56a]. Within [Lin60]. Without [AFM68, Ung68]. Word [BP64, FH63, Gre66b, Pet63, Sei67]. Word-Length [FH63]. Word-Recognition [BP64]. Words [Hig65]. Working [Fuc56]. Workshop [Ano66-35, Brz66, Fif66]. Worst [BdT65, Lyn69a, Wra62]. Worst-Case [BdT65, Lyn69a]. Woven [MMT66]. Write [Mat66, Waa68]. Writing [Har69a].

Year [Ano60-31]. Years [McL62]. Yielding [Cof68, Met62]. Yields [Whi69a].

ZAM [Luk63]. ZAM-41 [Luk63]. Zero [Hoa55].

References

Aldrich:1966:BTM

[AA66] W. H. Aldrich and R. L. Alonso. The “Braid” transformer memory. IEEE Transactions on Electronic Computers, EC-15(4): 502–508, August 1966. CO-DEN IEECA8. ISSN 0367-

                                                 7508. URL http://ieeexplore.
                                                 ieee.org/stamp/stamp.jsp?tp=&
                                                 arnumber=4038823.

Allen:1961:MMC

[ABC61] C. A. Allen, G. D. Bruce, and E. D. Councill. A 2.18microsecond megabit core storage unit. IRE Transactions on Electronic Computers, EC-10 (2):233–237, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219193.

Aberth:1969:MCL

[Abe69] O. Aberth. A multiple computer linkage. IEEE Transactions on Computers, C-18(12): 1132–1134, December 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671191.

Abhyankar:1958:MSP

[Abh58] Shreeram Abhyankar. Minimal “Sum of Products of Sums” expressions of Boolean functions. IRE Transactions on Electronic Computers, EC-7(4):268–276, December 1958. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222658.

Abhyankar:1959:AME

[Abh59] Shreeram Abhyankar. Absolute minimal expressions of Boolean functions. IRE Transactions on Electronic Computers, EC-8(1):3–8, March 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222754.

Alonso:1963:SAL

[ABSH63] R. L. Alonso, H. Blair-Smith, and A. L. Hopkins. Some aspects of the logical design of a control computer: A case study. IEEE Transactions on Electronic Computers, EC-12(6):687–697, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038019.

Amar:1967:DLC

[AC67] V. Amar and N. Condulmari. Diagnosis of large combinational networks. IEEE Transactions on Electronic Computers, EC-16(5): 675–680, October 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039162.

Armstrong:1960:MAR

[ACGN60] D. B. Armstrong, T. H. Crowley, U. F. Gianola, and E. E. Newhall. Magnetic analogs of relay contact networks for logic. IRE Transactions on Electronic Computers, EC-9(1):30–35, March 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221601.

Amarel:1964:MGN

[ACW64] S. Amarel, G. Cooke, and R. O. Winder. Majority gate networks. IEEE Transactions on Electronic Computers, EC-13(1): 4–13, February 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038070.

Anderson:1963:MTD

[AD63] D. J. Anderson and D. L. Dietmeyer. A magnetic ternary device. IEEE Transactions on Electronic Computers, EC-12(6): 911–914, December 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038040.

Abate:1969:OPD

[AD69] J. Abate and H. Dubner. Optimizing the performance of a drum-like storage. IEEE Transactions on Computers, C-18(11): 992–997, November 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671161.

Adey:1962:RBC

[Ade62] W. R. Adey. Review: Sensory Communication: Contributions to the Symposium on Principles of Sensory Communication, by W. A. Rosenblith. IRE Transactions on Electronic Computers, EC-11(3): 425, June 1962. CODEN IRELAO. ISSN 0367-9950.

Aoki:1961:BSU

[AE61] M. Aoki and G. Estrin. Bilateral switching using nonsymmetric elements. IRE Transactions on Electronic Computers, EC-10(1):42–50, March 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219150.

Armstrong:1968:RAS

[AFM68] D. B. Armstrong, A. D. Friedman, and P. R. Menon. Realization of asynchronous sequential circuits without inserted delay elements. IEEE Transactions on Computers, C-17(2):129–134, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687302.

Armstrong:1969:DAC

[AFM69] D. B. Armstrong, A. D. Friedman, and P. R. Menon. Design of asynchronous circuits assuming unbounded gate delays. IEEE Transactions on Computers, C-18(12):1110–1120, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671187.

Allen:1968:MTM

[AG68] C. M. Allen and D. D. Givone. A minimization technique for multiple-valued logic systems. IEEE Transactions on Computers, C-17 (2):182–184, February 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687309.

Aggarwal:1964:STO

[Agg64] G. K. Aggarwal. Simulation of third-order systems by a single operational amplifier. IEEE Transactions on Electronic Computers, EC-13(6):752, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038314.

Aggarwal:1965:EFG

[Agg65] G. K. Aggarwal. Effect of finite gain on one-amplifier circuits. IEEE Transactions on Electronic Computers, EC-14(3):491–492, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038475.

Aufenkamp:1957:ASM

[AH57] D. D. Aufenkamp and F. E.

Hohn. Analysis of sequential machines. IRE Transactions on Electronic Computers, EC-6(4): 276–285, December 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222039.

Allinnick:1962:MCA

[AH62] R. C. Allinnick and J. L. Haynes.

Magnetic core access switches. IRE Transactions on Electronic Computers, EC-11(3):352–368, June 1962. CODEN IRELAO. ISSN 0367-9950.

Arnold:1963:APS

[AH63] R. F. Arnold and M. A. Harrison. Algebraic properties of symmetric and partially symmetric Boolean functions. IEEE Transactions on Electronic Computers, EC-12(3):244–251, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037858.

Aho:1968:RPL

[Aho68] A. V. Aho. R68-27 programming languages for automata. IEEE Transactions on Computers, C-17(6):606, June 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687408.

Ahrons:1965:CSL

[Ahr65a] R. W. Ahrons. Calculations of speed of ladder network for superconductive associative memories. IEEE Transactions on Electronic Computers, EC-14(2):267–270, April 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038418.

Ahrons:1965:RBP

[Ahr65b] R. W. Ahrons. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-14 (2):275–280, April 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038421.

Aaronson:1960:MUD

[AJ60] D. A. Aaronson and D. B. James. Magnetostrictive ultrasonic delay lines for a PCM communication system. IRE Transactions on Electronic Computers, EC-9(3): 329–332, September 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219854.

Aus:1969:TLI

[AK69] H. M. Aus and G. A. Korn. Tablelookup/interpolation function generation for fixed-point digital computations. IEEE Transactions on Computers, C-18(8):745–749, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671351.

Akers:1961:AMM

[Ake61a] Sheldon B. Akers. On the algebraic manipulation of majority logic. IRE Transactions on Electronic Computers, EC-10 (4):779, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219289.

Akers:1961:TTM

[Ake61b] Sheldon B. Akers. A truth table method for the synthesis of combinational logic. IRE Transactions on Electronic Computers, EC-10(4):604–615, December 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219266.

Akers:1965:RBP

[Ake65a] S. B. Akers. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-14(4):667–671, August 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038529.

Akers:1965:DAM

[Ake65b] Sheldon B. Akers. A diagrammatic approach to multilevel logic synthesis. IEEE Transactions on Electronic Computers, EC-14 (2):174–181, April 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038396.

Akers:1965:CG

[Ake65c] Sheldon B. Akers. On the construction of (d, k) graphs. IEEE Transactions on Electronic Computers, EC-14(3):488, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038473.

Akers:1966:RBP

[Ake66] S. B. Akers. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-15(6): 953–962, December 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038949.

Akers:1967:MLP

[Ake67] Sheldon B. Akers. A modification of Lee’s path connection algorithm. IEEE Transactions on Electronic Computers, EC-16(1):97–98, February 1967.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038994.

Akers:1968:MIM

[Ake68] S. B. Akers, Jr. On maximum inversion with minimum inverters. IEEE Transactions on Computers, C-17(2):134–135, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687303.

Aleksander:1967:ANP

[Ale67] I. Aleksander. Array networks for a parallel adder and its control. IEEE Transactions on Electronic Computers, EC-16(2):226–229, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039037.

Alonso:1962:RBD

[Alo62] R. L. Alonso. Review: Tunnel Diode Storage Using Current Sensing, by E. R. Beck, D. A. Savitt, and A. E. Whiteside. IRE Transactions on Electronic Computers, EC-11(3): 422, June 1962. CODEN IRELAO. ISSN 0367-9950.

Alonso:1964:RRM

[Alo64] R. L. Alonso. R64-30 the rope memory — A permanent storage device. IEEE Transactions on Electronic Computers, EC-13(3):319, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038179.

Alrich:1955:EDE

[Alr55] John C. Alrich. Engineering description of the ElectroData digital computer. IRE Transactions on Electronic Computers, EC-4(1):1–10, March 1955. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407946.

Angelo:1968:STM

[ALS68] E. J. Angelo, Jr., J. Logan, and K. W. Sussman. The separation technique: A method for simulating transistors to aid integrated circuit design. IEEE Transactions on Computers, C-17(2):113–116, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687300.

Atkin:1962:IPD

[AM62] J. Atkin and N. B. Marple. Information processing by data interrogation. IRE Transactions on Electronic Computers, EC-11 (2):181–187, April 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219351.

Aaron:1967:NDD

[AM67a] M. R. Aaron and S. K. Mitra. A note on the design of digital-to-analog converters. IEEE Transactions on Electronic Computers, EC-16(5):685–686, October 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039166.

Aaron:1967:SRD

[AM67b] M. R. Aaron and S. K. Mitra. Synthesis of resistive digitalto-analog conversion ladders for arbitrary codes with fixed positive weights. IEEE Transactions on Electronic Computers, EC-16 (3):277–281, June 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039065.

Amarel:1964:PN

[Ama64a] S. Amarel. PTGEC news. IEEE Transactions on Electronic Computers, EC-13(1):81, February 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038089.

Amarel:1964:RAF

[Ama64b] S. Amarel. R64-50 on the analysis of functional symmetry. IEEE Transactions on Electronic Computers, EC-13(4):485, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038230.

Amari:1967:TAP

[Ama67] Shunichi Amari. A theory of adaptive pattern classifiers. IEEE Transactions on Electronic Computers, EC-16(3):299–307, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039068.

Amodei:1964:HSA

[Amo64] J. J. Amodei. High-speed adders and comparators using transistors and tunnel dioes. IEEE Transactions on Electronic Computers, EC-13(5):563–575, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038249.

Amoroso:1966:TEH

[Amo66] Frank Amoroso. Timing extraction at high bit densities. IEEE Transactions on Electronic Computers, EC-15(3):383–385, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038781.

Anderson:1953:EDL

[And53] J. R. Anderson. Electrical delay lines for digital computer applications. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(2):5–13, June 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Anderson:1956:NTF

[And56] John R. Anderson. A new type of ferroelectric shift register. IRE Transactions on Electronic Computers, EC-5(4):184–191, December 1956. CODEN IRELAO. ISSN 0367-9950.

Anderson:1962:MES

[And62] T. C. Anderson. A method for evaluating Stieltjes integrals on the analog computer. IRE Transactions on Electronic Computers, EC-11(4):552–555, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219395.

Anderson:1963:RMI

[And63] T. C. Anderson. R63-93 multiple integrals on a non-repetitive analog computer. IEEE Transactions on Electronic Computers, EC-12(5):582, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037999.

Anderson:1964:CGN

[And64] J. P. Anderson. Computer group news. IEEE Transactions on Electronic Computers, EC-13 (6):792, December 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038322.

Anderson:1966:AI

[And66] J. P. Anderson. Author index.

IEEE Transactions on Electronic Computers, EC-15(4):700, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038853.

Anderson:1967:RBP

[And67a] J. P. Anderson. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-16 (3):375–380, June 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039086.

Andrews:1967:EN

[And67b] J. Andrews. Editor’s notice. IEEE Transactions on Electronic Computers, EC-16(5):549, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039141.

Andrews:1968:HSA

[And68] H. Andrews. A high-speed algorithm for the computer generation of Fourier transforms. IEEE Transactions on Computers, C-17(4):373–375, April 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687348. See comments

[Mac69].

Anonymous:1953:Ca

[Ano53a] Anonymous. Contributors. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(3):12, ???? 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407734.

Anonymous:1953:Cb

[Ano53b] Anonymous. Contributors. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(4):27–28, December 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407691.

Anonymous:1953:E

[Ano53c] Anonymous. Editorial. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(4):1, December 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407684.

Anonymous:1953:NPS

[Ano53d] Anonymous. Notice to prospective speakers. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(4):1, December 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407685.

Anonymous:1953:RSa

[Ano53e] Anonymous. Review section.

Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(3):13–27, ???? 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407735.

Anonymous:1953:RSb

[Ano53f] Anonymous. Review section.

Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(4):29–47, December 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407692.

Anonymous:1955:CIP

[Ano55a] Anonymous. Constitution for IRE Professional Group on Electronic Computers. IRE Transactions on Electronic Computers, EC-4(3):88–92, September 1955. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407905.

Anonymous:1955:Cb

[Ano55b] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-4(1):32–33, March 1955. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407953.

Anonymous:1955:Cc

[Ano55c] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-4(2):75, June 1955. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407897.

Anonymous:1955:Cd

[Ano55d] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-4(3):119–120, September 1955. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5407913.

Anonymous:1955:Ce

[Ano55e] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-4(4):159–160, December 1955. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219488.

Anonymous:1955:Ca

[Ano55f] Anonymous. Correction. IRE Transactions on Electronic Computers, EC-4(1):20, March 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407949.

Anonymous:1955:PNa

[Ano55g] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-4(1):32, March 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407952.

Anonymous:1955:PNb

[Ano55h] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-4(2):76, June 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407898.

Anonymous:1955:PNc

[Ano55i] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-4(3):119, September 1955. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5407912.

Anonymous:1955:PNd

[Ano55j] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-4(4):160, December 1955. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219489.

Anonymous:1955:RCLa

[Ano55k] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-4(1):39–47, March 1955. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407955.

Anonymous:1955:RCLb

[Ano55l] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-4(2):77–85, June 1955. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407899.

Anonymous:1955:RCLc

[Ano55m] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-4(3):121–131, September 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407914.

Anonymous:1955:RCLd

[Ano55n] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-4(4):161–173, December 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219490.

Anonymous:1955:WSS

[Ano55o] Anonymous. Where should I send my manuscript? IRE Transactions on Electronic Computers, EC-4(3):87, September 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407904.

Anonymous:1956:Cb

[Ano56a] Anonymous.  Contents.  IRE

Transactions on Electronic Computers, EC-5(1):61–62, March 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219795.

Anonymous:1956:Ca

[Ano56b] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-5(1):35, March 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219790.

Anonymous:1956:Cc

[Ano56c] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-5(2):94–95, June 1956.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219807.

Anonymous:1956:Ce

[Ano56d] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-5(3):168, September 1956. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219941.

Anonymous:1956:Cf

[Ano56e] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-5(4):256–257, December 1956. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219961.

Anonymous:1956:Cd

[Ano56f] Anonymous. Correction. IRE Transactions on Electronic Computers, EC-5(3):167, September 1956. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219940.

Anonymous:1956:IIT

[Ano56g] Anonymous. Index to IRE Transactions on Electronic Computers. IRE Transactions on Electronic Computers, EC-5(4): 1–4, December 1956. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219964.

Anonymous:1956:PNa

[Ano56h] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-5(1):42, March 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219792.

Anonymous:1956:PNb

[Ano56i] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-5(2):95–96, June 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219808.

Anonymous:1956:PNc

[Ano56j] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-5(3):169, September 1956. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219942.

Anonymous:1956:PNd

[Ano56k] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-5(4):257, December 1956. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219962.

Anonymous:1956:PPA

[Ano56l] Anonymous. PGEC papers awards for 1955. IRE Transactions on Electronic Computers, EC-5(2):63–64, June 1956. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219800.

Anonymous:1956:RCLa

[Ano56m] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-5(1):47–59, March 1956. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219794.

Anonymous:1956:RCLb

[Ano56n] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-5(2):97–109, June 1956. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219809.

Anonymous:1956:RCLc

[Ano56o] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-5(3):171–181, September 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219943.

Anonymous:1956:RCLd

[Ano56p] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-5(4):259–273, December 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219963.

Anonymous:1956:SIC

[Ano56q] Anonymous. Symposium on the impact of computers on science and society. IRE Transactions on Electronic Computers, EC-5(3): 142–158, September 1956. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219938.

Anonymous:1956:SDM

[Ano56r] Anonymous. Symposium: The design of machines to simulate the behavior of the human brain. IRE Transactions on Electronic Computers, EC-5(4): 240–255, December 1956. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219960.

Anonymous:1957:Ca

[Ano57a] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-6(1):60–61, March 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221566.

Anonymous:1957:Cb

[Ano57b] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-6(2):124–125, June 1957. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221588.

Anonymous:1957:Cd

[Ano57c] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-6(3):205–207, September 1957. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222026.

Anonymous:1957:Ce

[Ano57d] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-6(4):288, December 1957. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222043.

Anonymous:1957:Cc

[Ano57e] Anonymous. Correction. IRE Transactions on Electronic Computers, EC-6(3):202, September 1957. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222023.

Anonymous:1957:IIT

[Ano57f] Anonymous. Index to IRE Transactions on Electronic Computers. IRE Transactions on Electronic Computers, EC-6(4): 1–4, December 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222046.

Anonymous:1957:PNa

[Ano57g] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-6(1):61–62, March 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221567.

Anonymous:1957:PNb

[Ano57h] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-6(2):126–127, June 1957. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221589.

Anonymous:1957:PNc

[Ano57i] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-6(3):208, September 1957. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222027.

Anonymous:1957:PNd

[Ano57j] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-6(4):289, December 1957. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222044.

Anonymous:1957:REC

[Ano57k] Anonymous. Review of electronic computer progress during 1956. IRE Transactions on Electronic Computers, EC-6(1):55–60, March 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221565.

Anonymous:1957:RCLa

[Ano57l] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-6(1):63–69, March 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221568.

Anonymous:1957:RCLb

[Ano57m] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-6(2):129–141, June 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221590.

Anonymous:1957:RCLc

[Ano57n] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-6(3):209–229, September 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222028.

Anonymous:1957:RCLd

[Ano57o] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-6(4):291–307, December 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222045.

Anonymous:1957:WDS

[Ano57p] Anonymous. Ways of developing Soviet computer production. IRE Transactions on Electronic Computers, EC-6(1):37–49, March 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221563.

Anonymous:1958:CCa

[Ano58a] Anonymous. The Chairman’s column. IRE Transactions on Electronic Computers, EC-7(3): 189, September 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222572.

Anonymous:1958:Cb

[Ano58b] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-7(1):73–74, March 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222101.

Anonymous:1958:Cd

[Ano58c] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-7(2):183–184, June 1958. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222533.

Anonymous:1958:Cg

[Ano58d] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-7(3):252–253, September 1958. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222587.

Anonymous:1958:Ci

[Ano58e] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-7(4):326–327, December 1958. CODEN IRELAO.

ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222669.

Anonymous:1958:Ca

[Ano58f] Anonymous. Correction. IRE Transactions on Electronic Computers, EC-7(1):60, March 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222098.

Anonymous:1958:Cc

[Ano58g] Anonymous. Correction. IRE Transactions on Electronic Computers, EC-7(2):122, June 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222521.

Anonymous:1958:Ch

[Ano58h] Anonymous. Correction. IRE Transactions on Electronic Computers, EC-7(4):324, December 1958. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222667.

Anonymous:1958:Ce

[Ano58i] Anonymous. Corrections. IRE Transactions on Electronic Computers, EC-7(3):250, September 1958. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222584.

Anonymous:1958:Cf

[Ano58j] Anonymous. Corrections. IRE Transactions on Electronic Computers, EC-7(3):250, September 1958. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222585.

Anonymous:1958:HT

[Ano58k] Anonymous. Howard E. Tompkins. IRE Transactions on Electronic Computers, EC-7(4): 261, December 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222655.

Anonymous:1958:IIT

[Ano58l] Anonymous. Index to IRE Transactions on Electronic Computers. IRE Transactions on Electronic Computers, EC-7(4): 331–334, December 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222672.

Anonymous:1958:PNa

[Ano58m] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-7(1):75–76, March 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222102.

Anonymous:1958:PNb

[Ano58n] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-7(2):188, June 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222535.

Anonymous:1958:PNc

[Ano58o] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-7(3):259, September 1958. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5222589.

Anonymous:1958:PNd

[Ano58p] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-7(4):327, December 1958. CODEN IRELAO. ISSN

                                        0367-9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222670.

Anonymous:1958:QD

[Ano58q] Anonymous. Questions and discussion. IRE Transactions on Electronic Computers, EC-7 (2):150–155, June 1958. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222527.

Anonymous:1958:RB

[Ano58r] Anonymous. Reviews and bibliographies. IRE Transactions on Electronic Computers, EC-7(1):1, March 1958. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5222087.

Anonymous:1958:RCL

[Ano58s] Anonymous. Reviews of current literature. IRE Transactions on Electronic Computers, EC-7(1):77–89, March 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222103.

Anonymous:1958:SSEa

[Ano58t] Anonymous. SENEWS science education subcommittee newsletter. IRE Transactions on Electronic Computers, EC-7(2):185– 188, June 1958. CODEN IRE-LAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222534.

Anonymous:1958:SSEb

[Ano58u] Anonymous. SENEWS science education subcommittee newsletter. IRE Transactions on Electronic Computers, EC-7(3):254– 258, September 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222588.

Anonymous:1958:SSEc

[Ano58v] Anonymous. SENEWS science education subcommittee newsletter. IRE Transactions on Electronic Computers, EC-7(4):328– 330, December 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222671.

Anonymous:1958:WHW

[Ano58w] Anonymous. Willis H. Ware.

IRE Transactions on Electronic Computers, EC-7(3):190, September 1958. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222573.

Anonymous:1959:IAC

[Ano59a] Anonymous. 1959 index to abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-8(4): 521–531, December 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222071.

Anonymous:1959:ACCa

[Ano59b] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-8(1):69–79, March 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222765.

Anonymous:1959:ACCb

[Ano59c] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-8(2):236–247, June 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219533.

Anonymous:1959:ACCc

[Ano59d] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-8(3):409–420, September 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222705.

Anonymous:1959:ACCd

[Ano59e] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-8(4):507–520, December 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222070.

Anonymous:1959:AIa

[Ano59f] Anonymous. Author index. IRE Transactions on Electronic Computers, EC-8(1):82, March 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222767.

Anonymous:1959:AIb

[Ano59g] Anonymous. Author index. IRE Transactions on Electronic Computers, EC-8(2):250, June 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219535.

Anonymous:1959:AIc

[Ano59h] Anonymous. Author index.

IRE Transactions on Electronic Computers, EC-8(3):423, September 1959. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222707.

Anonymous:1959:CCa

[Ano59i] Anonymous. The Chairman’s column. IRE Transactions on Electronic Computers, EC-8(1):1– 2, March 1959. CODEN IRE-LAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222753.

Anonymous:1959:CCb

[Ano59j] Anonymous. The Chairman’s column. IRE Transactions on Electronic Computers, EC-8(2):90–91, June 1959. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219507.

Anonymous:1959:CCc

[Ano59k] Anonymous. The Chairman’s column. IRE Transactions on Electronic Computers, EC-8(3): 261, September 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222679.

Anonymous:1959:Ca

[Ano59l] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-8(1):83–84, March 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222768.

Anonymous:1959:Cb

[Ano59m] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-8(2):231–235, June 1959. CODEN IRELAO. ISSN

                                        0367-9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219532.

Anonymous:1959:Cc

[Ano59n] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-8(3):402–406, September 1959. CODEN IRELAO.

ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222703.

Anonymous:1959:Cd

[Ano59o] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-8(4):500–502, December 1959. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222068.

Anonymous:1959:IIT

[Ano59p] Anonymous. Index to IRE Transactions on Electronic Computers. IRE Transactions on Electronic Computers, EC-8(4): 1–6, December 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222074.

Anonymous:1959:IA

[Ano59q] Anonymous. Information for authors. IRE Transactions on Electronic Computers, EC-8(2): 253–260, June 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219537.

Anonymous:1959:PN

[Ano59r] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-8(1):87, March 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222770.

Anonymous:1959:PNNa

[Ano59s] Anonymous. PGEC news and notices. IRE Transactions on Electronic Computers, EC-8(2): 251–252, June 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219536.

Anonymous:1959:PNNb

[Ano59t] Anonymous. PGEC news and notices. IRE Transactions on Electronic Computers, EC-8(3): 430, September 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222709.

Anonymous:1959:PNNc

[Ano59u] Anonymous. PGEC news and notices. IRE Transactions on Electronic Computers, EC-8(4): 532, December 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222072.

Anonymous:1959:RBPa

[Ano59v] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-8(3):407– 408, September 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222704.

Anonymous:1959:RBPb

[Ano59w] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-8(4): 503–506, December 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222069.

Anonymous:1959:RE

[Ano59x] Anonymous. Richard O. Endres.

IRE Transactions on Electronic Computers, EC-8(2):89, June 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219506.

Anonymous:1959:Sa

[Ano59y] Anonymous. SENEWS. IRE Transactions on Electronic Computers, EC-8(1):85–86, March 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222769.

Anonymous:1959:Sb

[Ano59z] Anonymous. SENEWS. IRE Transactions on Electronic Computers, EC-8(3):424–429, September 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222708.

Anonymous:1959:Sc

[Ano59-27] Anonymous. SENEWS. IRE Transactions on Electronic Computers, EC-8(4):533–534, December 1959. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222073.

Anonymous:1959:SIa

[Ano59-28] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-8(1):79–82, March 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222766.

Anonymous:1959:SIb

[Ano59-29] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-8(2):248–249, June 1959. CODEN IRELAO. ISSN

                                        0367-9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219534.

Anonymous:1959:SIc

[Ano59-30] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-8(3):421–423, September 1959. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5222706.

Anonymous:1960:ACCa

[Ano60a] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-9(1):134–148, March 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221615.

Anonymous:1960:ACCb

[Ano60b] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-9(2):275–288, June 1960.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219837.

Anonymous:1960:ACCc

[Ano60c] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-9(3):389–404, September 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219867.

Anonymous:1960:ACCd

[Ano60d] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-9(4):531–545, December 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219900.

Anonymous:1960:ACE

[Ano60e] Anonymous. Arnold A. Cohen: Elected Chairman, PGEC, 1960–1961. IRE Transactions on Electronic Computers, EC-9(1):1, March 1960. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5221596.

Anonymous:1960:AIa

[Ano60f] Anonymous. Author index. IRE Transactions on Electronic Computers, EC-9(1):151, March 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221617.

Anonymous:1960:AIb

[Ano60g] Anonymous. Author index. IRE Transactions on Electronic Computers, EC-9(2):291–292, June 1960. CODEN IRELAO. ISSN

                                        0367-9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219839.

Anonymous:1960:AIc

[Ano60h] Anonymous. Author index. IRE Transactions on Electronic Computers, EC-9(3):404–405, September 1960. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219868.

Anonymous:1960:Ca

[Ano60i] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-9(1):125–129, March 1960. CODEN IRELAO. ISSN

                                        0367-9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221613.

Anonymous:1960:Cc

[Ano60j] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-9(2):267–269, June 1960. CODEN IRELAO. ISSN

                                        0367-9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219835.

Anonymous:1960:Cd

[Ano60k] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-9(3):374–376, September 1960. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219865.

Anonymous:1960:Cf

[Ano60l] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-9(4):511–514, December 1960. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219898.

Anonymous:1960:Cb

[Ano60m] Anonymous. Correction. IRE Transactions on Electronic Computers, EC-9(2):261, June 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219827.

Anonymous:1960:Ce

[Ano60n] Anonymous. Correction. IRE Transactions on Electronic Computers, EC-9(4):509, December 1960. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219895.

Anonymous:1960:IIT

[Ano60o] Anonymous. Index to IRE Transactions on Electronic Computers. IRE Transactions on Electronic Computers, EC-9(4): 1–9, December 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219904.

Anonymous:1960:IA

[Ano60p] Anonymous. Information for authors. IRE Transactions on Electronic Computers, EC-9(4):nil5, December 1960. CODEN IRELAO. ISSN 0367-9950.

Anonymous:1960:Na

[Ano60q] Anonymous.  Notices.  IRE

Transactions on Electronic Computers, EC-9(2):294, June 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219841.

Anonymous:1960:Nb

[Ano60r] Anonymous. Notices. IRE Transactions on Electronic Computers, EC-9(3):411–413, September 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219871.

Anonymous:1960:Nc

[Ano60s] Anonymous. Notices. IRE Transactions on Electronic Computers, EC-9(4):574–578, December 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219903.

Anonymous:1960:PNb

[Ano60t] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-9(3):409–410, September 1960. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219870.

Anonymous:1960:PNc

[Ano60u] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-9(4):573, December 1960. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219902.

Anonymous:1960:PNN

[Ano60v] Anonymous. PGEC news and notices. IRE Transactions on Electronic Computers, EC-9(1): 152–153, March 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221618.

Anonymous:1960:PNa

[Ano60w] Anonymous. PGEC notices. IRE Transactions on Electronic Computers, EC-9(2):293, June 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219840.

Anonymous:1960:RBPa

[Ano60x] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-9(1): 130–133, March 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221614.

Anonymous:1960:RBPb

[Ano60y] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-9(2): 270–274, June 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219836.

Anonymous:1960:RBPc

[Ano60z] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-9(3):377– 388, September 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219866.

Anonymous:1960:RBPd

[Ano60-27] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-9(4): 515–530, December 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219899.

Anonymous:1960:SIa

[Ano60-28] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-9(1):148–151, March 1960. CODEN IRELAO. ISSN

                                        0367-9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221616.

Anonymous:1960:SIb

[Ano60-29] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-9(2):288–291, June 1960. CODEN IRELAO. ISSN

                                        0367-9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219838.

Anonymous:1960:SIc

[Ano60-30] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-9(3):405–408, September 1960. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219869.

Anonymous:1960:TYC

[Ano60-31] Anonymous. Two-year cumulative index for abstracts of current computer literature 1959– 1960. IRE Transactions on Electronic Computers, EC-9(4): 547–572, December 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219901.

Anonymous:1961:ACCa

[Ano61a] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-10(1):118–136, March 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219170.

Anonymous:1961:ACCb

[Ano61b] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-10(2):316–332, June 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219209.

Anonymous:1961:ACCc

[Ano61c] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-10(3):555–571, September 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219252.

Anonymous:1961:ACCd

[Ano61d] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-10(4):797–843, December 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219297.

Anonymous:1961:AIa

[Ano61e] Anonymous. Author index. IRE Transactions on Electronic Computers, EC-10(1):137, March 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219171.

Anonymous:1961:AIb

[Ano61f] Anonymous. Author index. IRE Transactions on Electronic Computers, EC-10(2):335–336, June 1961. CODEN IRELAO. ISSN

                                        0367-9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219211.

Anonymous:1961:AIc

[Ano61g] Anonymous. Author index. IRE Transactions on Electronic Computers, EC-10(3):574, September 1961. CODEN IRELAO.

ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219254.

Anonymous:1961:Cb

[Ano61h] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-10(1):102–104, March 1961. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219168.

Anonymous:1961:Cc

[Ano61i] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-10(2):292–295, June 1961. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219207.

Anonymous:1961:Ce

[Ano61j] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-10(3):535–540, September 1961. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219250.

Anonymous:1961:Cf

[Ano61k] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-10(4):784–788, December 1961. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219295.

Anonymous:1961:Ca

[Ano61l] Anonymous. Correction. IRE Transactions on Electronic Computers, EC-10(1):62, March 1961.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219153.

Anonymous:1961:Cd

[Ano61m] Anonymous. Correction. IRE Transactions on Electronic Computers, EC-10(3):461, September 1961. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219235.

Anonymous:1961:Ea

[Ano61n] Anonymous.  Editorial.  IRE

Transactions on Electronic Computers, EC-10(2):150, June 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219182.

Anonymous:1961:Eb

[Ano61o] Anonymous. Editorial. IRE Transactions on Electronic Computers, EC-10(3):345, September 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219221.

Anonymous:1961:IIT

[Ano61p] Anonymous. Index to IRE Transactions on Electronic Computers. IRE Transactions on Electronic Computers, EC-10(4): 849–859, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219300.

Anonymous:1961:IAa

[Ano61q] Anonymous. Information for authors. IRE Transactions on Electronic Computers, EC-10(2):343, June 1961. CODEN IRELAO. ISSN 0367-9950.

Anonymous:1961:IAb

[Ano61r] Anonymous. Information for authors. IRE Transactions on Electronic Computers, EC-10(3):585, September 1961. CODEN IRELAO. ISSN 0367-9950.

Anonymous:1961:NEC

[Ano61s] Anonymous. New Editor-in-Chief of IRETEC. IRE Transactions on Electronic Computers, EC-10(2):149, April 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219181.

Anonymous:1961:Na

[Ano61t] Anonymous. Notices. IRE Transactions on Electronic Computers, EC-10(1):143–147, March 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219174.

Anonymous:1961:Nb

[Ano61u] Anonymous. Notices. IRE Transactions on Electronic Computers, EC-10(2):341–342, June 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219213.

Anonymous:1961:Nc

[Ano61v] Anonymous. Notices. IRE Transactions on Electronic Computers, EC-10(3):579–584, September 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219256.

Anonymous:1961:Nd

[Ano61w] Anonymous. Notices. IRE Transactions on Electronic Computers, EC-10(4):847–848, December 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219299.

Anonymous:1961:PNa

[Ano61x] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-10(1):142–143, March 1961. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219173.

Anonymous:1961:PNb

[Ano61y] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-10(2):337–340, June 1961. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219212.

Anonymous:1961:PNc

[Ano61z] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-10(3):575–578, September 1961. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219255.

Anonymous:1961:PNd

[Ano61-27] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-10(4):845–846, December 1961. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219298.

Anonymous:1961:RBPa

[Ano61-28] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-10(1): 105–117, March 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219169.

Anonymous:1961:RBPb

[Ano61-29] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-10(2): 296–315, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219208.

Anonymous:1961:RBPc

[Ano61-30] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-10(3):541– 554, September 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219251.

Anonymous:1961:RBPd

[Ano61-31] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-10(4): 789–796, December 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219296.

Anonymous:1961:SIa

[Ano61-32] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-10(1):138–141, March 1961. CODEN IRELAO. ISSN

                                        0367-9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219172.

Anonymous:1961:SIb

[Ano61-33] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-10(2):332–335, June 1961. CODEN IRELAO. ISSN

                                        0367-9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219210.

Anonymous:1961:SIc

[Ano61-34] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-10(3):571–574, September 1961. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219253.

Anonymous:1962:ACCa

[Ano62a] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-11(1):105–115, February 1962.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219332.

Anonymous:1962:ACCb

[Ano62b] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-11(2):305–316, April 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219371.

Anonymous:1962:ACC

[Ano62c] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-11(3):427–??, June 1962. CODEN IRELAO. ISSN 0367-9950.

Anonymous:1962:ACCc

[Ano62d] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-11(4):591–603, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219411.

Anonymous:1962:ACCd

[Ano62e] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-11(5):727–737, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219446.

Anonymous:1962:ACCe

[Ano62f] Anonymous. Abstracts of current computer literature. IRE Transactions on Electronic Computers, EC-11(6):809–832, December 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219471.

Anonymous:1962:AN

[Ano62g] Anonymous. AFIPS news. IRE Transactions on Electronic Computers, EC-11(5):740–741, October 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219450.

Anonymous:1962:AIa

[Ano62h] Anonymous. Author index. IRE Transactions on Electronic Computers, EC-11(1):118, February 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219334.

Anonymous:1962:AIb

[Ano62i] Anonymous. Author index. IRE Transactions on Electronic Computers, EC-11(2):318, April 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219373.

Anonymous:1962:AIc

[Ano62j] Anonymous. Author index. IRE Transactions on Electronic Computers, EC-11(4):606, August 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219413.

Anonymous:1962:AId

[Ano62k] Anonymous. Author index.

IRE Transactions on Electronic Computers, EC-11(5):739, October 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219448.

Anonymous:1962:Ca

[Ano62l] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-11(1):90–92, February 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219330.

Anonymous:1962:Cb

[Ano62m] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-11(2):291–294, April 1962. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219369.

Anonymous:1962:Cd

[Ano62n] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-11(4):580–582, August 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219409.

Anonymous:1962:Cf

[Ano62o] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-11(5):715–716, October 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219444.

Anonymous:1962:Cg

[Ano62p] Anonymous. Contributors. IRE Transactions on Electronic Computers, EC-11(6):797, December 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219469.

Anonymous:1962:Ce

[Ano62q] Anonymous. Correction. IRE Transactions on Electronic Computers, EC-11(5):714, October 1962. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219443.

Anonymous:1962:INa

[Ano62r] Anonymous. IFIP news. IRE Transactions on Electronic Computers, EC-11(1):120, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219336.

Anonymous:1962:INb

[Ano62s] Anonymous. IFIP news. IRE Transactions on Electronic Computers, EC-11(6):833, December 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219473.

Anonymous:1962:IIT

[Ano62t] Anonymous. Index to IRE Transactions on Electronic Computers. IRE Transactions on Electronic Computers, EC-11(6):

836–845, December 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219475.

Anonymous:1962:Na

[Ano62u] Anonymous. Notices. IRE Transactions on Electronic Computers, EC-11(1):120–122, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219337.

Anonymous:1962:Nb

[Ano62v] Anonymous. Notices. IRE Transactions on Electronic Computers, EC-11(2):321–322, April 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219376.

Anonymous:1962:Nc

[Ano62w] Anonymous. Notices. IRE Transactions on Electronic Computers, EC-11(4):607–609, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219415.

Anonymous:1962:Nd

[Ano62x] Anonymous. Notices. IRE Transactions on Electronic Computers, EC-11(5):741–742, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219451.

Anonymous:1962:Ne

[Ano62y] Anonymous. Notices. IRE Transactions on Electronic Computers, EC-11(6):834–835, December 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219474.

Anonymous:1962:O

[Ano62z] Anonymous. Obituary. IRE Transactions on Electronic Computers, EC-11(2):320, April 1962. CODEN IRELAO. ISSN 0367-9950.

Anonymous:1962:PNa

[Ano62-27] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-11(1):119, February 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219335.

Anonymous:1962:PNb

[Ano62-28] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-11(2):319, April 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219374.

Anonymous:1962:PNc

[Ano62-29] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-11(4):607, August 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219414.

Anonymous:1962:PNd

[Ano62-30] Anonymous. PGEC news. IRE Transactions on Electronic Computers, EC-11(5):740, October 1962. CODEN IRELAO.

ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219449.

Anonymous:1962:PNe

[Ano62-31] Anonymous. PGEC news.

IRE Transactions on Electronic Computers, EC-11(6):833, December 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219472.

Anonymous:1962:PIS

[Ano62-32] Anonymous. Proposed IRE standards for analog computers. IRE Transactions on Electronic Computers, EC-11(1):67–79, February 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219319.

Anonymous:1962:RBPa

[Ano62-33] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-11(1): 93–104, February 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219331.

Anonymous:1962:RBPb

[Ano62-34] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-11(2): 295–303, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219370.

Anonymous:1962:RBPc

[Ano62-35] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-11(4): 583–590, August 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219410.

Anonymous:1962:RBPd

[Ano62-36] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-11(5): 717–726, October 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219445.

Anonymous:1962:RBPe

[Ano62-37] Anonymous. Reviews of books and papers in the computer field. IRE Transactions on Electronic Computers, EC-11(6): 798–807, December 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219470.

Anonymous:1962:SAH

[Ano62-38] Anonymous. Special analoghybrid computer issue. IRE Transactions on Electronic Computers, EC-11(1):1, February 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219307.

Anonymous:1962:SIa

[Ano62-39] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-11(1):116–117, February 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219333.

Anonymous:1962:SIb

[Ano62-40] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-11(2):316–318, April 1962. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219372.

Anonymous:1962:SIc

[Ano62-41] Anonymous. Subject index.

IRE Transactions on Electronic Computers, EC-11(4):604–605, August 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219412.

Anonymous:1962:SId

[Ano62-42] Anonymous. Subject index. IRE Transactions on Electronic Computers, EC-11(5):737–739, October 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219447.

Anonymous:1962:WAC

[Ano62-43] Anonymous. Walter L. Anderson, Chairman, 1962–1963. IRE Transactions on Electronic Computers, EC-11(3):323, June 1962. CODEN IRELAO. ISSN 0367-9950.

Anonymous:1963:ACCa

[Ano63a] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(1):39–54, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037791.

Anonymous:1963:ACCb

[Ano63b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(2):159–175, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037842.

Anonymous:1963:ACCc

[Ano63c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(3):337–351, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037895.

Anonymous:1963:ACCd

[Ano63d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(4):425–436, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037929.

Anonymous:1963:ACCe

[Ano63e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(4):425–439, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037947.

Anonymous:1963:ACCf

[Ano63f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(5):585–599, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038002.

Anonymous:1963:ACCg

[Ano63g] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(6):937–961, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038059.

Anonymous:1963:CPa

[Ano63h] Anonymous. Call for papers. IEEE Transactions on Electronic Computers, EC-12(1):1, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037757.

Anonymous:1963:CPb

[Ano63i] Anonymous. Call for papers. IEEE Transactions on Electronic Computers, EC-12(3):182, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037850.

Anonymous:1963:CPc

[Ano63j] Anonymous. Call for papers. IEEE Transactions on Electronic Computers, EC-12(4):355, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037906.

Anonymous:1963:Ca

[Ano63k] Anonymous. Contributors. IEEE Transactions on Electronic Computers, EC-12(2):149, April 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037821.

Anonymous:1963:Cd

[Ano63l] Anonymous. Contributors. IEEE Transactions on Electronic Computers, EC-12(4):415, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037927.

Anonymous:1963:Cb

[Ano63m] Anonymous. Correction. IEEE Transactions on Electronic Computers, EC-12(3):232, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037856.

Anonymous:1963:Cc

[Ano63n] Anonymous. Correction. IEEE Transactions on Electronic Computers, EC-12(4):400, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037914.

Anonymous:1963:Ce

[Ano63o] Anonymous. Correction. IEEE Transactions on Electronic Computers, EC-12(5):550, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037971.

Anonymous:1963:ITE

[Ano63p] Anonymous. IEEE Transactions on Electronic Computers. IEEE Transactions on Electronic Computers, EC-12(6):965–965–i, December 1963. CODEN IEECA8. ISSN 0367-7508.

Anonymous:1963:IN

[Ano63q] Anonymous. IFIP news. IEEE Transactions on Electronic Computers, EC-12(1):56, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037793.

Anonymous:1963:IAa

[Ano63r] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-12(1): 59–b, February 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037796.

Anonymous:1963:IAb

[Ano63s] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-12(2):179–b, April 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037846.

Anonymous:1963:IAc

[Ano63t] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-12(3):354–c, June 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037901.

Anonymous:1963:IAd

[Ano63u] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-12 (4):442–a, August 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037951.

Anonymous:1963:IAe

[Ano63v] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-12(5): 606–a, October 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038005.

Anonymous:1963:IAf

[Ano63w] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-12(6): 965–j, December 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038064.

Anonymous:1963:MPTa

[Ano63x] Anonymous. Membership in the Professional Technical Group on Electronic Computers. IEEE Transactions on Electronic Computers, EC-12(3):354, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037898.

Anonymous:1963:MPTb

[Ano63y] Anonymous. Membership in the Professional Technical Group on Electronic Computers. IEEE Transactions on Electronic Computers, EC-12(4):442, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037950.

Anonymous:1963:MPTc

[Ano63z] Anonymous. Membership in the Professional Technical Group on Electronic Computers. IEEE Transactions on Electronic Computers, EC-12(6):964, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038062.

Anonymous:1963:Na

[Ano63-27] Anonymous. Notices. IEEE Transactions on Electronic Computers, EC-12(1):57–59, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037794.

Anonymous:1963:Nb

[Ano63-28] Anonymous. Notices. IEEE Transactions on Electronic Computers, EC-12(2):177–179, April 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037844.

Anonymous:1963:Nc

[Ano63-29] Anonymous. Notices. IEEE Transactions on Electronic Computers, EC-12(3):353, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037897.

Anonymous:1963:Nd

[Ano63-30] Anonymous. Notices. IEEE Transactions on Electronic Computers, EC-12(4):441, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037949.

Anonymous:1963:Ne

[Ano63-31] Anonymous. Notices. IEEE Transactions on Electronic Computers, EC-12(5):601–606, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038004.

Anonymous:1963:Nf

[Ano63-32] Anonymous. Notices. IEEE Transactions on Electronic Computers, EC-12(6):963, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038061.

Anonymous:1963:PNc

[Ano63-33] Anonymous. PGEC news. IEEE Transactions on Electronic Computers, EC-12(3):352, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037896.

Anonymous:1963:PNa

[Ano63-34] Anonymous. PTGEC news. IEEE Transactions on Electronic Computers, EC-12(1):55–56, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037792.

Anonymous:1963:PNb

[Ano63-35] Anonymous. PTGEC news. IEEE Transactions on Electronic Computers, EC-12(2):176, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037843.

Anonymous:1963:PNd

[Ano63-36] Anonymous. PTGEC news. IEEE Transactions on Electronic Computers, EC-12(4):440, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037948.

Anonymous:1963:PNe

[Ano63-37] Anonymous. PTGEC news. IEEE Transactions on Electronic Computers, EC-12(5):600–601, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038003.

Anonymous:1963:PNf

[Ano63-38] Anonymous. PTGEC news. IEEE Transactions on Electronic Computers, EC-12(6):962, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038060.

Anonymous:1963:PP

[Ano63-39] Anonymous. PTGEC people. IEEE Transactions on Electronic Computers, EC-12(1): 59–59–a, February 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037795.

Anonymous:1963:RER

[Ano63-40] Anonymous. R63-92 on the effectiveness of receptors in recognition systems. IEEE Transactions on Electronic Computers, EC-12(5):582, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037998.

Anonymous:1964:ACCa

[Ano64a] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(1):67–80, February 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038093.

Anonymous:1964:ACCb

[Ano64b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(2):165–179, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038133.

Anonymous:1964:ACCc

[Ano64c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(3):323–340, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038188.

Anonymous:1964:ACCd

[Ano64d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(4):487–501, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038233.

Anonymous:1964:ACCe

[Ano64e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(5):651–668, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038290.

Anonymous:1964:ACCf

[Ano64f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(6):763–791, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038333.

Anonymous:1964:CGNa

[Ano64g] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-13(3): 341–342, June 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038189.

Anonymous:1964:CGNb

[Ano64h] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-13(4): 502–503, August 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038234.

Anonymous:1964:CGNc

[Ano64i] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-13(5):669, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038291.

Anonymous:1964:CGNd

[Ano64j] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-13(6): 792, December 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038334.

Anonymous:1964:Ca

[Ano64k] Anonymous. Contributors. IEEE Transactions on Electronic Computers, EC-13(2):158–159, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038121.

Anonymous:1964:Cb

[Ano64l] Anonymous. Contributors. IEEE Transactions on Electronic Computers, EC-13(5):643–644, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038274.

Anonymous:1964:IAf

[Ano64m] Anonymous. Index author. IEEE Transactions on Electronic Computers, EC-13(6):793–794–j, December 1964. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038335.

Anonymous:1964:IAa

[Ano64n] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-13(1): 82–c, February 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038098.

Anonymous:1964:IAb

[Ano64o] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-13(2):182–a, April 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038136.

Anonymous:1964:IAc

[Ano64p] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-13(3):342–a, June 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038190.

Anonymous:1964:IAd

[Ano64q] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-13 (4):503–d, August 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038238.

Anonymous:1964:IAe

[Ano64r] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-13(5): 672–a, October 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038293.

Anonymous:1964:IAg

[Ano64s] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-13(6): 794–k, December 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038336.

Anonymous:1964:KWU

[Ano64t] Anonymous. Keith W. Uncapher new Chairman of the IEEE Computer Group. IEEE Transactions on Electronic Computers, EC-13(3):183, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038141.

Anonymous:1964:NAE

[Ano64u] Anonymous. New Associate Editor for logic and switching theory. IEEE Transactions on Electronic Computers, EC-13(5):505, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038243.

Anonymous:1964:Na

[Ano64v] Anonymous.  Notices.  IEEE

Transactions on Electronic Computers, EC-13(1):82, February 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038095.

Anonymous:1964:Nb

[Ano64w] Anonymous.  Notices.  IEEE

Transactions on Electronic Computers, EC-13(2):181–182, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038135.

Anonymous:1964:Nc

[Ano64x] Anonymous. Notices. IEEE Transactions on Electronic Computers, EC-13(5):670–672, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038292.

Anonymous:1964:PNa

[Ano64y] Anonymous. PTGEC news. IEEE Transactions on Electronic Computers, EC-13(1):81, February 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038094.

Anonymous:1964:PNb

[Ano64z] Anonymous. PTGEC news. IEEE Transactions on Electronic Computers, EC-13(2):180–181, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038134.

Anonymous:1964:SIC

[Ano64-27] Anonymous. The special issue on computer languages editorial. IEEE Transactions on Electronic Computers, EC-13(4):343–344, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038195.

Anonymous:1965:ACCa

[Ano65a] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-14(1):95–109, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038376.

Anonymous:1965:ACCb

[Ano65b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-14(2):281–298, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038439.

Anonymous:1965:ACCc

[Ano65c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-14(3):503–521, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038492.

Anonymous:1965:ACCd

[Ano65d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-14(4):651–662, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038526.

Anonymous:1965:ACCe

[Ano65e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-14(5):747–761, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038581.

Anonymous:1965:ACCf

[Ano65f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-14(6):969–980, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038632.

Anonymous:1965:AI

[Ano65g] Anonymous. Author index. IEEE Transactions on Electronic Computers, EC-14(2):294, April 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038423.

Anonymous:1965:CP

[Ano65h] Anonymous. Call for papers. IEEE Transactions on Electronic Computers, EC-14(6):767, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038593.

Anonymous:1965:CGNa

[Ano65i] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-14(1): 110–117, February 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038377.

Anonymous:1965:CGNb

[Ano65j] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-14(2): 299–306, April 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038440.

Anonymous:1965:CGNc

[Ano65k] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-14(3):522, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038493.

Anonymous:1965:CGNd

[Ano65l] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-14(4): 672–675, August 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038540.

Anonymous:1965:CGNe

[Ano65m] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-14(5):762, October 1965. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038582.

Anonymous:1965:CGNf

[Ano65n] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-14(6): 1008, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038646.

Anonymous:1965:Ca

[Ano65o] Anonymous. Contributors. IEEE Transactions on Electronic Computers, EC-14(4):648–650, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038525.

Anonymous:1965:Cc

[Ano65p] Anonymous. Contributors. IEEE Transactions on Electronic Computers, EC-14(6):966–968, December 1965. CODEN IEECA8. ISSN 0367-7508. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4038631.

Anonymous:1965:Cb

[Ano65q] Anonymous. Correction. IEEE Transactions on Electronic Computers, EC-14(6):797, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038598.

Anonymous:1965:ICG

[Ano65r] Anonymous. IEEE Computer Group constitution and bylaws preface. IEEE Transactions on Electronic Computers, EC-14(1):1, February 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038341.

Anonymous:1965:ITE

[Ano65s] Anonymous. IEEE Transactions on Electronic Computers annual index 1965 volume EC-14. IEEE Transactions on Electronic Computers, EC-14(6):1017–b–1017–m, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038653.

Anonymous:1965:IASa

[Ano65t] Anonymous. Information about the Society. IEEE Transactions on Electronic Computers, EC-14(1):119–119–b, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038379.

Anonymous:1965:IASb

[Ano65u] Anonymous. Information about the Society. IEEE Transactions on Electronic Computers, EC-14(2):307–307–f, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038441.

Anonymous:1965:IASc

[Ano65v] Anonymous. Information about the Society. IEEE Transactions on Electronic Computers, EC-14 (3):523–531, June 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038494.

Anonymous:1965:IASd

[Ano65w] Anonymous. Information about the Society. IEEE Transactions on Electronic Computers, EC-14(4):677–677–d, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038542.

Anonymous:1965:IASe

[Ano65x] Anonymous. Information about the Society. IEEE Transactions on Electronic Computers, EC-14(5):763–763–b, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038583.

Anonymous:1965:IASf

[Ano65y] Anonymous. Information about the Society. IEEE Transactions on Electronic Computers, EC-14(6):1009–1017, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038647.

Anonymous:1965:IAa

[Ano65z] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-14(1): 119–f, February 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038384.

Anonymous:1965:IAb

[Ano65-27] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-14(2):307–h, April 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038443.

Anonymous:1965:IAc

[Ano65-28] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-14(3):531–d, June 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038498.

Anonymous:1965:IAd

[Ano65-29] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-14 (4):677–h, August 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038546.

Anonymous:1965:IAe

[Ano65-30] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-14 (5):763–f, October 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038588.

Anonymous:1965:IAf

[Ano65-31] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-14(6): 1017–n, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038654.

Anonymous:1965:NAE

[Ano65-32] Anonymous. New Associate Editor for analog and hybrid computers. IEEE Transactions on Electronic Computers, EC-14(2):123, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038390.

Anonymous:1965:NEC

[Ano65-33] Anonymous. New Editor-in-Chief.

IEEE Transactions on Electronic Computers, EC-14(4):533, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038503.

Anonymous:1966:ACCa

[Ano66a] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-15(1):133–148, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038700.

Anonymous:1966:ACCb

[Ano66b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-15(2):275–294, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038749.

Anonymous:1966:ACCc

[Ano66c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-15(3):395–410, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038797.

Anonymous:1966:ACCd

[Ano66d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-15(4):683–700, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038861.

Anonymous:1966:ACCe

[Ano66e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-15(5):817–827, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038897.

Anonymous:1966:ACCf

[Ano66f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-15(6):963–999, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038967.

Anonymous:1966:CGNa

[Ano66g] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-15(1): 149–150, February 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038701.

Anonymous:1966:CGNb

[Ano66h] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-15(2): 295–296, April 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038750.

Anonymous:1966:CGNc

[Ano66i] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-15(3):412, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038799.

Anonymous:1966:CGNd

[Ano66j] Anonymous. Computer group news. IEEE Transactions on Electronic Computers, EC-15(4):682, August 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038860.

Anonymous:1966:Ca

[Ano66k] Anonymous. Contributors. IEEE Transactions on Electronic Computers, EC-15(2):263–264, April 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038730.

Anonymous:1966:Cb

[Ano66l] Anonymous. Contributors. IEEE Transactions on Electronic Computers, EC-15(5):814–816, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038896.

Anonymous:1966:E

[Ano66m] Anonymous. Editorial. IEEE Transactions on Electronic Computers, EC-15(1):1, February 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038659.

Anonymous:1966:EN

[Ano66n] Anonymous. Editor’s notice. IEEE Transactions on Electronic Computers, EC-15(3):303, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038762.

Anonymous:1966:IPC

[Ano66o] Anonymous. IEEE Proceedings computer issue. IEEE Transactions on Electronic Computers, EC-15(4):419, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038812.

Anonymous:1966:ITE

[Ano66p] Anonymous. IEEE Transactions on Electronic Computers annual index 1966 volume EC-15. IEEE Transactions on Electronic Computers, EC-15(6):999–b–999–q, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038969.

Anonymous:1966:ICB

[Ano66q] Anonymous. Implementation of communications-based EDP information systems. IEEE Transactions on Electronic Computers, EC-15(2):271–272, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038746.

Anonymous:1966:IASa

[Ano66r] Anonymous. Information about the Society. IEEE Transactions on Electronic Computers, EC-15(1):151–155, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038702.

Anonymous:1966:IASb

[Ano66s] Anonymous. Information about the Society. IEEE Transactions on Electronic Computers, EC-15 (2):297–299, April 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038751.

Anonymous:1966:IASc

[Ano66t] Anonymous. Information about the Society. IEEE Transactions on Electronic Computers, EC-15 (3):415–417, June 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038802.

Anonymous:1966:IASd

[Ano66u] Anonymous. Information about the Society. IEEE Transactions on Electronic Computers, EC-15(4):701–713, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038862.

Anonymous:1966:IASe

[Ano66v] Anonymous. Information about the Society. IEEE Transactions on Electronic Computers, EC-15(5):843–847, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038916.

Anonymous:1966:IAa

[Ano66w] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-15(1): 155–b, February 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038706.

Anonymous:1966:IAb

[Ano66x] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-15(2):299–d, April 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038756.

Anonymous:1966:IAc

[Ano66y] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-15(3):417–d, June 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038807.

Anonymous:1966:IAd

[Ano66z] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-15 (4):713–b, August 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038870.

Anonymous:1966:IAe

[Ano66-27] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-15 (5):847–d, October 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038922.

Anonymous:1966:IAf

[Ano66-28] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-15(6): 999–r, December 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038971.

Anonymous:1966:N

[Ano66-29] Anonymous. Notices. IEEE Transactions on Electronic Computers, EC-15(3):413, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038800.

Anonymous:1966:ORA

[Ano66-30] Anonymous. On the organization of reliable automata. IEEE Transactions on Electronic Computers, EC-15(3):411, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038798.

Anonymous:1966:RRTb

[Ano66-31] Anonymous. R66-15 A reduction technique for prime implicant tables. IEEE Transactions on Electronic Computers, EC-15(5):841, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038914.

Anonymous:1966:RRTa

[Ano66-32] Anonymous. The role of “Reliability” in total systems design. IEEE Transactions on Electronic Computers, EC-15(2):272–273, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038747.

Anonymous:1966:SAS

[Ano66-33] Anonymous. Sixth annual symposium on switching circuit theory and logical design. IEEE Transactions on Electronic Computers, EC-15(2):274, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038748.

Anonymous:1966:SCI

[Ano66-34] Anonymous. Special computer issue. IEEE Transactions on Electronic Computers, EC-15(3):301, June 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038761.

Anonymous:1966:SRW

[Ano66-35] Anonymous. Summary report of a workshop on programming languages for command and control. IEEE Transactions on Electronic Computers, EC-15(1):

131–132, February 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038699.

Anonymous:1967:IIT

[Ano67a] Anonymous. 1967 index IEEE Transactions on Electronic Computers vol. EC-16. IEEE Transactions on Electronic Computers, EC-16(6):912–a–912–t, December 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039219.

Anonymous:1967:ACCa

[Ano67b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-16(1):115–132, February 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039013.

Anonymous:1967:ACCb

[Ano67c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-16(2):243–260, April 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039058.

Anonymous:1967:ACCc

[Ano67d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-16(3):381–396, June 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039097.

Anonymous:1967:ACCd

[Ano67e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-16(6):881–912, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039218.

Anonymous:1967:C

[Ano67f] Anonymous. Contributors. IEEE Transactions on Electronic Computers, EC-16(5):690–694, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039170.

Anonymous:1967:IAa

[Ano67g] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-16(1): 132–a, February 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039014.

Anonymous:1967:IAb

[Ano67h] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-16(2):260–a, April 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039059.

Anonymous:1967:IAc

[Ano67i] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-16(3):396–a, June 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039098.

Anonymous:1967:IAd

[Ano67j] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-16 (4):547–b, August 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039136.

Anonymous:1967:IAe

[Ano67k] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-16(5): 716–a, October 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039184.

Anonymous:1967:IAf

[Ano67l] Anonymous. Information for authors. IEEE Transactions on Electronic Computers, EC-16(6): 912–u, December 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039221.

Anonymous:1968:ACCa

[Ano68a] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17 (1):99–108, January 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008884.

Anonymous:1968:ACCb

[Ano68b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17 (2):193–204, February 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687318.

Anonymous:1968:ACCc

[Ano68c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17 (3):285–297, March 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687334.

Anonymous:1968:ACCd

[Ano68d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17(4):411–420, April 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687367.

Anonymous:1968:ACCe

[Ano68e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17(5):511–519, May 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687383.

Anonymous:1968:ACCf

[Ano68f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17(6):609–620, June 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687413.

Anonymous:1968:ACCg

[Ano68g] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17(7):705–717, July 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687436.

Anonymous:1968:ACCh

[Ano68h] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17 (8):805–817, August 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687459.

Anonymous:1968:ACCi

[Ano68i] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17(9):903–921, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687485.

Anonymous:1968:ACCj

[Ano68j] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17(10):1005–1024, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687253.

Anonymous:1968:ACCk

[Ano68k] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17 (11):1101–1120, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687277.

Anonymous:1968:ACCl

[Ano68l] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-17 (12):1183–1230, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687294.

Anonymous:1968:ATA

[Ano68m] Anonymous. Announcement transactions available in microfiche form. IEEE Transactions on Computers, C-17(11):1026, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1968:AI

[Ano68n] Anonymous. Author index. IEEE Transactions on Computers, C-17(1):112, January 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008886.

Anonymous:1968:CID

[Ano68o] Anonymous. A capacitor-IGFET digital-to-analog converter. IEEE Transactions on Computers, C-17(10):999–1000, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687249.

Anonymous:1968:Ca

[Ano68p] Anonymous. Contributors. IEEE Transactions on Computers, C-17 (1):95–96, January 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008882.

Anonymous:1968:Cb

[Ano68q] Anonymous. Contributors. IEEE Transactions on Computers, C-17 (2):189–190, February 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1968:Cc

[Ano68r] Anonymous. Contributors. IEEE Transactions on Computers, C-17 (4):403–404, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1968:Cd

[Ano68s] Anonymous. Contributors. IEEE Transactions on Computers, C-17 (5):509–510, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1968:Ce

[Ano68t] Anonymous. Contributors. IEEE Transactions on Computers, C-17 (6):604–605, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1968:Cf

[Ano68u] Anonymous. Contributors. IEEE Transactions on Computers, C-17 (7):703–704, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1968:Cg

[Ano68v] Anonymous. Contributors. IEEE Transactions on Computers, C-17 (8):803–804, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1968:Ch

[Ano68w] Anonymous. Contributors. IEEE Transactions on Computers, C-17 (9):901–902, September 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1968:Ci

[Ano68x] Anonymous. Contributors. IEEE Transactions on Computers, C-17(11):1096–1097, November 1968.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1968:Cj

[Ano68y] Anonymous. Contributors. IEEE Transactions on Computers, C-17(12):1178–1179, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1968:ENa

[Ano68z] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-17(1):1, January 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008861.

Anonymous:1968:ENb

[Ano68-27] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-17(9):821, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687468.

Anonymous:1968:ENc

[Ano68-28] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-17(11):1025, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687258.

Anonymous:1968:IAa

[Ano68-29] Anonymous. Information for authors. IEEE Transactions on Computers, C-17(2):204–a, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687319.

Anonymous:1968:IAb

[Ano68-30] Anonymous. Information for authors. IEEE Transactions on Computers, C-17(3):300–a, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687337.

Anonymous:1968:IAc

[Ano68-31] Anonymous. Information for authors. IEEE Transactions on Computers, C-17(4):420–a, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687368.

Anonymous:1968:IAd

[Ano68-32] Anonymous. Information for authors. IEEE Transactions on Computers, C-17(5):524–b, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687393.

Anonymous:1968:IAe

[Ano68-33] Anonymous. Information for authors. IEEE Transactions on Computers, C-17(6):620–b, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687415.

Anonymous:1968:IAf

[Ano68-34] Anonymous. Information for authors. IEEE Transactions on Computers, C-17(7):720–a, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687441.

Anonymous:1968:IAg

[Ano68-35] Anonymous. Information for authors. IEEE Transactions on Computers, C-17(8):820–a, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687464.

Anonymous:1968:IAh

[Ano68-36] Anonymous. Information for authors. IEEE Transactions on Computers, C-17(9):924–a, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687490.

Anonymous:1968:IAi

[Ano68-37] Anonymous. Information for authors. IEEE Transactions on Computers, C-17(10):1024–a, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687254.

Anonymous:1968:IAj

[Ano68-38] Anonymous. Information for authors. IEEE Transactions on Computers, C-17(11):1120–a, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687278.

Anonymous:1968:IAk

[Ano68-39] Anonymous. Information for authors. IEEE Transactions on Computers, C-17(12):1230–c, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687297.

Anonymous:1968:LPR

[Ano68-40] Anonymous. A light pen for remote time-shared graphic consoles. IEEE Transactions on Computers, C-17(8):799–802, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687456.

Anonymous:1968:RSM

[Ano68-41] Anonymous. Realization of sequential machines. IEEE Transactions on Computers, C-17(12): 1177, December 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687289.

Anonymous:1968:RBP

[Ano68-42] Anonymous. Reviews of books and papers in the computer field. IEEE Transactions on Computers, C-17(1):97–98, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008883.

Anonymous:1968:SCS

[Ano68-43] Anonymous. SAHYB-2: A continuous system simulation language compatible with FORTRAN-IV. IEEE Transactions on Computers, C-17(2):187–188, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687311.

Anonymous:1968:SI

[Ano68-44] Anonymous. Subject index. IEEE Transactions on Computers, C-17(1):108–112, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008885.

Anonymous:1968:TCA

[Ano68-45] Anonymous. Two complete axiom systems for the extended language of regular expressions. IEEE Transactions on Computers, C-17(7):700–701, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687431.

Anonymous:1968:VRC

[Ano68-46] Anonymous. Variable-radix counters based on feedback shiftregisters. IEEE Transactions on Computers, C-17(10):1000– 1002, October 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687250.

Anonymous:1969:ICG

[Ano69a] Anonymous. 1969 IEEE Computer Group Conference. IEEE Transactions on Computers, C-18 (11):986, November 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671159.

Anonymous:1969:IIT

[Ano69b] Anonymous. 1969 index IEEE Transactions on Computers vol. C-18. IEEE Transactions on Computers, C-18(12):1222–a–1222– k, December 1969. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671201.

Anonymous:1969:ACCa

[Ano69c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-18 (1):85–104, January 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671130.

Anonymous:1969:ACCb

[Ano69d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-18 (2):185–203, February 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671218.

Anonymous:1969:ACCc

[Ano69e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-18 (3):285–302, March 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671244.

Anonymous:1969:ACCd

[Ano69f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-18(4):385–399, April 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671266.

Anonymous:1969:ACCe

[Ano69g] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-18(5):479–498, May 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671289.

Anonymous:1969:ACCf

[Ano69h] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-18(6):574–592, June 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671314.

Anonymous:1969:ACCg

[Ano69i] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-18(7):671–688, July 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671338.

Anonymous:1969:ACCh

[Ano69j] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-18 (8):765–783, August 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671363.

Anonymous:1969:ACCi

[Ano69k] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-18(9):867–884, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671383.

Anonymous:1969:ACCk

[Ano69l] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-18 (11):1073–1090, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671177.

Anonymous:1969:ACCl

[Ano69m] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-18 (12):1147–1222, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671200.

Anonymous:1969:ACCj

[Ano69n] Anonymous. Abstracts of the current computer literature. IEEE Transactions on Computers, C-18 (10):966–984, October 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671154.

Anonymous:1969:AVC

[Ano69o] Anonymous. Attenuationless voltage or current propagation on a line with ohmic losses. IEEE Transactions on Computers, C-18 (3):280–281, March 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671239.

Anonymous:1969:CPC

[Ano69p] Anonymous. Comment on “Pattern Classification Design by Linear Programming”. IEEE Transactions on Computers, C-18(4):378–379, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671260.

Anonymous:1969:Ca

[Ano69q] Anonymous. Contributors. IEEE Transactions on Computers, C-18 (1):81–82, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:Cb

[Ano69r] Anonymous. Contributors. IEEE Transactions on Computers, C-18 (2):184, February 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:Cc

[Ano69s] Anonymous. Contributors. IEEE Transactions on Computers, C-18 (4):382–383, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:Cd

[Ano69t] Anonymous. Contributors. IEEE Transactions on Computers, C-18 (5):475–476, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:Ce

[Ano69u] Anonymous. Contributors. IEEE Transactions on Computers, C-18 (6):570–571, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:Cf

[Ano69v] Anonymous. Contributors. IEEE Transactions on Computers, C-18 (7):666–667, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:Cg

[Ano69w] Anonymous. Contributors. IEEE Transactions on Computers, C-18 (8):761–762, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:Ch

[Ano69x] Anonymous. Contributors. IEEE Transactions on Computers, C-18 (10):961–962, October 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:Ci

[Ano69y] Anonymous. Contributors. IEEE Transactions on Computers, C-18(11):1069–1072, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:Cj

[Ano69z] Anonymous. Contributors. IEEE Transactions on Computers, C-18 (12):1145, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:DE

[Ano69-27] Anonymous. Digital engineering. IEEE Transactions on Computers, C-18(12):1144, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671197.

Anonymous:1969:ENa

[Ano69-28] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-18(9):785, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671368.

Anonymous:1969:ENb

[Ano69-29] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-18(12):1091, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671184.

Anonymous:1969:Fa

[Ano69-30] Anonymous. Foreword. IEEE Transactions on Computers, C-18(8):689, August 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671342.

Anonymous:1969:Fb

[Ano69-31] Anonymous. Foreword. IEEE Transactions on Computers, C-18 (11):985, November 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671158.

Anonymous:1969:ITCa

[Ano69-32] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-18(3):c1–304–b, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:ITCb

[Ano69-33] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-18(4):c1–399–c, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:ITCc

[Ano69-34] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-18(5):c1–498–d, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:ITCd

[Ano69-35] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-18(6):c1–592–d, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:ITCe

[Ano69-36] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-18(7):c1–688–b, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:ITCf

[Ano69-37] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-18(8):c1–783–c, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:ITCg

[Ano69-38] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-18(9):c1–884–b, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:ITCh

[Ano69-39] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-18(10):c1–984–b, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:ITCi

[Ano69-40] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-18(12):c1–1222–n, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1969:IAa

[Ano69-41] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(1):104–a, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671131.

Anonymous:1969:IAb

[Ano69-42] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(2):204–a, February 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671221.

Anonymous:1969:IAc

[Ano69-43] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(3):304–a, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671247.

Anonymous:1969:IAd

[Ano69-44] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(4):399–b, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671268.

Anonymous:1969:IAe

[Ano69-45] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(5):498–c, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671292.

Anonymous:1969:IAf

[Ano69-46] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(6):592–c, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671317.

Anonymous:1969:IAg

[Ano69-47] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(7):688–a, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671339.

Anonymous:1969:IAh

[Ano69-48] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(8):783–b, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671365.

Anonymous:1969:IAi

[Ano69-49] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(9):884–a, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671384.

Anonymous:1969:IAj

[Ano69-50] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(10):984–a, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671155.

Anonymous:1969:IAk

[Ano69-51] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(11):1090–c, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671180.

Anonymous:1969:IAl

[Ano69-52] Anonymous. Information for authors. IEEE Transactions on Computers, C-18(12):1222–m, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671203.

Anonymous:1969:AMF

[Ano69-53] Anonymous. On asynchronous machines with flip-flops. IEEE Transactions on Computers, C-18(5):473, May 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671282.

Anonymous:1969:DMC

[Ano69-54] Anonymous. On the determination of the maximum compatibility classes. IEEE Transactions on Computers, C-18(7):665, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671330.

Anonymous:1969:FD

[Ano69-55] Anonymous. T. F. Dwyer.

IEEE Transactions on Computers, C-18(8):760, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671356.

Allen:1966:HDG

[AP66] M. W. Allen and R. J. Potter.

A hardware device for generalized mapping functions. IEEE Transactions on Electronic Computers, EC-15(1):118–120, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038681.

Allen:1963:CEM

[APP^{+}63] M. W. Allen, T. Pearcey, J. P.

Penny, G. A. Rose, and J. G. Sanderson. CIRRUS, an economical multiprogram computer with microprogram control. IEEE Transactions on Electronic Computers, EC-12(6):663–671, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038016.

Arango:1966:TIT

[APVS66] H. Arango, M. Pascual, M. E.

Valentinuzzi, and J. Santos. Threshold implementation of ternary systems. IEEE Transactions on Electronic Computers, EC-15(4): 661–662, August 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038842.

Arant:1957:TST

[Ara57] Gene W. Arant. A time-sequential tabular analysis of flip-flop logical operation. IRE Transactions on Electronic Computers, EC-6(2):72–74, June 1957. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221574.

Arbib:1966:IAS

[Arb66] M. A. Arbib. Information about the Society. IEEE Transactions on Electronic Computers, EC-15(5):843–847, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038901.

Arden:1963:RSS

[Ard63] D. N. Arden. R63-55 states of sequential machines whose logical elements involve delay. IEEE Transactions on Electronic Computers, EC-12(3):335, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037892.

Arden:1965:RBC

[Ard65] B. W. Arden. R65-12 bounded context translation. IEEE Transactions on Electronic Computers, EC-14(1):93, February 1965.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038372.

Armstrong:1962:EAI

[Arm62a] D. B. Armstrong. On the efficient assignment of internal codes to sequential machines. IRE Transactions on Electronic Computers, EC-11(5):611–622, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219422.

Armstrong:1962:PAA

[Arm62b] D. B. Armstrong. A programmed algorithm for assigning internal codes to sequential machines. IRE Transactions on Electronic Computers, EC-11(4): 466–472, August 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219385.

Armstrong:1962:RBS

[Arm62c] D. B. Armstrong. Review: On the State Assignment Problem for Sequential Machines II, by R. E.Stearns and J. Hartmanis. IRE Transactions on Electronic Computers, EC-11(3):419, June 1962. CODEN IRELAO. ISSN 0367-9950.

Armstrong:1965:RIR

[Arm65] D. B. Armstrong. R65-14 interwoven redundant logic. IEEE Transactions on Electronic Computers, EC-14(1):94, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038374.

Armstrong:1966:FNM

[Arm66] D. B. Armstrong. On finding a nearly minimal set of fault detection tests for combinational logic nets. IEEE Transactions on Electronic Computers, EC-15 (1):66–73, February 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038667.

Arthur:1961:GMS

[Art61] M. E. Arthur. Geometric mapping of switching functions. IRE Transactions on Electronic Computers, EC-10(4):631–637, December 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219269.

Arango:1966:FCP

[AS66] H. Arango and J. Santos. A fast carry-propagation circuit for base 3 signed non redundant arithmetic. IEEE Transactions on Electronic Computers, EC-15 (2):254–255, April 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038725.

Arango:1968:TCD

[ASC68] H. Arango, J. Santos, and A. Chacur. Ternary cyclo-decompositions. IEEE Transactions on Computers, C-17 (12):1175–1176, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687288.

Ashley:1965:RBP

[Ash65] J. R. Ashley. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-14(6):1003–1007, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038635.

Ashenhurst:1966:AI

[Ash66] R. L. Ashenhurst. Annual index 1966. IEEE Transactions on Electronic Computers, EC-15(6):1001–1016, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038953.

Ataka:1964:BTT

[Ata64] Hikosaburo Ataka. A basic theorem on threshold devices. IEEE Transactions on Electronic Computers, EC-13(5):631, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038267.

Atkins:1968:HRD

[Atk68] D. E. Atkins. Higher-radix division using estimates of the divisor and partial remainders. IEEE Transactions on Computers, C-17(10):925–934, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687237. See comments

[Whi69b].

Atrubin:1965:ODR

[Atr65] A. J. Atrubin. A one-dimensional real-time iterative multiplier. IEEE Transactions on Electronic Computers, EC-14(3):394–399, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038457.

Aufenkamp:1958:ASM

[Auf58] D. D. Aufenkamp. Analysis of sequential machines II. IRE Transactions on Electronic Computers, EC-7(4):299–306, December 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222663.

Aufenkamp:1962:RBM

[Auf62] D. D. Aufenkamp. Review: On Methods for Realizing a Finite Automaton Whose Cyclical Nature is Determined by the Variation of the Input State, by M. A. Aizerman, L. A. Gusev, L. I. Rozonoer, I. M. Smirnova and A. A. Tal’. IRE Transactions on Electronic Computers, EC-11(3):419, June 1962. CODEN IRELAO. ISSN 0367-9950.

Avizienis:1961:SDN

[Avi61] Algirdas Avižienis. Signed-digit number representations for fast parallel arithmetic. IRE Transactions on Electronic Computers, EC-10(3):389–400, September 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219227.

Avizienis:1963:RBP

[Avi63] A. Avizienis. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-12(1):30–38, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037773.

Astrahan:1956:RAD

[AW56] M. M. Astrahan and L. R. Walters. Reliability of an air defense computing system: Marginal checking and maintenance programming. IRE Transactions on Electronic Computers, EC-5(4): 233–237, December 1956. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219958.

Anacker:1967:PEC

[AW67] Wilhelm Anacker and Chu Ping Wang. Performance evaluation of computing systems with memory hierarchies. IEEE Transactions on Electronic Computers, EC-16(6):764–773, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039193.

Axelrod:1963:RBP

[Axe63] M. S. Axelrod. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-12 (3):330–336, June 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037883.

Ablow:1968:IDT

[AYT68] C. M. Ablow, M. Yoeli, and J. Turner. Irreducible decompositions of transformation graphs by assignment techniques. IEEE Transactions on Computers, C-17(4):325–329, April 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687343.

Azgapetian:1968:CAP

[Azg68] V. Azgapetian. Comment on “An Analog Photoresistive Multiplier”. IEEE Transactions on Computers, C-17(2):188, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687312. See [Mas67].

Baker:1957:IAP

[Bak57] W. R. G. Baker. The IRE “Affiliate” plan — A new venture in engineering society structure and service. IRE Transactions on Electronic Computers, EC-6(2):71, June 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221573.

Burns:1961:CCS

[BAL61] L. L. Burns, G. A. Alphonse, and G. W. Leck. Coincident-current superconductive memory. IRE Transactions on Electronic Computers, EC-10(3):438–446, September 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219232.

Baldwin:1962:CET

[Bal62a] J. A. Baldwin. Circuits employing toroidal magnetic cores as analogs of multipath cores. IRE Transactions on Electronic Computers, EC-11(2):218–223, April 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219355.

Baldwin:1962:FRT

[Bal62b] J. A. Baldwin. Flux reversal in three-rung laddics. IRE Transactions on Electronic Computers, EC-11(5):664–676, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219429.

Banerji:1963:PN

[Ban63] R. B. Banerji. PTGEC news.

IEEE Transactions on Electronic Computers, EC-12(5):600, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037989.

Banerji:1965:IAS

[Ban65] R. B. Banerji. Information about the Society. IEEE Transactions on Electronic Computers, EC-14 (3):523–531, June 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038484.

Bartee:1961:CDM

[Bar61] Thomas C. Bartee. Computer design of multiple-output logical networks. IRE Transactions on Electronic Computers, EC-10 (1):21–30, March 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219148.

Barnes:1963:CEG

[Bar63] George H. Barnes. Comment on “An Electronic Generator of Random Numbers”. IEEE Transactions on Electronic Computers, EC-12(1):22, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037766. See [DS62].

Bareiss:1964:ITE

[Bar64] E. H. Bareiss. IEEE Transactions on Electronic Computers. IEEE Transactions on Electronic Computers, EC-13(6):793–803, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038324.

Barus:1966:EMS

[Bar66] Carl Barus. An easily mechanized scheme for an adaptive pattern recognizer. IEEE Transactions on Electronic Computers, EC-15(3):385–387, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038783.

Bartee:1967:C

[Bar67] T. C. Bartee. Contributors. IEEE Transactions on Electronic Computers, EC-16(2):232–233, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039043.

Barnes:1968:RST

[Bar68] B. B. Barnes. R68-26 some techniques for accuracy improvement in analog computation. IEEE Transactions on Computers, C-17(5):523–524, May 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687391.

Bashkow:1963:C

[Bas63a] T. R. Bashkow. Contributors. IEEE Transactions on Electronic Computers, EC-12(6): 923–930, December 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038046.

Bashkow:1963:RDA

[Bas63b] T. R. Bashkow. R63-106 the D 825 automatic operating and scheduling program. IEEE Transactions on Electronic Computers, EC-12(6):935–936, December 1963.

CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038057.

Bashkow:1964:SCA

[Bas64] Theodore R. Bashkow. A sequential circuit for algebraic statement translation. IEEE Transactions on Electronic Computers, EC-13(2):102–105, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038106.

Batcher:1965:NSS

[Bat65] Kenneth E. Batcher. On the number of stable states in a NOR network. IEEE Transactions on Electronic Computers, EC-14(6):931–932, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038617.

Bauer:1958:ART

[Bau58] Walter F. Bauer. Aspects of real-time simulation. IRE Transactions on Electronic Computers, EC-7(2):134–136, June 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222524.

Bauer:1964:CGN

[Bau64] W. F. Bauer. Computer group news. IEEE Transactions on Electronic Computers, EC-13(5):669, October 1964. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038279.

Bartik:1956:SCC

[BB56a] W. J. Bartik and T. H. Bonn.

A small coincident-current magnetic memory. IRE Transactions on Electronic Computers, EC-5(2):73–78, June 1956. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219802.

Booth:1956:LCT

[BB56b] G. W. Booth and T. P. Bothwell. Logic circuits for a transistor digital computer. IRE Transactions on Electronic Computers, EC-5(3):132–138, September 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219933.

Betts:1961:FTC

[BB61] R. Betts and G. Bishop. Ferrite toroid core circuit analysis. IRE Transactions on Electronic Computers, EC-10(1):51– 56, March 1961. CODEN IRE-LAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219151.

Bledsoe:1962:MMT

[BB62] W. W. Bledsoe and C. L. Bisson.

Memory matrices for the n-tuple pattern recognition method. IRE Transactions on Electronic Computers, EC-11(3):414, June 1962. CODEN IRELAO. ISSN 0367-9950.

Banerji:1969:SDR

[BB69] D. K. Banerji and J. A. Brzozowski. Sign detection in residue number systems. IEEE Transactions on Computers, C-18(4):313–320, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671251;   http://
ieeexplore.ieee.org/xpl/tocresult.
jsp?isnumber=35034.

Brooks:1959:PDB

[BBB59] Frederick P. Brooks, Jr., Gerrit A. Blaauw, and Werner Buchholz. Processing data in bits and pieces. IRE Transactions on Electronic Computers, EC-8 (2):118–124, June 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219512. This paper

contains on page 121 the first published reference to the term “byte”. An article of the same title appears in “Information Processing, Proceedings of the International Conference on Information Processing, UNESCO, Paris, 15–20 June 1959”, pp. 375–381, 1959. From [Buc62, page 40]: “Byte denotes a group of bits used to encode a character, or the number of bits transmitted in parallel to and from input-output units. A term other than character is used here because a given character may be represented in different applications by more than one code, and different codes may use different numbers of bits (i.e., different byte sizes). In input-output transmission the grouping of bits may be completely arbitrary and have no relation to actual characters. (The term is coined from bite, but respelled to avoid accidental mutation to bit.)”.

Barnes:1968:IIC

[BBK^{+}68] G. H. Barnes, R. M. Brown, M. Kato, D. J. Kuck, D. L. Slotnick, and R. A. Stokes. The ILLIAC IV computer. IEEE Transactions on Computers, C-17(8):746–757, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687448.

Bates:1957:ACA

[BBP57] M. R. Bates, D. H. Bock, and F. D. Powell. Analog computer applications in predictor design. IRE Transactions on Electronic Computers, EC-6(3): 143–153, September 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222011.

Bartee:1965:DAG

[BC65] T. C. Bartee and D. J. Chapman.

Design of an accumulator for a general purpose computer. IEEE Transactions on Electronic Computers, EC-14(4):570–574, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038508.

Bargainer:1966:DNC

[BC66] J. D. Bargainer and C. L. Coates. Decimal numbers for checking summability. IEEE Transactions on Electronic Computers, EC-15(3):372, June 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038774.

Bargainer:1968:MMT

[BC68] J. D. Bargainer, Jr. and C. L. Coates. Minimal multiplexed threshold gate realizations. IEEE Transactions on Computers, C-17(6):566–578, June 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687401.

Bennion:1961:BSA

[BCE61] D. R. Bennion, H. D. Crane, and D. C. Engelbart. A bibliographical sketch of all-magnetic logic schemes. IRE Transactions on Electronic Computers, EC-10 (2):203–206, June 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219190.

Brzozowski:1964:LAS

[BD64] J. A. Brzozowski and W. A. Davis. On the linearity of autonomous sequential machines. IEEE Transactions on Electronic Computers, EC-13(6):673–679, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038298.

Bongenaar:1965:WCC

[BdT65] W. Bongenaar and N. C. de Troye.

Worst-case considerations in designing logical circuits. IEEE Transactions on Electronic Computers, EC-14(4):590–599, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038511.

Baer:1969:BMP

[BE69] J.-L. E. Baer and G. Estrin. Bounds for maxium parallelism in a bilogic graph model of computations. IEEE Transactions on Computers, C-18(11):1012–1014, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671165.

Bedrij:1962:CSA

[Bed62] O. J. Bedrij. Carry-select adder. IRE Transactions on Electronic Computers, EC-11(3):340– 346, June 1962. CODEN IRELAO. ISSN 0367-9950.

Bedrosian:1964:RRC

[Bed64] S. D. Bedrosian. R64-19 (1)

the role of checkout-status control of weapon systems. IEEE Transactions on Electronic Computers, EC-13(2):164, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038132.

Beister:1968:IFT

[Bei68] J. Beister. On the implementation of failure-tolerant counters. IEEE Transactions on Computers, C-17(9):885–886, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687476.

Bekey:1959:GIA

[Bek59] George A. Bekey. Generalized integration on the analog computer. IRE Transactions on Electronic Computers, EC-8 (2):210–217, June 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219524.

Bekey:1965:RMS

[Bek65] G. A. Bekey. R65-11 methods for solving engineering problems using analog computers. IEEE Transactions on Electronic Computers, EC-14(1):92–93, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038371.

Bekey:1967:RAC

[Bek67] G. A. Bekey. R67-30 analog computer methods for parameter optimization. IEEE Transactions on Electronic Computers, EC-16(2):242, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039056.

Bennion:1960:NMS

[Ben60] D. R. Bennion. A note on magnetic shift registers. IRE Transactions on Electronic Computers, EC-9(2):262, June 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219828.

Ben:1969:RHA

[Ben69a] A. Ben. R69-21 hybrid assumed mode solution of non-linear partial differential equations. IEEE Transactions on Computers, C-18(7):670, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671337.

Benham:1969:RTT

[Ben69b] R. D. Benham. R69-1 transport time-delay simulation for transmission line representation. IEEE Transactions on Computers, C-18(1):83, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671128.

Berlin:1958:SVS

[Ber58] R. D. Berlin. Synthesis of Nvalued switching circuits. IRE Transactions on Electronic Computers, EC-7(1):52–56, March 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222096.

Bergman:1960:TDL

[Ber60a] R. H. Bergman. Tunnel diode logic circuits. IRE Transactions on Electronic Computers, EC-9(4): 430–438, December 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219881.

Berwin:1960:FPA

[Ber60b] Ted W. Berwin. Frequencyto-period-to-analog computer for flowrate measurement. IRE Transactions on Electronic Computers, EC-9(1):62–71, March 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221606.

Bernstein:1966:APP

[Ber66a] A. J. Bernstein. Analysis of programs for parallel processing. IEEE Transactions on Electronic Computers, EC-15(5):757–763, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038883.

Bernstein:1966:CGN

[Ber66b] A. J. Bernstein. Computer group news. IEEE Transactions on Electronic Computers, EC-15(4):682, August 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038850.

Beyer:1963:UWC

[Bey63] W. A. Beyer. Uniqueness of weighted code representations, II.

IEEE Transactions on Electronic Computers, EC-12(2):137, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037816.

Beyer:1964:UWC

[Bey64] W. A. Beyer. Uniqueness of weighted code representations, III. IEEE Transactions on Electronic Computers, EC-13(2): 153–154, April 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038114.

Bianchini:1961:IVA

[BF61] R. Bianchini and C. Freiman.

On internal variable assignments for sequential switching circuits. IRE Transactions on Electronic Computers, EC-10(1):95–96, March 1961. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219161.

Bashkow:1963:PSD

[BFK63] T. R. Bashkow, J. Friets, and A. Karson. A programming system for detection and diagnosis of machine malfunctions. IEEE Transactions on Electronic Computers, EC-12(1):10–17, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037760.

Ball:1966:C

[BFL^{+}66a] N. A. Ball, H. Q. Foster, W. H.

Long, I. E. Sutherland, and R. L.

Wigington. Correction. IEEE Transactions on Electronic Computers, EC-15(5):749, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038881.

Ball:1966:SMC

[BFL^{+}66b] N. A. Ball, H. Q. Foster, W. H. Long, I. E. Sutherland, and R. L. Wigington. A shared memory computer display system. IEEE Transactions on Electronic Computers, EC-15(5):750–756, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038882.

Bay:1956:HSF

[BG56a] Z. Bay and N. T. Grisamore. Highspeed flip-flops for the millimicrosecond region. IRE Transactions on Electronic Computers, EC-5(3):121–125, September 1956. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219931.

Bay:1956:PGH

[BG56b] Z. Bay and N. T. Grisamore. Pulse generator and high-speed memory circuit. IRE Transactions on Electronic Computers, EC-5(4): 213–218, December 1956. CODEN IRELAO. ISSN 0367-9950.

Baker:1966:PPC

[BGR66] R. H. Baker, D. H. Galvin, and R. W. Rasche. Pulse powered circuits. IEEE Transactions on Electronic Computers, EC-15 (3):321–323, June 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038766.

Backus:1964:F

[BH64] J. W. Backus and W. P. Heising. Fortran. IEEE Transactions on Electronic Computers, EC-13(4):382–385, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038201.

Bernstein:1965:ACS

[BH65a] A. Bernstein and J. Hopcroft.

Annual cumulative subject index. IEEE Transactions on Electronic Computers, EC-14(6):981– 1000, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038633.

Bingulac:1965:ACG

[BH65b] S. P. Bingulac and E. A. Humo.

Analog computer generation of Bessel functions of arbitrary order. IEEE Transactions on Electronic Computers, EC-14(6): 886–889, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038609.

Bergland:1967:DRT

[BH67] G. D. Bergland and H. W.

Hale. Digital real-time spectral analysis. IEEE Transactions on Electronic Computers, EC-16 (2):180–185, April 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039026.

Bryant:1962:CFS

[BHK62] P. R. Bryant, F. G. Heath, and R. D. Killick. Counting with feedback shift registers by means of a jump technique. IRE Transactions on Electronic Computers, EC-11(2):285–286, April 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219364.

Brooks:1957:EMC

[BHNW57] F. P. Brooks, A. L. Hopkins, P. G. Neumann, and W. V. Wright. An experiment in musical composition. IRE Transactions on Electronic Computers, EC-6(3):175–182, September 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222016.

Barrett:1961:CCP

[BHRS61] W. A. Barrett, F. B. Humphrey, J. A. Ruff, and H. L. Stadler. A card-changeable permanentmagnet-twistor memory of large capacity. IRE Transactions on Electronic Computers, EC-10(3): 451–461, September 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219234.

Bickart:1960:CEF

[Bic60] Theodore A. Bickart. Computation of an expression for the frequency spectrum of a twosided signal. IRE Transactions on Electronic Computers, EC-9 (2):263–264, June 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219831.

Biega:1968:RAS

[Bie68] B. C. Biega. R68-7 analog simulation of ferroresonant system including analysis of hysteresis loop. IEEE Transactions on Computers, C-17(2):192, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687317.

Bingulac:1968:RAA

[Bin68] S. P. Bingulac. R68-38 accurate analog computer generation of Bessel functions for large ranges. IEEE Transactions on Computers, C-17(8):819, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687462.

Birta:1967:CNN

[Bir67] L. G. Birta. Comment on “Nonobservable or Noncontrollable Analog Schemes for the Solution of a Special Class of Differential Equations”. IEEE Transactions on Electronic Computers, EC-16(4):516–517, August 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039122. See [Mar66b].

Bittmann:1959:TFM

[Bit59] Eric E. Bittmann. Thin-film memories. IRE Transactions on Electronic Computers, EC-8(2):92–97, June 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219508.

Bittmann:1962:RBD

[Bit62] E. E. Bittmann. Review: Proposal for Demonstration of Domain Wall Storage and Logic, by D. O. Smith. IRE Transactions on Electronic Computers, EC-11(3): 423, June 1962. CODEN IRELAO. ISSN 0367-9950.

Bittner:1966:RCP

[Bit66] B. P. Bittner. R66-87 computer programming. IEEE Transactions on Electronic Computers, EC-15(6):961, December 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038965.

Brown:1964:CC

[BJST64] R. M. Brown, R. D. Jenks, J. E. Stifle, and R. L. Trogdon. The CSX-1 computer. IEEE Transactions on Electronic Computers, EC-13(3):247–250, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038149.

Bryant:1962:CNF

[BK62a] P. R. Bryant and R. D. Killick.

Correction to “Nonlinear Feedback Shift Registers”. IRE Transactions on Electronic Computers, EC-11(5):714, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219442.

Bryant:1962:NFS

[BK62b] P. R. Bryant and R. D. Killick. Nonlinear feedback shift registers. IRE Transactions on Electronic Computers, EC-11(3):410– 411, June 1962. CODEN IRELAO. ISSN 0367-9950.

Bernstein:1966:IAS

[BL66] A. J. Bernstein and P. M. Lewis. Information about the Society. IEEE Transactions on Electronic Computers, EC-15(4):701–713, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038854.

Blachman:1956:WTD

[Bla56a] N. M. Blachman. On the wiring of two-dimensional multiplecoincidence magnetic memories. IRE Transactions on Electronic Computers, EC-5(1):19–21, March 1956. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219787.

Blachman:1956:RIA

[Bla56b] N. M. Blachman. A report on the International Analogy Computation Meeting. IRE Transactions on Electronic Computers, EC-5(1):36–41, March 1956. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219791.

Blachman:1956:CE

[Bla56c] Nelson M. Blachman. Change of editorship. IRE Transactions on Electronic Computers, EC-5(4):183, December 1956. CODEN IRELAO. ISSN 0367-9950.

Blachman:1956:SAD

[Bla56d] Nelson M. Blachman. Some automatic digital computers in Western Europe. IRE Transactions on Electronic Computers, EC-5(3): 158–166, September 1956. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219939.

Blankenbaker:1958:LMP

[Bla58] John V. Blankenbaker. Logically micro-programmed computers. IRE Transactions on Electronic Computers, EC-7(2): 103–109, June 1958. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222518.

Blache:1962:ATM

[Bla62] Yves J. M. Blache. Automatic time-mark generation for analog computer phase-plane trajectories. IRE Transactions on Electronic Computers, EC-11(4): 571–572, August 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219400.

Bledsoe:1961:BLS

[Ble61a] W. W. Bledsoe. A basic limitation on the speed of digital computers. IRE Transactions on Electronic Computers, EC-10 (3):530, September 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219243.

Bledsoe:1961:FRT

[Ble61b] W. W. Bledsoe. Further results on the N-tuple pattern recognition method. IRE Transactions on Electronic Computers, EC-10(1):96, March 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219162.

Bledsoe:1964:RAS

[Ble64] W. W. Bledsoe. R64-61 the automatic speech recognition system for conversational sound. IEEE Transactions on Electronic Computers, EC-13(5):648, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038283.

Bledsoe:1966:RPR

[Ble66] W. W. Bledsoe. R66-64 pattern recognition studies in the biomedical sciences. IEEE Transactions on Electronic Computers, EC-15(5):839, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038909.

Brzozowski:1963:SFG

[BM63] J. A. Brzozowski and E. J. Mc-Cluskey. Signal flow graph techniques for sequential circuit state diagrams. IEEE Transactions on Electronic Computers, EC-12(2):67–76, April 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037802.

Brzozowski:1964:ASF

[BM64] J. A. Brzozowski and E. J. Mc-Cluskey. About signal flow graph techniques for sequential circuits. IEEE Transactions on Electronic Computers, EC-13(2):154, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038115.

Burner:1969:PDC

[BMRS69] H. B. Burner, R. P. Million, O. W. Rechard, and J. S. Sobolewski. A programmable data concentrator for a large computing system. IEEE Transactions on Computers, C-18(11):1030–1038, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671169.

Bush:1959:PTA

[BO59] L. Bush and P. Orlando. A perturbation technique for analog computers. IRE Transactions on Electronic Computers, EC-8(2):218–221, June 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219525.

Bobrow:1965:RTC

[Bob65] D. G. Bobrow. R65-25 training a computer to assign descriptors to documents: Experiments in automatic indexing. IEEE Transactions on Electronic Computers, EC-14(2):278, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038433.

Bohn:1960:PPM

[Boh60] E. V. Bohn. A pulse position modulation analog computer. IRE Transactions on Electronic Computers, EC-9(2):256–261, June 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219826.

Bolton:1967:RSL

[Bol67] G. R. Bolton. R67-40 satellite lifetime program. IEEE Transactions on Electronic Computers, EC-16(3):379–380, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039093.

Bonner:1966:PRT

[Bon66] R. E. Bonner. Pattern recognition with three added requirements. IEEE Transactions on Electronic Computers, EC-15(5): 770–781, October 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038885.

Booth:1962:VFM

[Boo62] Theodore M. Booth. The vertex-frame method for obtaining minimal proposition-letter formulas. IRE Transactions on Electronic Computers, EC-11(2): 144–154, April 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219347.

Booth:1964:RCF

[Boo64] T. L. Booth. R64-49 computation with finite fields. IEEE Transactions on Electronic Computers, EC-13(4):484–485, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038229.

Booth:1967:C

[Boo67] T. L. Booth. Contributors. IEEE Transactions on Electronic Computers, EC-16(1):106–107, February 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038997.

Booth:1968:SPR

[Boo68] T. L. Booth. Statistical properties of random digital sequences. IEEE Transactions on Computers, C-17(5):452–461, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687374.

Born:1969:TTS

[Bor69] R. C. Born. Transformation of ternary switching functions to completely symmetric ternary switching functions. IEEE Transactions on Computers, C-18(6):559–560, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671304.

Bouchet:1968:AMM

[Bou68] A. Bouchet. An algebraic method for minimizing the number of states in an incomplete sequential machine. IEEE Transactions on Computers, C-17(8):795–798, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687455.

Bowdon:1966:N

[Bow66] E. K. Bowdon. Notices. IEEE Transactions on Electronic Computers, EC-15(2):295–296, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038739.

Bowdon:1967:RPA

[Bow67] E. K. Bowdon. R67-9 PERT as an aid to logic design. IEEE Transactions on Electronic Computers, EC-16(1):110, February 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039005.

Bowdon:1969:PAN

[Bow69] E. K. Bowdon, Sr. Priority assignment in a network of computers. IEEE Transactions on Computers, C-18(11):1021–1026, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671167.

Bloch:1959:MCL

[BP59] E. Bloch and R. C. Paulsen. Magnetic core logic in a high-speed card-to-tape converter. IRE Transactions on Electronic Computers, EC-8(2):169–181, June 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219518.

Brzozowski:1963:CSM

[BP63] J. A. Brzozowski and J. F. Poage. On the construction of sequential machines from regular expressions. IEEE Transactions on Electronic Computers, EC-12 (4):402–403, August 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037916.

Brick:1964:MWR

[BP64] Donald B. Brick and George G.

Pick. Microsecond word-recognition system. IEEE Transactions on Electronic Computers, EC-13(1): 57–59, February 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038083.

Bolgiano:1969:RPT

[BP69] L. P. Bolgiano, Jr. and M. J. Piovoso. Relationship of Poisson transform to Laguerre expansions. IEEE Transactions on Computers, C-18(12):1128–1131, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671189.

Boulton:1969:PCL

[BR69] P. I. P. Boulton and P. A. Reid.

A process-control language. IEEE Transactions on Computers, C-18 (11):1049–1053, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671172.

Braun:1958:DCC

[Bra58] Edward L. Braun. Digital computers in continuous control systems. IRE Transactions on Electronic Computers, EC-7 (2):123–128, June 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222522.

Bray:1959:EOS

[Bra59] T. E. Bray. An electro-optical shift register. IRE Transactions on Electronic Computers, EC-8 (2):113–117, June 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219511.

Bradshaw:1963:LRU

[Bra63] John A. Bradshaw. Letter recognition using a captive scan. IEEE Transactions on Electronic Computers, EC-12(1):26, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037770.

Braverman:1964:AI

[Bra64] D. Braverman. Author index. IEEE Transactions on Electronic Computers, EC-13(3):340, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038174.

Brady:1966:CHS

[Bra66] Allen H. Brady. The conjectured highest scoring machines for Rado’s Σ(k) for the value k = 4. IEEE Transactions on Electronic Computers, EC-15(5):802–803, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038890.

Brcic:1965:IFD

[Brc65] I. Brcic. “Ideally Fast” decimal counters with bistables. IEEE Transactions on Electronic Computers, EC-14(5):733–737, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038565.

Brearley:1965:IIS

[Bre65a] H. C. Brearley. ILLIAC II — A short description and annotated bibliography. IEEE Transactions on Electronic Computers, EC-14(3):399–403, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038458.

Breuer:1965:ITN

[Bre65b] Melvin A. Breuer. Implementation of threshold nets by integer linear programming. IEEE Transactions on Electronic Computers, EC-14(6):950–952, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038624.

Bremerman:1967:RCT

[Bre67] H. Bremerman. R67-16 computers and thought. IEEE Transactions on Electronic Computers, EC-16(1):113–114, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039012.

Brigham:1961:SPB

[Bri61] Robert C. Brigham. Some properties of binary counters with feedback. IRE Transactions on Electronic Computers, EC-10(4): 699–701, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219275.

Bridgman:1963:CAC

[Bri63] Arthur Bridgman. Comments on “Analog Computer Simulation of a Frequency Detector”. IEEE Transactions on Electronic Computers, EC-12(6):911, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038039.

Bridgman:1966:CES

[Bri66] Arthur Bridgman. Comments on “Experimental Study of a New Method of Time Delay for Analog Computers”. IEEE Transactions on Electronic Computers, EC-15(4):667–669, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038846. See [Wie65a].

Bristol:1967:CTE

[Bri67] Edgar H. Bristol. Compensation for truncation errors in accumulative addition using random methods. IEEE Transactions on Electronic Computers, EC-16(4):518, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039125.

Brown:1955:REC

[Bro55a] David R. Brown. Review of electronic computer progress during 1954. IRE Transactions on Electronic Computers, EC-4 (1):33–38, March 1955. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407954.

Brown:1955:SNL

[Bro55b] R. M. Brown. Some notes on logical binary counters. IRE Transactions on Electronic Computers, EC-4(2):67–69, June 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407894.

Broadbent:1960:TMF

[Bro60a] Kent D. Broadbent. A thin magnetic film shift register. IRE Transactions on Electronic Computers, EC-9(3):321–323, September 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219852.

Broughton:1960:FMO

[Bro60b] M. B. Broughton. A feedback method for obtaining a synchro output signal proportional to input angle θ for large θ. IRE Transactions on Electronic Computers, EC-9(3):359–362, September 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219859.

Brown:1960:EDC

[Bro60c] David T. Brown. Error detecting and correcting binary codes for arithmetic operations. IRE Transactions on Electronic Computers, EC-9(3):333–337, September 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219855.

Brown:1963:NET

[Bro63a] Frank M. Brown. A nodeelimination theorem for Boolean matrices. IEEE Transactions on Electronic Computers, EC-12 (2):133–134, April 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037811.

Brown:1963:RED

[Bro63b] R. G. Brown. R63-88 equitable distribution. IEEE Transactions on Electronic Computers, EC-12(5):581, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037994.

Brown:1964:LCR

[Bro64] Richard M. Brown. On-line computer recognition of handprinted characters. IEEE Transactions on Electronic Computers, EC-13(6):750–752, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038313.

Brown:1965:CTS

[Bro65a] Frank M. Brown. Code transformation in sequential machines. IEEE Transactions on Electronic Computers, EC-14(6):822–829, December 1965. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038601.

Brown:1965:ESL

[Bro65b] Richard M. Brown. An experimental study of an on-line mancomputer system. IEEE Transactions on Electronic Computers, EC-14(1):82–85, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038360.

Brown:1966:FSO

[Bro66] J. Reese Brown. First-and secondorder ferrite memory core characteristics and their relationship to system performance. IEEE Transactions on Electronic Computers, EC-15(4):485–501, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038822.

Brown:1968:OMI

[Bro68] F. M. Brown. The origin of the method of iterated consensus. IEEE Transactions on Computers, C-17(8):802, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687457.

Brown:1969:RFE

[Bro69a] B. Brown. R69-26 further experimental data on the behavior of programs in a paging environment. IEEE Transactions on Computers, C-18(9):865, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671380.

Brown:1969:CCP

[Bro69b] F. M. Brown. Comment on “Canonical Programming of Nonlinear and Time-Varying Differential Equations”. IEEE Transactions on Computers, C-18(6):566, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671306.

Bruns:1957:IES

[Bru57] Robert A. Bruns. An integral-error-squared method for evaluating analog computer components. IRE Transactions on Electronic Computers, EC-6(1):35–36, March 1957. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221561.

Brzozowski:1962:SRE

[Brz62] Janusz A. Brzozowski. A survey of regular expressions and their applications. IRE Transactions on Electronic Computers, EC-11(3):324– 335, June 1962. CODEN IRELAO. ISSN 0367-9950.

Brzozowski:1963:ACC

[Brz63] J. A. Brzozowski. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(6):937–961, December 1963. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038048.

Brzozowski:1964:PN

[Brz64a] J. A. Brzozowski. PTGEC news. IEEE Transactions on Electronic Computers, EC-13(2): 180–181, April 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038126.

Brzozowski:1964:RES

[Brz64b] J. A. Brzozowski. Regular expressions from sequential circuits. IEEE Transactions on Electronic Computers, EC-13(6): 741–744, December 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038308.

Brzozowski:1965:REL

[Brz65a] J. A. Brzozowski. Regular expressions for linear sequential circuits. IEEE Transactions on Electronic Computers, EC-14(2): 148–156, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038393.

Brzozowski:1965:SPR

[Brz65b] J. A. Brzozowski. Some problems in relay circuit design. IEEE Transactions on Electronic Computers, EC-14(4):630–634, August 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038518.

Brzozowski:1966:WR

[Brz66] J. A. Brzozowski. Workshop reports. IEEE Transactions on Electronic Computers, EC-15(1): 131–132, February 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038689.

Brzozowski:1967:AI

[Brz67] J. A. Brzozowski. Author index. IEEE Transactions on Electronic Computers, EC-16(1):132, February 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039001.

Blattner:1959:FML

[BS59] D. J. Blattner and F. Sterzer. Fast microwave logic circuits. IRE Transactions on Electronic Computers, EC-8(3):297–301, September 1959. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5222686.

Barkouki:1963:TEE

[BS63] M. F. Barkouki and I. Stein. Theoretical and experimental evaluation of RZ and NRZ recording characteristics. IEEE Transactions on Electronic Computers, EC-12(2):92–100, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037804.

Beddoes:1964:SCT

[BS64] M. P. Beddoes and C. A. T.

Salama. Series connected tunneldiode multilevel register. IEEE Transactions on Electronic Computers, EC-13(3):308–309, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038166.

Brousil:1967:TLN

[BS67] James K. Brousil and David R.

Smith. A threshold logic network for shape invariance. IEEE Transactions on Electronic Computers, EC-16(6):818–828, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039198.

Born:1968:TSF

[BS68a] R. C. Born and A. K. Scidmore. Transformation of switching functions to completely symmetric switching functions. IEEE Transactions on Computers, C-17(6):596–599, June 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687404.

Brzozowski:1968:DAS

[BS68b] Janusz A. Brzozowski and Shanker Singh. Definite asynchronous sequential circuits. IEEE Transactions on Computers, C-17(1): 18–26, January 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008864.

Bashkow:1967:SDF

[BSK67] Theodore R. Bashkow, Azra Sasson, and Arnold Kronfeld. System design of a FORTRAN machine. IEEE Transactions on Electronic Computers, EC-16(4):485–499, August 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039113.

Brown:1961:IEC

[BTW61] W. G. Brown, J. Tierney, and R. Wasserman. Improvement of electronic-computer reliability through the use of redundancy. IRE Transactions on Electronic Computers, EC-10(3): 407–416, September 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219229.

Buchholz:1962:PCS

[Buc62] Werner Buchholz, editor. Planning a Computer System: Project Stretch. McGraw-Hill, New York, NY, USA, 1962. xvii + 322 pp. LCCN 1876. URL http:/

                                        /ed-thelen.org/comp-hist/IBM-

7030-Planning-McJones.pdf. This

important book is the primary description of the influential IBM 7030 Stretch computer, written by its architects. See also a detailed critical review [Str62].

Budak:1966:ITD

[Bud66] Aram Budak. An improved time delay simulation. IEEE Transactions on Electronic Computers, EC-15(1):122, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038684.

Buelow:1960:ICS

[Bue60] F. K. Buelow. Improvements to current switching. IRE Transactions on Electronic Computers, EC-9(4):415–418, December 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219878.

Burns:1963:ACC

[Bur63] L. L. Burns. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(5):585–596, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037986.

Butcher:1964:PSU

[But64] I. R. Butcher. A prewired storage unit. IEEE Transactions on Electronic Computers, EC-13 (2):106–111, April 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038107.

Burke:1965:C

[BV65] R. E. Burke and J. G. Van Bosse. Correction. IEEE Transactions on Electronic Computers, EC-14(1):61, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038350.

Burke:1965:NC

[BvB65] R. E. Burke and J. G. van Bosse. NAND–AND circuits. IEEE Transactions on Electronic Computers, EC-14(1):63–65, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038352.

Becker:1961:TLC

[BW61] C. L. Becker and J. V. Wait. Two-level correlation on an analog computer. IRE Transactions on Electronic Computers, EC-10(4):752–758, December 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219284. See correction

[BW62].

Becker:1962:CTL

[BW62] C. L. Becker and J. V. Wait. Correction to “Two-Level Correlation of an Analog Computer”. IRE Transactions on Electronic Computers, EC-11(4):578–579, August 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219408. See [BW61].

Brayton:1963:AEC

[BW63] R. Brayton and R. Willoughby.

An analysis of the effect of component tolerances on the amplification of the balanced-pair tunnel-diode circuit. IEEE Transactions on Electronic Computers, EC-12(3):269–274, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037862.

Bobrow:1964:LPE

[BW64] Daniel G. Bobrow and Joseph Weizenbaum. List processing and extension of language facility by embedding. IEEE Transactions on Electronic Computers, EC-13(4):395–400, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038203.

Byrne:1965:N

[Byr65] E. R. Byrne. Notices. IEEE Transactions on Electronic Computers, EC-14(3):522, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038483.

Cadzow:1968:SND

[Cad68] J. A. Cadzow. Synthesis of nonlinear decision boundaries by cascaded threshold gates. IEEE Transactions on Computers, C-17 (12):1165–1172, December 1968.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687286.

Cahn:1953:AAC

[Cah53] Lee Cahn. Accuracy of an analog computer. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(4): 12–18, December 1953. CO-DEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407689.

Campeau:1957:SAD

[Cam57] Joseph O. Campeau. The synthesis and analysis of digital systems by Boolean matrices. IRE Transactions on Electronic Computers, EC-6(4):231–241, December 1957. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222033.

Cameron:1964:GMT

[Cam64] S. H. Cameron. The generation of minimal threshold nets by an integer program. IEEE Transactions on Electronic Computers, EC-13(3):299–302, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038159.

Campeau:1969:BOC

[Cam69] J. O. Campeau. The block-oriented computer. IEEE Transactions on Computers, C-18(8):706–718, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671346.

Canaday:1966:RCG

[Can66] R. H. Canaday. R66-88 computer graphics-ten unsolved problems. IEEE Transactions on Electronic Computers, EC-15(6): 961–962, December 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038966.

Canaday:1967:RLW

[Can67a] R. H. Canaday. R67-26 the Lincoln Wand. IEEE Transactions on Electronic Computers, EC-16(2):240, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039052.

Canaday:1967:RBP

[Can67b] R. H. Canaday. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-16(6):877–880, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039211.

Carter:1960:NCS

[Car60] I. P. V. Carter. A new core switch for magnetic matrix stores and other purposes. IRE Transactions on Electronic Computers, EC-9(2):176–191, June 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219817.

Carr:1964:RMT

[Car64] W. N. Carr. R64-81 300-Mc tunnel-diode logic circuits. IEEE Transactions on Electronic Computers, EC-13(6):762, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038332.

Carling:1965:RBP

[Car65] L. N. Carling. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-14(5):743–746, October 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038570.

Carroll:1968:RSM

[Car68] C. C. Carroll. R68-40 sequential machines and automata theory. IEEE Transactions on Computers, C-17(9):922–923, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687486.

Catt:1966:TLT

[Cat66] Ivor Catt. Time loss through gating of asynchronous logic signal pulses. IEEE Transactions on Electronic Computers, EC-15 (1):108–111, February 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038676.

Catt:1967:CND

[Cat67] Ivor Catt. Crosstalk (noise) in digital systems. IEEE Transactions on Electronic Computers, EC-16(6):743–763, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039192.

Choudhury:1962:MCS

[CB62] A. K. Choudhury and M. S. Basu. A mechanized chart for simplification of switching functions. IRE Transactions on Electronic Computers, EC-11(5):713–714, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219441.

Chien:1963:CSD

[CB63] R. T. Chien and R. J. Barbetta. A class of simple decoding networks for Marcus matrix switches. IEEE Transactions on Electronic Computers, EC-12(5):551–553, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037973.

Choudhury:1966:CDT

[CD66a] A. K. Choudhury and S. R. Das.

Comment on “Detection of Totally Symmetric Boolean Functions”. IEEE Transactions on Electronic Computers, EC-15(5):813, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038895. See [She65d].

Choudhury:1966:CIN

[CD66b] A. K. Choudhury and S. R. Das. Computing irredundant normal forms from abbreviated presence functions. IEEE Transactions on Electronic Computers, EC-15(3):387, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038784.

Cohn:1965:IML

[CE65] M. Cohn and S. Even. Identification and minimization of linear machines. IEEE Transactions on Electronic Computers, EC-14(3):367–376, June 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038454.

Cohn:1969:DSR

[CE69a] M. Cohn and S. Even. The design of shift register generators for finite sequences. IEEE Transactions on Computers, C-18(7):660–662, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671328.

Cohn:1969:GCC

[CE69b] M. Cohn and S. Even. A Gray code counter. IEEE Transactions on Computers, C-18(7):662–664, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671329.

Cantor:1962:LEF

[CET62] D. Cantor, G. Estrin, and R. Turn.

Logarithmic and exponential function evaluation in a variable structure digital computer. IRE Transactions on Electronic Computers, EC-11(2):155–164, April 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219348.

Cardillo:1968:SSP

[CF68] G. P. Cardillo and King-Sun Fu.

On suboptimal sequential pattern recognition. IEEE Transactions on Computers, C-17(8):789–792, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687453.

Cox:1964:QE

[CG64] J. R. Cox and D. H. Glaeser. A quantising encoder. IEEE Transactions on Electronic Computers, EC-13(3):250–254, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038150.

Crane:1965:BPD

[CG65] B. A. Crane and J. A. Githens.

Bulk processing in distributed logic memory. IEEE Transactions on Electronic Computers, EC-14(2):186–196, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038398.

Chao:1959:GRT

[Cha59a] S. C. Chao. A generalized resistor-transistor logic circuit and some applications. IRE Transactions on Electronic Computers, EC-8(1):8–12, March 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222755.

Chao:1959:GCT

[Cha59b] Stanley K. Chao. A glow counting tube read-out technique and its application. IRE Transactions on Electronic Computers, EC-8(3):317–320, September 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222690.

Chao:1960:GRC

[Cha60] S. C. Chao. Generalized RTL circuits — supplementary. IRE Transactions on Electronic Computers, EC-9(3):371–372, September 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219862.

Chao:1964:DDT

[Cha64] R. Y. Chao. A digital decoding technique. IEEE Transactions on Electronic Computers, EC-13(6):745–746, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038310.

Chaitin:1965:ITF

[Cha65a] Gregory J. Chaitin. An improvement on a theorem of E. F. Moore. IEEE Transactions on Electronic Computers, EC-14 (3):466–467, June 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038467.

Chang:1965:ASO

[Cha65b] Herbert Y. Chang. An algorithm for selecting an optimum set of diagnostic tests. IEEE Transactions on Electronic Computers, EC-14(5):706–711, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038556.

Chang:1966:RRT

[Cha66] D. M. Y. Chang. The role of “Reliability” in total systems design. IEEE Transactions on Electronic Computers, EC-15 (2):272–273, April 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038733.

Chan:1967:RBP

[Cha67] Shu-Park Chan. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-16 (1):107–114, February 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038998.

Chander:1968:DMT

[Cha68] Suresh Chander. On Dolotta– McCluskey technique. IEEE Transactions on Computers, C-17(1):95, January 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008881.

Chan:1969:SSD

[Cha69] Shu-Kwan Chan. The serial solution of the diffusion equation using nonstandard hybrid techniques. IEEE Transactions on Computers, C-18(9):786–799, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671369.

Chen:1958:MCP

[Che58] Mao-Chao Chen. A magnetic core parallel adder. IRE Transactions on Electronic Computers, EC-7(4):262–264, December 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222656.

Cheney:1961:ADC

[Che61a] Philip W. Cheney. Analog-to-digital conversion with threshold detectors. IRE Transactions on Electronic Computers, EC-10 (1):100–101, March 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219167.

Cheney:1961:DCB

[Che61b] Philip Warren Cheney. A digital correlator based on the residue number system. IRE Transactions on Electronic Computers, EC-10(1):63–70, March 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219154.

Chen:1968:CPR

[Che68a] Chi-Hau Chen. Computer processing of radiation sensor PCM data. IEEE Transactions on Computers, C-17(10):987–989, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687245.

Chen:1968:CAM

[Che68b] K. A. Chen. Computer aided memory design using transmission line models. IEEE Transactions on Computers, C-17(7):640–648, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687420.

Chen:1968:UDL

[Che68c] Tsai Hwa Chen. The use of delay lines in reading a Manchester code. IEEE Transactions on Computers, C-17(9):827–845, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687470.

Chien:1959:OME

[Chi59] Robert T. Chien. Orthogonal matrices, error-correcting codes and load-sharing matrix switches. IRE Transactions on Electronic Computers, EC-8(3): 400, September 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222701.

Chirlian:1963:FLR

[Chi63] P. M. Chirlian. The fabrication of low resistances at liquid helium temperatures. IEEE Transactions on Electronic Computers, EC-12(2):148, April 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037820.

Chow:1957:OCR

[Cho57] C. K. Chow. An optimum character recognition system using decision functions. IRE Transactions on Electronic Computers, EC-6(4):247–254, December 1957. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222035.

Chow:1959:COC

[Cho59] C. K. Chow. Comments on optimum character recognition systems. IRE Transactions on Electronic Computers, EC-8(2):230, June 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219530.

Chow:1960:TDD

[Cho60] W. F. Chow. Tunnel diode digital circuitry. IRE Transactions on Electronic Computers, EC-9(3):295–301, September 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219848.

Chow:1962:RMU

[Cho62] C. K. Chow. A recognition method using neighbor dependence. IRE Transactions on Electronic Computers, EC-11(5): 683–690, October 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219431.

Cho:1963:MTA

[Cho63a] Yohan Cho. A method of theoretical analysis of high-speed junction diode logic circuits. IEEE Transactions on Electronic Computers, EC-12(5):492–502, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037963.

Chow:1963:ERC

[Cho63b] C. K. Chow. An experimental result on character recognition. IEEE Transactions on Electronic Computers, EC-12(1):25, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037769.

Chow:1963:N

[Cho63c] C. K. Chow. Notices. IEEE Transactions on Electronic Computers, EC-12(6):963, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038050.

Cho:1964:AI

[Cho64] Yohan Cho. Author index. IEEE Transactions on Electronic Computers, EC-13(4):501, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038223.

Choudhury:1965:ACA

[Cho65a] A. K. Choudhury. Annual cumulative author index. IEEE Transactions on Electronic Computers, EC-14(6):1000–1002, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038634.

Chow:1965:SIT

[Cho65b] C. K. Chow. Statistical independence and threshold functions. IEEE Transactions on Electronic Computers, EC-14(1): 66–68, February 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038354.

Choudhury:1966:ACC

[Cho66] A. K. Choudhury. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-15(6):963–975, December 1966. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038950.

Chow:1967:PWC

[Cho67] Woo F. Chow. Plated wire contentaddressable memories with bitsteering technique. IEEE Transactions on Electronic Computers, EC-16(5):642–652, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039155.

Christopherson:1961:MSD

[Chr61] Warren A. Christopherson. Matrix switch and drive system for a low-cost magnetic-core memory. IRE Transactions on Electronic Computers, EC-10(2): 238–246, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219194.

Chu:1961:GTQ

[Chu61] J. T. Chu. A generalization of a theorem of Quine for simplifying truth functions. IRE Transactions on Electronic Computers, EC-10(2):165–168, June 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219185.

Chu:1965:DRA

[Chu65a] Yaohan Chu. A destructive-readout associative memory. IEEE Transactions on Electronic Computers, EC-14(4):600–605, August 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038512.

Chu:1965:DEP

[Chu65b] Yaohan Chu. Direct execution of programs in floating code by address interpretation. IEEE Transactions on Electronic Computers, EC-14(3):417–422, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038460.

Chu:1966:CSW

[Chu66a] W. W. Chu. Computer simulation of waveform distortions in digital magnetic recordings. IEEE Transactions on Electronic Computers, EC-15(3):328–336, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038769.

Chu:1966:C

[Chu66b] W. W. Chu. Correction. IEEE Transactions on Electronic Computers, EC-15(3):327, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038768.

Chu:1967:CSE

[Chu67a] Wesley W. Chu. A computer simulation of electrical loss and loading effect in magnetic recording. IEEE Transactions on Electronic Computers, EC-16(4):430–434, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039106.

Chu:1967:MMD

[Chu67b] Wesley W. Chu. A mathematical model for diagnosing system failures. IEEE Transactions on Electronic Computers, EC-16 (3):327–331, June 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039071.

Chu:1969:OFA

[Chu69a] W. W. Chu. Optimal file allocation in a multiple computer system. IEEE Transactions on Computers, C-18(10):885–889, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671135.

Chuang:1969:TLC

[Chu69b] Y. H. Chuang. Transition logic circuits and a synthesis method. IEEE Transactions on Computers, C-18(2):154–168, February 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671211.

Chan:1965:C

[CK65a] S. K. Chan and E. F. Kurtz.

Correction. IEEE Transactions on Electronic Computers, EC-14(2):247, April 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038407.

Chan:1965:UAC

[CK65b] S. K. Chan and E. F. Kurtz. On using an analog computer to study hydrodynamic stability. IEEE Transactions on Electronic Computers, EC-14(2):233–238, April 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038404.

Coates:1962:SPR

[CKL62] C. L. Coates, R. B. Kirchner, and P. M. Lewis, II. A simplified procedure for the realization of linearly-separable switching functions. IRE Transactions on Electronic Computers, EC-11 (4):447–458, August 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219383.

Caldwell:1960:PAD

[CKLP60] W. F. Caldwell, G. A. Korn, V. R. Latorre, and G. R. Peterson. A precision amplitudedistribution amplifier. IRE Transactions on Electronic Computers, EC-9(2):252–255, June 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219825.

Cohn:1961:AMD

[CL61] M. Cohn and R. Lindaman. Axiomatic majority-decision logic. IRE Transactions on Electronic Computers, EC-10(1):17–21, March 1961. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219147.

Coates:1963:RPT

[CL63] C. L. Coates and P. M. Lewis. A realization procedure for threshold gate networks. IEEE Transactions on Electronic Computers, EC-12(5):454–461, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037957.

Coates:1964:DTG

[CL64] C. L. Coates and P. M. Lewis.

DONUT: A threshold gate computer. IEEE Transactions on Electronic Computers, EC-13(3): 240–247, June 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038148.

Clark:1956:FAM

[Cla56] John R. Clark. Fourier analysis by machine methods. IRE Transactions on Electronic Computers, EC-5(3):141, September 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219937.

Clague:1967:NCS

[Cla67] Dennis John Clague. New classes of synchronous codes. IEEE Transactions on Electronic Computers, EC-16(3):290–298, June 1967.

CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039067.

Carleton:1964:FEF

[CLS64] J. T. Carleton, P. E. Lego, and R. M. Suarez. A FOR-TRAN extension to facilitate proposal preparation. IEEE Transactions on Electronic Computers, EC-13(4):456–462, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038211.

Clymer:1959:OAS

[Cly59] A. Ben Clymer. Operational analog simulation of the vibration of a beam and a rectangular multicellular structure. IRE Transactions on Electronic Computers, EC-8(3):381–391, September 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222699.

Clymer:1965:RSS

[Cly65] A. B. Clymer. R65-82 the SIOUX system and hybrid block diagrams. IEEE Transactions on Electronic Computers, EC-14(6): 1006–1007, December 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038644.

Chang:1959:MFS

[CM59] H. Chang and A. G. Milnes. Magnetic fields of square-loop thin films of oblate spheroidal geometry. IRE Transactions on Electronic Computers, EC-8(4): 458–464, December 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222058.

Chang:1960:MFT

[CM60] H. Chang and A. G. Milnes.

Magnetic fields of twistors represented by confocal hollow prolate spheroids. IRE Transactions on Electronic Computers, EC-9(2):199–207, June 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219819.

Connolly:1961:ACC

[CM61] T. W. Connolly and K. S. Miller.

Analog computation of covariance matrices. IRE Transactions on Electronic Computers, EC-10 (3):533, September 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219248.

Carr:1962:BCT

[CM62a] W. N. Carr and A. G. Milnes.

Bias-controlled tunnel-pair logic circuits. IRE Transactions on Electronic Computers, EC-11(6): 773–779, December 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219462.

Chirlian:1962:OAU

[CM62b] P. M. Chirlian and V. A. Marsocci. Operational amplifiers using controlled superconductors. IRE Transactions on Electronic Computers, EC-11(1):6–9, February 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219309.

See note [CM63].

Chirlian:1963:NRO

[CM63] P. M. Chirlian and V. A. Marsocci. Note on a review of “Operational Amplifiers Using Controlled Superconductors”. IEEE Transactions on Electronic Computers, EC-12(2):137, April 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037815. See [CM62b].

Chang:1965:CIN

[CM65] D. M. Y. Chang and T. H. Mott. Computing irredundant normal forms from abbreviated presence functions. IEEE Transactions on Electronic Computers, EC-14(3):335–342, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038450.

Coffman:1968:MUM

[CM68] E. G. Coffman and A. C. McKellar. On the motion of an unbounded, Markov queue in random access storage. IEEE Transactions on Computers, C-17(6):600–603, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687406.

Casey:1966:RPC

[CN66] R. Casey and G. Nagy. Recognition of printed Chinese characters. IEEE Transactions on Electronic Computers, EC-15(1):91–101, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038670.

Casey:1968:ARM

[CN68] R. G. Casey and G. Nagy. An autonomous reading machine. IEEE Transactions on Computers, C-17(5):492–503, May 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687379.

Coates:1963:RDL

[Coa63] C. L. Coates. R63-56 disjunctively linear logic nets. IEEE Transactions on Electronic Computers, EC-12(3):335–336, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037893.

Coates:1964:ACC

[Coa64] C. L. Coates. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(6):763–791, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038321.

Coffman:1968:SPM

[Cof68] E. G. Coffman. A simple probability model yielding performance bounds for modular memory systems. IEEE Transactions on Computers, C-17(1):86–89, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008878.

Coffman:1969:TBR

[Cof69a] E. G. Coffman, Jr. On the tradeoff between response and preemption costs in a foregroundbackground computer service discipline. IEEE Transactions on Computers, C-18(10):942–947, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671143.

Coffman:1969:RDP

[Cof69b] E. G. Coffman, Jr. R69-27 demand paging in perspective. IEEE Transactions on Computers, C-18(9):865–866, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671381.

Cohn:1962:ICF

[Coh62] Martin Cohn. Inconsistent canonical forms of switching functions. IRE Transactions on Electronic Computers, EC-11(2):

284–285, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219363.

Cohler:1963:SI

[Coh63] E. U. Cohler. Subject index. IEEE Transactions on Electronic Computers, EC-12(4):437–438, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037930.

Cohn:1964:TLA

[Coh64] Martin Cohn. A theorem on linear automata. IEEE Transactions on Electronic Computers, EC-13(1):52–53, February 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038077.

Coleman:1961:OFL

[Col61] Robert P. Coleman. Orthogonal functions for the logical design of switching circuits. IRE Transactions on Electronic Computers, EC-10(3):379–383, September 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219225.

Colley:1967:CTP

[Col67] L. E. Colley. Computation of timephase displacements of binary linear sequence generators. IEEE Transactions on Electronic Computers, EC-16(3):357–359, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039077. See comment

[Dou68].

Cole:1969:RTC

[Col69a] S. N. Cole. Real-time computation by n-dimensional iterative arrays of finite-state machines. IEEE Transactions on Computers, C-18(4):349–365, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671256.

Coles:1969:RAQ

[Col69b] L. S. Coles. R69-30 automatic question-answering of English-like questions about simple diagrams. IEEE Transactions on Computers, C-18(10):963, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671150.

Connelly:1962:RTA

[Con62] Mark E. Connelly. Real-time analog-digital computation. IRE Transactions on Electronic Computers, EC-11(1):31–41, February 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219313.

See comments [VS62a, VS62b].

Connolly:1966:CCH

[Con66] John B. Connolly. Cross coupling in high speed digital systems. IEEE Transactions on Electronic Computers, EC-15(3):323–327, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038767.

Connelly:1967:RCT

[Con67a] M. E. Connelly. R67-57 circle test evaluation of a method of compensating hybrid computing error by predicted integral. IEEE Transactions on Electronic Computers, EC-16(5):713–714, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039177.

Connolly:1967:RAD

[Con67b] T. W. Connolly. R67-62 analog and digital computation of Fourier series and integrals. IEEE Transactions on Electronic Computers, EC-16(5):715–716, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039182.

Connelly:1968:AR

[Con68] M. E. Connelly. Author’s reply.

IEEE Transactions on Computers, C-17(2):188, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687313.

Cooper:1963:MCB

[Coo63a] Paul W. Cooper. The multiplecategory Bayes decision procedure. IEEE Transactions on Electronic Computers, EC-12(1):18, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037761.

Cooper:1963:MEU

[Coo63b] Paul W. Cooper. Multivariate extension of univariate distributions. IEEE Transactions on Electronic Computers, EC-12(5): 572–573, October 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037982.

Cooper:1964:APR

[Coo64a] David B. Cooper. Adaptive pattern recognition and signal detection using stochastic approximation. IEEE Transactions on Electronic Computers, EC-13(3):306–307, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038164.

Cooperman:1964:MTD

[Coo64b] Michael Cooperman. 300-Mc tunnel-diode logic circuits. IEEE Transactions on Electronic Computers, EC-13(1):18–26, February 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038072.

Cooper:1966:NAH

[Coo66] Paul W. Cooper. A note on an adaptive hypersphere decision boundary. IEEE Transactions on Electronic Computers, EC-15(6):948–949, December 1966.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038947.

Copes:1968:ROG

[Cop68] E. J. Copes. R68-35 optimal generation of arbitrary functions. IEEE Transactions on Computers, C-17(7):719–720, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687440.

Coraluppi:1964:SDD

[Cor64] Giorgio Coraluppi. On signal deterioration in digital computers. IEEE Transactions on Electronic Computers, EC-13(3):303, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038161.

Costanza:1963:ICL

[Cos63] James L. Costanza. The identification of combinational logic switching structure. IEEE Transactions on Electronic Computers, EC-12(2):132–133, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037810.

Cosgriff:1969:MLC

[Cos69] R. L. Cosgriff. Multiplex logic circuits. IEEE Transactions on Computers, C-18(6):552–556, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671302.

Couleur:1958:BBD

[Cou58] John F. Couleur. BIDEC — A binary-to-decimal or decimalto-binary converter. IRE Transactions on Electronic Computers, EC-7(4):313–316, December 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222665.

Cover:1965:GSP

[Cov65] Thomas M. Cover. Geometrical and statistical properties of systems of linear inequalities with applications in pattern recognition. IEEE Transactions on Electronic Computers, EC-14(3):326–334, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038449.

Cowell:1963:AI

[Cow63] W. R. Cowell. Author index. IEEE Transactions on Electronic Computers, EC-12(2):175, April 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037825.

Cowgill:1964:LEB

[Cow64] D. Cowgill. Logic equations for a built-in square root method. IEEE Transactions on Electronic Computers, EC-13(2):156–157, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038119.

Cox:1961:SAD

[Cox61] Jerome R. Cox. A squaring analogdigital converter. IRE Transactions on Electronic Computers, EC-10(1):98–100, March 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219166.

Chaplin:1966:TDB

[CP66] Mildred Chaplin and W. Penney.

Two-dimensional binary arrays. IEEE Transactions on Electronic Computers, EC-15(1):121, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038683.

Colussi:1968:AOT

[CP68] S. P. Colussi and G. V. Pallottino. An approach to optimum tolerance adaptive threshold elements. IEEE Transactions on Computers, C-17(7):684–691, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687425.

Calabi:1962:IMN

[CR62] L. Calabi and J. A. Riley. Inessentiality in minimal networks and formulas. IRE Transactions on Electronic Computers, EC-11(5): 711–713, October 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219440.

Crane:1960:N

[Cra60a] H. D. Crane. The neuristor. IRE Transactions on Electronic Computers, EC-9(3):370– 371, September 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219861.

Crane:1960:SDU

[Cra60b] H. D. Crane. Sequence detection using all-magnetic circuits. IRE Transactions on Electronic Computers, EC-9(2): 155–160, June 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219814.

Crane:1961:DAMb

[Cra61] H. D. Crane. Design of an allmagnetic computing system: Part II — logical design. IRE Transactions on Electronic Computers, EC-10(2):221–232, June 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219192.

Cramer:1964:RSP

[Cra64] M. L. Cramer. R64-37 some principles and circuit techniques for controlling machine tools from a central digital computer. IEEE Transactions on Electronic Computers, EC-13(3):322, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038186.

Crane:1968:PFA

[Cra68] B. A. Crane. Path finding with associative memory. IEEE Transactions on Computers, C-17(7):691–693, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687426.

Croy:1960:IAD

[Cro60] John E. Croy. Improved arrangement of a decimal multiplier. IRE Transactions on Electronic Computers, EC-9(2):263, June 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219830.

Croy:1961:RTM

[Cro61] John E. Croy. Rapid technique of manual or machine binary-to-decimal integer conversion using decimal radix arithmetic. IRE Transactions on Electronic Computers, EC-10(4): 777, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219286.

Crowther:1964:RDL

[Cro64] T. S. Crowther. R64-31 A dynamic large signal model for a singledomain thin magnetic film inductor. IEEE Transactions on Electronic Computers, EC-13(3):319, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038180.

Crocker:1968:RSR

[Cro68] C. F. Crocker. R68-17 synthesis of resistive digital-to-analog conversion ladders for arbitrary codes with fixed positive weights. IEEE Transactions on Computers, C-17(4):409–410, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687365.

Castanias:1958:RCP

[CS58] R. P. Castanias and J. E. Sherman. Review of computer progress in 1957. IRE Transactions on Electronic Computers, EC-7(1):65–72, March 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222100.

Chernak:1964:SEC

[CS64] J. Chernak and C. F. Simone. Simple equivalent circuits for a step recovery diode and a switching transistor. IEEE Transactions on Electronic Computers, EC-13(5):623–625, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038263.

Crofut:1966:DTD

[CS66] W. A. Crofut and M. R. Sottile. Design techniques of a delay-line content-addressed memory. IEEE Transactions on Electronic Computers, EC-15(4):529–534, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038826.

Connolly:1967:FEC

[CS67] J. B. Connolly and W. G. Schmidt.

Failure-erasure circuitry: A duplicative technique of failuremasking systems. IEEE Transactions on Electronic Computers, EC-16(1):82–85, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038986.

Coffman:1968:RWM

[CS68] E. G. Coffman, Jr. and M. S.

Schmookler. A random-walk model of a queue storage problem. IEEE Transactions on Computers, C-17(11):1093–1095, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687270.

Cunningham:1954:TDN

[Cun54] W. J. Cunningham. Time-delay networks for an analog computer. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-3(4):16–18, December 1954. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Curl:1968:RMC

[Cur68a] F. G. Curl. R68-14 microprogrammed control in problemoriented languages. IEEE Transactions on Computers, C-17(4): 407–408, April 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687362.

Curtis:1968:PSC

[Cur68b] H. A. Curtis. Polylinear sequential circuit realizations of finite automata. IEEE Transactions on Computers, C-17(3): 251–259, March 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687326.

Curtis:1969:SPR

[Cur69] H. A. Curtis. Systematic procedures for realizing synchronous sequential machines using flipflop memory: Part I. IEEE Transactions on Computers, C-18 (12):1121–1127, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671188.

Cutlip:1968:CDP

[Cut68] W. F. Cutlip. On the cascade decomposition of prefix automata. IEEE Transactions on Computers, C-17(1):94–95, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008880.

Crane:1961:DAMa

[CV61] H. D. Crane and E. K. Van De Riet. Design of an all-magnetic computing system: Part I — circuit design. IRE Transactions on Electronic Computers, EC-10 (2):207–220, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219191.

Combet:1965:CBT

[CVV65] M. Combet, H. Van Zonneveld, and L. Verbeek. Computation of the base two logarithm of binary numbers. IEEE Transactions on Electronic Computers, EC-14(6):863–867, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038605.

Char:1968:RDG

[CW68] B. F. Char and C. S. Wells. R68-42 on designing generalized file records for management information systems. IEEE Transactions on Computers, C-17(9):924, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687488.

Cohen:1969:TGB

[CW69] S. Cohen and R. O. Winder.

Threshold gate building blocks. IEEE Transactions on Computers, C-18(9):816–823, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671371.

Dabadghao:1969:MFF

[Dab69] S. V. Dabadghao. A method for finding feedback partitions for sequential machines. IEEE Transactions on Computers, C-18(5):465–467, May 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671280.

Daggett:1959:DBC

[Dag59] D. H. Daggett. Decimal–binary conversions in CORDIC. IRE Transactions on Electronic Computers, EC-8(3):335–339, September 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222694.

Daggett:1965:CGN

[Dag65] N. L. Daggett. Computer group news. IEEE Transactions on Electronic Computers, EC-14(6): 1008, December 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038636.

Daley:1965:N

[Dal65] R. C. Daley. Notices. IEEE Transactions on Electronic Computers, EC-14(6):1008, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038637.

Danielsson:1966:BM

[Dan66] P. E. Danielsson. Boolean memories. IEEE Transactions on Electronic Computers, EC-15(1): 29–35, February 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038664.

Darms:1968:RHS

[Dar68] D. Darms. R68-29 hydro system optimization model. IEEE Transactions on Computers, C-17(6):607, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687410.

Das:1969:MDM

[Das69] S. K. Das. A method of decision making in pattern recognition. IEEE Transactions on Computers, C-18(4):329–333, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671253.

Dauer:1966:IP

[Dau66] Frederick W. Dauer. Impact printing. IEEE Transactions on Electronic Computers, EC-15(5): 794–798, October 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038887.

Davidson:1957:SCC

[Dav57a] Charles H. Davidson. A survey of the characteristics of currently used bistable multivibrators. IRE Transactions on Electronic Computers, EC-6(2):121, June 1957.

CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221584.

Davies:1957:SFT

[Dav57b] D. W. Davies. Switching functions of three variables. IRE Transactions on Electronic Computers, EC-6(4):265–275, December 1957. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222038.

David:1962:RBV

[Dav62] E. E. David, Jr. Review: Digitalto-Voice Conversion, by E. L. Ragland. IRE Transactions on Electronic Computers, EC-11(3): 424, June 1962. CODEN IRELAO. ISSN 0367-9950.

David:1963:MPT

[Dav63a] E. E. David. Membership in the Professional Technical Group on Electronic Computers. IEEE Transactions on Electronic Computers, EC-12(3):354, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037889.

Davies:1963:CFA

[Dav63b] D. W. Davies. An 11-cryotron full adder. IEEE Transactions on Electronic Computers, EC-12(3):319, June 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037873.

Davies:1965:LSP

[Dav65] D. W. Davies. Longest “Separated”

paths and loops in an N cube. IEEE Transactions on Electronic Computers, EC-14(2):261, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038415.

Davidow:1966:IAS

[Dav66] W. H. Davidow. Information about the Society. IEEE Transactions on Electronic Computers, EC-15(1):151–155, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038695.

Davis:1967:AAI

[Dav67] Wayne A. Davis. An approach to the assignment of input codes. IEEE Transactions on Electronic Computers, EC-16 (4):435–442, August 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039107.

Davis:1968:SSR

[Dav68] W. A. Davis. Single shift-register realizations for sequential machines. IEEE Transactions on Computers, C-17(5):421–431, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687371.

Davidson:1969:AND

[Dav69a] E. S. Davidson. An algorithm for NAND decomposition under network constraints. IEEE Transactions on Computers, C-18(12): 1098–1109, December 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671186.

Davis:1969:CVC

[Dav69b] F. C. Davis. Comment on “A Variable Counter Design Technique”. IEEE Transactions on Computers, C-18(3):281, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671240. See [Jon69a,

Tae69, Ren68].

Davis:1969:IIP

[Dav69c] R. L. Davis. The ILLIAC IV processing element. IEEE Transactions on Computers, C-18(9): 800–816, September 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671370.

Davis:1969:LDU

[Dav69d] W. A. Davis. Logical design using shift registers. IEEE Transactions on Computers, C-18(10): 958–960, October 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671147.

Davis:1966:LSM

[DB66] W. A. Davis and J. A. Brzozowski. On the linearity of sequential machines. IEEE Transactions on Electronic Computers, EC-15(1):21–29, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038663.

DeSarkar:1969:SIS

[DBC69] S. C. De Sarkar, A. K. Basu, and A. K. Choudhury. Simplification of incompletely specified flow tables with the help of prime closed sets. IEEE Transactions on Computers, C-18(10):953–956, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671145.

Das:1965:MTE

[DC65] S. R. Das and A. K. Choudhury. Maxterm type expressions of switching functions and their prime implicants. IEEE Transactions on Electronic Computers, EC-14(6):920–923, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038613.

Duley:1968:DSD

[DD68] J. R. Duley and D. L. Dietmeyer. A digital system design language (DDL). IEEE Transactions on Computers, C-17(9):850–861, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687472.

Duley:1969:TDD

[DD69] J. R. Duley and D. L. Dietmeyer. Translation of a DDL digital system specification to Boolean equations. IEEE Transactions on Computers, C-18(4):305–313, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671250.

deRennaeSouza:1968:NEN

[de 68a] Celso de Renna e Souza. A note on embedding nonlinear machines. IEEE Transactions on Computers, C-17(9):894–896, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687480.

deRennaeSouza:1968:RML

[de 68b] Celso de Renna e Souza. R 68-44 mathematical logic. IEEE Transactions on Computers, C-17 (10):1003, October 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687251.

deRennaeSouza:1969:TSR

[de 69] Celso de Renna e Souza. A theorem on the state reduction of synthesized stochastic machines. IEEE Transactions on Computers, C-18(5):473–474, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671283.

Deiters:1965:ODD

[Dei65] Robert M. Deiters. Optimum design of a diode squarer by applying the criterion of square root of the integral of per cent error squared. IEEE Transactions on Electronic Computers, EC-14(3):456–463, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038465.

Dunn:1955:DCU

[DEL55] W. H. Dunn, C. Eldert, and P. V. Levonian. A digital computer for use in an operational flight trainer. IRE Transactions on Electronic Computers, EC-4(2):55–63, June 1955. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407892.

Deland:1962:SBS

[Del62] E. C. Deland. Simulation of a biological system on an analog computer. IRE Transactions on Electronic Computers, EC-11 (1):17–25, February 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219311.

Dennis:1963:AN

[Den63] J. B. Dennis. AFIPS news. IEEE Transactions on Electronic Computers, EC-12(5):600–601, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037990.

Dennis:1965:CGN

[Den65] J. B. Dennis. Computer group news. IEEE Transactions on Electronic Computers, EC-14(3):522, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038482.

Dennis:1967:RIR

[Den67] J. B. Dennis. R67-11 index register allocation. IEEE Transactions on Electronic Computers, EC-16(1):111, February 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039007.

Denham:1969:RAF

[Den69a] W. F. Denham. R69-33 applications of functional optimization techniques for the serial hybrid computer solution of partial differential equations. IEEE Transactions on Computers, C-18(10):965, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671153.

Denning:1969:ECB

[Den69b] P. J. Denning. Equipment configuration in balanced computer systems. IEEE Transactions on Computers, C-18(11): 1008–1012, November 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671164.

Denning:1969:RDS

[Den69c] P. J. Denning. R69-28 dynamic storage allocation systems. IEEE Transactions on Computers, C-18(9):866, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671382.

Deo:1966:SDC

[Deo66] Narsingh Deo. The self-diagnosability of a computer. IEEE Transactions on Electronic Computers, EC-15(5):799, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038888.

Deo:1968:GPR

[Deo68] N. Deo. Generalized parallel redundancy in digital computers. IEEE Transactions on Computers, C-17(6):600, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&

                                        arnumber=1687405. See comment

[Gir69].

Deo:1969:AR

[Deo69] N. Deo. Author’s reply. IEEE Transactions on Computers, C-18(1):80, January 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671125.

Dertouzos:1964:AST

[Der64] Michael L. Dertouzos. An approach to single-threshold-element synthesis. IEEE Transactions on Electronic Computers, EC-13(5): 519–528, October 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038245.

Dertouzos:1965:CGN

[Der65] M. L. Dertouzos. Computer group news. IEEE Transactions on Electronic Computers, EC-14(4):672–675, August 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038530.

Dertouzos:1966:AI

[Der66] M. L. Dertouzos. Author index. IEEE Transactions on Electronic Computers, EC-15(2):294, April 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038737.

Dertouzos:1967:PLG

[Der67] Michael L. Dertouzos. PHASE-PLOT: An on-line graphical display technique. IEEE Transactions on Electronic Computers, EC-16(2):203–209, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039029.

Duda:1966:PCI

[DF66] R. O. Duda and H. Fossum. Pattern classification by iteratively determined linear and piecewise linear discriminant functions. IEEE Transactions on Electronic Computers, EC-15(2):220–232, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038716. See comments

[Kno67a].

Dertouzos:1967:MCT

[DF67] Michael L. Dertouzos and Zachary Fluhr. Minimization and convexity in threshold logic. IEEE Transactions on Electronic Computers, EC-16(2):212–215, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039031.

David:1968:HSF

[DF68a] C. A. David and B. Feldman.

High-speed fixed memories using large-scale integrated resistor matrices. IEEE Transactions on Computers, C-17(8):721–728, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687445.

Doty:1968:TL

[DF68b] K. Doty and H. Frank. A theorem on linearity. IEEE Transactions on Computers, C-17(3): 270–272, March 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687329.

Duffy:1961:FGU

[DG61] R. M. Duffy and C. P. Gilbert. A function generator using coldcathode selector tubes. IRE Transactions on Electronic Computers, EC-10(1):71–77, March 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219155.

Douce:1969:EAD

[DH69] J. L. Douce and T. J. Healy. Evaluation of the amplitude distribution of quasi-Gaussian signals obtained from pseudorandom noise. IEEE Transactions on Computers, C-18(8):749–752, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671352.

Diamantides:1956:MES

[Dia56] Nick D. Diamantides. A multipurpose electronic switch for analog computer simulation and autocorrelation applications. IRE Transactions on Electronic Computers, EC-5(4):197–202, December 1956. CO-DEN IRELAO. ISSN 0367-9950.

Dietmeyer:1963:CPN

[Die63] Donald L. Dietmeyer. Conversion from positive to negative and imaginary radix. IEEE Transactions on Electronic Computers, EC-12(1):20–22, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037765.

Dietmeyer:1968:BPO

[Die68] D. L. Dietmeyer. Bounds on the period of oscillatory activity in randomly interconnected networks of neuron-like elements. IEEE Transactions on Computers, C-17(6):578–591, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687402.

Daly:1963:HSA

[DK63] W. G. Daly and J. F. Kruy.

A high-speed arithmetic unit using tunnel diodes. IEEE Transactions on Electronic Computers, EC-12(5):503–511, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037964.

Dertouzos:1969:LSB

[DKP69] M. L. Dertouzos, M. E. Kaliski, and K. P. Polzen. On-line simulation of block-diagram systems. IEEE Transactions on Computers, C-18(4):333–342, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671254.

deRennaeSouza:1969:RAD

[dL69] Celso de Renna e Souza and R. Jeffrey Leake. Relationships among distinct models and notions of equivalence for stochastic finitestate systems. IEEE Transactions on Computers, C-18(7):633–641, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671323.

Dolotta:1964:CIS

[DM64] T. A. Dolotta and E. J. McCluskey. The coding of internal states of sequential circuits. IEEE Transactions on Electronic Computers, EC-13(5):549–562, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038248.

dHoop:1966:SPS

[dM66] H. d’Hoop and R. Monterosso. SAHYB — A program for simulation of analog and hybrid computers. IEEE Transactions on Electronic Computers, EC-15 (3):381–382, June 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038778.

Davidson:1968:CAS

[DM68] E. S. Davidson and G. Metze. Comments on “An Algorithm for Synthesis of Multiple-Output Combinational Logic”. IEEE Transactions on Computers, C-17(11): 1091–1092, November 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687269. See [SD68,

Has69].

Davidson:1969:AR

[DM69] E. S. Davidson and G. Metze.

Authors’ reply. IEEE Transactions on Computers, C-18(9): 863–864, September 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671379.

Dolotta:1964:C

[Dol64] T. A. Dolotta. Contributors. IEEE Transactions on Electronic Computers, EC-13(3):313–315, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038170.

Domenico:1957:STS

[Dom57] R. J. Domenico. Simulation of transistor switching circuits on the IBM 704. IRE Transactions on Electronic Computers, EC-6(4):242–247, December 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222034.

Donegan:1965:IAS

[Don65a] J. J. Donegan. Information about the Society. IEEE Transactions on Electronic Computers, EC-14(6):1009–1017, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038638.

Donegan:1965:RST

[Don65b] J. J. Donegan. R65-15 the 4102-S space track program. IEEE Transactions on Electronic Computers, EC-14(1):94, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038375.

Donegan:1967:RDT

[Don67] J. J. Donegan. R67-12 digital training devices. IEEE Transactions on Electronic Computers, EC-16(1):111, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039008.

Dor:1967:GLB

[Dor67] Neville M. Dor. Guide to the length of buffer storage required for random (Poisson) input and constant output rates. IEEE Transactions on Electronic Computers, EC-16(5):683–684, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039165.

Dorrough:1969:MAA

[Dor69] D. C. Dorrough. A methodical approach to analyzing and synthesizing a self-repairing computer. IEEE Transactions on Computers, C-18(1):22–42, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671115.

Douce:1968:CCT

[Dou68] J. L. Douce. Comment on “Computation of Time-Phase Displacements of Binary Linear Sequence Generators”. IEEE Transactions on Computers, C-17(4):402, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687356. See [Col67].

Dow:1957:ACE

[Dow57] Paul C. Dow. An analysis of certain errors in electronic differential analyzers. I. Bandwidth limitations. IRE Transactions on Electronic Computers, EC-6(4): 255–260, December 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222036.

Dow:1958:ACE

[Dow58] Paul C. Dow. An analysis of certain errors in electronic differential analyzers II — capacitor dielectric absorption. IRE Transactions on Electronic Computers, EC-7(1):17–22, March 1958.

CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222090.

Dunworth:1969:ECB

[DR69] A. Dunworth and J. I. Roche. The error characteristics of the binary rate multiplier. IEEE Transactions on Computers, C-18 (8):741–745, August 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671350.

Dreyfus:1967:RIO

[Dre67] S. Dreyfus. R67-10 introduction and overview of the Multics system. IEEE Transactions on Electronic Computers, EC-16(1):110–111, February 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039006.

Drucker:1969:COR

[Dru69] H. Drucker. Computer optimization of recognition networks. IEEE Transactions on Computers, C-18(10):918–923, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671140.

Deering:1961:ICT

[DS61] C. S. Deering and C. B. Shelman. An incremental computer technique for solving coordinaterotation equations. IRE Transactions on Electronic Computers, EC-10(4):748–751, December 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219283.

Dillard:1962:EGR

[DS62] G. M. Dillard and R. E. Simmons. An electronic generator of random numbers. IRE Transactions on Electronic Computers, EC-11(2):284, April 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219362. See comments

[Bar63].

Dietmeyer:1965:COF

[DS65] D. L. Dietmeyer and P. R. Schneider. A computer-oriented factoring algorithm for NOR logic design. IEEE Transactions on Electronic Computers, EC-14(6): 868–874, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038606.

Dietmeyer:1967:ISR

[DS67a] Donald L. Dietmeyer and Peter R.

Schneider. Identification of symmetry, redundancy and equivalence of Boolean functions. IEEE Transactions on Electronic Computers, EC-16(6):804–817, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039197.

Dwyer:1967:DAC

[DS67b] Thomas F. Dwyer and Rabah Shahbender. Device applications of charge rectifiers. IEEE Transactions on Electronic Computers, EC-16(6):723–731, December 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039190.

Dietmeyer:1969:LDA

[DS69] D. L. Dietmeyer and Yueh-Hsung Su. Logic design automation of fan-in limited NAND networks. IEEE Transactions on Computers, C-18(1):11–22, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671114.

Dunnet:1965:RMM

[Dun65] W. J. Dunnet. R65-69 mathematical models for predicting pulse characteristics in digital logic systems. IEEE Transactions on Electronic Computers, EC-14(5):745–746, October 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038579.

Dvorak:1968:TRC

[Dvo68] V. Dvorak. A two-rail cascade synthesis of Boolean functions. IEEE Transactions on Computers, C-17(6):592–596, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687403.

Dick:1967:ADC

[DW67] D. E. Dick and H. J. Wertz. Analog and digital computation of Fourier series and integrals. IEEE Transactions on Electronic Computers, EC-16(1):8–13, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038978.

Dwyer:1969:CFT

[Dwy69] T. F. Dwyer. Comments on “Fault Testing and Diagnosis in Combinational Digital Circuits”. IEEE Transactions on Computers, C-18(8):760, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671354. See [Kau68].

Earnshaw:1964:DTD

[Ear64] J. B. Earnshaw. Design for a tunnel diode-transistor store with nondestructive read-out of information. IEEE Transactions on Electronic Computers, EC-13(6): 710–722, December 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038303.

Earley:1968:RUT

[Ear68] J. Earley. R68-46 use of transition matrices in compiling. IEEE Transactions on Computers, C-17 (11):1098, November 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687272.

Easley:1958:TCD

[Eas58] James W. Easley. Transistor characteristics for direct-coupled transistor logic circuits. IRE Transactions on Electronic Computers, EC-7(1):6–16, March 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222089.

Estrin:1963:PPR

[EBTB63] G. Estrin, B. Bussell, R. Turn, and J. Bibb. Parallel processing in a restructurable computer system. IEEE Transactions on Electronic Computers, EC-12(6): 747–755, December 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038024.

Eckes:1962:NAC

[Eck62] Harry R. Eckes. New analogcomputer setup for a bistable element permitting symmetrical triggering. IRE Transactions on Electronic Computers, EC-11(5):709, October 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219437.

Eckes:1965:FMC

[Eck65] H. R. Eckes. A fast modecontrol switch for iterative differential analyzers. IEEE Transactions on Electronic Computers, EC-14(6):946–950, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038623.

Eriksson:1967:TPB

[ED67] Henrik Eriksson and Per E. Danielsson. Two problems on Boolean memories. IEEE Transactions on Electronic Computers, EC-16 (3):364–365, June 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039079.

Edwards:1961:HST

[Edw61] D. B. G. Edwards. High-speed transistorized adders. IRE Transactions on Electronic Computers, EC-10(3):532, September 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219247.

Eisenstadt:1961:ABF

[EG61] B. M. Eisenstadt and B. Gold.

Autocorrelations for Boolean functions of noiselike periodic sequences. IRE Transactions on Electronic Computers, EC-10(3): 383–388, September 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219226.

Estrin:1956:NHS

[EGP56] G. Estrin, B. Gilchrist, and J. H.

Pomerene. A note on high-speed digital multiplication. IRE Transactions on Electronic Computers, EC-5(3):140, September 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219936.

Einhorn:1961:USA

[Ein61] Sidney N. Einhorn. The use of the simplex algorithm in the mechanization of Boolean switching functions by means of magnetic cores. IRE Transactions on Electronic Computers, EC-10(4): 615–622, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219267. See comments

[Min62, Ein62].

Einhorn:1962:ARC

[Ein62] S. N. Einhorn. Author’s reply [to comments on “The Use of the Simplex Algorithm in the Mechanization of Boolean Switching Functions by Means of Magnetic Cores”]. IRE Transactions on Electronic Computers, EC-11(4):574, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219403. See [Ein61,

Min62].

Even:1967:MMS

[EKP67] S. Even, I. Kohavi, and A. Paz. On minimal modulo 2 sums of products for switching functions. IEEE Transactions on Electronic Computers, EC-16(5):671–674, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039160.

Elfant:1962:ADF

[Elf62] R. F. Elfant. Approximations for the demagnetizing factors of hollow cylinders represented by confocal hollow prolate spheroids. IRE Transactions on Electronic Computers, EC-11 (6):789, December 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219464.

Ellis:1965:SCL

[Ell65] D. T. Ellis. A synthesis of combinational logic with NAND or NOR elements. IEEE Transactions on Electronic Computers, EC-14(5):701–705, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038555.

Elspas:1960:SCS

[Els60] Bernard Elspas. Self-complementary symmetry types of Boolean functions. IRE Transactions on Electronic Computers, EC-9(2): 264–266, June 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219832.

Elspas:1963:PP

[Els63a] B. Elspas. PTGEC people. IEEE Transactions on Electronic Computers, EC-12(1):59, February 1963.

CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037781.

Elspas:1963:RCL

[Els63b] B. Elspas. R63-74 coding for logical operations. IEEE Transactions on Electronic Computers, EC-12(4):422–423, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037944.

Elspas:1964:RAC

[Els64] B. Elspas. R64-35 an application of coding theory to a file address problem. IEEE Transactions on Electronic Computers, EC-13(3):321, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038184.

Elspas:1969:RCG

[Els69] B. Elspas. R69-22 computation over Galois fields using shiftregisters. IEEE Transactions on Computers, C-18(8):763, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671359.

Even:1966:TPC

[EM66] S. Even and A. R. Meyer. Test for planarity of a circuit given by an expression. IEEE Transactions on Electronic Computers, EC-15(3):372–375, June 1966.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038775.

Even:1969:SBE

[EM69] S. Even and A. R. Meyer. Sequential Boolean equations. IEEE Transactions on Computers, C-18 (3):230–240, March 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671229.

Embley:1966:N

[Emb66] R. W. Embley. Notices. IEEE Transactions on Electronic Computers, EC-15(3):413, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038792.

Edwards:1961:ACF

[ENB61] H. H. Edwards, V. L. Newhouse, and J. W. Bremer. Analysis of a crossed-film cryoton shift register. IRE Transactions on Electronic Computers, EC-10 (2):285–287, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219201.

Endres:1959:CC

[End59] Richard O. Endres. The Chairman’s column. IRE Transactions on Electronic Computers, EC-8 (4):431, December 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222053.

Engelbart:1961:CGT

[Eng61a] Douglas C. Engelbart. Correction: “Games that Teach the Fundamentals of Computer Operation”. IRE Transactions on Electronic Computers, EC-10(2):291, June 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219206. See [Eng61b].

Engelbart:1961:GTF

[Eng61b] Douglas C. Engelbart. Games that teach the fundamentals of computer operation. IRE Transactions on Electronic Computers, EC-10(1):31–41, March 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219149. See correction

[Eng61a].

Epley:1963:PN

[Epl63] D. L. Epley. PTGEC news. IEEE Transactions on Electronic Computers, EC-12(6):962, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038049.

Epley:1966:RBP

[Epl66] D. L. Epley. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-15(4):676–682, August 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038849.

Epley:1968:RMC

[Epl68] D. L. Epley. R68-5 models of computational systems-cyclic to acyclic graph transformations. IEEE Transactions on Computers, C-17(2):191–192, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687315.

Epstein:1958:SEC

[Eps58] George Epstein. Synthesis of electronic circuits for symmetric functions. IRE Transactions on Electronic Computers, EC-7 (1):57–60, March 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222097.

Epstein:1968:CPT

[Eps68] B. M. Epstein. A 48-channel PCM tape data-acquisition system. IEEE Transactions on Computers, C-17(3):259–267, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687327.

Elspas:1964:BRM

[ES64] B. Elspas and R. A. Short. A bound on the run measure of switching functions. IEEE Transactions on Electronic Computers, EC-13(1):1–4, February 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038069.

Estrin:1963:AAC

[ET63] G. Estrin and R. Turn. Automatic assignment of computations in a variable structure computer system. IEEE Transactions on Electronic Computers, EC-12(6):755–773, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038025.

Even:1963:SBC

[Eve63] Shimon Even. Snake in the box codes. IEEE Transactions on Electronic Computers, EC-12(1):18, February 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037762.

Even:1964:RNR

[Eve64] Shimon Even. Rational numbers and regular events. IEEE Transactions on Electronic Computers, EC-13(6):740–741, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038307.

Even:1965:CMS

[Eve65a] Shimon Even. Comments on the minimization of stochastic machines. IEEE Transactions on Electronic Computers, EC-14(4): 634–637, August 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038519.

Even:1965:ILA

[Eve65b] Shimon Even. On information lossless automata of finite order. IEEE Transactions on Electronic Computers, EC-14(4):561–569, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038507.

Even:1966:AI

[Eve66] S. Even. Author index. IEEE Transactions on Electronic Computers, EC-15(1):148, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038692.

Elbourn:1953:DCT

[EW53] Robert D. Elbourn and Richard P.

Witt. Dynamic circuit techniques used in SEAC and DYSEAC. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(1):2–9, March 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Fantauzzi:1968:NMC

[Fan68] G. Fantauzzi. NORNAND Maitra cascades. IEEE Transactions on Computers, C-17(11):1074– 1080, November 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687265.

Farrell:1962:PGL

[Far62] James L. Farrell. Pulse generator with logarithmic spacing. IRE Transactions on Electronic Computers, EC-11(4): 531–534, August 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219393.

Farrand:1968:REA

[Far68] W. A. Farrand. R68-45 electrically alterable digital differential analyzer. IEEE Transactions on Computers, C-17(10):1004, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687252.

Farmer:1969:DCU

[Far69] N. A. Farmer. A digital comparator for use with computer displays. IEEE Transactions on Computers, C-18(3):269–270, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671234.

Fu:1967:DPA

[FCC67] King-Sun Fu, Y. T. Chien, and Gerald P. Cardillo. A dynamic programming approach to sequential pattern recognition. IEEE Transactions on Electronic Computers, EC-16(6):790–803, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039196.

Fan:1963:AMO

[FDBW63] G. Fan, E. Donath, E. S. Barrekette, and A. Wirgin. Analysis of a magneto-optic readout system. IEEE Transactions on Electronic Computers, EC-12(1):3–9, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037759.

Feigenbaum:1961:SCC

[Fei61] Edward A. Feigenbaum. Soviet cybernetics and computer sciences — 1960. IRE Transactions on Electronic Computers, EC-10(4):759–776, December 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219285.

Fein:1963:RP

[Fei63a] Louis Fein. Renaming the PTGEC.

IEEE Transactions on Electronic Computers, EC-12(2):136, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037814.

Fein:1963:STC

[Fei63b] Louis Fein. Scientific and technical conventions. IEEE Transactions on Electronic Computers, EC-12(2):141–142, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037818.

Feldman:1963:RSI

[Fel63] J. Feldman. R63-89 on the solution of an information retrieval problem. IEEE Transactions on Electronic Computers, EC-12(5):581, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037995.

Fellgett:1965:LDA

[Fel65] Peter Fellgett. Logical design of analog-to-digital converters. IEEE Transactions on Electronic Computers, EC-14(5):740–741, October 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038567.

Fenwick:1969:BMO

[Fen69] P. M. Fenwick. Binary multiplication with overlapped addition cycles. IEEE Transactions on Computers, C-18(1):71–74, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671120.

Ferrari:1967:DMU

[Fer67] Domenico Ferrari. A division method using a parallel multiplier. IEEE Transactions on Electronic Computers, EC-16(2): 224–226, April 1967. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039036. See correction

[Fer69].

Ferrari:1968:FCP

[Fer68] D. Ferrari. Fast carry-propagation iterative networks. IEEE Transactions on Computers, C-17(2): 136–145, February 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687304.

Ferrari:1969:CDM

[Fer69] D. Ferrari. Correction to “A Division Method Using a Parallel Multiplier”. IEEE Transactions on Computers, C-18(10):960, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671148. See [Fer67].

Firschein:1963:ASD

[FF63a] O. Firschein and M. Fischler.

Automatic subclass determination for pattern-recognition applications. IEEE Transactions on Electronic Computers, EC-12 (2):137–141, April 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037817.

Fukinuki:1963:SBL

[FF63b] T. Fukinuki and M. Fujinaka. A stable basic logic circuit. IEEE Transactions on Electronic Computers, EC-12(3):318–319, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037872.

Firschein:1967:CPR

[FF67] O. Firschein and M. A. Fischler. Comment on “A Pattern Recognition Experiment with Near-Optimum Results”. IEEE Transactions on Electronic Computers, EC-16(2):231, April 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039039. See [RH66,

Kno67a].

Flores:1960:ORS

[FG60] I. Flores and L. Grey. Optimization of reference signals for character recognition systems. IRE Transactions on Electronic Computers, EC-9(1):54–61, March 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221605.

Frei:1961:MRM

[FG61] E. H. Frei and J. Goldberg. A method for resolving multiple responses in a parallel search file. IRE Transactions on Electronic Computers, EC-10(4):718– 722, December 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219279.

Freeman:1964:AJP

[FG64] H. Freeman and L. Garder. Apictorial jigsaw puzzles: The computer solution of a problem in pattern recognition. IEEE Transactions on Electronic Computers, EC-13(2):118–127, April 1964.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038109.

Ferrari:1969:CSS

[FG69] D. Ferrari and A. Grasselli. A cellular structure for sequential networks. IEEE Transactions on Computers, C-18(10):947–953, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671144.

Friedman:1969:UST

[FGU69] A. D. Friedman, R. L. Graham, and J. D. Ullman. Universal single transition time asynchronous state assignments. IEEE Transactions on Computers, C-18(6):541–547, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671300.

Fogarty:1962:FSO

[FH62] L. E. Fogarty and R. M. Howe.

Flight simulation of orbital and re-entry vehicles. IRE Transactions on Electronic Computers, EC-11(4):555–563, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219396.

Flynn:1963:VFL

[FH63] M. J. Flynn and D. S. Henderson. Variable field-length data manipulation in a fixed wordlength memory. IEEE Transactions on Electronic Computers, EC-12(5):512–516, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037965.

Fukunaga:1965:DTR

[FI65] K. Fukunaga and T. Ito. A design theory of recognition functions in self-organizing systems. IEEE Transactions on Electronic Computers, EC-14(1):44–52, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038348.

Fialkowski:1966:COS

[Fia66] Konrad R. Fialkowski. The 2’s complement and other semi-systematic binary codes. IEEE Transactions on Electronic Computers, EC-15(4):603–605, August 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038834.

Fielder:1964:SAU

[Fie64] Daniel C. Fielder. On Shannon’s almost uniform distribution. IEEE Transactions on Electronic Computers, EC-13(1):53–54, February 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038078.

Fife:1966:WR

[Fif66] D. W. Fife. Workshop reports. IEEE Transactions on Electronic Computers, EC-15(3):411, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038790.

Fife:1968:RSM

[Fif68a] D. W. Fife. R68-10 SODAS and a methodology for system design. IEEE Transactions on Computers, C-17(4):405, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687358.

Fife:1968:RCS

[Fif68b] D. W. Fife. R68-47 computer scheduling methods and their countermeasures. IEEE Transactions on Computers, C-17(11): 1098–1099, November 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687273.

Filippazzi:1967:NAP

[Fil67] F. Filippazzi. A new approach to permanent memory. IEEE Transactions on Electronic Computers, EC-16(3):370–371, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039082.

Findler:1961:CMG

[Fin61] N. V. Findler. Computer model of gambling and bluffing. IRE Transactions on Electronic Computers, EC-10(1):97–98, March 1961.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219165.

Firschein:1963:MPT

[Fir63a] O. Firschein. Membership in the Professional Technical Group on Electronic Computers. IEEE Transactions on Electronic Computers, EC-12(6):964, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038051.

Firschein:1963:REM

[Fir63b] O. Firschein. R63-70 experiments on machine learning to recognize. IEEE Transactions on Electronic Computers, EC-12 (4):420–421, August 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037940.

Firschein:1964:ACC

[Fir64a] O. Firschein. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(3):323–336, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038172.

Firschein:1964:RBP

[Fir64b] O. Firschein. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-13(1):62–65, February 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038085.

Firschein:1965:RCP

[Fir65] O. Firschein. R65-29 A computer program which ’understands’. IEEE Transactions on Electronic Computers, EC-14(2):279–280, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038437.

Fischmann:1957:GCC

[Fis57] Arieh F. Fischmann. A gray code counter. IRE Transactions on Electronic Computers, EC-6(2):120, June 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221582.

Fischler:1963:RLP

[Fis63] M. A. Fischler. R63-31 A “Logical Pattern” recognition program. IEEE Transactions on Electronic Computers, EC-12(2): 154–155, April 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037831.

Fischler:1964:RAJ

[Fis64] M. A. Fischler. R64-77 apictorial jigsaw puzzles: The computer solution of a problem in pattem recognition. IEEE Transactions on Electronic Computers, EC-13(6):760, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038328.

Fischer:1966:RCM

[Fis66] M. A. Fischer. R66-86 A chess mating combinations program. IEEE Transactions on Electronic Computers, EC-15(6):961, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038964.

Fitzhugh:1963:RAP

[Fit63] R. Fitzhugh. R63-40 an active pulse transmission line simulating nerve axon. IEEE Transactions on Electronic Computers, EC-12(2):158, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037840.

Frank:1965:RGD

[FK65] C. J. Frank and A. F. Kaupe. R65-21 GIER — A Danish computer of medium size. IEEE Transactions on Electronic Computers, EC-14(2):277, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038429.

Fuschillo:1966:MMT

[FK66] N. Fuschillo and J. Kroboth. Multifunction monolithic thin-film compatible DTL logic circuits for data processing. IEEE Transactions on Electronic Computers, EC-15(4):640–646, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038838.

Fukunaga:1969:CBR

[FK69] K. Fukunaga and T. F. Krile. Calculation of Bayes’ recognition error for two multivariate Gaussian distributions. IEEE Transactions on Computers, C-18(3):220–229, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671228.

Freeman:1967:AST

[FL67] Herbert Freeman and Philippe P.

Loutrel. An algorithm for the solution of the two-dimensional “Hidden-Line” problem. IEEE Transactions on Electronic Computers, EC-16(6):784–790, December 1967. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4039195.

Flehinger:1963:PN

[Fle63a] B. J. Flehinger. PTGEC news.

IEEE Transactions on Electronic Computers, EC-12(1):55, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037777.

Flehinger:1963:SI

[Fle63b] B. J. Flehinger. Subject index. IEEE Transactions on Electronic Computers, EC-12(1): 51–54, February 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037775.

Fleck:1967:SCD

[Fle67] A. C. Fleck. On the strong connectedness of the direct product. IEEE Transactions on Electronic Computers, EC-16(1):90, February 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038990.

Flores:1956:RNS

[Flo56] Ivan Flores. Reflected number systems. IRE Transactions on Electronic Computers, EC-5(2):79–82, June 1956. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219803.

Flores:1958:OCR

[Flo58] Ivan Flores. An optimum character recognition system using decision functions. IRE Transactions on Electronic Computers, EC-7(2):180, June 1958. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222530.

Flores:1962:RBS

[Flo62] I. Flores. Review: Computer Synthesis of Character Recognition Systems, by D. N. Freeman. IRE Transactions on Electronic Computers, EC-11(3):424, June 1962. CODEN IRELAO. ISSN 0367-9950.

Floyd:1964:SI

[Flo64a] R. W. Floyd. Summary of the issue. IEEE Transactions on Electronic Computers, EC-13(4):345, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038196.

Floyd:1964:SPL

[Flo64b] Robert W. Floyd. The syntax of programming languages — A survey. IEEE Transactions on Electronic Computers, EC-13(4):346–353, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038197.

Flores:1968:RIP

[Flo68] I. Flores. R68-37 intercommunication of processors and memory. IEEE Transactions on Computers, C-17(8):818–819, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687461.

Friedman:1968:SAS

[FM68] A. D. Friedman and P. R. Menon.

Synthesis of asynchronous sequential circuits with multiple-input changes. IEEE Transactions on Computers, C-17(6):559–566, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687400.

Fogel:1957:HCF

[Fog57] Lawrence J. Fogel. The human computer in flight control. IRE Transactions on Electronic Computers, EC-6(3):195–202, September 1957. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222022.

Fogarty:1967:RGP

[Fog67] L. E. Fogarty. R67-28 A generalpurpose analog translational trajectory program for orbiting and reentry vehicles. IEEE Transactions on Electronic Computers, EC-16(2):241, April 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039054.

Foster:1968:DPA

[Fos68a] C. C. Foster. Determination of priority in associative memories. IEEE Transactions on Computers, C-17(8):788–789, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687452.

Foster:1968:CBP

[Fos68b] K. A. Foster. Comments on “Basic Properties and a Construction Method for Fail-Safe Logical Systems”. IEEE Transactions on Computers, C-17(7):702, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687433. See [MK67].

Freeman:1954:TSA

[FP54] H. Freeman and E. Parsons. Timesharing analog multiplier (TSAM).

Transactions of the I.R.E. Professional Group on Electronic Computers, EC-3(1):11–17, March 1954. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Fischler:1962:TRA

[FP62] M. A. Fischler and E. A. Poe.

Threshold realization of arithmetic circuits. IRE Transactions on Electronic Computers, EC-11 (2):287–288, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219366.

Flores:1958:MSW

[FR58] I. Flores and F. Ragonese. A method for synthesizing the waveform generated by a character, printed in magnetic ink, in passing beneath a magnetic reading head. IRE Transactions on Electronic Computers, EC-7(4): 277–282, December 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222659.

Frankel:1957:LDS

[Fra57] Stanley P. Frankel. The logical design of a simple general purpose computer. IRE Transactions on Electronic Computers, EC-6(1):5–14, March 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221555.

Frankel:1958:MLC

[Fra58] S. P. Frankel. On the minimum logical complexity required for a general purpose computer. IRE Transactions on Electronic Computers, EC-7(4):282–285, December 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222660.

Frankel:1959:LDM

[Fra59] Stanley P. Frankel. A logic design for a microwave computer. IRE Transactions on Electronic Computers, EC-8(3): 271–276, September 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222683.

Frank:1964:ROS

[Fra64a] John E. Frank. The role of operational simulation to communications system design. IEEE Transactions on Electronic Computers, EC-13(5):640, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038271.

Frankel:1964:RLO

[Fra64b] S. Frankel. R64-36 the logical organization of the PB 440 microprogrammable computer. IEEE Transactions on Electronic Computers, EC-13(3):321, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038185.

Frazer:1965:CGN

[Fra65] W. D. Frazer. Computer group news. IEEE Transactions on Electronic Computers, EC-14(2): 299–306, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038425.

Fraser:1969:IBR

[Fra69] D. Fraser. Incrementing a bitreversed integer. IEEE Transactions on Computers, C-18(1): 74, January 1969. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671121.

Freeman:1961:CSC

[Fre61a] D. N. Freeman. Computer synthesis of character-recognition systems. IRE Transactions on Electronic Computers, EC-10(4): 735–747, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219282.

Freeman:1961:EAG

[Fre61b] Herbert Freeman. On the encoding of arbitrary geometric configurations. IRE Transactions on Electronic Computers, EC-10 (2):260–268, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219197.

Freeman:1967:CMS

[Fre67a] Herbert Freeman. Calculation of mean shift for a binary multiplier using 2, 3, or 4 bits at a time. IEEE Transactions on Electronic Computers, EC-16(6): 864–866, December 1967. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039205.

Freiman:1967:ACC

[Fre67b] C. V. Freiman. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-16(5):695–705, October 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039171.

Freedman:1968:RLE

[Fre68] H. A. Freedman. R68-25 an online editor. IEEE Transactions on Computers, C-17(5):523, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687390.

Friedes:1966:C

[Fri66a] A. Friedes. Corrections. IEEE Transactions on Electronic Computers, EC-15(1):101, February 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038671.

Friedes:1966:CPS

[Fri66b] Albert Friedes. Critical pairs and set systems. IEEE Transactions on Electronic Computers, EC-15(1):102–104, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038672.

Friedman:1966:FAS

[Fri66c] Arthur D. Friedman. Feedback in asynchronous sequential circuits. IEEE Transactions on Electronic Computers, EC-15(5): 740–749, October 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038880.

Friedman:1966:FSS

[Fri66d] Arthur D. Friedman. Feedback in synchronous sequential switching circuits. IEEE Transactions on Electronic Computers, EC-15 (3):354–367, June 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038771.

Friedman:1966:DG

[Fri66e] Henry D. Friedman. A design for (d, k) graphs. IEEE Transactions on Electronic Computers, EC-15(2):253–254, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038724.

Friedman:1967:FDR

[Fri67a] A. D. Friedman. Fault detection in redundant circuits. IEEE Transactions on Electronic Computers, EC-16(1):99–100, February 1967.

CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038995.

Friedman:1967:CMS

[Fri67b] Arthur D. Friedman. Comment on “A Method for the Selection of Prime Implicants”. IEEE Transactions on Electronic Computers, EC-16(2):221–222, April 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039034. See [Luc66a].

Frola:1965:NPC

[Fro65] F. R. Frola. A note on the problem of classifying vertices of the n cube into m categories. IEEE Transactions on Electronic Computers, EC-14(5):729–730, October 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038563.

Fullwood:1959:ABT

[FS59] R. Fullwood and P. Spiegel. An accurate bidirectional transmission gate using semiconductor diodes. IRE Transactions on Electronic Computers, EC-8 (4):498, December 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222065.

Fife:1965:TCD

[FS65] D. W. Fife and J. L. Smith. Transmission capacity of disk storage systems with concurrent arm positioning. IEEE Transactions on Electronic Computers, EC-14 (4):575–582, August 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038509.

Fischler:1968:ATS

[FT68] M. A. Fischler and M. Tannenbaum.

Assumptions in the threshold synthesis of symmetric switching functions. IEEE Transactions on Computers, C-17(3):273–279, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687331.

Fuchs:1956:WTR

[Fuc56] Alexander Fuchs. Working time in repetitive analog computers. IRE Transactions on Electronic Computers, EC-5(2):94, June 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219806.

Fuchs:1957:NAD

[Fuc57] Alexander Fuchs. A note on the accuracy of differential analyzers. IRE Transactions on Electronic Computers, EC-6(1):36, March 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221562.

Fukunaga:1964:TNA

[Fuk64] Keinosuke Fukunaga. A theory of nonlinear autonomous sequential nets using z transforms. IEEE Transactions on Electronic Computers, EC-13(3):310–312, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038168.

Frank:1966:IRS

[FY66] H. Frank and S. S. Yau. Improving reliability of a sequential machine by error-correcting state assignments. IEEE Transactions on Electronic Computers, EC-15(1):111–113, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038677.

Friedman:1969:CMU

[FY69a] T. D. Friedman and S. C. Yang. Correction to “Methods Used in an Automatic Logic Design Generator (ALERT)”. IEEE Transactions on Computers, C-18(12): 1142–1143, December 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671196. See [FY69b].

Friedman:1969:MUA

[FY69b] T. D. Friedman and Sih-Chin Yang. Methods used in an automatic logic design generator (ALERT). IEEE Transactions on Computers, C-18(7):593–614, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671320. See correction

[FY69a].

Gillespie:1958:ASM

[GA58] R. G. Gillespie and D. D.

Aufenkamp. On the analysis of sequential machines. IRE Transactions on Electronic Computers, EC-7(2):119–122, June 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222520.

Gabelman:1962:SBF

[Gab62] Irving J. Gabelman. The synthesis of Boolean functions using a single threshold element. IRE Transactions on Electronic Computers, EC-11(5):639–642, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219424.

Gabelman:1964:PTS

[Gab64a] Irving J. Gabelman. Properties and transformations of single threshold element functions. IEEE Transactions on Electronic Computers, EC-13(6):680–684, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038299.

Gabor:1964:RSP

[Gab64b] A. Gabor. R64-80 signal-processing for increased bit densities in digital magnetic recording. IEEE Transactions on Electronic Computers, EC-13(6):761–762, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038331.

Gabelman:1965:ATE

[Gab65] Irving J. Gabelman. An algorithm for threshold element analysis. IEEE Transactions on Electronic Computers, EC-14(4): 623–625, August 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038515.

Gabor:1967:ACS

[Gab67] Andrew Gabor. Adaptive coding for self-clocking recording. IEEE Transactions on Electronic Computers, EC-16(6):866–868, December 1967. CODEN IEECA8. ISSN 0367-7508. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4039206.

Gaffney:1968:SDM

[Gaf68] J. E. Gaffney, Jr. Sequential decision-making device for information-processing applications. IEEE Transactions on Computers, C-17(12):1151–1156, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687284.

Gaines:1968:RRP

[Gai68] B. R. Gaines. R68-18 random pulse machines. IEEE Transactions on Computers, C-17(4):410, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687366.

Galli:1962:SSL

[Gal62] E. J. Galli. The Stenowriter — A system for the lexical processing of Stenotypy. IRE Transactions on Electronic Computers, EC-11(2):187–199, April 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219352.

Gamblin:1968:AHS

[Gam68] R. L. Gamblin. An analysis of high-speed, linear-passive binary, read-only stores. IEEE Transactions on Computers, C-17(9):889–893, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687478.

Gange:1964:ACC

[Gan64] R. A. Gange. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(5):651–663, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038276.

Garner:1958:GPC

[Gar58] Harvey L. Garner. Generalized parity checking. IRE Transactions on Electronic Computers, EC-7(3):207–213, September 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222577.

Garner:1959:RNS

[Gar59a] Harvey L. Garner. The residue number system. IRE Transactions on Electronic Computers, EC-8(2):140–147, June 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219515.

Garner:1959:RMS

[Gar59b] Harvey L. Garner. A ring model for the study of multiplication for complement codes. IRE Transactions on Electronic Computers, EC-8(1):25–30, March 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222757.

Garner:1965:RID

[Gar65] H. L. Garner. R65-22 improving digital computer performance using residue number theory. IEEE Transactions on Electronic Computers, EC-14(2):277, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=4038430;   http://
ieeexplore.ieee.org/xpl/RecentIssue.
jsp?punumber=4037753;   http:
//ieeexplore.ieee.org/xpl/tocresult.
jsp?isnumber=4038385.

Garner:1966:ECA

[Gar66] Harvey L. Garner. Error codes for arithmetic operations. IEEE Transactions on Electronic Computers, EC-15(5):763–770, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038884.

Garwick:1969:RUS

[Gar69] J. V. Garwick. R69-32 A universal syntax-directed top-down analyzer. IEEE Transactions on Computers, C-18(10):964–965, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671152.

Gaston:1963:STL

[Gas63] C. A. Gaston. A simple test for linear separability. IEEE Transactions on Electronic Computers, EC-12(2):134–135, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037812.

Gaskill:1964:VPO

[Gas64] Roger A. Gaskill. A versatile problem-oriented language for engineers. IEEE Transactions on Electronic Computers, EC-13(4): 415–421, August 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038206.

Gaskill:1965:RPD

[Gas65] R. A. Gaskill. R65-44 PACTO-LUS — A digital analog simulator program for the IBM 1620. IEEE Transactions on Electronic Computers, EC-14(3):502, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038490.

Gaskill:1966:RDS

[Gas66] R. A. Gaskill. R66-60 digital simulation languages: A critique and A guide. IEEE Transactions on Electronic Computers, EC-15(5):836, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038905.

Gauss:1960:CA

[Gau60] E. J. Gauss. Concerning abstracts. IRE Transactions on Electronic Computers, EC-9(1):125, March 1960. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5221612.

Gaudiosi:1964:PMV

[Gau64] J. V. Gaudiosi. Proposed magnetic variation analog computer for use in flight simulators. IEEE Transactions on Electronic Computers, EC-13(3):312, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038169.

Glaser:1955:LVQ

[GB55] E. M. Glaser and H. Blasbalg. A logarithmic voltage quantizer. IRE Transactions on Electronic Computers, EC-4(4): 150–155, December 1955. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219485.

Ghosh:1969:MSG

[GBC69] S. Ghosh, D. Basu, and A. K.

Choudhury. Multigate synthesis of general Boolean functions by threshold logic elements. IEEE Transactions on Computers, C-18(5):451–456, May 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671277.

Graham:1967:RFF

[GD67] Peter J. Graham and Raymond J.

Distler. RST flip-flop input equations. IEEE Transactions on Electronic Computers, EC-16(4): 443–445, August 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039108. See comments

[Sur68].

Green:1962:ABA

[GDD62] C. Green, H. D’Hoop, and A. Debroux. APACHE-A breakthrough in analog computing. IRE Transactions on Electronic Computers, EC-11(5):699–706, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219433.

Gear:1963:AI

[Gea63] C. W. Gear. Author index. IEEE Transactions on Electronic Computers, EC-12(3):351, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037886.

Gear:1965:RIS

[Gea65] C. W. Gear. R65-56 IBM System 360 engineering. IEEE Transactions on Electronic Computers, EC-14(4):670–671, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038537.

Gelernter:1961:NSR

[Gel61] Herbert Gelernter. A note on the system requirements of a digital computer for the manipulation of list structures. IRE Transactions on Electronic Computers, EC-10(3):484–489, September 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219237.

Geller:1962:OIJ

[Gel62] Sidney B. Geller. Obtaining the internal junction characteristics of a transistor for use in analog simulation. IRE Transactions on Electronic Computers, EC-11(5):709–710, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219438.

Germeroth:1960:CTB

[Ger60] Joel H. Germeroth. Casting out threes in binary numbers. IRE Transactions on Electronic Computers, EC-9(3): 373, September 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219863.

Gerace:1963:MCC

[Ger63] G. B. Gerace. Microprogrammed control for computing systems. IEEE Transactions on Electronic Computers, EC-12(6):733–747, December 1963. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038023.

Gerace:1968:DSD

[Ger68] G. B. Gerace. Digital system design automation — a method for designing a digital system as a sequential network system. IEEE Transactions on Computers, C-17(11):1044–1061, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687262.

Giloi:1968:CMI

[GG68] W. Giloi and H. Grebe. Construction of multistep integration formulas for simulation purposes. IEEE Transactions on Computers, C-17(12):1121–1131, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687281.

Gerace:1969:SML

[GG69] G. B. Gerace and G. Gestri. Sequential machines with less delay elements than feedback paths.

IEEE Transactions on Computers, C-18(2):132–144, February 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671209.

Grinich:1959:PTS

[GH59] V. H. Grinich and I. Haas. P-N-π-N triode switching applications. IRE Transactions on Electronic Computers, EC-8 (2):108–113, June 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219510.

Ghiron:1962:RMR

[Ghi62] Hugo Ghiron. Rules to manipulate regular expressions of finite automata. IRE Transactions on Electronic Computers, EC-11(4):574–575, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219404.

Gram:1963:GDC

[GHI^{+}63] C. Gram, O. Hestvik, H. Isaksson, P. T. Jacobsen, J. Jensen, P. Naur, B. S. Petersen, and B. Svejgaard. GIER — A Danish computer of medium size. IEEE Transactions on Electronic Computers, EC-12(6):629–650, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038014.

Gordon:1966:SCP

[GHL^{+}66] B. B. Gordon, R. W. House, A. P. Lechler, L. D. Nelson, and T. Rado. Simplification of the covering problem for multiple output logical networks. IEEE Transactions on Electronic Computers, EC-15(6):891–897, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038932.

Groner:1967:LCC

[GHR67] Gabriel F. Groner, J. F. Heafner, and T. W. Robinson. On-line computer classification of handprinted Chinese characters as a translation aid. IEEE Transactions on Electronic Computers, EC-16(6):856–860, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039203.

Gianola:1964:C

[Gia64] U. G. Gianola. Contributors. IEEE Transactions on Electronic Computers, EC-13(4):477–479, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038219.

Gilbert:1959:DPV

[Gil59a] Edward O. Gilbert. The design of position and velocity servos for multiplying and function generation. IRE Transactions on Electronic Computers, EC-8(3):391–399, September 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222700.

Gilbert:1959:LSA

[Gil59b] Elmer G. Gilbert. Linear system approximation by differential analyzer simulation of orthonormal approximation functions. IRE Transactions on Electronic Computers, EC-8(2):204–209, June 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219523.

Gill:1959:SSD

[Gil59c] Arthur Gill. Systematic scaling for digital differential analyzers. IRE Transactions on Electronic Computers, EC-8(4): 486–489, December 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222063.

Gill:1960:CEF

[Gil60a] Arthur Gill. Characterizing experiments for finite-memory binary automata. IRE Transactions on Electronic Computers, EC-9(4):469–471, December 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219887.

Gill:1960:MCN

[Gil60b] Arthur Gill. Minimization of contact networks subject to reliability specifications. IRE Transactions on Electronic Computers, EC-9(1):122–123, March 1960.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221610.

Gill:1960:OOS

[Gil60c] Arthur Gill. The optimal organization of serial memory transfers. IRE Transactions on Electronic Computers, EC-9(1):12–15, March 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221598.

Gill:1961:CFS

[Gil61a] Arthur Gill. Cascaded finitestate machines. IRE Transactions on Electronic Computers, EC-10(3):366–370, September 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219223.

Gill:1961:NMD

[Gil61b] Arthur Gill. A note on Moore’s Distinguishability Theorem. IRE Transactions on Electronic Computers, EC-10(2):290–291, June 1961. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219205.

Gilbert:1962:RBI

[Gil62] E. N. Gilbert. Review: Transmission of Information, by R. M. Fano. IRE Transactions on Electronic Computers, EC-11(3):424, June 1962. CODEN IRELAO. ISSN 0367-9950.

Gilbert:1963:PN

[Gil63a] E. N. Gilbert. PTGEC news. IEEE Transactions on Electronic Computers, EC-12(2):176, April 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037826.

Gill:1963:RSD

[Gil63b] A. Gill. R63-34 some dangers in state reduction of sequential machines. IEEE Transactions on Electronic Computers, EC-12(2):156, April 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037834.

Gill:1964:ALS

[Gil64a] Arthur Gill. Analysis of linear sequential circuits by confluence sets. IEEE Transactions on Electronic Computers, EC-13 (3):226–231, June 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038146.

Gillis:1964:TAH

[Gil64b] John E. Gillis. A technique for achieving high bit packing density on magnetic tape. IEEE Transactions on Electronic Computers, EC-13(2):112–117, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038108.

Gill:1965:RBP

[Gil65a] A. Gill. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-14(3):497–502, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038478.

Gill:1965:BMS

[Gil65b] Arthur Gill. On the bound to the memory of a sequential machine. IEEE Transactions on Electronic Computers, EC-14 (3):464–466, June 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038466.

Gilbert:1966:IAS

[Gil66a] E. G. Gilbert. Information about the Society. IEEE Transactions on Electronic Computers, EC-15 (3):415–417, June 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038793.

Gill:1966:ACC

[Gil66b] A. Gill. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-15(2):275–288, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038735.

Gill:1966:SPT

[Gil66c] Arthur Gill. On the series-to-parallel transformation of linear sequential circuits. IEEE Transactions on Electronic Computers, EC-15(1):107–108, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038675.

Gilligan:1966:HSM

[Gil66d] T. J. Gilligan. 21/2D high speed memory systems — past, present, and future. IEEE Transactions on Electronic Computers, EC-15(4):475–485, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038821.

Gilliland:1966:SSA

[Gil66e] M. C. Gilliland. A spectral stability analysis of finite difference operators. IEEE Transactions on Electronic Computers, EC-15(6):849–854, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038927.

Gilson:1966:C

[Gil66f] R. P. Gilson. Correction. IEEE Transactions on Electronic Computers, EC-15(6):925, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038936.

Gilson:1966:SRA

[Gil66g] Rodman P. Gilson. Some results of amplitude distribution experiments on shift register generated pseudo-random noise. IEEE Transactions on Electronic Computers, EC-15(6):926–927, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038937.

Gilbert:1969:RNH

[Gil69] E. O. Gilbert. R69-11 A new high performance computer dc amplifier. IEEE Transactions on Computers, C-18(5):477, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671287.

Gimpel:1965:MPB

[Gim65a] James F. Gimpel. A method of producing a Boolean function having an arbitrarily prescribed prime implicant table. IEEE Transactions on Electronic Computers, EC-14(3):485–488, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038472.

Gimpel:1965:RTP

[Gim65b] James F. Gimpel. A reduction technique for prime implicant tables. IEEE Transactions on Electronic Computers, EC-14(4):535–541, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038504.

Gimpel:1967:MTN

[Gim67] James F. Gimpel. The minimization of TANT networks. IEEE Transactions on Electronic Computers, EC-16(1):18–38, February 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038980.

Ginsburg:1959:SMS

[Gin59a] Seymour Ginsburg. Synthesis of minimal-state machines. IRE Transactions on Electronic Computers, EC-8(4):441–449, December 1959. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5222056.

Ginsburg:1959:STM

[Gin59b] Seymour Ginsburg. A synthesis technique for minimal state sequential machines. IRE Transactions on Electronic Computers, EC-8(1):13–24, March 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222756.

Ginsburg:1959:TRG

[Gin59c] Seymour Ginsburg. A technique for the reduction of a given machine to a minimal-state machine. IRE Transactions on Electronic Computers, EC-8(3): 346–355, September 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222696.

Ginsburg:1962:EAM

[Gin62] Seymour Ginsburg. Examples of abstract machines. IRE Transactions on Electronic Computers, EC-11(2):132–135, April 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219345.

Ginsburg:1963:RSF

[Gin63] S. Ginsburg. R63-73 signal flow graph techniques for sequential circuit state diagrams. IEEE Transactions on Electronic Computers, EC-12(4):422, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037943.

Ginzburg:1966:ASP

[Gin66] A. Ginzburg. About some properties of definite, reverse-definite and related automata. IEEE Transactions on Electronic Computers, EC-15(5):806–810, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038893.

Giroux:1966:EBP

[Gir66] Paul H. Giroux. Estimates for best placement of voters in a triplicated logic network. IEEE Transactions on Electronic Computers, EC-15(3):382, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038779.

Giroux:1969:CGP

[Gir69] P. H. Giroux. Comment on “Generalized Parallel Redundancy in Digital Computers”. IEEE Transactions on Computers, C-18(1):80, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671124. See [Deo68].

Giguere:1959:TPC

[GJN59] W. J. Giguere, J. H. Jamison, and J. C. Noll. Transistor pulse circuits for 160-MC clock rates. IRE Transactions on Electronic Computers, EC-8(4):432–438, December 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222054.

Goldstick:1962:DMS

[GK62] G. H. Goldstick and E. F. Klein. Design of memory sense amplifiers. IRE Transactions on Electronic Computers, EC-11 (2):236–253, April 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219357.

Gaines:1965:ICM

[GL65a] R. S. Gaines and C. V. Lee. An improved cell memory. IEEE Transactions on Electronic Computers, EC-14(1):72–75, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038356.

Grasselli:1965:MMN

[GL65b] A. Grasselli and F. Luccio. A method for minimizing the number of internal states in incompletely specified sequential networks. IEEE Transactions on Electronic Computers, EC-14(3):350–359, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038452.

Goldstein:1967:SSS

[GL67] Larry Joel Goldstein and Stephen W.

Leibholz. On the synthesis of signal switching networks with transient blocking. IEEE Transactions on Electronic Computers, EC-16(5):637–641, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039154.

Gleman:1968:RED

[Gle68] R. Gleman. R68-31 the effect of digital compensation for computation delay in a hybrid loop. IEEE Transactions on Computers, C-17(6):608, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687412.

Glinski:1962:MDL

[Gli62] G. S. Glinski. Magnetostrictive delay lines. IRE Transactions on Electronic Computers, EC-11(1):86, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219326.

Glinski:1965:N

[Gli65] G. S. Glinski. Notices. IEEE Transactions on Electronic Computers, EC-14(5):762, October 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038575.

Glinski:1966:AI

[Gli66a] G. S. Glinski. Author index. IEEE Transactions on Electronic Computers, EC-15(5):832, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038899.

Glinski:1966:RUA

[Gli66b] G. S. Glinski. R66-84 the use of an analogue computer in some operational research problems. IEEE Transactions on Electronic Computers, EC-15(6): 959–960, December 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038962.

Glucksman:1965:PPC

[Glu65] Herbert A. Glucksman. A parapropagation pattern classifier. IEEE Transactions on Electronic Computers, EC-14(3):434–443, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038463.

Glucksman:1966:ILS

[Glu66] H. Glucksman. On the improvement of a linear separation by extending the adaptive process with a stricter criterion. IEEE Transactions on Electronic Computers, EC-15(6):941–944, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038945.

Goldstick:1962:DCC

[GM62] G. H. Goldstick and D. G.

MacKie. Design of computer circuits using linear programming techniques. IRE Transactions on Electronic Computers, EC-11 (4):518–530, August 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219392.

Geller:1963:MPS

[GM63a] S. B. Geller and P. A. Mantek. A 100 megapulse per second binary counter with impedance steering. IEEE Transactions on Electronic Computers, EC-12(5):568, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037979.

Gregory:1963:SC

[GM63b] J. Gregory and R. McReynolds.

The SOLOMON computer. IEEE Transactions on Electronic Computers, EC-12(6):774–781, December 1963. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038026.

Gilbert:1965:ECA

[GM65] C. P. Gilbert and B. J. J. McHugh. Economical circuits for the analog solution of algebraic equations. IEEE Transactions on Electronic Computers, EC-14(2):217–229, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038402.

Geller:1961:GJT

[GMB61] S. B. Geller, P. A. Mantek, and D. R. Boyle. A general junction-transistor equivalent circuit for use in large-signal switching analysis. IRE Transactions on Electronic Computers, EC-10 (4):670–679, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219272.

Goto:1960:EDH

[GMN^{+}60] E. Goto, K. Murata, K. Nakazawa, K. Nakagawa, T. Moto-Oka, Y. Matsuoka, Y. Ishibashi, H. Ishida, T. Soma, and E. Wada. Esaki diode high-speed logical circuits. IRE Transactions on Electronic Computers, EC-9(1):25–29, March 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221600.

Graham:1959:DLE

[GMPS59] M. Graham, G. L. Miller, H. R. Pate, and R. Spinrad. The design of a large electrostatic memory. IRE Transactions on Electronic Computers, EC-8(4): 479–485, December 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222062.

Gray:1959:SOM

[GNV59] H. J. Gray, H. H. Nishino, and A. L. Vivatson. System organization of a multiple-cockpit digital operational flight trainer. IRE Transactions on Electronic Computers, EC-8(3):326–330, September 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222692.

Gupta:1963:MST

[GO63] S. C. Gupta and J. J. O’Donnell.

Manipulation of state transition matrices of finite Markov chains. IEEE Transactions on Electronic Computers, EC-12(1): 19–20, February 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037764.

Goering:1964:RCH

[Goe64] Orville Goering. Radix conversion in a hexadecimal machine. IEEE Transactions on Electronic Computers, EC-13(2):154–155, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038116.

Golay:1957:LBB

[Gol57] Marcel J. E. Golay. The logic of bidirectional binary counters. IRE Transactions on Electronic Computers, EC-6(1):1–4, March 1957. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221554.

Goldstick:1960:CSN

[Gol60] G. H. Goldstick. Comparison of saturated and nonsaturated switching circuit techniques. IRE Transactions on Electronic Computers, EC-9(2):161–175, June 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219815.

Goldberg:1963:RBP

[Gol63] J. Goldberg. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-12(4):416–423, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037928.

Goldman:1964:PCM

[Gol64] Herbert D. Goldman. Protecting core memory circuits with error correcting cyclic codes. IEEE Transactions on Electronic Computers, EC-13(3):303–304, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038162.

Golay:1969:HPP

[Gol69] M. J. E. Golay. Hexagonal parallel pattern transformations. IEEE Transactions on Computers, C-18(8):733–740, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671349.

Gonzalez:1963:MIC

[Gon63] Rodolfo Gonzalez. A multilayer iterative circuit computer. IEEE Transactions on Electronic Computers, EC-12(6):781–790, December 1963. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038027.

Gonzalez:1965:DAC

[Gon65] Guillermo Gonzalez. Delay approximations for correlation measurements using analog computers. IEEE Transactions on Electronic Computers, EC-14(4): 606–617, August 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038513.

Goode:1955:PSA

[Goo55] H. H. Goode. PGEC student activities and education in computers. IRE Transactions on Electronic Computers, EC-4(2):49–51, June 1955. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407890.

Goodstein:1964:TDA

[Goo64] L. P. Goodstein. Tunnel-diode amplitude comparator. IEEE Transactions on Electronic Computers, EC-13(3):309–310, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038167.

Gose:1965:STN

[Gos65] Earl E. Gose. A synthesis technique for networks consisting of logical functions feeding a linear summation element. IEEE Transactions on Electronic Computers, EC-14(2):254–256, April 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038412.

Goto:1964:NLG

[Got64] Eiichi Goto. A note on logical gain. IEEE Transactions on Electronic Computers, EC-13(5): 606–608, October 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038255.

Gould:1958:NCN

[Gou58] Roderick Gould. A note on contact networks for switching functions of four variables. IRE Transactions on Electronic Computers, EC-7(3):196–198, September 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222575.

Gilchrist:1955:FCL

[GPW55] Bruce Gilchrist, J. H. Pomerene, and S. Y. Wong. Fast carry logic for digital computers. IRE Transactions on Electronic Computers, EC-4(4):133–136, December 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219482.

Gray:1964:TCI

[GR64] Harry J. Gray and Henry Ruston. On techniques for coping with the information explosion. IEEE Transactions on Electronic Computers, EC-13(2):156, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038118.

Golten:1967:UHC

[GR67] J. W. Golten and David Rees. The use of hybrid computing in the analysis of steel rolling. IEEE Transactions on Electronic Computers, EC-16(6):717–722, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039189.

Goldman:1968:CSC

[GR68] H. D. Goldman and J. Rom. Considering solder connections, does triplicated majority voting apply to integrated circuits. IEEE Transactions on Computers, C-17(10): 990–992, October 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687246.

Graham:1956:IMW

[Gra56] Martin Graham. An improved method for Williams storage. IRE Transactions on Electronic Computers, EC-5(3):140, September 1956. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219935.

Gray:1958:IMA

[Gra58] Harry J. Gray. Investigations of magnetic amplifiers with feedback. IRE Transactions on Electronic Computers, EC-7(3):213– 217, September 1958. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222578.

Grado:1960:SEA

[Gra60a] Gilbert R. Grado. A solution to the Euler angle transformation equations. IRE Transactions on Electronic Computers, EC-9(3):362–369, September 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219860.

Gray:1960:EAN

[Gra60b] R. L. Gray. An electrically alterable nondestructive twistor memory. IRE Transactions on Electronic Computers, EC-9(4): 451–455, December 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219883.

Grasselli:1962:DPM

[Gra62a] A. Grasselli. The design of programmodifiable micro-programmed control units. IRE Transactions on Electronic Computers, EC-11(3): 336–339, June 1962. CODEN IRE-LAO. ISSN 0367-9950.

Grasselli:1962:CUS

[Gra62b] Antonio Grasselli. Control units for sequencing complex asynchronous operations. IRE Transactions on Electronic Computers, EC-11 (4):483–493, August 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219387.

Grau:1964:RAP

[Gra64a] A. A. Grau. R64-63 application of pushdown-store machines. IEEE Transactions on Electronic Computers, EC-13(5):649, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038285.

Gray:1964:RPS

[Gra64b] H. J. Gray. R64-18 predicting signal degeneration and gate compatibility in logic circuits. IEEE Transactions on Electronic Computers, EC-13(2):163–164, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038131.

Gray:1964:FSL

[Gra64c] Harry J. Gray. Fields in strip-lines for film memory application. IEEE Transactions on Electronic Computers, EC-13(5):576–580, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038250.

Grasselli:1966:MCP

[Gra66] A. Grasselli. Minimal closed partitions for incompletely specified flow tables. IEEE Transactions on Electronic Computers, EC-15(2):245–249, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038722.

Graham:1969:ECB

[Gra69] M. Graham. Error correction in batch-fabricated memories. IEEE Transactions on Computers, C-18(6):566–567, June 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671307.

Greenwood:1956:RCM

[Gre56] Donald T. Greenwood. The representation of constraints by means of an electronic differential analyzer. IRE Transactions on Electronic Computers, EC-5(3): 111–113, September 1956. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219929.

Green:1961:CLS

[Gre61] Bert F. Green. Computer languages for symbol manipulation. IRE Transactions on Electronic Computers, EC-10(4): 729–735, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219281.

Greenwood:1962:CHO

[Gre62] Donald T. Greenwood. A comparison of higher-order difference methods in the solution of beamvibration problems. IRE Transactions on Electronic Computers, EC-11(1):9–17, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219310.

Green:1963:REH

[Gre63] C. Green. R63-38 experience with hybrid computation. IEEE Transactions on Electronic Computers, EC-12(2):157, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037838.

Greibach:1964:N

[Gre64a] S. A. Greibach. Notices. IEEE Transactions on Electronic Computers, EC-13(6):792, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038323.

Greibach:1964:RCS

[Gre64b] S. A. Greibach. R64-17 the computer-stored thesaurus and its use in concept processing. IEEE Transactions on Electronic Computers, EC-13(2):163, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038130.

Greenwood:1966:RTN

[Gre66a] D. T. Greenwood. R66-80 transient neutron distribution solutions by compressed and realtime computer complexes. IEEE Transactions on Electronic Computers, EC-15(6):958, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038958.

Gregory:1966:DAU

[Gre66b] Robert T. Gregory. On the design of the arithmetic unit of a fixed-word-length computer from the standpoint of computational accuracy. IEEE Transactions on Electronic Computers, EC-15 (2):255–257, April 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038726.

Greenwood:1968:RMM

[Gre68a] D. T. Greenwood. R68-16 A mathematical model for the investigation of three-dimensional fields with asymmetric boundaries. IEEE Transactions on Computers, C-17(4):409, April 1968.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687364.

Greer:1968:CMS

[Gre68b] D. L. Greer. Characterization of the magnetic second-harmonic analog memory. IEEE Transactions on Computers, C-17(6):551–558, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687399.

Greenwood:1969:RHC

[Gre69a] D. T. Greenwood. R69-8 hybrid computer integration of partial differential equations by use of an assumed sum separation of variables. IEEE Transactions on Computers, C-18(4):384, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671265.

Greibach:1969:RMO

[Gre69b] S. A. Greibach. R69-18 multitape one-way nonwriting automata. IEEE Transactions on Computers, C-18(7):669, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671334.

Greibach:1969:RRT

[Gre69c] S. A. Greibach. R69-23 the reduction of tape reversals for offline one-tape Turing machines. IEEE Transactions on Computers, C-18(8):763, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671360.

Gray:1955:DMA

[GRT55] H. J. Gray, M. Rubinoff, and J. Tompkins. A diode multiplexer for analog voltages. IRE Transactions on Electronic Computers, EC-4(2):64–66, June 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407893.

Grisamore:1958:LDU

[GRU58] N. T. Grisamore, L. S. Rotolo, and G. U. Uyehara. Logical design using the Stroke function. IRE Transactions on Electronic Computers, EC-7(2):181–183, June 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222532.

Gano:1958:TGA

[GS58] J. J. Gano and G. F. Sandy. Thermistors for the gradual application of heater voltage to thermionic tubes. IRE Transactions on Electronic Computers, EC-7(1):61–64, March 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222099.

Goldberg:1965:ACL

[GS65] J. Goldberg and R. A. Short.

Antiparallel control logic. IEEE Transactions on Electronic Computers, EC-14(3):383–393, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038456.

Gamblin:1968:HST

[GT68] R. L. Gamblin and C. J. Tunis. A high-speed threshold memory element. IEEE Transactions on Computers, C-17(9):893–894, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687479.

Guffin:1962:CSL

[Guf62] Ronald M. Guffin. A computer for solving linear simultaneous equations using the residue number system. IRE Transactions on Electronic Computers, EC-11 (2):164–173, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219349.

Gunning:1958:CPI

[Gun58] William F. Gunning. Computers in process industry control. IRE Transactions on Electronic Computers, EC-7(2): 129–133, June 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222523.

Gurzi:1965:EBP

[Gur65] Karyl J. Gurzi. Estimates for best placement of voters in a triplicated logic network. IEEE Transactions on Electronic Computers, EC-14(5):711–717, October 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038557.

Hadlock:1967:FMS

[Had67] Frank Hadlock. On finding a minimal set of diagnostic tests. IEEE Transactions on Electronic Computers, EC-16(5):674–675, October 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039161.

Hagelbarger:1956:SSE

[Hag56] D. W. Hagelbarger. SEER, A SEquence Extrapolating Robot. IRE Transactions on Electronic Computers, EC-5(1):1–7, March 1956. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219783.

Hall:1969:MTR

[Hal69] K. S. Hall. Modified twisted-ring counter circuit. IEEE Transactions on Computers, C-18(6):568, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671308.

Hamilton:1958:TPG

[Ham58] Douglas J. Hamilton. A transistor pulse generator for digital systems. IRE Transactions on Electronic Computers, EC-7(3): 244–249, September 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222583.

Hamilton:1960:CBA

[Ham60] Douglas J. Hamilton. Current build-up in avalanche transistors with resistance loads. IRE Transactions on Electronic Computers, EC-9(4):456–460, December 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219884.

Hammel:1965:RLC

[Ham65a] D. Hammel. R65-54 the logic of computer arithmetic. IEEE Transactions on Electronic Computers, EC-14(4):670, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038535.

Hamming:1965:C

[Ham65b] R. W. Hamming. Contributors.

IEEE Transactions on Electronic Computers, EC-14(5):742, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038569.

Handler:1962:TMP

[Han62] Howard Handler. A technique for measuring the phase margin of an operational amplifier. IRE Transactions on Electronic Computers, EC-11(4):571, August 1962.

CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219399.

Handler:1963:MPS

[Han63a] Howard Handler. Measurement of phase shift. IEEE Transactions on Electronic Computers, EC-12(3):324–325, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037881.

Hanson:1963:TTL

[Han63b] W. H. Hanson. Ternary threshold logic. IEEE Transactions on Electronic Computers, EC-12 (3):191–197, June 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037853.

Hanson:1963:TLS

[Han63c] W. H. Hanson. Threshold-logic synthesis by algebraic methods. IEEE Transactions on Electronic Computers, EC-12(4):401–402, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037915.

Hanlon:1966:CAA

[Han66] A. G. Hanlon. Content-addressable and associative memory systems a survey. IEEE Transactions on Electronic Computers, EC-15 (4):509–521, August 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038824.

Handler:1967:MCS

[Han67] Howard Handler. Monte Carlo solution of partial differential equations using a hybrid computer. IEEE Transactions on Electronic Computers, EC-16(5):603–610, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039150.

Hannauer:1968:RDA

[Han68] G. Hannauer. R68-51 the design of an automatic patching system. IEEE Transactions on Computers, C-17(12):1179–1181, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687291.

Harris:1956:PVR

[Har56] J. N. Harris. A programmed variable-rate counter for generating the sine function. IRE Transactions on Electronic Computers, EC-5(1):21–26, March 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219788.

Harris:1957:ADM

[Har57] Bernard Harris. An algorithm for determining minimal representations of a logic function. IRE Transactions on Electronic Computers, EC-6(2):103–108, June 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221580.

Harris:1958:DCT

[Har58] J. R. Harris. Direct-coupled transistor logic circuitry. IRE Transactions on Electronic Computers, EC-7(1):2–6, March 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222088.

Hargrave:1961:MDL

[Har61a] Lee E. Hargrave. A magnetostrictive delay-line shift register. IRE Transactions on Electronic Computers, EC-10(4):702– 708, December 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219276.

Hartmanis:1961:SAP

[Har61b] J. Hartmanis. On the state assignment problem for sequential machines. I. IRE Transactions on Electronic Computers, EC-10 (2):157–165, June 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219184.

Hartmanis:1962:MAC

[Har62] Juris Hartmanis. Maximal autonomous clocks of sequential machines. IRE Transactions on Electronic Computers, EC-11(1): 83–86, February 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219325.

Harman:1963:NFC

[Har63a] M. G. Harman. A new form of cryotron logical circuitry. IEEE Transactions on Electronic Computers, EC-12(5):568–570, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037980.

Harrison:1963:ACC

[Har63b] M. A. Harrison. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(2):159–172, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037823.

Harrison:1963:NEC

[Har63c] Michael A. Harrison. The number of equivalence classes of Boolean functions under groups containing negation. IEEE Transactions on Electronic Computers, EC-12(5):559–561, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037976.

Hartmanis:1963:MPT

[Har63d] J. Hartmanis. Membership in the Professional Technical Group on Electronic Computers. IEEE Transactions on Electronic Computers, EC-12(1):60, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037782.

Hartmanis:1963:ESM

[Har63e] Juris Hartmanis. The equivalence of sequential machine models. IEEE Transactions on Electronic Computers, EC-12(1):18–19, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037763.

Harrison:1964:RUD

[Har64] Michael A. Harrison. A remark on uniform distribution. IEEE Transactions on Electronic Computers, EC-13(5):630–631, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038266.

Haring:1965:SI

[Har65a] D. R. Haring. Subject index. IEEE Transactions on Electronic Computers, EC-14(4):663–666, August 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038528.

Hartmanis:1965:TTL

[Har65b] Juris Hartmanis. Two tests for the linearity of sequential machines. IEEE Transactions on Electronic Computers, EC-14(6): 781–786, December 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038595.

Haring:1966:SAS

[Har66a] D. R. Haring. Sixth annual symposium on switching circuit theory and logical design. IEEE Transactions on Electronic Computers, EC-15(2):274, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038734.

Haring:1966:SI

[Har66b] D. R. Haring. Subject index. IEEE Transactions on Electronic Computers, EC-15(5):828–831, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038898.

Haring:1966:MTB

[Har66c] Donald R. Haring. Multi-threshold building blocks. IEEE Transactions on Electronic Computers, EC-15(4):662–663, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038843.

Haring:1966:MTT

[Har66d] Donald R. Haring. Multi-threshold threshold elements. IEEE Transactions on Electronic Computers, EC-15(1):45–65, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038666.

Hartmanis:1966:MFR

[Har66e] J. Hartmanis. Minimal feedback realizations of sequential machines. IEEE Transactions on Electronic Computers, EC-15(6): 931–933, December 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038940.

Harrison:1967:RDN

[Har67a] Michael A. Harrison. A remark on determining the number of states of a sequential machine. IEEE Transactions on Electronic Computers, EC-16(5):689–690, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039169.

Hartmanis:1967:ACC

[Har67b] J. Hartmanis. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-16(1):115–128, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038999.

Haring:1968:SRR

[Har68a] D. R. Haring. On shift-register realizations of sequential machines and finite-state universal sequential machines. IEEE Transactions on Computers, C-17(4):309–312, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687341.

Haring:1968:TIR

[Har68b] D. R. Haring. A technique for improving the reliability of certain classes of threshold elements.

IEEE Transactions on Computers, C-17(10):997–998, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687248.

Harley:1968:LTT

[Har68c] J. Harley. The linear transformer tree. IEEE Transactions on Computers, C-17(8):738–746, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687447.

Harman:1968:ADI

[Har68d] M. G. Harman. An attempt to design an improved multiplication system. IEEE Transactions on Computers, C-17(11):1090, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687268.

Harrison:1968:ESA

[Har68e] Michael A. Harrison. On equivalence of state assignments. IEEE Transactions on Computers, C-17 (1):55–57, January 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008869.

Harrison:1969:RTW

[Har69a] M. A. Harrison. R69-14 translator writing systems. IEEE Transactions on Computers, C-18(6):572–573, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671312.

Harrison:1969:RDS

[Har69b] M. A. Harrison. R69-9 deterministic stack automata and the quotient operator. IEEE Transactions on Computers, C-18(5):477, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671285.

Hastings:1966:RCB

[Has66] C. W. Hastings. R66-78 computation of the base two logarithm of binary number. IEEE Transactions on Electronic Computers, EC-15(6):956–957, December 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038956.

Haspel:1969:RCA

[Has69] C. H. Haspel. Remarks on “Comments on ’An Algorithm for Synthesis of Multiple-Output Combinational Logic”’. IEEE Transactions on Computers, C-18(9): 863, September 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671378. See [SD68,

DM68].

Hausner:1962:CPT

[Hau62a] Arthur Hausner. Correction: “Parametric Techniques for Eliminating Division and Treating Singularities in Computer Solutions of Ordinary Differential”. IRE Transactions on Electronic Computers, EC-11(4):570, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219398.

Hausner:1962:PTE

[Hau62b] Arthur Hausner. Parametric techniques for eliminating division and treating singularities in computer solutions of ordinary differential equations. IRE Transactions on Electronic Computers, EC-11(1):42–45, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219314.

Hausner:1965:ACT

[Hau65a] Arthur Hausner. Analog computer techniques for problems in complex variables. IEEE Transactions on Electronic Computers, EC-14(6):898–908, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038611.

Hausner:1965:LE

[Hau65b] Arthur Hausner. Letter to the editor. IEEE Transactions on Electronic Computers, EC-14(5):741, October 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038568.

Hausner:1965:SLE

[Hau65c] Arthur Hausner. The solution of Lagrange’s equations by analog computation. IEEE Transactions on Electronic Computers, EC-14(1):53–61, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038349.

Hauser:1967:VI

[Hau67] A. A. Hauser. Vector integrators. IEEE Transactions on Electronic Computers, EC-16(1): 80–82, February 1967. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038985.

Hausner:1969:RSP

[Hau69] A. Hausner. R69-2 the synthesis and processing of signals with discontinuities in the time domain. IEEE Transactions on Computers, C-18(1):83–84, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671129.

Hawryszkiewycz:1967:MCP

[Haw67] Igor T. Hawryszkiewycz. Microprogrammed control in problemoriented languages. IEEE Transactions on Electronic Computers, EC-16(5):652–658, October 1967.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039156.

Haynes:1952:MMM

[Hay52] M. K. Haynes. Multidimensional magnetic memory selection systems. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-1(??):25–29, December 1952. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Haynes:1961:LCU

[Hay61] John L. Haynes. Logic circuits using square-loop magnetic devices: A survey. IRE Transactions on Electronic Computers, EC-10(2):191–203, June 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219189.

Hays:1969:AR

[Hay69a] G. G. Hays. Author’s reply.

IEEE Transactions on Computers, C-18(9):862, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671377.

Hays:1969:CAD

[Hay69b] G. G. Hays. Computer-aided design: Simulation of digital design logic. IEEE Transactions on Computers, C-18(1):1–10, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&

                                        arnumber=1671113. See comment

[Mck69].

Hazeltine:1964:CIM

[Haz64] Barrett Hazeltine. Construction of irredundant multilevel switching functions. IEEE Transactions on Electronic Computers, EC-13(4):468–470, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038214.

Hazeltine:1965:EAS

[Haz65] Barrett Hazeltine. Encoding of asynchronous sequential circuits. IEEE Transactions on Electronic Computers, EC-14(5): 727–729, October 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038562.

Hoagland:1960:HDD

[HB60] A. S. Hoagland and G. C. Bacon. High density digital magnetic recording techniques. IRE Transactions on Electronic Computers, EC-9(1):2–11, March 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221597.

Howe:1968:LHT

[HC68] A. B. Howe and C. L. Coates. Logic hazards in threshold networks. IEEE Transactions on Computers, C-17(3):238–251, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687325.

Hadlock:1969:RSM

[HC69] F. O. Hadlock and C. L. Coates.

Realization of sequential machines with threshold elements. IEEE Transactions on Computers, C-18(5):428–439, May 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671275.

Haring:1967:RPM

[HD67] Donald R. Haring and Richard J.

Diephuis. A realization procedure for multithreshold threshold elements. IEEE Transactions on Electronic Computers, EC-16(6):828–835, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039199.

Hellerman:1966:CRS

[HDW66] L. Hellerman, W. L. Duda, and S. Winograd. Continuity and realizability of sequence transformations. IEEE Transactions on Electronic Computers, EC-15 (4):560–569, August 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038829.

Heath:1956:RAD

[Hea56] Harold F. Heath. Reliability of an air defense computing system: Component development. IRE Transactions on Electronic Computers, EC-5(4): 224–226, December 1956. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219956.

Heijn:1963:RSC

[Hei63] H. J. Heijn. R63-39 the Solomon computer. IEEE Transactions on Electronic Computers, EC-12(2):158, April 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037839. See comments

[SBM63].

Hellerman:1955:IIN

[Hel55] H. Hellerman. On the input impedance network error in operational amplifiers. IRE Transactions on Electronic Computers, EC-4(3):118, September 1955. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407910.

Hellerman:1963:CTV

[Hel63] Leo Hellerman. A catalog of three-variable or-invert and andinvert logical circuits. IEEE Transactions on Electronic Computers, EC-12(3):198–223, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037854.

Hellerman:1966:ASM

[Hel66a] H. Hellerman. On the average speed of a multiple-module storage system. IEEE Transactions on Electronic Computers, EC-15(4):670, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038847.

Hellerman:1966:PPA

[Hel66b] H. Hellerman. Parallel processing of algebraic expressions. IEEE Transactions on Electronic Computers, EC-15(1):82–91, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038669.

Hendrickson:1960:FHA

[Hen60] Herbert C. Hendrickson. Fast high-accuracy binary parallel addition. IRE Transactions on Electronic Computers, EC-9(4): 465–469, December 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219886.

Henry:1962:ACS

[Hen62] Eugene W. Henry. Analogcomputer simulation of a frequency detector. IRE Transactions on Electronic Computers, EC-11(1):80, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219320.

Henderson:1964:SNW

[Hen64a] Keith W. Henderson. Some notes on the Walsh functions. IEEE Transactions on Electronic Computers, EC-13(1):50–52, February 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038076.

Hennie:1964:RRF

[Hen64b] F. C. Hennie. R64-48 reduction of feedback loops in sequential circuits and carry leads in iterative networks. IEEE Transactions on Electronic Computers, EC-13(4):484, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038228.

Hennie:1966:LTM

[Hen66] F. C. Hennie. On-line Turing machine computations. IEEE Transactions on Electronic Computers, EC-15(1):35–44, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038665.

Henry:1967:RNG

[Hen67] R. F. Henry. Random noise generation by hybrid computer. IEEE Transactions on Electronic Computers, EC-16(6):872–873, December 1967. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4039209.

Hepner:1965:IMS

[Hep65] Charles F. Hepner. Improved methods of simulating time delays. IEEE Transactions on Electronic Computers, EC-14(2):239–243, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038405.

Hermann:1962:SSG

[Her62] P. J. Hermann. Simulation of steam generation in a heat exchanger. IRE Transactions on Electronic Computers, EC-11(1): 53–57, February 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219316.

Herriot:1965:RA

[Her65] J. G. Herriot. R65-26 ALGOL.

IEEE Transactions on Electronic Computers, EC-14(2):278, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038434.

Herrmann:1966:ATF

[Her66] O. Herrmann. About transfer functions with approximately constant time-delay. IEEE Transactions on Electronic Computers, EC-15(3):385, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038782.

Herzog:1967:ACC

[Her67] B. Herzog. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-16(3):381–392, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039087.

Horn:1957:SVN

[HF57] R. E. Horn and V. G. Fauque. Synthesis of vector networks. IRE Transactions on Electronic Computers, EC-6(4):261–265, December 1957. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5222037.

Hausner:1966:CAP

[HF66] A. Hausner and Cita M. Furlani. Chebyshev all-pass approximants for time-delay simulation. IEEE Transactions on Electronic Computers, EC-15(3):314–321, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038765.

Henrichon:1969:NPP

[HF69] E. G. Henrichon, Jr. and King-Sun Fu. A nonparametric partitioning procedure for pattern classification. IEEE Transactions on Computers, C-18(7):614–624, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671321.

Howe:1957:TRA

[HG57] R. M. Howe and E. G. Gilbert. Trigonometric resolution in analog computers by means of multiplier elements. IRE Transactions on Electronic Computers, EC-6(2):86–92, June 1957.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221576.

Howe:1953:SLD

[HH53] Carl E. Howe and Robert M.

Howe. Solution of linear differential equations with variable coefficients by the electronic differential analyzer. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(4):3–8, December 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407687.

Hirsch:1957:CTS

[HH57] C. J. Hirsch and F. C. Hallden.

Computing techniques for the sampling parametric computer. IRE Transactions on Electronic Computers, EC-6(2):108–119, June 1957. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221581.

Horton:1962:CMR

[HH62] W. P. Horton and L. E. Hargrave, Jr. Comment on “A Magnetostrictive Register”. IRE Transactions on Electronic Computers, EC-11(3): 410, June 1962. CODEN IRELAO. ISSN 0367-9950.

Hellerman:1968:CSU

[HH68] L. Hellerman and G. E. Hoernes. Control storage use in implementing an associative memory for a time-shared processor.

IEEE Transactions on Computers, C-17(12):1144–1151, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687283.

Highleyman:1961:AMC

[Hig61a] W. H. Highleyman. An analog method for character recognition. IRE Transactions on Electronic Computers, EC-10(3): 502–512, September 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219239.

Highleyman:1961:FCT

[Hig61b] W. H. Highleyman. Further comments on the N-tuple pattern recognition method. IRE Transactions on Electronic Computers, EC-10(1):97, March 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219164.

Highleyman:1961:NLS

[Hig61c] W. H. Highleyman. A note on linear separation. IRE Transactions on Electronic Computers, EC-10(4):777–778, December 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219287.

Highleyman:1961:NOP

[Hig61d] W. H. Highleyman. A note on optimum pattern recognition systems. IRE Transactions on Electronic Computers, EC-10 (2):287–288, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219202.

Highleyman:1963:DCR

[Hig63a] W. H. Highleyman. Data for character recognition studies. IEEE Transactions on Electronic Computers, EC-12(2):135–136, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037813.

Highleyman:1963:RAR

[Hig63b] W. H. Highleyman. R63-91 automatic reading machine for telegraph service. IEEE Transactions on Electronic Computers, EC-12(5):581–582, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037997.

Highleyman:1965:RSA

[Hig65] W. H. Highleyman. R65-58 A system for automatic recognition of handwritten words. IEEE Transactions on Electronic Computers, EC-14(4):671, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038539.

Higashi:1966:TFR

[Hig66] Paul Higashi. A thin-film rod memory for the NCR 315 RMC computer. IEEE Transactions on Electronic Computers, EC-15 (4):459–467, August 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038819.

Hilberg:1966:SMR

[Hil66] Wolfgang Hilberg. Simultaneous multiple response in associative memories and readout of the detector matrix. IEEE Transactions on Electronic Computers, EC-15(1):117–118, February 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038680.

Hinkelman:1953:UTP

[HK53] Thomas D. Hinkelman and Max H. Kraus. The UNIVAC tube program. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(3):8–12, ???? 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407733.

Highleyman:1960:CCR

[HK60] W. H. Highleyman and L. A. Kamentsky. Comments on a character recognition method of Bledsoe and Browning. IRE Transactions on Electronic Computers, EC-9(2):263, June 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219829.

Ho:1965:ALI

[HK65] Y-C. Ho and R. L. Kashyap. An algorithm for linear inequalities and its applications. IEEE Transactions on Electronic Computers, EC-14(5):683–688, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038553.

Hall:1969:ASH

[HK69] C. R. Hall and S. J. Kahne. Automated scaling for hybrid computers. IEEE Transactions on Computers, C-18(5):416–423, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671273.

Hampton:1963:HAD

[HKM63] R. Hampton, G. A. Korn, and B. Mitchell. Hybrid analog-digital random-noise generation. IEEE Transactions on Electronic Computers, EC-12(4):412–413, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037925.

Hamer:1957:MEA

[HL57] H. Hamer and A. Lupinski.

A method for evaluating amplifier phase shift at low frequencies. IRE Transactions on Electronic Computers, EC-6(3): 203, September 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222024.

Hawkins:1963:PCO

[HM63] J. K. Hawkins and C. J. Munsey. A parallel computer organization and mechanizations. IEEE Transactions on Electronic Computers, EC-12(3):251–262, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037859.

Hamming:1965:NLB

[HM65a] R. W. Hamming and W. L. Mammel. A note on the location of the binary point in a computing machine. IEEE Transactions on Electronic Computers, EC-14(2):260–261, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038414.

Hopcroft:1965:SMT

[HM65b] J. E. Hopcroft and R. L. Mattson. Synthesis of minimal threshold logic networks. IEEE Transactions on Electronic Computers, EC-14(4):552–560, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038506.

Hannigan:1967:RST

[HM67a] Joseph M. Hannigan and Charles G. Masters. Redundant system test point allocation and mission reliability estimation procedures. IEEE Transactions on Electronic Computers, EC-16(5):591–596, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039148.

Hawkins:1967:PLC

[HM67b] J. K. Hawkins and C. J. Munsey. Parallel logic with charge storage techniques. IEEE Transactions on Electronic Computers, EC-16(4):507–508, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039116.

Hilibrand:1959:SPD

[HMSG59] J. Hilibrand, C. W. Mueller, C. F.

Stocker, and R. D. Gold. Semiconductor parametric diodes in microwave computers. IRE Transactions on Electronic Computers, EC-8(3):287–297, September 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222685.

Haring:1967:TMS

[HO67] Donald R. Haring and Dennis Ohori. A tabular method for the synthesis of multithreshold threshold elements. IEEE Transactions on Electronic Computers, EC-16(2):216–220, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039032.

Hoagland:1955:LRS

[Hoa55] A. S. Hoagland. A logical reading system for nonreturn-to-zero magnetic recording. IRE Transactions on Electronic Computers, EC-4(3):93–95, September 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407906.

Hoagland:1962:RBR

[Hoa62] A. S. Hoagland. Review: Card Random Access Memory (CRAM): Functions and Use, by L. Bloom, I. Pardo, W. Keating, and E. Mayne. IRE Transactions on Electronic Computers, EC-11(3):422, June 1962. CODEN IRELAO. ISSN 0367-9950.

Hoagland:1964:ACC

[Hoa64] A. S. Hoagland. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(1):67–78, February 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038086.

Hobbs:1966:PFS

[Hob66] L. C. Hobbs. Present and future state-of-the-art in computer memories. IEEE Transactions on Electronic Computers, EC-15 (4):534–550, August 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038827.

Hockney:1962:IAG

[Hoc62] Roger Hockney. An intersection algorithm giving all irredundant normal forms from a prime implicant list. IRE Transactions on Electronic Computers, EC-11 (2):289–290, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219368.

Hollander:1961:DOS

[Hol61a] Gerhard L. Hollander. Drum organization for strobe addressing. IRE Transactions on Electronic Computers, EC-10(4):722– 729, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219280.

Holst:1961:BSC

[Hol61b] Per A. Holst. Bibliography on switching circuits and logical algebra. IRE Transactions on Electronic Computers, EC-10(4): 638–661, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219270.

Holick:1963:ANF

[Hol63a] A. Holick. Analysis of noncatastrophic failures in digital guidance systems. IEEE Transactions on Electronic Computers, EC-12(4):365–371, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037909.

Hollander:1963:EFF

[Hol63b] Gerhard L. Hollander. Effect of fanin and fan-out limitations on the value of computer circuits. IEEE Transactions on Electronic Computers, EC-12(3):317–318, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037871.

Holst:1963:DAE

[Hol63c] Per Asbjorn Holst. Dynamic accuracy and error in analog computations. IEEE Transactions on Electronic Computers, EC-12 (3):313–316, June 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037870.

Hopner:1964:HDB

[Hop64] Emil Hopner. High-density binary recording using nonsaturation techniques. IEEE Transactions on Electronic Computers, EC-13(3):255–261, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038151.

Hopcroft:1966:SI

[Hop66] J. Hopcroft. Subject index. IEEE Transactions on Electronic Computers, EC-15(3):406–409, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038788.

Hopcroft:1967:RFP

[Hop67a] J. E. Hopcroft. R67-24 formal properties of grammars. IEEE Transactions on Electronic Computers, EC-16(2):238–239, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039050.

Hopcroft:1967:RBP

[Hop67b] J. E. Hopcroft. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-16 (5):710–716, October 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039174.

Horna:1963:MLS

[Hor63] Otakar A. Horna. Majoritylogic synthesis. IEEE Transactions on Electronic Computers, EC-12(2):131–132, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037809.

Horna:1965:FME

[Hor65a] Otakar A. Horna. Figure of merit of electronic switching devices. IEEE Transactions on Electronic Computers, EC-14(4):643–646, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038522.

Horna:1965:GSM

[Hor65b] Otakar A. Horna. A geometric synthesis method of three-input majority logic networks. IEEE Transactions on Electronic Computers, EC-14(3):475–481, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038470.

Hornbuckle:1968:RIC

[Hor68] G. D. Hornbuckle. R68-13 an introduction to computer graphic terminals. IEEE Transactions on Computers, C-17(4):407, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687361.

Horst:1969:RDM

[Hor69] R. L. Horst. R69-12 design method for digital dead beat controller. IEEE Transactions on Computers, C-18(5):478, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671288.

Howe:1956:RNF

[How56] Robert M. Howe. Representation of nonlinear functions by means of operational amplifiers. IRE Transactions on Electronic Computers, EC-5(4):203–206, December 1956. CO-DEN IRELAO. ISSN 0367-9950.

Howard:1967:RAI

[How67] B. E. Howard. R67-61 APE II — A modem analog/hybrid computer for laboratory instruction. IEEE Transactions on Electronic Computers, EC-16(5):715, October 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039181.

Harris:1964:TDD

[HP64a] H. H. Harris and W. D. Pricer. Tunnel diode delay-line memory. IEEE Transactions on Electronic Computers, EC-13(3):269–272, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038153.

Homer:1964:CCS

[HP64b] E. G. Homer and W. Palmer. Comparison of computational speeds of digital differential analyzers and general purpose computers. IEEE Transactions on Electronic Computers, EC-13(3):307–308, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038165.

Hruz:1969:UTB

[Hru69] B. Hruz. Unateness test of a Boolean function and two general synthesis methods using threshold logic elements. IEEE Transactions on Computers, C-18(2): 122–131, February 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671208.

Heijn:1961:PCP

[HS61] H. J. Heijn and J. C. Selman. The Philips Computer PASCAL. IRE Transactions on Electronic Computers, EC-10 (2):175–183, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219187.

Hartmanis:1963:SFE

[HS63a] J. Hartmanis and R. E. Stearns. A study of feedback and errors in sequential machines. IEEE Transactions on Electronic Computers, EC-12(3):223–232, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037855.

Hsiao:1963:CDS

[HS63b] M. Y. Hsiao and F. F. Sellers. The carry-dependent sum adder. IEEE Transactions on Electronic Computers, EC-12(3):265–268, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037861.

Harris:1964:DAF

[HS64a] P. E. Harris and B. E. Simmons. DC accuracy in a fast boxcar circuit via a comparator. IEEE Transactions on Electronic Computers, EC-13(3):285–288, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038156.

Hsiao:1964:SPT

[HS64b] M. Y. Hsiao and K. Y. Sih. Serial-to-parallel transformation of linear-feedback shift-register circuits. IEEE Transactions on Electronic Computers, EC-13(6): 738–740, December 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038306.

Hardie:1967:DUF

[HS67a] Fred H. Hardie and Robert J.

Suhocki. Design and use of fault simulation for Saturn computer design. IEEE Transactions on Electronic Computers, EC-16(4):412–429, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039105.

Haring:1967:RSS

[HS67b] Donald R. Haring and Alfred K.

Susskind. On realizations of synchronous sequential machines. IEEE Transactions on Electronic Computers, EC-16(5):686–687, October 1967. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4039167.

Hornbuckle:1969:DSG

[HS69a] G. D. Hornbuckle and R. N.

Spann. Diagnosis of single-gate failures in combinational circuits. IEEE Transactions on Computers, C-18(3):216–220, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671227.

Huttenhoff:1969:AUC

[HS69b] J. H. Huttenhoff and R. R. Shively. Arithmetic unit of a computing element in a global, highly parallel computer. IEEE Transactions on Computers, C-18(8):695–698, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671344.

Hwa:1969:ART

[HS69c] H. R. Hwa and C. L. Sheng. An approach for the realization of threshold functions of order r. IEEE Transactions on Computers, C-18(10):923–939, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671141.

Hohn:1957:TN

[HSA57] F. E. Hohn, S. Seshu, and D. D. Aufenkamp. The theory of nets. IRE Transactions on Electronic Computers, EC-6(3): 154–161, September 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222012.

Hsiao:1962:DSU

[Hsi62] M. Y. Hsiao. A decoder scheme using magnetic-core matrices for systematic codes. IRE Transactions on Electronic Computers, EC-11(4):575–576, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219405.

Hsia:1967:SCR

[Hsi67] Yukun Hsia. Solenoid-coupled READ-ONLY memory. IEEE Transactions on Electronic Computers, EC-16(3):344–350, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039074.

Hsiao:1968:SCE

[Hsi68] M. Y. Hsiao. Single-channel error correction in an f-channel system. IEEE Transactions on Computers, C-17(10):935–943, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687238.

Hughes:1967:FFT

[Hug67] Gordon F. Hughes. Feed-forward threshold logic nets for digital switching and pattern recognition. IEEE Transactions on Electronic Computers, EC-16(4): 463–472, August 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039111.

Humphrey:1963:REA

[Hum63] W. S. Humphrey. R63-57 on the efficient assignment of internal codes to sequential machines. IEEE Transactions on Electronic Computers, EC-12(3):336, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037894.

Huskey:1954:RS

[Hus54] H. D. Huskey. Review section.

Transactions of the I.R.E. Professional Group on Electronic Computers, EC-3(1):23–39, March 1954. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Hause:1969:VLC

[HW69] A. D. Hause and D. R. Weller. A variable-length code for an incremental display system. IEEE Transactions on Computers, C-18(7):643–644, July 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671325.

Ibarra:1967:EFS

[Iba67] Oscar H. Ibarra. On the equivalence of finite-state sequential machine models. IEEE Transactions on Electronic Computers, EC-16(1):88–90, February 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038988.

Isaac:1963:DSR

[ID63] M. G. Isaac and V. T. De-Buono. Deterministic and stochastic response of linear time variable systems. IEEE Transactions on Electronic Computers, EC-12(5):532–540, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037969.

Ihm:1963:RCF

[Ihm63] P. Ihm. R63-72 cluster formation and diagnostic significance in psychiatric symptom evaluation. IEEE Transactions on Electronic Computers, EC-12(4):422, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037942.

Ishida:1965:LNU

[IS65] H. Ishida and R. M. Stewart.

A learning network using adaptive threshold elements. IEEE Transactions on Electronic Computers, EC-14(3):481–485, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038471.

Isberg:1963:RMI

[Isb63] C. A. Isberg. R63-69 multiple inputoutput link in computer systems. IEEE Transactions on Electronic Computers, EC-12(4):420, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037939.

Ide:1967:EIN

[IT67] E. R. Ide and Cyril J. Tunis. An experimental investigation of a nonsupervised adaptive algorithm. IEEE Transactions on Electronic Computers, EC-16(6):860–864, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039204.

Ito:1969:NCS

[Ito69] T. Ito. Note on a class of statistical recognition functions.

IEEE Transactions on Computers, C-18(1):76–79, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671123.

Jackson:1967:RMI

[JA67] Charles L. Jackson and Robert A. Ankerlin. A rapid method for the identification of the type of a four-variable Boolean function. IEEE Transactions on Electronic Computers, EC-16(6): 870–871, December 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039208.

Jacobsohn:1964:SFM

[Jac64a] D. Jacobsohn. A suggestion for a fast multiplier. IEEE Transactions on Electronic Computers, EC-13(6):754, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038318.

Jacobsohn:1964:SOD

[Jac64b] David Jacobsohn. A self-organizing drum. IEEE Transactions on Electronic Computers, EC-13(3):302, June 1964. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038160.

Jackson:1968:RMS

[Jac68] C. L. Jackson. R68-50 multiprogramming system performance measurement and analysis. IEEE Transactions on Computers, C-17 (11):1100, November 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687276.

Jarvis:1963:EIH

[Jar63] D. B. Jarvis. The effects of interconnections on high-speed logic circuits. IEEE Transactions on Electronic Computers, EC-12(5): 476–487, October 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037961.

Jones:1969:TSD

[JC69] J. W. Jones, Jr. and C. C. Carroll. A timed-shared digital filter realization. IEEE Transactions on Computers, C-18(11): 1027–1030, November 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671168.

Jarvis:1960:TCS

[JMW60] D. B. Jarvis, L. P. Morgan, and J. A. Weaver. Transistor current switching and routing techniques. IRE Transactions on Electronic Computers, EC-9(3): 302–308, September 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219849.

Johnson:1968:ASRa

[JO68a] D. L. Johnson and K. H. O’Keefe. The application of shift registers to secondary state assignment: Part I. IEEE Transactions on Computers, C-17(10):954–965, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687241.

Johnson:1968:ASRb

[JO68b] D. L. Johnson and K. H. O’Keefe. The application of shift registers to secondary state assignment: Part II. IEEE Transactions on Computers, C-17(10):966–977, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687242.

Johnston:1961:CMD

[Joh61] R. C. Johnston. Cryosar memory design. IRE Transactions on Electronic Computers, EC-10(4): 712–717, December 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219278.

Johnson:1963:N

[Joh63] E. O. Johnson. Notices. IEEE Transactions on Electronic Computers, EC-12(4):441, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037934.

Jones:1964:RBP

[Jon64] E. D. Jones. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-13(4):480–486, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038220.

Jones:1969:ACV

[Jon69a] R. Jones. Another comment on “A Variable Counter Design Technique”. IEEE Transactions on Computers, C-18(3): 281–282, March 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671242. See [Ren68,

Dav69b, Tae69].

Jones:1969:FAE

[Jon69b] R. M. Jones. Factors affecting the efficiency of a virtual memory. IEEE Transactions on Computers, C-18(11):1004–1008, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671163.

Josephs:1966:ESV

[Jos66] H. C. Josephs. The effect of spatial variations in voltage and temperature in emitter coupled logic systems. IEEE Transactions on Electronic Computers, EC-15(2):239–242, April 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038719.

Joseph:1967:ILS

[Jos67] Earl C. Joseph. Impact of large scale integration on aerospace computers. IEEE Transactions on Electronic Computers, EC-16(5): 558–561, October 1967. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039144.

Janoski:1966:CPA

[JSS66] R. M. Janoski, R. L. Schaefer, and J. J. Skiles. COBLOC — A program for all-digital simulation of a hybrid computer. IEEE Transactions on Electronic Computers, EC-15(1):74–82, February 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038668.

Kaenel:1961:HSA

[Kae61] R. A. Kaenel. High-speed analogto-digital converters utilizing tunnel diodes. IRE Transactions on Electronic Computers, EC-10 (2):273–284, June 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219199.

Kaenel:1962:HFT

[Kae62] R. A. Kaenel. Hysteresisfree tunnel-diode amplitude comparator. IRE Transactions on Electronic Computers, EC-11(2): 286–287, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219365.

Kahne:1968:SFC

[Kah68] S. J. Kahne. Sensitivity-function calculation in linear systems using time-shared analog integration. IEEE Transactions on Computers, C-17(4):375–379, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687349.

Kain:1967:SC

[Kai67] Richard Y. Kain. Synthesis of UP– DOWN counters. IEEE Transactions on Electronic Computers, EC-16(2):146–151, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039021.

Kalbfell:1957:EAM

[Kal57] David C. Kalbfell. An electronic analog multiplier. IRE Transactions on Electronic Computers, EC-6(2):100–103, June 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221579.

Kamat:1960:TMS

[Kam60a] D. S. Kamat. Two modulatorswitches for computer applications. IRE Transactions on Electronic Computers, EC-9(1): 123–124, March 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221611.

Kampe:1960:DGP

[Kam60b] Thomas W. Kampe. The design of a general-purpose microprogramcontrolled computer with elementary structure. IRE Transactions on Electronic Computers, EC-9(2):208–213, June 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219820.

Kamentsky:1961:STM

[Kam61] L. A. Kamentsky. The simulation of three machines which read rows of handwritten arabic numbers. IRE Transactions on Electronic Computers, EC-10(3): 489–501, September 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219238.

Karush:1955:SMS

[Kar55] William Karush. Stability of a method of smoothing in a digital control computer. IRE Transactions on Electronic Computers, EC-4(1):26–31, March 1955. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407951.

Karp:1963:RCC

[Kar63a] R. M. Karp. R63-32 computational chains and the simplification of computer programs. IEEE Transactions on Electronic Computers, EC-12(2):155, April 1963.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037832.

Karplus:1963:MPT

[Kar63b] W. J. Karplus. Membership in the Professional Technical Group on Electronic Computers. IEEE Transactions on Electronic Computers, EC-12(4):442, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037935.

Karp:1964:STS

[Kar64a] Richard M. Karp. Some techniques of state assignment for synchronous sequential machines. IEEE Transactions on Electronic Computers, EC-13(5):507–518, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038244.

Karplus:1964:HCT

[Kar64b] Walter J. Karplus. A hybrid computer technique for treating nonlinear partial differential equations. IEEE Transactions on Electronic Computers, EC-13(5): 597–605, October 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038254.

Karplus:1967:IIT

[Kar67] W. J. Karplus. 1967 index IEEE Transactions on Electronic Computers vol. EC-16. IEEE Transactions on Electronic Computers, EC-16(6):913–932, December 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039215.

Karplus:1969:RQA

[Kar69] W. J. Karplus. R69-4 quasianalogue discrete simulating mediums. IEEE Transactions on Computers, C-18(2):204, February 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671220.

Kasischke:1962:ACT

[Kas62] F. W. Kasischke. Analogcomputer technique for simulation of dynamic pressure in reentry studies. IRE Transactions on Electronic Computers, EC-11 (1):80–82, February 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219321.

Kaszerman:1963:NST

[Kas63] Philip Kaszerman. A nonlinearsummation threshold device. IEEE Transactions on Electronic Computers, EC-12(6):914–915, December 1963. CODEN IEECA8. ISSN 0367-7508. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4038041.

Kashyap:1966:SSF

[Kas66] R. L. Kashyap. Synthesis of switching functions by threshold elements. IEEE Transactions on Electronic Computers, EC-15(4):619–628, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038836.

Kateley:1966:C

[Kat66] J. Kateley. Contributors. IEEE Transactions on Electronic Computers, EC-15(3):388–389, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038785.

Kato:1968:RSA

[Kat68] M. Kato. R68-21 system architecture for large scale integration. IEEE Transactions on Computers, C-17(5):521, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687386.

Kautz:1958:UDE

[Kau58] W. H. Kautz. Unit-distance errorchecking codes. IRE Transactions on Electronic Computers, EC-7(2):179–180, June 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222529.

Kautz:1960:CWC

[Kau60] William H. Kautz. Constantweight counters and decoding trees. IRE Transactions on Electronic Computers, EC-9(2): 231–244, June 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219823.

Kautz:1961:RSS

[Kau61] William H. Kautz. The realization of symmetric switching functions with linear-input logical elements. IRE Transactions on Electronic Computers, EC-10(3): 371–378, September 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219224.

Kautz:1962:RBP

[Kau62] W. H. Kautz. Review: Some Properties of Binary Counters With Feedback, by R. C. Brigham. IRE Transactions on Electronic Computers, EC-11(3):419, June 1962. CODEN IRELAO. ISSN 0367-9950.

Kaufman:1964:ACC

[Kau64] B. A. Kaufman. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(4):487–498, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038221.

Kaufman:1966:SIH

[Kau66a] Bruce A. Kaufman. The special issue on high-speed memories. IEEE Transactions on Electronic Computers, EC-15(4):420–422, August 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038813.

Kautz:1966:ACC

[Kau66b] W. H. Kautz. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-15(3):395–406, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038787.

Kautz:1966:SAP

[Kau66c] William H. Kautz. A survey and assessment of progress in switching theory and logical design in the Soviet Union. IEEE Transactions on Electronic Computers, EC-15(2):164–204, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038713.

Kaufman:1967:FWM

[Kau67a] Alvin B. Kaufman. Ferromagnetic wire memory. IEEE Transactions on Electronic Computers, EC-16(1):86–88, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038987.

Kaupp:1967:CMT

[Kau67b] H. R. Kaupp. Characteristics of microstrip transmission lines. IEEE Transactions on Electronic Computers, EC-16(2):185–193, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039027.

Kautz:1967:CTA

[Kau67c] William H. Kautz. A cellular threshold array. IEEE Transactions on Electronic Computers, EC-16(5):680–682, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039163.

Kautz:1968:FTD

[Kau68] W. H. Kautz. Fault testing and diagnosis in combinational digital circuits. IEEE Transactions on Computers, C-17(4):352–366, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687346. See comments

[Dwy69].

Kautz:1969:AR

[Kau69a] W. H. Kautz. Author’s reply. IEEE Transactions on Computers, C-18(8):760, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671355.

Kautz:1969:CLM

[Kau69b] W. H. Kautz. Cellular logicin-memory arrays. IEEE Transactions on Computers, C-18(8): 719–727, August 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671347.

KazMin:1965:ACC

[Kaz65] A. I. Kaz’Min. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-14(5):747–757, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038571.

Kundu:1964:TMD

[KB64] P. Kundu and S. Banerji. Transistorized multiplier and divider and its applications. IEEE Transactions on Electronic Computers, EC-13(3):288–295, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038157.

Kovach:1954:AMU

[KC54] L. D. Kovach and W. Comley.

An analog multiplier using thyrite. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-3(2):42–45, June 1954. CO-DEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Kovach:1958:NTF

[KC58] L. D. Kovach and W. Comley. Nonlinear transfer functions with thyrite. IRE Transactions on Electronic Computers, EC-7(2):91–97, June 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222516.

Kovach:1960:NSS

[KC60] L. D. Kovach and W. Comley. A new, solid-state, nonlinear analog component. IRE Transactions on Electronic Computers, EC-9(4):496–503, December 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219892.

Keir:1962:DOD

[KCT62] Y. A. Keir, P. W. Cheney, and M. Tannenbaum. Division and overflow detection in residue number systems. IRE Transactions on Electronic Computers, EC-11 (4):501–507, August 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219389.

Keir:1964:APV

[Kei64] Yates A. Keir. Algebraic properties of 3-valued compositions. IEEE Transactions on Electronic Computers, EC-13(5):635–639, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038270.

Keirstead:1968:RLD

[Kei68] R. E. Keirstead. R68-8 language directed computer design. IEEE Transactions on Computers, C-17(3):298, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687335.

Keller:1962:ACR

[Kel62] Robert E. Keller. An analog computer realization of the Euclidean tools. IRE Transactions on Electronic Computers, EC-11 (4):564–570, August 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219397.

Keller:1963:RAH

[Kel63] R. E. Keller. R63-94 application of hybrid analog and digital techniques in the automatic map compilation system. IEEE Transactions on Electronic Computers, EC-12(5):582–583, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038000.

Kellerman:1968:FLN

[Kel68] E. Kellerman. A formula for logical network cost. IEEE Transactions on Computers, C-17(9): 881–884, September 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687475.

Kilburn:1962:OLS

[KELS62] T. Kilburn, D. B. G. Edwards, M. J. Lanigan, and F. H. Sumner. One-level storage system. IRE Transactions on Electronic Computers, EC-11 (2):223–235, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219356.

Kerfoot:1956:TCA

[Ker56] Branch P. Kerfoot. Transistors in current-analog computing. IRE Transactions on Electronic Computers, EC-5(2):86–93, June 1956. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219805.

Kaupe:1966:RPP

[KF66] A. F. Kaupe and C. J. Frank. R66-63 parallel processing of algebraic expressions. IEEE Transactions on Electronic Computers, EC-15(5):837–839, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038908.

Kaufman:1963:HSD

[KH63] B. A. Kaufman and J. S. Hammond. A high-speed direct-coupled magnetic memory sense amplifier employing tunnel-diode discriminators. IEEE Transactions on Electronic Computers, EC-12 (3):282–295, June 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037865.

Kriz:1963:DCM

[KI63] T. A. Kriz and K. Ishii. Directional couplers as microwave logic. IEEE Transactions on Electronic Computers, EC-12(6): 915–916, December 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038042.

Kriz:1965:MRF

[KI65] T. A. Kriz and K. Ishii. Microwave readout of ferrite memory storage. IEEE Transactions on Electronic Computers, EC-14(1):75–76, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038357.

Kriz:1966:ENM

[KI66] T. A. Kriz and T. K. Ishii. An experimental nondestructive microwave ferrite core readout using stripline. IEEE Transactions on Electronic Computers, EC-15(4):428–435, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038815.

Kiyono:1962:ADC

[KII62] T. Kiyono, K. Ikeda, and H. Ichiki.

Analog-to-digital converter utilizing an Esaki diode stack. IRE Transactions on Electronic Computers, EC-11(6):791–792, December 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219466.

Kilmer:1959:IAE

[Kil59] William L. Kilmer. An idealized over-all error-correcting digital computer having only an error-detecting combinational part. IRE Transactions on Electronic Computers, EC-8(3):321– 325, September 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222691.

Kilmer:1962:ISN

[Kil62] William Kilmer. Iterative switching networks composed of combinational cells. IRE Transactions on Electronic Computers, EC-11(2):123–131, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219344.

Kime:1966:CGN

[Kim66a] C. R. Kime. Computer group news. IEEE Transactions on Electronic Computers, EC-15(1):149, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038693.

Kime:1966:OCE

[Kim66b] Charles R. Kime. An organization for checking experiments on sequential circuits. IEEE Transactions on Electronic Computers, EC-15(1):113–115, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038678.

Kime:1967:ACC

[Kim67] C. R. Kime. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-16(4):531–543, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039130.

Kintner:1959:DPL

[Kin59] Paul M. Kintner. Dual-polarity logic as a design tool. IRE Transactions on Electronic Computers, EC-8(2):227–228, June 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219527.

King:1966:DAD

[Kin66a] R. E. King. A dual amplifier differentiator. IEEE Transactions on Electronic Computers, EC-15(1):120–121, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038682.

King:1966:AIN

[Kin66b] William F. King. Analysis of iterative NOR autonomous sequential machines. IEEE Transactions on Electronic Computers, EC-15(4):569–577, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038830.

Kirsch:1964:CIE

[Kir64a] Russell A. Kirsch. Computer interpretation of English text and picture patterns. IEEE Transactions on Electronic Computers, EC-13(4):363–376, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038199.

Kirshner:1964:RCD

[Kir64b] H. J. Kirshner. R64-33 A computer driven simulation environment for air traffic control studies. IEEE Transactions on Electronic Computers, EC-13(3):320, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038182.

Kohavi:1968:VLD

[KK68] I. Kohavi and Z. Kohavi. Variablelength distinguishing sequences and their application to the design of fault-detection experiments. IEEE Transactions on Computers, C-17(8):792–795, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687454.

Krause:1958:DAC

[KL58] C. A. Krause and R. R. Lowe. Design of AC computing amplifiers using transistors. IRE Transactions on Electronic Computers, EC-7(3):191–195, September 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222574.

Kohavi:1967:DSM

[KL67] Zvi Kohavi and Pierre Lavallee. Design of sequential machines with fault-detection capabilities. IEEE Transactions on Electronic Computers, EC-16(4):473–484, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039112.

Klein:1953:ABC

[Kle53] Rudolph J. Klein. Automatic beam current stabilization for Williams tube memories. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(4):8–11, December 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407688.

Kleinrock:1964:RAC

[Kle64] L. Kleinrock. R64-34 analysis of computing-load assignment in a MultiProcessor computer. IEEE Transactions on Electronic Computers, EC-13(3):320, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038183.

Kleinrock:1969:RSD

[Kle69] L. Kleinrock. R69-5 the statistics of discrete-event simulation. IEEE Transactions on Computers, C-18(3):303, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671245.

Klir:1966:C

[Kli66] J. Klir. Contributors. IEEE Transactions on Electronic Computers, EC-15(4):670–675, August 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038848.

Klir:1967:NBB

[Kli67] Jiri Klir. A note on the basic block diagrams of finite automata from the engineering point of view. IEEE Transactions on Electronic Computers, EC-16(2):223–224, April 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039035.

Klir:1969:RFL

[Kli69] G. J. Klir. R69-3 A formula for logical network cost. IEEE Transactions on Computers, C-18(2):204, February 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671219.

Klock:1961:FNN

[Klo61] Harold F. Klock. Functional notation for NOR and NAND networks. IRE Transactions on Electronic Computers, EC-10 (4):778, December 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219288.

Kouvaras:1968:DSS

[KLP68] N. Kouvaras, D. Lagoyannis, and L. Ponticopoulos. A digital system of simultaneous addition of several binary numbers. IEEE Transactions on Computers, C-17 (10):992–997, October 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687247.

Kautz:1968:CIA

[KLW68] W. H. Kautz, K. N. Levitt, and A. Waksman. Cellular interconnection arrays. IEEE Transactions on Computers, C-17(5):443–451, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687373.

Kovatch:1961:HEA

[KM61] G. Kovatch and W. E. Meserve.

The Hall-Effect analog multiplier. IRE Transactions on Electronic Computers, EC-10(3):512– 515, September 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219240.

Knuth:1964:SSL

[KM64a] D. E. Knuth and J. L. McNeley. SOL — A symbolic language for general purpose systems simulation. IEEE Transactions on Electronic Computers, EC-13(4):401–408, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038204.

Knuth:1964:FDS

[KM64b] Donald E. Knuth and J. L. Mc-Neley. A formal definition of SOL. IEEE Transactions on Electronic Computers, EC-13(4): 409–414, August 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038205.

Kermani:1969:IDD

[KMM69] M. B. Kermani, M. H. Mickle, and L. P. McNamee. Identification of disjunctively decomposable logic functions employing a Karnaugh map. IEEE Transactions on Computers, C-18(3):277–279, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671238.

Knoll:1967:CPR

[Kno67a] Arnold L. Knoll. Comments on“A Pattern Recognition Experiment with Near-Optimum Results” and “Pattern classification by iteratively determined linear and piecewise linear discriminant functions”. IEEE Transactions on Electronic Computers, EC-16(4):519, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039126. See [RH66,

FF67] and [DF66].

Knoll:1967:SAD

[Kno67b] Arnold L. Knoll. Spectrum analysis of digital magnetic recording waveforms. IEEE Transactions on Electronic Computers, EC-16(6):732–743, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039191.

Knowlton:1968:CHS

[Kno68] K. C. Knowlton. A combination hardware-software debugging system. IEEE Transactions on Computers, C-17(1):84–86, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008877.

Knoll:1969:ECL

[Kno69] A. L. Knoll. Experiments with “Characteristic Loci” for recognition of handprinted characters. IEEE Transactions on Computers, C-18(4):366–372, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671257.

Knudsen:1965:RAS

[Knu65a] H. K. Knudsen. R65-80 application of state variable techniques in analog computer programming. IEEE Transactions on Electronic Computers, EC-14(6):1005–1006, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038642.

Knudsen:1965:SDD

[Knu65b] Harold K. Knudsen. The scaling of digital differential analyzers. IEEE Transactions on Electronic Computers, EC-14(4):583–590, August 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038510.

Knudsen:1967:AI

[Knu67] H. K. Knudsen. Author index. IEEE Transactions on Electronic Computers, EC-16(4):547, August 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039132.

Kudielka:1966:CSF

[KO66] V. Kudielka and P. Oliva. Complete sets of functions of two and three binary variables. IEEE Transactions on Electronic Computers, EC-15(6):930–931, December 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038939.

Koester:1963:SI

[Koe63] C. J. Koester. Subject index. IEEE Transactions on Electronic Computers, EC-12(5):597–599, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037987.

Kohr:1963:MDD

[Koh63] Robert H. Kohr. A method for the determination of a differential equation model for simple nonlinear systems. IEEE Transactions on Electronic Computers, EC-12(4):394–400, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037913.

Kohavi:1964:NOM

[Koh64a] Z. Kohavi. New officers and members of the Board. IEEE Transactions on Electronic Computers, EC-13(3):185–192, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038143.

Kohavi:1964:SSA

[Koh64b] Zvi Kohavi. Secondary state assignment for sequential machines. IEEE Transactions on Electronic Computers, EC-13(3):193–203, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038144.

Kohavi:1965:AI

[Koh65a] Z. Kohavi. Author index. IEEE Transactions on Electronic Computers, EC-14(4):663, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038527.

Kohavi:1965:ROD

[Koh65b] Zvi Kohavi. Reduction of output dependency in sequential machines. IEEE Transactions on Electronic Computers, EC-14(6): 932–934, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038618.

Konigsberg:1960:DAM

[Kon60] R. L. Konigsberg. DC amplifier misalignment in computing systems. IRE Transactions on Electronic Computers, EC-9(3): 352–358, September 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219858.

Konkle:1964:SI

[Kon64] K. H. Konkle. Subject index. IEEE Transactions on Electronic Computers, EC-13(1):78–80, February 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038087.

Konkle:1968:ACP

[Kon68] K. H. Konkle. An analog comparator as a pseudo-light pen for computer displays. IEEE Transactions on Computers, C-17(1):54–55, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008868.

Kopf:1969:RSO

[Kop69] J. Kopf. R69-25 A system organization for resource allocation. IEEE Transactions on Computers, C-18(8):764, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671362.

Korn:1960:ACR

[Kor60] Granino A. Korn. Analogcomputer representation of inelastic stops. IRE Transactions on Electronic Computers, EC-9(1):121, March 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221608.

Korn:1962:EDE

[Kor62a] Granino A. Korn. Exact design equations for operational amplifiers with four-terminal computing networks. IRE Transactions on Electronic Computers, EC-11(1):82–83, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219324.

Korn:1962:NAV

[Kor62b] Granino A. Korn. Network approximations of variable time delays require caution. IRE Transactions on Electronic Computers, EC-11(1):82, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219322.

Korn:1963:POA

[Kor63a] G. A. Korn. Performance of operational amplifiers with electronic mode switching. IEEE Transactions on Electronic Computers, EC-12(3):310–312, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037869.

Korn:1963:AHS

[Kor63b] Granino A. Korn. Analog/ hybrid storage and pulse modulation. IEEE Transactions on Electronic Computers, EC-12(4): 411–412, August 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037924.

Korn:1964:EPM

[Kor64] Granino A. Korn. Enforcing Pontryagin’s maximum principle by continuous steepest descent. IEEE Transactions on Electronic Computers, EC-13(4):475–476, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038218.

Korn:1965:RPH

[Kor65] G. A. Korn. R65-41 progress of hybrid computation at united aircraft research laboratories. IEEE Transactions on Electronic Computers, EC-14(3):501, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038487.

Korn:1967:RHC

[Kor67a] G. A. Korn. R67-29 hybrid computer solutions of partialdifferential equations by Monte Carlo methods. IEEE Transactions on Electronic Computers, EC-16(2):241–242, April 1967.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039055.

Korn:1967:RSC

[Kor67b] G. A. Korn. R67-63 stochastic computing. IEEE Transactions on Electronic Computers, EC-16(5):716, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039183.

Korn:1967:G

[Kor67c] Israel Korn. On (d, k) graphs.

IEEE Transactions on Electronic Computers, EC-16(1):90, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038989.

Kostyshyn:1962:HAS

[Kos62] Bohdan Kostyshyn. A harmonic analysis of saturation recording in a magnetic medium. IRE Transactions on Electronic Computers, EC-11(2):253–263, April 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219358.

Kovatch:1963:C

[Kov63] G. Kovatch. Contents. IEEE Transactions on Electronic Computers, EC-12(6):965–974, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038052.

Kovatch:1964:N

[Kov64] G. Kovatch. Notices. IEEE Transactions on Electronic Computers, EC-13(5):670–672, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038280.

Kovach:1966:RAC

[Kov66] L. D. Kovach. R66-67 analog computer techniques for problems in complex variables. IEEE Transactions on Electronic Computers, EC-15(5):840, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038911.

Keen:1967:ODE

[KP67] Arthur W. Keen and Jacqueline L. Peters. Operational differentiators employing positive feedback. IEEE Transactions on Electronic Computers, EC-16(3):371–372, June 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039083.

Krause:1967:TDC

[Kra67] A. E. Krause. Tunnel diode-charge storage diode shift register. IEEE Transactions on Electronic Computers, EC-16(4):506–507, August 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039115.

Krishnamurthy:1963:CMD

[Kri63] E. V. Krishnamurthy. On computer multiplication and division using binary logarithms. IEEE Transactions on Electronic Computers, EC-12(3):319–320, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037874.

Krishnamurthy:1964:SAE

[Kri64] E. V. Krishnamurthy. A simple algorithm for evaluating positive values of the function x^{y}. IEEE Transactions on Electronic Computers, EC-13(1):55, February 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038080.

Krishnan:1966:TOB

[Kri66] T. Krishnan. On the threshold order of a Boolean function. IEEE Transactions on Electronic Computers, EC-15(3):369–372, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038773.

Kroll:1959:PPS

[Kro59] Norman Kroll. Properties of propagating structures with variable parameter elements. IRE Transactions on Electronic Computers, EC-8(3):307, September 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222688.

Kruy:1963:AI

[Kru63] J. F. Kruy. Author index. IEEE Transactions on Electronic Computers, EC-12(4):439, August 1963.

CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037931.

Kruy:1964:RTD

[Kru64] J. F. Kruy. R64-79 tunnel diode storage system with nondestructive read-out. IEEE Transactions on Electronic Computers, EC-13(6):761, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038330.

Kettel:1961:AAM

[KS61] E. Kettel and W. Schneider. An accurate analog multiplier and divider. IRE Transactions on Electronic Computers, EC-10 (2):269–272, June 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219198.

Kazmierczak:1963:ASP

[KS63] H. Kazmierczak and K. Steinbuch. Adaptive systems in pattern recognition. IEEE Transactions on Electronic Computers, EC-12(6):822–835, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038030.

Kella:1967:NAD

[KS67] J. Kella and A. Shani. A note on the accuracy of digital differential analyzers. IEEE Transactions on Electronic Computers, EC-16(2):230, April 1967.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039038.

Kella:1968:RCD

[KS68] J. Kella and A. Shani. On the reversibility of computations in a digital differential analyzer. IEEE Transactions on Computers, C-17(3):283–284, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687333.

Kiessling:1965:LSC

[KT65] C. E. Kiessling and C. J. Tunis. Linearly separable codes for adaptive threshold networks. IEEE Transactions on Electronic Computers, EC-14(6):935–936, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038619.

Kuck:1966:RCS

[Kuc66] D. J. Kuck. R66-59 computer software: Programming systems for digital computers. IEEE Transactions on Electronic Computers, EC-15(5):836, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038904.

Kuck:1968:IIS

[Kuc68a] D. J. Kuck. ILLIAC IV software and application programming. IEEE Transactions on Computers, C-17(8):758–770, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687449.

Kuck:1968:RGC

[Kuc68b] D. J. Kuck. R68-23 the Greenblatt chess program. IEEE Transactions on Computers, C-17(5):522, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687388.

Kudlich:1964:RGM

[Kud64] R. A. Kudlich. R64-65 the GUS multicomputer system. IEEE Transactions on Electronic Computers, EC-13(5):649, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038287.

Kump:1962:MCS

[Kum62] H. J. Kump. The magnetic configuration of stylus recording. IRE Transactions on Electronic Computers, EC-11(2): 263–273, April 1962. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219359.

Kuo:1968:SNE

[Kuo68] M. C. Y. Kuo. Solution of nonlinear equations. IEEE Transactions on Computers, C-17(9): 897–898, September 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687481. See comments

[Ogb69, Vic69b, Tho69, Tay69].

Kuo:1969:AR

[Kuo69] M. C. Y. Kuo. Author’s reply.

IEEE Transactions on Computers, C-18(12):1140–1142, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671195.

Kurek:1964:TEC

[Kur64] Nicholas B. Kurek. Transistor equivalent circuit in operational amplifier configuration. IEEE Transactions on Electronic Computers, EC-13(6):752–753, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038315.

Kaplan:1965:CAT

[KW65] K. R. Kaplan and R. O. Winder.

Chebyshev approximation and threshold functions. IEEE Transactions on Electronic Computers, EC-14(2):250–252, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038410.

Kline:1969:TLD

[KW69a] R. M. Kline and D. F. Wann.

Threshold logic design of pulsetype sequential networks. IEEE Transactions on Computers, C-18(5):459–465, May 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671279.

Koczela:1969:DHP

[KW69b] L. J. Koczela and G. Y. Wang. The design of a highly parallel computer organization. IEEE Transactions on Computers, C-18(6):520–529, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671297.

Leiner:1954:SOD

[LA54] A. L. Leiner and S. N. Alexander. System organization of the DYSEAC. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-3(1):1–10, March 1954. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Leake:1968:DSG

[LA68] R. J. Leake and H. L. Althaus. DDA scaling graph. IEEE Transactions on Computers, C-17(1):81–84, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008876.

Lambert:1959:NRM

[Lam59] L. M. Lambert. Nondestructive readout of metallic-tape computer cores. IRE Transactions on Electronic Computers, EC-8(4): 470–474, December 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222060.

Lamson:1964:DAD

[Lam64] Roger C. Lamson. A division algorithm for a digital differential analyzer. IEEE Transactions on Electronic Computers, EC-13(1):54–55, February 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038079.

Landauer:1963:BTU

[Lan63a] Walter I. Landauer. The balanced tree and its utilization in information retrieval. IEEE Transactions on Electronic Computers, EC-12(6):863–871, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038033.

Langefors:1963:DDP

[Lan63b] B. Langefors. The D21 Data Processing System by Svenska Aeroplan Aktiebolaget, Sweden. IEEE Transactions on Electronic Computers, EC-12(6):650–662, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038015.

Landsverk:1964:FCC

[Lan64a] O. Landsverk. A fast coincident current magnetic core memory. IEEE Transactions on Electronic Computers, EC-13(5):580–585, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038251.

Landweber:1964:DPP

[Lan64b] Peter S. Landweber. Decision problems of phrase-structure grammars. IEEE Transactions on Electronic Computers, EC-13(4): 354–362, August 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038198.

Langdon:1968:AAC

[Lan68] G. G. Langdon, Jr. Analysis of asynchronous circuits under different delay assumptions. IEEE Transactions on Computers, C-17(12):1131–1143, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687282.

Langdom:1969:CSM

[Lan69a] G. G. Langdom, Jr. Correction to “Subtraction by Minuend Complementation”. IEEE Transactions on Computers, C-18(8):760, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671357. See [Lan69b].

Langdon:1969:SMC

[Lan69b] G. G. Langdon, Jr. Subtraction by minuend complementation. IEEE Transactions on Computers, C-18(1):74–76, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671122. See correction

[Lan69a].

Langdon:1969:DFA

[Lan69c] O. G. Langdon, Jr. Delay-free asynchronous circuits with constrained line delays. IEEE Transactions on Computers, C-18(2): 175–181, February 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671213.

Larkin:1961:AMD

[Lar61] John P. Larkin. Axiomatic majority-decision logic. IRE Transactions on Electronic Computers, EC-10(3):530, September 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219244.

Larrowe:1966:RTM

[Lar66a] V. L. Larrowe. R66-69 A theory and method for correlation analysis of nonstationary signals. IEEE Transactions on Electronic Computers, EC-15(5):841, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038913.

Larrowe:1966:BPQ

[Lar66b] Vernon L. Larrowe. Band-pass quadrature filters. IEEE Transactions on Electronic Computers, EC-15(5):726–731, October 1966.

CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038878.

Larrowe:1969:RAC

[Lar69] V. L. Larrowe. R69-20 an analog card capacity correlator as applied to OCR. IEEE Transactions on Computers, C-18(7):669–670, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671336.

Lauer:1968:RIP

[Lau68] H. C. Lauer. R68-11 intercommunication of processors and memory. IEEE Transactions on Computers, C-17(4):405–406, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687359.

Lavallee:1965:NCL

[Lav65a] Pierre Lavallee. Nonstable cycle and level sets for linear sequential machines. IEEE Transactions on Electronic Computers, EC-14(6):957–959, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038627.

Lavallee:1965:SNG

[Lav65b] Pierre Lavallee. Some new group theoretic properties of singular linear sequential machines. IEEE Transactions on Electronic Computers, EC-14(6):

959–961, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038628.

Lawler:1961:MSC

[Law61] E. L. Lawler. Minimization of switching circuits subject to reliability conditions. IRE Transactions on Electronic Computers, EC-10(4):781–782, December 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219293. See correction

[Law62b].

Lawer:1962:EAM

[Law62a] E. L. Lawer. Electrical assemblies with a minimum number of interconnections. IRE Transactions on Electronic Computers, EC-11(1):86–88, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219327.

Lawler:1962:CMS

[Law62b] E. L. Lawler. Correction to “Minimization of Switching Circuits Subject to Reliability Conditions”. IRE Transactions on Electronic Computers, EC-11(2):289, April 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219367. See [Law61].

Lawler:1967:RAL

[Law67] E. L. Lawler. R67-8 an algorithm for linear inequalities and its applications. IEEE Transactions on Electronic Computers, EC-16(1):110, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039004.

Lawson:1968:PLO

[Law68] H. W. Lawson, Jr. Programminglanguage-oriented instruction streams. IEEE Transactions on Computers, C-17(5):476–485, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687377.

Lehman:1960:NSC

[LB60] M. Lehman and N. Burla. A note on the simultaneous-carry-generation system for high-speed adders. IRE Transactions on Electronic Computers, EC-9(4): 510, December 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219896.

Lehman:1961:STH

[LB61] M. Lehman and N. Burla. Skip techniques for high-speed carrypropagation in binary arithmetic units. IRE Transactions on Electronic Computers, EC-10(4): 691–698, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219274.

Light:1966:AAR

[LBB66] L. Light, J. Badger, and D. Barnes.

An automatic acoustic ray tracing computer. IEEE Transactions on Electronic Computers, EC-15(5):719–725, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038877.

Lewin:1965:FRC

[LBG65] M. H. Lewin, H. R. Beelitz, and J. Guarracini. Fixed resistorcard memory. IEEE Transactions on Electronic Computers, EC-14(3):428–434, June 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038462.

Libaw:1953:PDC

[LC53] William H. Libaw and Leonard J.

Craig. A photoelectric decimalcoded shaft digitizer. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(3):1–4, ???? 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407731.

Lee:1959:HSA

[LC59] R. C. Lee and F. B. Cox. A high-speed analog-digital computer for simulation. IRE Transactions on Electronic Computers, EC-8(2):186–197, June 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219520.

Lewis:1963:STM

[LC63a] P. M. Lewis and C. L. Coates. Is switching theory mathematics or engineering? IEEE Transactions on Electronic Computers, EC-12(3):320–321, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037876.

Lewis:1963:RLF

[LC63b] P. M. Lewis and C. L. Coates. Realization of logical functions by a network of threshold components with specified sensitivity. IEEE Transactions on Electronic Computers, EC-12(5):443–454, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037956.

Lippel:1957:MOC

[LE57] B. Lippel and I. J. Epstein. A method for obtaining complete digital coding chains. IRE Transactions on Electronic Computers, EC-6(2):121, June 1957. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221583.

Lechner:1963:SI

[Lec63a] R. J. Lechner. Subject index. IEEE Transactions on Electronic Computers, EC-12(2):172–174, April 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037824.

Lechner:1963:TAS

[Lec63b] Robert J. Lechner. Transformations among switching function canonical forms. IEEE Transactions on Electronic Computers, EC-12(2):129–130, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037807.

Lechner:1964:RNC

[Lec64] R. J. Lechner. R64-52 on the number of classes of [n, k] switching networks. IEEE Transactions on Electronic Computers, EC-13(4):486, August 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038232.

Lechner:1967:CBE

[Lec67] Robert J. Lechner. A correspondence between equivalence classes of switching functions and group codes. IEEE Transactions on Electronic Computers, EC-16(5):621–624, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039152.

Ledley:1959:BME

[Led59a] Robert S. Ledley. Boolean matrix equations in digital circuit design. IRE Transactions on Electronic Computers, EC-8 (2):131–139, June 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219514.

Ledley:1959:SBE

[Led59b] Robert S. Ledley. Solution to Boolean equations. IRE Transactions on Electronic Computers, EC-8(2):230, June 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219531.

Ledley:1963:RSR

[Led63] R. S. Ledley. R63-16 the storage and retrieval of physiological and medical data in a modern hospital. IEEE Transactions on Electronic Computers, EC-12(1):36, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037785.

Lee:1961:APC

[Lee61] C. Y. Lee. An algorithm for path connections and its applications. IRE Transactions on Electronic Computers, EC-10(3): 346–365, September 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219222.

Lee:1962:NSP

[Lee62a] C. Y. Lee. A note on the n th shortest path problem. IRE Transactions on Electronic Computers, EC-11(4):572–573, August 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219401.

Lee:1962:ASC

[Lee62b] Fred Lee. An automatic selfchecking and fault-locating method. IRE Transactions on Electronic Computers, EC-11(5):649–654, October 1962. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219426.

Lee:1969:SLA

[Lee69] F. F. Lee. Study of “Look-Aside” memory. IEEE Transactions on Computers, C-18(11): 1062–1064, November 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671174.

Lehman:1957:HSD

[Leh57] M. Lehman. High-speed digital multiplication. IRE Transactions on Electronic Computers, EC-6(3):204–205, September 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222025.

Lehman:1961:SMS

[Leh61] M. Lehman. Serial matrix storage systems. IRE Transactions on Electronic Computers, EC-10 (2):247–252, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219195.

Lehman:1963:MAB

[Leh63a] M. Lehman. The minimization of assimilations in binary carrystorage arithmetic units. IEEE Transactions on Electronic Computers, EC-12(4):409–410, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037921.

Lehmer:1963:RDO

[Leh63b] D. H. Lehmer. R63–17 division and overflow detection in residue number systems. IEEE Transactions on Electronic Computers, EC-12(1):36–37, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=4037786;   http://
                                        ieeexplore.ieee.org/xpl/RecentIssue.
                                        jsp?punumber=4037753;   http:
                                        //ieeexplore.ieee.org/xpl/tocresult.
                                        jsp?isnumber=4037754.

Leichner:1957:FCM

[Lei57] Gene H. Leichner. A fast circulating memory. IRE Transactions on Electronic Computers, EC-6(2):123–124, June 1957. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221587.

Leibholz:1967:SEG

[Lei67] Stephen W. Leibholz. System effectiveness as a generalization of system availability. IEEE Transactions on Electronic Computers, EC-16(4):500–505, August 1967.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039114.

Lempel:1969:SFS

[Lem69] A. Lempel. On k-stable feedback shift registers. IEEE Transactions on Computers, C-18(7):652–660, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671327.

Lehman:1963:SNG

[LEN63a] M. Lehman, R. Eshed, and Z. Netter. Sabrac — A new generation serial computer. IEEE Transactions on Electronic Computers, EC-12(6):618–628, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038013.

Lenno:1963:N

[Len63b] W. T. Lenno. Notices. IEEE Transactions on Electronic Computers, EC-12(5):601–606, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037991.

Lentz:1963:AI

[Len63c] T. Lentz. Author index. IEEE Transactions on Electronic Computers, EC-12(5):599, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037988.

Lentz:1965:RSP

[Len65] T. Lentz. R65-43 A semipermanent memory utilizing correlation addressing. IEEE Transactions on Electronic Computers, EC-14(3):501–502, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038489.

Leondes:1963:N

[Leo63a] C. T. Leondes. Notices. IEEE Transactions on Electronic Computers, EC-12(1):57–59, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037780.

Leondes:1963:RND

[Leo63b] C. T. Leondes. R63-41 A nonlinear digital optimizing program for process control systems. IEEE Transactions on Electronic Computers, EC-12(2):158, April 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037841.

Lerner:1965:C

[Ler65a] S. B. Lerner. Contributors. IEEE Transactions on Electronic Computers, EC-14(2):272–274, April 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038420.

Lerner:1965:HCA

[Ler65b] Stuart B. Lerner. Hazard correction in asynchronous sequential circuits. IEEE Transactions on Electronic Computers, EC-14 (2):265–267, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038417.

Lesh:1964:RRT

[Les64] H. F. Lesh. R64-38 the DES-1 — A real-time digital simulation computer. IEEE Transactions on Electronic Computers, EC-13(3):322, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038187.

Levien:1963:RBP

[Lev63a] R. E. Levien. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-12(2):150–158, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037822.

Levy:1963:RSS

[Lev63b] S. Levy. R63-35 A study of some self-correcting sequential networks. IEEE Transactions on Electronic Computers, EC-12(2): 156–157, April 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037835.

Levy:1963:CEA

[Lev63c] Sanford M. Levy. A counter example to an assertion of Jackson. IEEE Transactions on Electronic Computers, EC-12(5):573, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037983.

Levine:1968:CPT

[Lev68] E. Levine. On the characterizing parameters of a threshold function. IEEE Transactions on Computers, C-17(7):696–697, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687429.

Lewis:1963:RIT

[Lew63a] P. M. Lewis. R63-33 introduction to the theory of finite state machines. IEEE Transactions on Electronic Computers, EC-12(2):155–156, April 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037833.

Lewis:1963:PPD

[Lew63b] Thomas B. Lewis. Primary processor and data storage equipment for the Orbiting Astronomical Observatory. IEEE Transactions on Electronic Computers, EC-12(6):677–687, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038018.

Lewin:1964:CGN

[Lew64] M. H. Lewin. Computer group news. IEEE Transactions on Electronic Computers, EC-13(3):341, June 1964. CODEN IEECA8.

ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038175.

Lewis:1965:RPR

[Lew65a] J. B. Lewis. R65-81 the philosophy and the result: Comcor’s Ci-5000 hybrid computing system. IEEE Transactions on Electronic Computers, EC-14(6):1006, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038643.

Lewis:1965:RCL

[Lew65b] P. M. Lewis. R65-30 comments on learning and adaptive machines for pattern recognition. IEEE Transactions on Electronic Computers, EC-14(2):280, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038438.

Lewis:1966:RTS

[Lew66a] J. B. Lewis. R66-83 A time-shared hybrid simulation facility. IEEE Transactions on Electronic Computers, EC-15(6):959, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038961.

Lewis:1966:LBN

[Lew66b] P. M. Lewis. A lower bound on the number of corrections required for convergence of the single threshold gate adaptive procedure. IEEE Transactions on Electronic Computers, EC-15(6):

933–935, December 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038941.

Lewin:1967:SI

[Lew67a] M. H. Lewin. Subject index. IEEE Transactions on Electronic Computers, EC-16(2):257–259, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039046.

Lewin:1967:PEK

[Lew67b] Morton H. Lewin. Portable electronic keyboard for computer input by telephone. IEEE Transactions on Electronic Computers, EC-16(3):332–334, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039072.

Lilamand:1956:TDM

[Lil56] M. Lejet Lilamand. A timedivision multiplier. IRE Transactions on Electronic Computers, EC-5(1):26–34, March 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219789.

Lindaman:1960:TDM

[Lin60] R. Lindaman. A theorem for deriving majority-logic networks within an augmented Boolean algebra. IRE Transactions on Electronic Computers, EC-9(3): 338–342, September 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219856.

Lincoln:1963:SMC

[Lin63] A. James Lincoln. A sequential magnetic-core storage system. IEEE Transactions on Electronic Computers, EC-12(6): 917–918, December 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038043.

Ling:1966:HSB

[Lin66] Huei Ling. High speed binary parallel adder. IEEE Transactions on Electronic Computers, EC-15(5):799–802, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038889.

Lippel:1955:DCA

[Lip55] Bernard Lippel. A decimal code for analog-to-digital conversion. IRE Transactions on Electronic Computers, EC-4(4): 158–159, December 1955. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219487.

Lippel:1964:NBW

[Lip64] B. Lippel. Negative-bit-weight codes and excess codes. IEEE Transactions on Electronic Computers, EC-13(3):304–306, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038163.

Liss:1963:TUT

[Lis63] Donald Liss. A test for unate truth functions. IEEE Transactions on Electronic Computers, EC-12(4):405, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037918.

Little:1968:RMC

[Lit68] W. D. Little. R68-30 Monte Carlo solution of partial differential equations using a hybrid computer. IEEE Transactions on Computers, C-17(6):607–608, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687411.

Liu:1963:DFS

[Liu63a] C. L. Liu. Determination of the final state of an automaton whose initial state is unknown. IEEE Transactions on Electronic Computers, EC-12(6):918–921, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038044.

Liu:1963:OFA

[Liu63b] C. L. Liu. k th-order finite automaton. IEEE Transactions on Electronic Computers, EC-12(5): 470–475, October 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037960.

Liu:1964:OFA

[Liu64a] C. L. Liu. k th-order finite automaton. IEEE Transactions on Electronic Computers, EC-13(5):642, October 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038273.

Liu:1964:PAD

[Liu64b] C. N. Liu. A programmed algorithm for designing multifont character recognition logics. IEEE Transactions on Electronic Computers, EC-13(5):586–593, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038252.

Liu:1965:DMT

[Liu65] Ming-Tsan Liu. On the dualmonotonicity of threshold functions. IEEE Transactions on Electronic Computers, EC-14(4): 625–627, August 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038516.

Liu:1966:SAP

[Liu66a] C. L. Liu. Some algebraic properties of multi-threshold functions. IEEE Transactions on Electronic Computers, EC-15(6): 935–938, December 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038943.

Liu:1966:SI

[Liu66b] C. L. Liu. Subject index. IEEE Transactions on Electronic Computers, EC-15(2):288–293, April 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038736.

Liu:1966:CGN

[Liu66c] C. N. Liu. Computer group news. IEEE Transactions on Electronic Computers, EC-15(3):412, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038791.

Lynch:1960:SAH

[LK60] J. T. Lynch and J. J. Karew. System application of hybrid logic circuitry. IRE Transactions on Electronic Computers, EC-9(4): 418–423, December 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219879.

Levy:1967:SUL

[LLMS67] Saul Y. Levy, Robert J. Linhardt, Henry S. Miller, and Robert D. Sidnam. System utilization of largescale integration. IEEE Transactions on Electronic Computers, EC-16(5):562–566, October 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039145.

Lawler:1969:MCM

[LLT69] E. L. Lawler, K. N. Levitt, and J. Turner. Module clustering to minimize delay in digital networks. IEEE Transactions on Computers, C-18(1):47–57, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671117.

Leaycraft:1962:CHS

[LM62] E. C. Leaycraft and E. H. Metan.

Characteristics of a high-speed multipath core for a coincident-current memory. IRE Transactions on Electronic Computers, EC-11(3):405– 409, June 1962. CODEN IRELAO. ISSN 0367-9950.

Leiner:1954:SDS

[LNSW54] A. L. Leiner, W. A. Notz, J. L.

Smith, and A. Weinberger. System design of the SEAC and DYSEAC. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-3(2):8–23, June 1954. CO-DEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Lo:1961:STD

[Lo61] Arthur W. Lo. Some thoughts on digital components and circuit techniques. IRE Transactions on Electronic Computers, EC-10(3):416–425, September 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219230.

Lo:1963:AN

[Lo63] A. W. Lo. AFIPS news. IEEE Transactions on Electronic Computers, EC-12(4):440, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037933.

Lo:1964:C

[Lo64] A. W. Lo. Contributors. IEEE Transactions on Electronic Computers, EC-13(1):60–61, February 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038084.

Loce:1967:SI

[Loc67] K. Loce. Subject index. IEEE Transactions on Electronic Computers, EC-16(3):393–396, June 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039088.

Loev:1964:UAT

[Loe64] D. Loev. “Uniform-Access” time — A needed concept. IEEE Transactions on Electronic Computers, EC-13(6):753, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038316.

Loomis:1965:CSD

[Loo65] H. H. Loomis. Completeness of sets of delayed-logic devices. IEEE Transactions on Electronic Computers, EC-14(2):157–172, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038394.

Loomis:1966:MRA

[Loo66] H. H. Loomis. The maximum rate accumulator. IEEE Transactions on Electronic Computers, EC-15(4):628–639, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038837.

Lorens:1964:IBF

[Lor64] C. S. Lorens. Invertible Boolean functions. IEEE Transactions on Electronic Computers, EC-13(5): 529–541, October 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038246.

Lotsch:1969:MFD

[Lot69] H. K. V. Lotsch. Magneticfield design considerations for a plated-wire memory. IEEE Transactions on Computers, C-18(10): 894–899, October 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671137.

Low:1962:RBM

[Low62] P. R. Low. Review: Testing of Micrologic Elements, by R. H. Norman and R. C. Anderson. IRE Transactions on Electronic Computers, EC-11(3):422, June 1962. CODEN IRE-LAO. ISSN 0367-9950.

Lovin-Roceric:1965:IAS

[LR65] M. Lovin-Roceric. Information about the Society. IEEE Transactions on Electronic Computers, EC-14(5):763–764, October 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038576.

Lawler:1960:UPF

[LS60] E. L. Lawler and G. A. Salton. The use of parenthesis-free notation for the automatic design of switching circuits. IRE Transactions on Electronic Computers, EC-9(3):342–352, September 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219857.

Liu:1966:EIM

[LS66] C. N. Liu and G. L. Shelton. An experimental investigation of a mixed-font print recognition system. IEEE Transactions on Electronic Computers, EC-15(6):916–925, December 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038935.

Lord:1963:DLP

[LTW63] P. A. Lord, C. J. Tunis, and H. L. Witter. A delay-line push-down list. IEEE Transactions on Electronic Computers, EC-12(6):872–874, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038034.

Lucal:1959:AOD

[Luc59] Harold M. Lucal. Arithmetic operations for digital computers using a modified reflected binary code. IRE Transactions on Electronic Computers, EC-8(4): 449–458, December 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222057.

Luccio:1966:MSP

[Luc66a] F. Luccio. A method for the selection of prime implicants. IEEE Transactions on Electronic Computers, EC-15(2):205–212, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038714. See comment

[Fri67b].

Luccio:1966:RNC

[Luc66b] F. Luccio. Reduction of the number of columns in flow table minimization. IEEE Transactions on Electronic Computers, EC-15(5):803–805, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038891.

Lucas:1969:AC

[Luc69a] P. Lucas. An accumulator chip.

IEEE Transactions on Computers, C-18(2):105–114, February 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671206.

Luccio:1969:EDP

[Luc69b] F. Luccio. Extending the definition of prime compatibility classes of states in incomplete sequential machine reduction. IEEE Transactions on Computers, C-18(6):537–540, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671299.

Lukaszewicz:1963:OLD

[Luk63] L. Lukaszewicz. Outline of the logical design of the ZAM-41 computer. IEEE Transactions on Electronic Computers, EC-12(6): 609–612, December 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038011.

Lukes:1967:OEO

[Luk67] Jaroslav H. Lukes. Oscillographic examination of the operation of function generators. IEEE Transactions on Electronic Computers, EC-16(2):133–139, April 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039019.

Lundh:1963:DIL

[Lun63] Yngvar Lundh. A digital integrator for on-line signal processing. IEEE Transactions on Electronic Computers, EC-12(1):26–28, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037771.

Ladd:1959:NRT

[LW59] D. W. Ladd and E. W. Wolf. A nonreal-time simulation of SAGE tracking and BOMARC guidance. IRE Transactions on Electronic Computers, EC-8(1):36–41, March 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222759.

Loomis:1965:CSL

[LW65] H. H. Loomis and R. H. Wyman.

On complete sets of logic primitives. IEEE Transactions on Electronic Computers, EC-14(2): 173–174, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038395.

Leiner:1961:UDC

[LWCL61] A. L. Leiner, A. Weinberger, C. Coleman, and H. Loberman. Using digital computers in the design and maintenance of new computers. IRE Transactions on Electronic Computers, EC-10(4): 680–690, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219273.

Levy:1964:NTS

[LWM64] S. Y. Levy, R. O. Winder, and T. H. Mott. A note on tributary switching networks. IEEE Transactions on Electronic Computers, EC-13(2):148–151, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038112.

Lynch:1969:WCA

[Lyn69a] W. T. Lynch. Worst-case analysis of a resistor memory matrix. IEEE Transactions on Computers, C-18(10):940–942, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671142.

Lynes:1969:HSD

[Lyn69b] D. J. Lynes. High-speed DC coupled digit detector. IEEE Transactions on Computers, C-18(1):43–47, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671116.

Metropolis:1958:SDC

[MA58] N. Metropolis and R. L. Ashenhurst. Significant digit computer arithmetic. IRE Transactions on Electronic Computers, EC-7(4):265–267, December 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222657.

Maling:1963:COP

[MA63a] K. Maling and E. L. Allen. A computer organization and programming system for automated maintenance. IEEE Transactions on Electronic Computers, EC-12(6):887–895, December 1963.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038036.

Metropolis:1963:BOU

[MA63b] N. Metropolis and R. L. Ashenhurst. Basic operations in an unnormalized arithmetic system. IEEE Transactions on Electronic Computers, EC-12(6): 896–904, December 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038037.

Miller:1963:GMS

[MA63c] W. F. Miller and R. Aschenbrenner. The GUS multicomputer system. IEEE Transactions on Electronic Computers, EC-12(6):671–676, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038017.

MacSorley:1962:RBA

[Mac62] O. L. MacSorley. Review: An Algorithm for Rapid Binary Division, by J. B. Wilson and R. S. Ledley. IRE Transactions on Electronic Computers, EC-11(3):420, June 1962. CODEN IRELAO. ISSN 0367-9950.

Macsorley:1964:RSC

[Mac64] O. L. Macsorley. R64-64 the Solomon computer. IEEE Transactions on Electronic Computers, EC-13(5):649, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038286.

Macleod:1969:CHS

[Mac69] I. D. G. Macleod. Comments on “A High-Speed Algorithm for the Computer Generation of Fourier Transforms”. IEEE Transactions on Computers, C-18(2):182, February 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671214. See [And68].

Madic:1962:RBI

[Mad62] Petar Madic. Review: An Incremental Computer Technique for Solving Coordinate-Rotation Equations, by C. S. Deering and C. B. Shelman. IRE Transactions on Electronic Computers, EC-11(3): 420, June 1962. CODEN IRELAO. ISSN 0367-9950.

Maitra:1962:CSN

[Mai62] K. K. Maitra. Cascaded switching networks of two-input flexible cells. IRE Transactions on Electronic Computers, EC-11 (2):136–143, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219346.

Maitra:1963:AI

[Mai63] K. K. Maitra. Author index. IEEE Transactions on Electronic Computers, EC-12(1):54, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037776.

Majerski:1967:DOD

[Maj67] Stanislaw Majerski. On determination of optimal distributions of carry skips in adders. IEEE Transactions on Electronic Computers, EC-16(1):45–58, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038982.

Miura:1957:NDF

[MAN57] T. Miura, H. Amemiya, and T. Numakura. A new diode function generator. IRE Transactions on Electronic Computers, EC-6(2):95–100, June 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221578.

Mandelbaum:1963:OSD

[Man63a] David Mandelbaum. Optimal scheduling of disk file data transfers. IEEE Transactions on Electronic Computers, EC-12(5):551, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037972.

Mann:1963:RDR

[Man63b] W. C. Mann. R63-75 design of a repairable redundant computer. IEEE Transactions on Electronic Computers, EC-12(4):423, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037945.

Mandelbaum:1964:MED

[Man64a] David Mandelbaum. A measure of efficiency of diagnostic tests upon sequential logic. IEEE Transactions on Electronic Computers, EC-13(5):630, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038265.

Mano:1964:MDL

[Man64b] M. Morris Mano. Multilevel diode logic circuits. IEEE Transactions on Electronic Computers, EC-13(5):620–623, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038262.

Manning:1966:SI

[Man66a] E. Manning. Subject index. IEEE Transactions on Electronic Computers, EC-15(1):145–147, February 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038691.

Manning:1966:CSDa

[Man66b] Eric Manning. On computer selfdiagnosis. Part I — experimental study of a processor. IEEE Transactions on Electronic Computers, EC-15(6):873–881, December 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038930.

Manning:1966:CSDb

[Man66c] Eric Manning. On computer selfdiagnosis. Part II — generalizations and design principles. IEEE Transactions on Electronic Computers, EC-15(6):882–890, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038931.

Mandelbaum:1967:CLS

[Man67a] David Mandelbaum. A comparison of linear sequential circuits and arithmetic sequences. IEEE Transactions on Electronic Computers, EC-16(2):151–157, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039022.

Manning:1967:RBP

[Man67b] E. Manning. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-16(4):526–530, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039129.

Manske:1968:CSN

[Man68] R. A. Manske. Computor simulation of narrowband systems. IEEE Transactions on Computers, C-17(4):301–308, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687340.

Maron:1954:LDF

[Mar54] M. E. Maron. Logic, discovery, and the foundations of computing machinery. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-3(2):2–7, June 1954. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Marcus:1956:DIS

[Mar56a] Mitchell P. Marcus. The detection and identification of symmetric switching functions with the use of tables of combinations. IRE Transactions on Electronic Computers, EC-5(4):237–239, December 1956. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219959.

Marsocci:1956:EAE

[Mar56b] Velio A. Marsocci. An error analysis of electronic analog computers. IRE Transactions on Electronic Computers, EC-5(4):207–212, December 1956. CODEN IRELAO. ISSN 0367-9950.

Marcus:1957:MPD

[Mar57] Mitchell P. Marcus. Minimization of the partially-developed transfer tree. IRE Transactions on Electronic Computers, EC-6(2):92–95, June 1957. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221577.

Marill:1962:CCS

[Mar62] Thomas Marill. Computational chains and the simplification of computer programs. IRE Transactions on Electronic Computers, EC-11(2):173–180, April 1962.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219350.

Marcus:1963:CBC

[Mar63a] Mitchell P. Marcus. Cascaded binary counters with feedback. IEEE Transactions on Electronic Computers, EC-12(4):361–364, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037908.

Marcus:1963:REH

[Mar63b] Mitchell P. Marcus. Relay essential hazards. IEEE Transactions on Electronic Computers, EC-12(4):405–407, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037919.

Marczynski:1965:RMS

[Mar65a] R. Marczynski. R65-83 A magnetically scanned magnetic tape transducer. IEEE Transactions on Electronic Computers, EC-14 (6):1007, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038645.

Marzocco:1965:CRH

[Mar65b] Frank N. Marzocco. Computer recognition of handwritten first names. IEEE Transactions on Electronic Computers, EC-14(2):210–217, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038401.

Marzollo:1966:EN

[Mar66a] A. Marzollo. Editor’s notice. IEEE Transactions on Electronic Computers, EC-15(5):715, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038875.

Marzollo:1966:NNA

[Mar66b] Angelo Marzollo. Nonobservable or noncontrollable analog schemes for the solution of a special class of differential equations. IEEE Transactions on Electronic Computers, EC-15(5):716–719, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038876. See comment

[Bir67].

Marsocci:1967:DEE

[Mar67a] Velio A. Marsocci. Dynamic errors in electronic analog computers. IEEE Transactions on Electronic Computers, EC-16(2): 210–212, April 1967. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039030.

Marzollo:1967:AR

[Mar67b] Angelo Marzollo. Author’s reply. IEEE Transactions on Electronic Computers, EC-16(4):517, August 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039123.

Markson:1968:RTC

[Mar68] E. E. Markson. R68-15 trajectory computation by a hybrid computer for the Apollo midcourse abort simulation. IEEE Transactions on Computers, C-17(4):408, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687363.

Marcus:1969:RFF

[Mar69a] M. P. Marcus. S-R-T flip-flop.

IEEE Transactions on Computers, C-18(6):568–569, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671309.

Marinos:1969:FLA

[Mar69b] P. N. Marinos. Fuzzy logic and its application to switching systems. IEEE Transactions on Computers, C-18(4):343–348, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671255.

Mason:1957:LCP

[Mas57] Robert M. Mason. The logical combination of punched paper tapes. IRE Transactions on Electronic Computers, EC-6(4): 285–286, December 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222040.

Masher:1960:DDT

[Mas60] Dale P. Masher. The design of diode-transistor NOR circuits. IRE Transactions on Electronic Computers, EC-9(1):15–24, March 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221599.

Massey:1966:NFM

[Mas66] James L. Massey. Note on finitememory sequential machines. IEEE Transactions on Electronic Computers, EC-15(4):658–659, August 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038840.

Massell:1967:RAP

[Mas67] E. Massell. R67-41 an analog photoresistive multiplier. IEEE Transactions on Electronic Computers, EC-16(3):380, June 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039094. See comment

[Azg68].

Mattson:1965:AI

[Mat65] R. L. Mattson. Author index. IEEE Transactions on Electronic Computers, EC-14(1):109, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038366.

Matick:1966:FNR

[Mat66] R. Matick. Fast nondestructive read, slow write memory device using thick magnetic films. IEEE Transactions on Electronic Computers, EC-15(4):435–441, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038816.

Matthews:1968:GPN

[Mat68] S. B. Matthews. Generation of pseudorandom noise having a Gaussian spectral density. IEEE Transactions on Computers, C-17(4):382–385, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687351.

Maurer:1967:DPR

[Mau67] Harold E. Maurer. A design procedure for radiation hardening certain NOR gate logic circuits. IEEE Transactions on Electronic Computers, EC-16(4):519–522, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039127.

Mayer:1960:PTB

[May60] R. P. Mayer. Progress toward block diagram standards. IRE Transactions on Electronic Computers, EC-9(2):266–267, June 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219834.

Mays:1964:EAP

[May64] C. Hugh Mays. Effects of adaptation parameters on convergence time and tolerance for adaptive threshold elements. IEEE Transactions on Electronic Computers, EC-13(4):465–468, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038213.

Mays:1965:BMT

[May65a] C. Hugh Mays. The boundary matrix of threshold functions. IEEE Transactions on Electronic Computers, EC-14(1):65–66, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038353.

Mays:1965:RAU

[May65b] C. Hugh Mays. The relationship of algorithms used with adjustable threshold elements to differential equations. IEEE Transactions on Electronic Computers, EC-14(1):62–63, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038351.

Morrill:1952:SEM

[MB52] C. D. Morrill and R. V. Baum. A stabilized electronic multiplier. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-1(??):52–59, December 1952. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Marsocci:1962:AOA

[MC62] V. A. Marsocci and P. M. Chirlian. On the accuracy of operational amplifiers utilizing controlled superconductors. IRE Transactions on Electronic Computers, EC-11(4):577, August 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219407.

Mott:1964:NPB

[MC64] H. Mott and C. C. Carroll. Numerical procedures for Boolean function minimization. IEEE Transactions on Electronic Computers, EC-13(4):470, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038215.

Mott:1965:FC

[MC65a] H. Mott and C. C. Carroll. A functional counter. IEEE Transactions on Electronic Computers, EC-14(4):646–647, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038523.

Mott:1965:VCE

[MC65b] H. Mott and C. C. Carroll. A versatile comparator for encoding devices. IEEE Transactions on Electronic Computers, EC-14 (3):423–427, June 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038461.

Muntz:1969:OPS

[MC69] R. R. Muntz and E. G. Coffman, Jr. Optimal preemptive scheduling on two-processor systems. IEEE Transactions on Computers, C-18 (11):1014–1020, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671166.

McAvoy:1968:LSD

[McA68] T. J. McAvoy. Least-square deadtime approximations. IEEE Transactions on Computers, C-17(2): 174–178, February 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687307.

McCulloch:1957:BC

[McC57] Warren S. McCulloch. Biological computers. IRE Transactions on Electronic Computers, EC-6(3):190–192, September 1957. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222019.

McCluskey:1958:ICS

[McC58] E. J. McCluskey, Jr. Iterative combinational switching networks — general design considerations. IRE Transactions on Electronic Computers, EC-7(4): 285–291, December 1958. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222661.

McCormick:1963:IPR

[McC63a] Bruce H. McCormick. The Illinois Pattern Recognition Computer — ILLIAC III. IEEE Transactions on Electronic Computers, EC-12(6):791–813, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038028.

Mccracken:1963:C

[Mcc63b] D. D. Mccracken. Contributors. IEEE Transactions on Electronic Computers, EC-12(5): 574–576, October 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037984.

Mcclure:1967:RSO

[Mcc67a] R. M. Mcclure. R67-25 A syntaxoriented translator. IEEE Transactions on Electronic Computers, EC-16(2):239, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039051.

Mccluskey:1967:C

[Mcc67b] E. J. Mccluskey. Contributors. IEEE Transactions on Electronic Computers, EC-16(4): 523–525, August 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039128.

McCary:1968:AAS

[McC68a] R. O. McCary. An approximation to the asymmetric strip-line coupling coefficient. IEEE Transactions on Computers, C-17(6):537–542, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687397.

Mcclure:1968:RDP

[Mcc68b] R. M. Mcclure. R68-36 dataless programming. IEEE Transactions on Computers, C-17(8):818, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687460.

Mcgee:1965:RUS

[Mcg65] W. C. Mcgee. R65-55 UNISIM — A simulation program for communications networks. IEEE Transactions on Electronic Computers, EC-14(4):670, August 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038536.

McGee:1968:RVM

[McG68] W. C. McGee. R68-49 virtual memory processes and sharing in Multics. IEEE Transactions on Computers, C-17(11):1099, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687275.

McGhee:1969:SAD

[McG69] R. B. McGhee. Some aids to the detection of hazards in combinational switching circuits. IEEE Transactions on Computers, C-18(6):561–565, June 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671305.

McKeeman:1967:RER

[McK67] W. M. McKeeman. Representation error for real numbers in binary computer arithmetic. IEEE Transactions on Electronic Computers, EC-16(5):682–683, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039164.

McKeeman:1968:RMI

[McK68] W. M. McKeeman. R68-28 A microprogrammed implementation of EULER on IBM System/360 Model 30. IEEE Transactions on Computers, C-17(6):607, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687409.

Mckay:1969:CCA

[Mck69] A. R. Mckay. Comment on “Computer-Aided Design: Simulation of Digital Design Logic”. IEEE Transactions on Computers, C-18(9):862, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671376. See [Hay69b].

McLeod:1962:TYC

[McL62] John McLeod. Ten years of computer simulation. IRE Transactions on Electronic Computers, EC-11(1):2–6, February 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219308.

McNaughton:1961:UTF

[McN61] Robert McNaughton. Unate truth functions. IRE Transactions on Electronic Computers, EC-10(1):1–6, March 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219145.

McQuillan:1962:DPM

[McQ62] J. D. R. McQuillan. The design problems of a megabit storage matrix for use in a high-speed computer. IRE Transactions on Electronic Computers, EC-11(3):390– 404, June 1962. CODEN IRELAO. ISSN 0367-9950.

Munson:1968:EHD

[MDH68] J. H. Munson, R. O. Duda, and P. E. Hart. Experiments with Highleyman’s data. IEEE Transactions on Computers, C-17(4):399–401, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687355.

Martin:1967:MCS

[ME67a] David E. Martin and Gerald Estrin. Models of computational systems-cyclic to acyclic graph transformations. IEEE Transactions on Electronic Computers, EC-16(1):70–79, February 1967.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038984.

Martin:1967:EMC

[ME67b] David F. Martin and Gerald Estrin. Experiments on models of computations and systems. IEEE Transactions on Electronic Computers, EC-16(1):59–69, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038983.

Martin:1969:PLC

[ME69] D. F. Martin and G. Estrin.

Path length computations on graph models of computations. IEEE Transactions on Computers, C-18(6):530–536, June 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671298.

Meagher:1959:HIM

[Mea59] R. E. Meagher. History and introduction — microwave techniques for computers. IRE Transactions on Electronic Computers, EC-8(3):263–265, September 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222681.

Meier:1963:C

[Mei63] D. A. Meier. Contributors. IEEE Transactions on Electronic Computers, EC-12(3):325–329, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037882.

Meier:1965:RBH

[Mei65] D. A. Meier. R65-45 A 105-bit high-speed ferrite memory systemdesign and operation. IEEE Transactions on Electronic Computers, EC-14(3):502, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038491.

Meissinger:1966:RSD

[Mei66] H. F. Meissinger. R66-66 sensitivity of dynamic systems to parameters which increase the order of mathematical models. IEEE Transactions on Electronic Computers, EC-15(5):839–840, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038910.

Meisel:1967:NIS

[Mei67a] W. S. Meisel. A note on internal state minimization in incompletely specified sequential networks. IEEE Transactions on Electronic Computers, EC-16(4):508–509, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039117.

Meissinger:1967:ACA

[Mei67b] H. F. Meissinger. Annual cumulative author index. IEEE Transactions on Electronic Computers, EC-16(6):910–912, December 1967.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039214.

Meisel:1968:NVT

[Mei68a] W. S. Meisel. Nets of variablethreshold threshold elements. IEEE Transactions on Computers, C-17(7):667–676, July 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687423.

Meisel:1968:VTT

[Mei68b] W. S. Meisel. Variable-threshold threshold elements. IEEE Transactions on Computers, C-17(7): 656–667, July 1968. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1687422.

Meisel:1969:PFM

[Mei69] W. S. Meisel. Potential functions in mathematical pattern recognition. IEEE Transactions on Computers, C-18(10):911–918, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671139.

Melas:1963:RED

[Mel63] C. M. Melas. R63-104 error detection, correction and control. IEEE Transactions on Electronic Computers, EC-12(6):935, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038055.

Mellen:1966:N

[Mel66] G. E. Mellen. Notices. IEEE Transactions on Electronic Computers, EC-15(1):149–150, February 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038694.

Menger:1961:CMT

[Men61a] Karl S. Menger. Correction to “A Modulo Two Adder for Three Inputs Using a Single Tunnel Diode”. IRE Transactions on Electronic Computers, EC-10 (4):781, December 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219292. See [Men61b].

Menger:1961:MTA

[Men61b] Karl S. Menger. A modulo two adder for three inputs using a single tunnel diode. IRE Transactions on Electronic Computers, EC-10(3):530–531, September 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219245. See correction

[Men61a].

Menger:1965:CCU

[Men65] Karl S. Menger. Characterization and cardinality of universal functions. IEEE Transactions on Electronic Computers, EC-14(5):720–721, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038559.

Menger:1969:TLN

[Men69a] K. S. Menger, Jr. A transform for logic networks. IEEE Transactions on Computers, C-18 (3):241–250, March 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671230.

Menn:1969:DLC

[Men69b] B. E. Menn. Digital letdown computer for vertical guidance. IEEE Transactions on Computers, C-18(11):1065–1068, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671175.

Meo:1965:DPM

[Meo65] Angelo R. Meo. On the determination of the ps maximal implicants of a switching function. IEEE Transactions on Electronic Computers, EC-14(6):830–840, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038602.

Meo:1966:MGN

[Meo66] Angelo Raffaele Meo. Majority gate networks. IEEE Transactions on Electronic Computers, EC-15(4):606–618, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038835.

Meo:1968:MTS

[Meo68] A. R. Meo. Modular tree structures. IEEE Transactions on Computers, C-17(5):432–442, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687372.

Merwin:1956:IEM

[Mer56] Richard E. Merwin. The IBM 705 EDPM memory system. IRE Transactions on Electronic Computers, EC-5(4):219–224, December 1956. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219955.

Merrill:1964:SPT

[Mer64a] Roy Merrill. Some properties of ternary threshold logic. IEEE Transactions on Electronic Computers, EC-13(5):632–635, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038269.

Merrill:1964:IDC

[Mer64b] Roy D. Merrill. Improving digital computer performance using residue number theory. IEEE Transactions on Electronic Computers, EC-13(2):93–101, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=4038105;   http://
                                        ieeexplore.ieee.org/xpl/RecentIssue.

jsp?punumber=4037753;   http:
//ieeexplore.ieee.org/xpl/tocresult.
jsp?isnumber=4038099.

Mermelstein:1965:RRT

[Mer65] P. Mermelstein. R65-27 real time recognition of hand-drawn characters. IEEE Transactions on Electronic Computers, EC-14(2):279, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038436.

Merrill:1966:TMP

[Mer66] Roy D. Merrill. A tabular minimization procedure for ternary switching functions. IEEE Transactions on Electronic Computers, EC-15(4):578–585, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038831.

Mermelstein:1967:C

[Mer67a] P. Mermelstein. Contributors.

IEEE Transactions on Electronic Computers, EC-16(3):373–375, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039085.

Merrill:1967:STS

[Mer67b] Roy D. Merrill. Symmetric ternary switching functions: Their detection and realization with threshold logic. IEEE Transactions on Electronic Computers, EC-16 (5):624–637, October 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039153.

Minnick:1963:SLS

[MES63] R. C. Minnick, B. Elspas, and R. A. Short. Symmetric latin squares. IEEE Transactions on Electronic Computers, EC-12 (2):130–131, April 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037808.

Metze:1962:CBD

[Met62] Gernot Metze. A class of binary divisions yielding minimally represented quotients. IRE Transactions on Electronic Computers, EC-11(6):761–764, December 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219460.

Metropolis:1965:RRB

[Met65a] N. Metropolis. Review: R65-35 Binary-Compatible Signed-Digit Arithmetic by Algirdas Avisienis. IEEE Transactions on Electronic Computers, EC-14(3):499, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/iel5/4037753/4038444/
                                        04038481.pdf.

Metropolis:1965:SI

[Met65b] N. Metropolis. Subject index. IEEE Transactions on Electronic Computers, EC-14(3):517–521, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038481.

Metze:1965:MSR

[Met65c] Gernot Metze. Minimal square rooting. IEEE Transactions on Electronic Computers, EC-14 (2):181–185, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038397.

Mezei:1961:MCE

[Mez61] J. E. Mezei. Minimal characterizing experiments for finite memory automata. IRE Transactions on Electronic Computers, EC-10(2):288, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219203.

Mow:1968:ARM

[MF68a] Chao-Wei Mow and King-Sun Fu.

An approach for the realization of multithreshold threshold elements. IEEE Transactions on Computers, C-17(1):32–46, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008866.

Mow:1968:GSD

[MF68b] Chao-Wei Mow and King-Sun Fu. Generation of self-dual and self-complementary dual functions. IEEE Transactions on Computers, C-17(1):57–66, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008870.

Mow:1968:ITC

[MF68c] Chao-Wei Mow and King-Sun Fu. Input tolerance considerations for multithreshold threshold elements. IEEE Transactions on Computers, C-17(1):46–54, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008867.

Mow:1969:LFT

[MF69] Chao-Wei Mow and King-Sun Fu. Loop-free threshold element structures. IEEE Transactions on Computers, C-18(3):257–267, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671232.

Mattson:1963:EIC

[MFF63] R. L. Mattson, O. Firschein, and M. Fischler. An experimental investigation of a class of pattern recognition synthesis algorithms. IEEE Transactions on Electronic Computers, EC-12(3):300–306, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037867.

Marill:1960:SRF

[MG60] T. Marill and D. M. Green. Statistical recognition functions and the design of pattern recognizers. IRE Transactions on Electronic Computers, EC-9(4):

472–477, December 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219888.

Mucciardi:1966:EPR

[MG66] A. N. Mucciardi and E. E. Gose.

Evolutionary pattern recognition in incomplete nonlinear multithreshold networks. IEEE Transactions on Electronic Computers, EC-15(2):257–261, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038727.

Mayer:1954:ODF

[MGN54] M. A. Mayer, B. M. Gordon, and R. N. Nicola. An operational digital feedback divider. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-3(1): 17–20, March 1954. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Maughmer:1958:SRP

[MH58] J. M. Maughmer and H. D. Huskey.

A study of refill phenomena in Williams’ tube memories. IRE Transactions on Electronic Computers, EC-7(1):23–31, March 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222091.

Miyata:1959:RRS

[MH59] J. J. Miyata and R. R. Hartel. The recording and reproduction of signals on magnetic medium using saturationtype recording. IRE Transactions on Electronic Computers, EC-8(2):159–169, June 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219517.

Miiller:1961:MGA

[Mii61] H. S. Miiller. Majority gates applied to simultaneous comparators. IRE Transactions on Electronic Computers, EC-10 (1):94–95, March 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219160.

Miiller:1964:RMR

[Mii64] H. S. Miiller. Resolving multiple responses in an associative memory. IEEE Transactions on Electronic Computers, EC-13(5):614–616, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038260.

Mikhailov:1965:AI

[Mik65a] N. N. Mikhailov. Author index. IEEE Transactions on Electronic Computers, EC-14(5):761, October 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038573.

Mikhailov:1965:SI

[Mik65b] N. N. Mikhailov. Subject index. IEEE Transactions on Electronic Computers, EC-14(5):

758–761, October 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038572.

Milch:1955:BSE

[Mil55] Alfred Milch. Bit storage via electro-optical feedback. IRE Transactions on Electronic Computers, EC-4(4):136–144, December 1955. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219483.

Miller:1957:NRM

[Mil57] George A. Miller. A note on the remarkable memory of man. IRE Transactions on Electronic Computers, EC-6(3):194– 195, September 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222021.

Miller:1958:FAS

[Mil58] Raymond E. Miller. Formal analysis and synthesis of bilateral switching networks. IRE Transactions on Electronic Computers, EC-7(3):231–244, September 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222582.

Miller:1965:AI

[Mil65] S. W. Miller. Author index. IEEE Transactions on Electronic Computers, EC-14(3):517, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038480.

Miller:1966:RBP

[Mil66] R. E. Miller. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-15(2):265–271, April 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038731.

Miller:1967:RAV

[Mil67] Edwin H. Miller. Reliability aspects of the variable instruction computer. IEEE Transactions on Electronic Computers, EC-16(5):596–602, October 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039149.

Minnick:1961:LIL

[Min61] Robert C. Minnick. Linearinput logic. IRE Transactions on Electronic Computers, EC-10(1):6–16, March 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219146.

Minnick:1962:CUS

[Min62] R. C. Minnick. Comments on “The Use of the Simplex Algorithm in the Mechanization of Boolean Switching Functions by Means of Magnetic Cores”. IRE Transactions on Electronic Computers, EC-11(4):573, August 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219402. See [Ein61,

Ein62].

Minnick:1963:RBP

[Min63] R. C. Minnick. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-12 (5):577–583, October 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037985.

Minnick:1964:CCL

[Min64] R. C. Minnick. Cutpoint cellular logic. IEEE Transactions on Electronic Computers, EC-13(6):685–698, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038300.

Minnick:1969:BDC

[Min69] R. C. Minnick. B69-1 digital computer system principles. IEEE Transactions on Computers, C-18(12):1146, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671199.

Misiurewicz:1966:CCM

[Mis66] Piotr Misiurewicz. Comment on “Counting with Majority-Logic Networks”. IEEE Transactions on Electronic Computers, EC-15(2):262, April 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038729. See [Pri65].

Mitchell:1962:RBM

[Mit62a] E. N. Mitchell. Review: Cryosar Memory Design, by R. C. Johnston. IRE Transactions on Electronic Computers, EC-11(3):422, June 1962. CODEN IRELAO. ISSN 0367-9950.

Mitchell:1962:CMD

[Mit62b] John N. Mitchell. Computer multiplication and division using binary logarithms. IRE Transactions on Electronic Computers, EC-11(4):512–517, August 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219391.

Miura:1966:REA

[Miu66] T. Miura. R66-82 error analysis of hybrid computer systems. IEEE Transactions on Electronic Computers, EC-15(6):959, December 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038960.

Miyata:1963:RAL

[Miy63] Fusachika Miyata. Realization of arbitrary logical functions using majority elements. IEEE Transactions on Electronic Computers, EC-12(3):183–191, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037852. See comment

[Miy64a].

Miyata:1964:CRA

[Miy64a] Fusachika Miyata. Comment on “Realization of Arbitrary Logical Functions Using Majority Elements”. IEEE Transactions on Electronic Computers, EC-13(2):148, April 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038111. See [Miy63].

Miyata:1964:EMC

[Miy64b] Fusachika Miyata. An extension of the method of Cohn and Lindaman. IEEE Transactions on Electronic Computers, EC-13(5):625–629, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038264.

Myhill:1961:SWR

[MK61] J. Myhill and W. H. Kautz.

On the size of weights required for linear-input switching functions. IRE Transactions on Electronic Computers, EC-10 (2):288–290, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219204.

Mine:1967:BPC

[MK67] Hisashi Mine and Yoshiaki Koga.

Basic properties and a construction method for fail-safe logical systems. IEEE Transactions on Electronic Computers, EC-16 (3):282–289, June 1967. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039066. See comments

[Fos68b].

Mine:1968:AR

[MK68] H. Mine and Y. Koga. Authors’ reply. IEEE Transactions on Computers, C-17(7):702, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687434.

Meisel:1969:HAS

[MK69] W. S. Meisel and R. S. Kashef. Hazards in asynchronous sequential circuits. IEEE Transactions on Computers, C-18(8):752–759, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671353.

Morgan:1961:DEA

[ML61] M. L. Morgan and J. C. Looney. Design of the ESIAC(R) algebraic computer. IRE Transactions on Electronic Computers, EC-10(3):524–529, September 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219242.

Mackay:1955:TC

[MM55] R. S. Mackay and R. MacIntyre. Ternary counters. IRE Transactions on Electronic Computers, EC-4(4):144–149, December 1955. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219484.

Mitchell:1966:CGN

[MMB66a] E. E. L. Mitchell, J. B. Mawson, and J. Bulger. Computer Group News Editor. IEEE Transactions on Electronic Computers, EC-15(3):303, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038763.

Mitchell:1966:GHS

[MMB66b] E. E. L. Mitchell, J. B. Mawson, and J. Bulger. A generalized hybrid simulation for an aerospace vehicle. IEEE Transactions on Electronic Computers, EC-15(3):304–313, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038764.

Martin:1960:ARP

[MMR60] R. J. Martin, N. A. Masnari, and J. E. Rowe. Analog representation of Poisson’s equation in two dimensions. IRE Transactions on Electronic Computers, EC-9(4):490–496, December 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219891.

Maeda:1966:WWM

[MMT66] H. Maeda, A. Matsushita, and M. Takashima. Woven wire memory for NDRO system. IEEE Transactions on Electronic Computers, EC-15(4):442–451, August 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038817.

Miura:1958:TCC

[MN58] T. Miura and M. Nagata. Theoretical consideration of computing errors of a slow type electronic analog computer. IRE Transactions on Electronic Computers, EC-7(4):306–312, December 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222664.

Mina:1961:SWG

[MN61] K. V. Mina and E. E. Newhall. A straightforward way of generating all Boolean functions of N variables using a single magnetic circuit. IRE Transactions on Electronic Computers, EC-10 (2):151–156, June 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219183.

Martin:1957:PMS

[MO57] W. L. Martin and S. R. Olson. PGEC membership survey. IRE Transactions on Electronic Computers, EC-6(1):49–55, March 1957. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221564.

Margolis:1963:RTV

[MO63] S. G. Margolis and J. J. O’Donnell. Rigorous treatments of variable time delays. IEEE Transactions on Electronic Computers, EC-12 (3):307–309, June 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037868.

Maxwell:1965:MDN

[MO65] L. M. Maxwell and E. E. Olander.

A method of determining the number of vertices contained in the SC network corresponding to a given SC transmission function. IEEE Transactions on Electronic Computers, EC-14(5):721–723, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038560.

Morgan:1959:BDM

[Mor59] Walter L. Morgan. Bibliography of digital magnetic circuits and materials. IRE Transactions on Electronic Computers, EC-8 (2):148–158, June 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219516.

Morreale:1967:PLA

[Mor67] Eugenio Morreale. Partitioned list algorithms for prime implicant determination from canonical forms. IEEE Transactions on Electronic Computers, EC-16 (5):611–620, October 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039151.

Mott:1960:DIN

[Mot60] Thomas H. Mott. Determination of the irredundant normal forms of a truth function by iterated consensus of the prime implicants. IRE Transactions on Electronic Computers, EC-9(2):245–252, June 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219824.

Mowle:1966:RBC

[Mow66] Frederic J. Mowle. Relations between P^{n} cycles and stable feedback shift registers. IEEE Transactions on Electronic Computers, EC-15(3):375–378, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038776.

Mowle:1969:RPP

[Mow69] F. J. Mowle. Readily programmable procedures for the analysis of nonlinear feedback shift registers. IEEE Transactions on Computers, C-18(9):824–829, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671372.

Mileto:1964:AVQ

[MP64] F. Mileto and G. Putzolu. Average values of quantities appearing in Boolean function minimization. IEEE Transactions on Electronic Computers, EC-13(2):87–92, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038104.

Mileto:1965:AVQ

[MP65] F. Mileto and G. Putzolu. Average values of quantities appearing in multiple output Boolean minimization. IEEE Transactions on Electronic Computers, EC-14(4):542–552, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038505.

Madich:1959:URD

[MPP59] P. Madich, J. Petrich, and N. Parezanovich. The use of a repetitive differential analyzer for finding roots of polynomial equations. IRE Transactions on Electronic Computers, EC-8 (2):182–185, June 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219519.

Munson:1959:ORS

[MR59] J. K. Munson and A. I. Rubin.

Optimization by random search on the analog computer. IRE Transactions on Electronic Computers, EC-8(2):200–203, June 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219522.

Martin:1965:SHS

[MR65] A. R. Martin and A. B. Rosenstein. A shiftrix for high-speed multiplication. IEEE Transactions on Electronic Computers, EC-14(4):639–643, August 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038521.

Mcghee:1966:SI

[MR66] R. B. Mcghee and I. S. Reed. Subject index. IEEE Transactions on Electronic Computers, EC-15(4):696–700, August 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038852.

Maurer:1968:HGC

[MR68] H. E. Maurer and R. C. Ricci. Horizons in guidance computer component technology. IEEE Transactions on Computers, C-17(7):621–634, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687418.

Mathis:1960:CPS

[MRS60] V. P. Mathis, H. Raillard, and J. J. Suran. Comparative performance of saturating and current-clamped high-frequency pulse circuits. IRE Transactions on Electronic Computers, EC-9(2):175, June 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219816.

Marcovitz:1958:ADR

[MS58] M. W. Marcovitz and E. Seif. Analytical design of resistorcoupled transistor logical circuits. IRE Transactions on Electronic Computers, EC-7(2): 109–119, June 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222519.

Melmed:1959:DSM

[MS59] A. Melmed and R. Shevlin.

Diode-steered magnetic-core memory. IRE Transactions on Electronic Computers, EC-8(4):474– 478, December 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222061.

Muth:1963:MOE

[MS63] V. O. Muth and A. K. Scidmore. A memory organization for an elementary list-processing computer. IEEE Transactions on Electronic Computers, EC-12(3):262–265, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037860.

Morrison:1966:CSF

[MS66] J. R. Morrison and D. E. Speliotis. Cobalt-substituted γ-Fe_{2}O_{3} as a high-density recording tape. IEEE Transactions on Electronic Computers, EC-15(5):782–793, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038886.

Massey:1968:ILS

[MS68] J. L. Massey and M. K. Sain.

Inverses of linear sequential circuits. IEEE Transactions on Computers, C-17(4):330–337, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687344.

Marcovitz:1969:IAS

[MS69] A. B. Marcovitz and C. M. Shub. An improved algorithm for the simplification of switching functions using unique identifiers on a Karnaugh map. IEEE Transactions on Computers, C-18(4):376–378, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671259.

Muroga:1959:PDC

[MT59] S. Muroga and K. Takashima. The Parametron Digital Computer MUSASINO-1. IRE Transactions on Electronic Computers, EC-8(3):308–316, September 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222689.

Muroga:1966:LBN

[MT66] S. Muroga and I. Toda. Lower bound of the number of threshold functions. IEEE Transactions on Electronic Computers, EC-15(5):805–806, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038892.

Marquitz:1968:IAC

[MT68] W. T. Marquitz and Y. Tokad.

On improving the analog computer solutions of linear systems. IEEE Transactions on Computers, C-17(3):268–270, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687328.

Miura:1967:HCS

[MTI67] Takeo Miura, Junji Tsuda, and Junzo Iwata. Hybrid computer solution of optimal control problems by the maximum principle. IEEE Transactions on Electronic Computers, EC-16(5):666–670, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039158. See comments

[Ros68].

Maki:1969:GDE

[MTS69] G. K. Maki, J. H. Tracey, and R. J.

Smith II. Generation of design equations in asynchronous sequential circuits. IEEE Transactions on Computers, C-18(5):467–472, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671281.

McCluskey:1959:NNI

[MU59] E. J. McCluskey and S. H. Unger.

A note on the number of internal variable assignments for sequential switching circuits. IRE Transactions on Electronic Computers, EC-8(4):439–440, December 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222055.

Mukhopadhyay:1963:DTP

[Muk63] Amarendra Mukhopadhyay. Detection of total or partial symmetry of a switching function with the use of decomposition charts. IEEE Transactions on Electronic Computers, EC-12(5):553–557, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037974.

Mukhopadhyay:1966:STS

[Muk66] A. Mukhopadhyay. Symmetric ternary switching functions. IEEE Transactions on Electronic Computers, EC-15(5):731–739, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038879.

Mukhopadhyay:1969:RUL

[Muk69a] A. Mukhopadhyay. R69-24 universal logic circuits and their modular realizations. IEEE Transactions on Computers, C-18(8):763–764, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671361.

Mukhopadhyay:1969:UCL

[Muk69b] A. Mukhopadhyay. Unate cellular logic. IEEE Transactions on Computers, C-18(2):114–121, February 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671207.

Muller:1954:ABA

[Mul54] D. E. Muller. Application of Boolean algebra to switching circuit design and to error detection. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-3(3):6–12, September 1954. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Muller:1956:CES

[Mul56] David E. Muller. Complexity in electronic switching circuits. IRE Transactions on Electronic Computers, EC-5(1):15–19, March 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219786.

Muller:1959:TTS

[Mul59a] D. E. Muller. Treatment of transition signals in electronic switching circuits by algebraic methods. IRE Transactions on Electronic Computers, EC-8(3): 401, September 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222702.

Mulligan:1959:FMS

[Mul59b] J. H. Mulligan. A figure of merit for single-pass data recording systems. IRE Transactions on Electronic Computers, EC-8 (1):48–54, March 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222761.

Mullin:1959:CRS

[Mul59c] A. A. Mullin. Combinational relay switching fields. IRE Transactions on Electronic Computers, EC-8(4):499, December 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222067.

Mullin:1959:SDS

[Mul59d] A. A. Mullin. Self-dual symmetric switching functions with a certain a-number constraint. IRE Transactions on Electronic Computers, EC-8(4):498–499, December 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222066. See correction

[Mul60].

Mullin:1960:CSD

[Mul60] A. A. Mullin. Correction to “Self-Dual Symmetric Switching Functions with a Certain a-Number Constraint”. IRE Transactions on Electronic Computers, EC-9(2):266, June 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219833. See [Mul59d].

Mullin:1963:N

[Mul63] A. A. Mullin. Notices. IEEE Transactions on Electronic Computers, EC-12(2):177–179, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037828.

Mullery:1964:POM

[Mul64] A. P. Mullery. A procedure-oriented machine language. IEEE Transactions on Electronic Computers, EC-13(4):449–455, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038210.

Munson:1967:SI

[Mun67] J. K. Munson. Subject index. IEEE Transactions on Electronic Computers, EC-16(4):544–547, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039131.

Murray:1955:VBS

[Mur55] D. B. Murray. A variable binary scaler. IRE Transactions on Electronic Computers, EC-4(2):70–74, June 1955. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407895.

Muroga:1965:GAS

[Mur65a] Saburo Muroga. Generation and asymmetry of self-dual threshold functions. IEEE Transactions on Electronic Computers, EC-14 (2):125–136, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038391.

Muroga:1965:LBN

[Mur65b] Saburo Muroga. Lower bounds of the number of threshold functions and a maximum weight. IEEE Transactions on Electronic Computers, EC-14(2):136–148, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038392.

Murphy:1965:RDP

[Mur65c] J. E. Murphy. R65-27 the Data Processing System Simulator (DPSS) (SP-1299/000/01). IEEE Transactions on Electronic Computers, EC-14(2):278–279, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038435.

Murrell:1968:RCS

[Mur68] T. A. Murrell. R68-22 current status of large scale integration technology. IEEE Transactions on Computers, C-17(5):521–522, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687387.

Malthaner:1955:CFM

[MV55] W. A. Malthaner and H. E. Vaughan. Control features of a magnetic-drum telephone office. IRE Transactions on Electronic Computers, EC-4(1):21–26, March 1955. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5407950.

Miller:1961:ICC

[MW61] K. S. Miller and J. B. Walsh.

Initial conditions in computer simulation. IRE Transactions on Electronic Computers, EC-10 (1):78–80, March 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219156.

Miller:1962:MLS

[MW62] H. S. Miller and R. O. Winder.

Majority-logic synthesis by geometric methods. IRE Transactions on Electronic Computers, EC-11(1):89–90, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219329.

Miller:1964:NTE

[MW64] R. E. Miller and S. Winograd.

On the number of transitions entering the states of a finite automaton. IEEE Transactions on Electronic Computers, EC-13(4): 463–464, August 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038212.

Majerski:1967:NGB

[MW67] Stanislaw Majerski and Michal Wiweger. NOR-gate binary adder with carry completion detection. IEEE Transactions on Electronic Computers, EC-16(1):90–92, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038991.

Miller:1968:SDC

[MW68] J. C. Miller and C. M. Wine. A simple display for characters and graphics. IEEE Transactions on Computers, C-17(5):470–475, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687376.

Mitchell:1968:DLL

[MWR68] R. H. Mitchell, T. Williams, and W. D. Ryan. A delay line and logic circuits utilizing charge-storage subharmonic parametric oscillators. IEEE Transactions on Computers, C-17(11):1037–1043, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687261.

McNaughton:1960:RES

[MY60] R. McNaughton and H. Yamada. Regular expressions and state graphs for automata. IRE Transactions on Electronic Computers, EC-9(1):39–47, March 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221603.

MacPherson:1961:SSC

[MY61] D. H. MacPherson and R. K. York. Semipermanent storage by capacitive coupling. IRE Transactions on Electronic Computers, EC-10(3):446–451, September 1961.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219233.

Myers:1957:EOT

[Mye57] George H. Myers. Extending the operational time of analog integrators. IRE Transactions on Electronic Computers, EC-6 (1):34–35, March 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221560.

Nadler:1963:ADC

[Nad63] Morton Nadler. An analogdigital character recognition system. IEEE Transactions on Electronic Computers, EC-12(6): 814–821, December 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038029.

Nadler:1965:RCR

[Nad65] M. Nadler. R65-70 the computer recognition of handwritten first names. IEEE Transactions on Electronic Computers, EC-14(5):746, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038580.

Nagy:1963:SAM

[Nag63] George Nagy. A survey of analog memory devices. IEEE Transactions on Electronic Computers, EC-12(4):388–393, August 1963.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037912.

Nambiar:1964:NWF

[Nam64] K. K. Nambiar. A note on the Walsh functions. IEEE Transactions on Electronic Computers, EC-13(5):631–632, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038268.

Narud:1960:SEP

[Nar60] Jan A. Narud. A secondaryemission pulse circuit, its analysis and application. IRE Transactions on Electronic Computers, EC-9(4):439–451, December 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219882.

Narasimhan:1961:MIS

[Nar61] R. Narasimhan. Minimizing incompletely specified sequential switching functions. IRE Transactions on Electronic Computers, EC-10(3):531–532, September 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219246.

Nash:1956:REC

[Nas56] John P. Nash. Review of electronic computer progress during 1955. IRE Transactions on Electronic Computers, EC-5(1):43–47, March 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219793.

Naslin:1967:ACC

[Nas67] P. Naslin. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-16(2):243–256, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039045.

Nathan:1957:DAD

[Nat57] Amos Nathan. Dynamic accuracy as a design criterion of linear electronic-analog differential analyzers. IRE Transactions on Electronic Computers, EC-6(2):74–86, June 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221575.

Nathan:1958:CEM

[Nat58] Amos Nathan. Computing and error matrices in linear differential analyzers. IRE Transactions on Electronic Computers, EC-7(1):32–33, March 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222092.

Nathan:1963:LNI

[Nat63] Amos Nathan. Linear and nonlinear interpolators. IEEE Transactions on Electronic Computers, EC-12(5):526–532, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037968.

Nathan:1964:CMG

[Nat64] Amos Nathan. Comments on “A Method of Generating Functions of Several Variables”. IEEE Transactions on Electronic Computers, EC-13(1):55–56, February 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038081. See [Wil63].

Nathan:1965:CM

[Nat65] Amos Nathan. The cascade multiplier. IEEE Transactions on Electronic Computers, EC-14 (2):243–247, April 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038406.

Nichols:1965:SAC

[NB65] A. J. Nichols and A. J. Bernstein. State assignments in combinational networks. IEEE Transactions on Electronic Computers, EC-14(3):343–349, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038451.

Neff:1960:EDL

[NBC60] G. W. Neff, S. A. Butler, and D. L. Critchlow. Esaki diode logic circuits. IRE Transactions on Electronic Computers, EC-9(4): 423–429, December 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219880.

Necula:1967:NPD

[Nec67] N. N. Necula. A numerical procedure for determination of the prime implicants of a Boolean function. IEEE Transactions on Electronic Computers, EC-16(5): 687–689, October 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039168.

Necula:1968:AMT

[Nec68a] N. N. Necula. An algorithm for multithreshold threshold element synthesis. IEEE Transactions on Computers, C-17(10):978–985, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687243.

Necula:1968:AAA

[Nec68b] N. N. Necula. An algorithm for the automatic approximate minimization of Boolean functions. IEEE Transactions on Computers, C-17(8):770–782, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687450.

Negrin:1964:SPT

[Neg64] Alan E. Negrin. Synthesis of practical three-input majority logic networks. IEEE Transactions on Electronic Computers, EC-13 (3):296–299, June 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038158.

Nelson:1954:ATU

[Nel54] E. C. Nelson. An algebraic theory for use in digital computer design. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-3(3):12–21, September 1954. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Nelson:1963:FET

[Nel63] Don J. Nelson. A fundamental error theory for analog computers. IEEE Transactions on Electronic Computers, EC-12(5):541–550, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037970. See comment

[Ott64].

Nelson:1964:RHA

[Nel64] D. J. Nelson. R64-76 A hybrid analog-digital parameter optimizer for Astrac II. IEEE Transactions on Electronic Computers, EC-13(6):760, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038327.

Netherwood:1958:LMD

[Net58] Douglas B. Netherwood. Logical machine design: A selected bibliography. IRE Transactions on Electronic Computers, EC-7 (2):155–178, June 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222528.

Netherwood:1959:LMD

[Net59a] Douglas B. Netherwood. Logical machine design II: A selected bibliography. IRE Transactions on Electronic Computers, EC-8(3):367–380, September 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222698.

Netherwood:1959:MSM

[Net59b] Douglas B. Netherwood. Minimal sequential machines. IRE Transactions on Electronic Computers, EC-8(3):339–345, September 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222695.

Neumann:1962:EDC

[Neu62a] Peter G. Neumann. Encoding and decoding for cyclic permutation codes. IRE Transactions on Electronic Computers, EC-11 (4):507–511, August 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219390.

Neumann:1962:LDN

[Neu62b] Peter G. Neumann. On the logical design of noiseless load-sharing matrix switches. IRE Transactions on Electronic Computers, EC-11(3): 369–373, June 1962. CODEN IRE-LAO. ISSN 0367-9950.

Newman:1964:RPP

[New64] E. G. Newman. R64-66 parallel processing in a restructurable computer system. IEEE Transactions on Electronic Computers, EC-13(5):649, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038288.

Newhall:1966:RMC

[New66a] E. E. Newhall. R66-25 magnetically controlled variable logic. IEEE Transactions on Electronic Computers, EC-15(2):270, April 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038743.

Newman:1966:RDM

[New66b] E. G. Newman. R66-62 A design for A multiple user multiprocessing system. IEEE Transactions on Electronic Computers, EC-15(5):837, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038907.

Newborn:1967:NBI

[New67] Monroe Newborn. Note on binary input-binary output finitememory sequential machines. IEEE Transactions on Electronic Computers, EC-16(4):514, August 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039119.

Newborn:1968:STB

[New68a] M. M. Newborn. A synthesis technique for binary input-binary output synchronous sequential Moore machines. IEEE Transactions on Computers, C-17(7):697–699, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687430.

Newborn:1968:MMB

[New68b] Monroe M. Newborn. Maximal memory binary input-binary output finite-memory sequential machines. IEEE Transactions on Computers, C-17(1):67–71, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008872.

Newman:1968:RMM

[New68c] W. M. Newman. R68-48 A multiprogramming monitor for small machines. IEEE Transactions on Computers, C-17(11):1099, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687274.

Narasimhan:1964:SFE

[NF64] R. Narasimhan and J. P. Fornango. Some further experiments in the parallel processing of pictures. IEEE Transactions on Electronic Computers, EC-13(6):748–750, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038312.

Nichols:1963:CAS

[Nic63] A. J. Nichols. Comments on Armstrong’s state assignment techniques. IEEE Transactions on Electronic Computers, EC-12(4): 407–409, August 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037920.

Nichols:1965:ACC

[Nic65a] A. J. Nichols. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-14(1):95–106, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038364.

Nichols:1965:MSR

[Nic65b] A. J. Nichols. Minimal shiftregister realizations of sequential machines. IEEE Transactions on Electronic Computers, EC-14(5):688–700, October 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038554.

Nienburg:1956:RAD

[Nie56] Raymond E. Nienburg. Reliability of an air defense computing system: Circuit design. IRE Transactions on Electronic Computers, EC-5(4):227–233, December 1956. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219957.

Nielsen:1968:REM

[Nie68] N. R. Nielsen. R68-12 an experimental model of System/360. IEEE Transactions on Computers, C-17(4):406–407, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687360.

Nieh:1969:RFA

[Nie69a] T. T. Nieh. R69-17 on finite automata with a time-variant structure. IEEE Transactions on Computers, C-18(7):668, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671333.

Nievergelt:1969:RPI

[Nie69b] J. Nievergelt. R69-13 perceptrons: An introduction to computational geometry. IEEE Transactions on Computers, C-18(6):572, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671311.

Nilson:1969:RFD

[Nil69] N. J. Nilson. R69-31 A formal deductive problem-solving system. IEEE Transactions on Computers, C-18(10):963–964, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671151.

Nisson:1960:AMD

[Nis60] C. J. Nisson. Analysis and methods for detection of some DC amplifier errors. IRE Transactions on Electronic Computers, EC-9 (1):121–122, March 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221609.

Nitzan:1964:FDS

[Nit64] David Nitzan. Flux divison in a saturable multipath core. IEEE Transactions on Electronic Computers, EC-13(3):272–277, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038154.

Nathan:1966:ODN

[NK66] A. Nathan and I. Korn. Optimal design of NLI squarers and sine generators. IEEE Transactions on Electronic Computers, EC-15(2):235–239, April 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038718.

Nathan:1957:DCS

[NM57] Amos Nathan and Yona Mahler. Demonstration of conditional stability on an analog computer. IRE Transactions on Electronic Computers, EC-6(4):287, December 1957. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5222041.

Nathan:1968:IVS

[NM68] A. Nathan and J. Molcho. Improved voltage selector and cascade multiplier circuits. IEEE Transactions on Computers, C-17(4):380–382, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687350.

Nolan:1962:RBL

[Nol62] J. F. Nolan. Review: Multi-Level Programming for a Real-Time System, by A. B. Shafritz, A. E. Miller, and K. Rose. IRE Transactions on Electronic Computers, EC-11(3): 426, June 1962. CODEN IRELAO. ISSN 0367-9950.

Nolan:1964:RGM

[Nol64] J. F. Nolan. R64-32 generalized multiprocessing and multiprogramming systems. IEEE Transactions on Electronic Computers, EC-13(3):319–320, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038181.

Newhouse:1956:HSS

[NP56] V. L. Newhouse and N. S. Prywes. High-speed shift registers using one core per bit. IRE Transactions on Electronic Computers, EC-5(3):114–120, September 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219930.

Newhall:1964:EIC

[NP64] E. E. Newhall and J. R. Perucca. Exploitation of initial conditions to achieve flux gain and nondestructive readout in balanced magnetic circuits. IEEE Transactions on Electronic Computers, EC-13(3):278–284, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038155.

Natarajan:1969:MAM

[NT69] N. K. Natarajan and P. A. V. Thomas. A multiaccess associative memory. IEEE Transactions on Computers, C-18(5):424–428, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671274.

Nelson:1964:SDA

[NW64] E. E. Nelson and H. F. Wolf. A sampled-data analog of neuron properties. IEEE Transactions on Electronic Computers, EC-13(6):744–745, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038309.

OBrien:1957:UDB

[O’B57] Joseph A. O’Brien. Unitdistance binary-decimal code translators. IRE Transactions on Electronic Computers, EC-6(2): 122–123, June 1957. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221585.

ODonnell:1964:AI

[O’D64] J. J. O’Donnell. Author index.

IEEE Transactions on Electronic Computers, EC-13(2):179, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038125.

OGrady:1967:CMC

[O’G67] Emmett P. O’Grady. Correlation method for computing sensitivity functions on a high-speed iterative analog computer. IEEE Transactions on Electronic Computers, EC-16(2):140–146, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039020.

Ogbuobiri:1969:CSN

[Ogb69] E. C. Ogbuobiri. Comment on “Solution of Nonlinear Equations”. IEEE Transactions on Computers, C-18(2):182–183, February 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671215. See [Kuo68,

Vic69b, Tho69, Tay69].

Oherren:1969:RRP

[O’h69] D. H. O’herren. Radar reflectivity plots-digital method. IEEE Transactions on Computers, C-18 (3):205–211, March 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671225.

Onyshkevych:1959:PPL

[OKL59] L. S. Onyshkevych, W. F. Kosonocky, and A. W. Lo. Parametric phaselocked oscillator — characteristics and applications to digital systems. IRE Transactions on Electronic Computers, EC-8(3): 277–286, September 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222684.

Olivares:1965:SNW

[Oli65] J. E. Olivares. Synthesis of nonconstant weight binary codes for minimum diode counters. IEEE Transactions on Electronic Computers, EC-14(1):69–72, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038355.

Olson:1965:RAM

[Ols65] C. D. Olson. R65-53 associative memory system implementation and characteristics. IEEE Transactions on Electronic Computers, EC-14(4):669–670, August 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038534.

ONeill:1967:EN

[O’N67a] L. A. O’Neill. Editor’s notice. IEEE Transactions on Electronic Computers, EC-16(1):1, February 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038976.

ONeill:1967:IAC

[O’N67b] Lawrence A. O’Neill. Iterative analog computation and the representation of signals. IEEE Transactions on Electronic Computers, EC-16(1):2–8, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038977.

ONeill:1969:RHS

[O’N69] L. A. O’Neill. R69-7 hybrid simulation of speech waveforms utilizing a Gaussian wave function representation. IEEE Transactions on Computers, C-18(4):384, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671264.

Ortel:1959:NLA

[Ort59] W. C. G. Ortel. Nanosecond logic by amplitude modulation at X band. IRE Transactions on Electronic Computers, EC-8(3): 265–271, September 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222682.

Otterman:1958:LNA

[Ott58] Joseph Otterman. On the loopand node-analysis approaches to the simulation of electrical networks. IRE Transactions on Electronic Computers, EC-7(3):

199–206, September 1958. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222576.

Otterman:1964:CFE

[Ott64] Joseph Otterman. Comment on “A Fundamental Error Theory for Analog Computers”. IEEE Transactions on Electronic Computers, EC-13(2):157, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038120. See [Nel63].

Overn:1963:SI

[Ove63] W. M. Overn. Subject index. IEEE Transactions on Electronic Computers, EC-12(3):349–351, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037885.

Patrick:1968:MMS

[PAB68] E. A. Patrick, D. R. Anderson, and F. K. Bechtel. Mapping multidimensional space to one dimension for computer output display. IEEE Transactions on Computers, C-17(10):949–953, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687240.

Page:1967:SI

[Pag67] C. V. Page. Subject index. IEEE Transactions on Electronic Computers, EC-16(1):128–131, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039000.

Paquette:1967:RHS

[Paq67] G. A. Paquette. R67-13 hybrid simulation of a helicopter. IEEE Transactions on Electronic Computers, EC-16(1):111–112, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039009.

Partridge:1962:LVR

[Par62] E. M. Partridge. A linear vector recognition function. IRE Transactions on Electronic Computers, EC-11(1):88, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219328.

Pasqualini:1967:FSI

[Pas67a] R. Pasqualini. Foreword to special issue on aerospace computers. IEEE Transactions on Electronic Computers, EC-16(5):550, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039142.

Pasqualini:1967:DCP

[Pas67b] Ronald Pasqualini. Design considerations for a parallel bit-organized MOS memory. IEEE Transactions on Electronic Computers, EC-16(5):551–557, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039143.

Paul:1967:RHC

[Pau67] R. J. A. Paul. R67-58 hybrid computers in the analysis of feedback control systems. IEEE Transactions on Electronic Computers, EC-16(5):714, October 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039178.

Prom:1956:JTS

[PC56] G. J. Prom and R. L. Crosby. Junction transistor switching circuits for high-speed digital computer applications. IRE Transactions on Electronic Computers, EC-5(4):192–196, December 1956. CO-DEN IRELAO. ISSN 0367-9950.

Parezanovic:1962:AIT

[PD62] N. Parezanovic and M. Dujmovic. Accuracy improvements of the tapped-potentiometer function generators. IRE Transactions on Electronic Computers, EC-11(1):63–66, February 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219318.

Perotto:1963:NMA

[Per63] Pier Giorgio Perotto. A new method for automatic character recognition. IEEE Transactions on Electronic Computers, EC-12(5):521–526, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037967.

Perlman:1968:SBS

[Per68] M. Perlman. The synthesis of binary sequence detectors. IEEE Transactions on Computers, C-17(9):873–880, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687474.

Petric:1961:NME

[Pet61] Jovan Petric. A new method of examining the stability of linear systems using a repetitive differential analyzer. IRE Transactions on Electronic Computers, EC-10(4):779–781, December 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219291.

Petrick:1963:RSE

[Pet63] S. R. Petrick. R63-71 some experiments in the generation of word and document associations. IEEE Transactions on Electronic Computers, EC-12(4):421, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037941.

Peterson:1964:N

[Pet64] W. W. Peterson. Notices. IEEE Transactions on Electronic Computers, EC-13(1):82, February 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038090.

Peters:1965:MSM

[Pet65] Charles J. Peters. A magnetically scanned magnetic tape transducer. IEEE Transactions on Electronic Computers, EC-14 (2):196–203, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038399.

Patrick:1969:CME

[PF69] E. A. Patrick and F. P. Fischer II. Cluster mapping with experimental computer graphics. IEEE Transactions on Computers, C-18(11):987–991, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671160.

Pfeiffer:1959:FQM

[Pfe59] Paul E. Pfeiffer. A four-quadrant multiplier using triangular waves, diodes, resistors, and operational amplifiers. IRE Transactions on Electronic Computers, EC-8 (2):222–227, June 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219526.

Pick:1964:SAN

[PGB64] G. G. Pick, S. B. Gray, and D. B. Brick. The solenoid array — A new computer element. IEEE Transactions on Electronic Computers, EC-13(1):27–35, February 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038073.

Picchia:1965:NTT

[Pic65] Walter Del Picchia. New type of transistorized digital analog converter. IEEE Transactions on Electronic Computers, EC-14(3):492, June 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038476.

Picchia:1966:MPF

[Pic66a] Walter Del Picchia. A multipurpose function generator. IEEE Transactions on Electronic Computers, EC-15(3):382–383, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038780.

Picchia:1966:TDA

[Pic66b] Walter Del Picchia. A transistorized digital analog converter with AC amplification. IEEE Transactions on Electronic Computers, EC-15(2):233–235, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038717.

Pierce:1963:PSD

[Pie63a] William H. Pierce. Predicting signal degeneration and gate compatibility in logic circuits. IEEE Transactions on Electronic Computers, EC-12(3):277–281, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037864.

Pierre:1963:TPT

[Pie63b] D. A. Pierre. Terminal properties of toroidal cores with circular cross sections. IEEE Transactions on Electronic Computers, EC-12(5):570–572, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037981.

Pierre:1963:OPD

[Pie63c] Donald A. Pierre. Optimization of pulse and digital circuits by use of the Lagrange multiplier technique. IEEE Transactions on Electronic Computers, EC-12(5):488–492, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037962.

Pierce:1965:SI

[Pie65] W. H. Pierce. Subject index. IEEE Transactions on Electronic Computers, EC-14(1):106–109, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038365.

Prywes:1961:ULH

[PLS61] N. S. Prywes, H. Lukoff, and J. Schwarz. UNIVAC–LARC highspeed circuitry: Case history in circuit optimization. IRE Transactions on Electronic Computers, EC-10(3):426–438, September 1961.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219231.

Pulvari:1958:SFM

[PM58] C. F. Pulvari and G. E. Mc-Duffie. Scanners for ferroelectric memory capacitors. IRE Transactions on Electronic Computers, EC-7(1):34–40, March 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222093.

Pohm:1960:MFM

[PM60] A. V. Pohm and E. N. Mitchell.

Magnetic film memories, A survey. IRE Transactions on Electronic Computers, EC-9(3):308– 314, September 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219850.

Pyne:1962:RRS

[PM62] I. B. Pyne and E. J. McCluskey. The reduction of redundancy in solving prime implicant tables. IRE Transactions on Electronic Computers, EC-11(4):473–482, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219386.

Preparata:1967:CAP

[PMC67] Franco P. Preparata, Gernot Metze, and Robert T. Chien. On the connection assignment problem of diagnosable systems. IEEE Transactions on Electronic Computers, EC-16(6):848–854, December 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039201.

Peeler:1967:ACS

[PMRC67] Donald L. Peeler, Paul H. Meredith, Louis M. Richards, and Christian O. Clark. Automatic checkout systems for Titan III and Apollo guidance computer programs. IEEE Transactions on Electronic Computers, EC-16(5):580–590, October 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039147.

Pohm:1964:CGN

[Poh64a] A. V. Pohm. Computer group news. IEEE Transactions on Electronic Computers, EC-13(4):502, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038224.

Pohm:1964:SI

[Poh64b] A. V. Pohm. Subject index. IEEE Transactions on Electronic Computers, EC-13(5):663–667, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038277.

Pohm:1966:MFS

[Poh66] A. V. Pohm. Magnetic film scratchpad memories. IEEE Transactions on Electronic Computers, EC-15(4):452–458, August 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038818.

Polimerou:1957:NMG

[Pol57] Lazarus G. Polimerou. A new method for generating a function of two independent variables. IRE Transactions on Electronic Computers, EC-6(3): 167–169, September 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222014.

Porter:1964:NCF

[Por64] Sigmund N. Porter. A new configuration for faster cryotron circuits. IEEE Transactions on Electronic Computers, EC-13(1): 56–57, February 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038082.

Powell:1959:DPV

[Pow59] R. V. Powell. Distributed parameter vibration with structural damping and noise excitation. IRE Transactions on Electronic Computers, EC-8(2):197–200, June 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219521.

Powers:1965:CNN

[Pow65] George L. Powers. Comment on NOR–NAND synthesis. IEEE Transactions on Electronic Computers, EC-14(4):648, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038524.

Powell:1969:PSD

[Pow69] T. J. Powell. A procedure for selecting diagnostic tests. IEEE Transactions on Computers, C-18 (2):168–175, February 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671212.

Paz:1968:CDR

[PP68] A. Paz and B. Peleg. On concatenative decompositions of regular events. IEEE Transactions on Computers, C-17(3):229–237, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687324.

Prather:1965:TC

[Pra65] Ronald E. Prather. On tree circuits. IEEE Transactions on Electronic Computers, EC-14(6): 841–851, December 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038603.

Prather:1966:NAE

[Pra66a] R. E. Prather. New Associate Editors and Reviews Editor. IEEE Transactions on Electronic Computers, EC-15(1):2, February 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038660.

Prather:1966:TVM

[Pra66b] R. E. Prather. Three variable multiple output tree circuits. IEEE Transactions on Electronic Computers, EC-15(1):3–13, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038661.

Pratt:1969:AFH

[Pra69] W. K. Pratt. An algorithm for a fast Hadamard matrix transform of order twelve. IEEE Transactions on Computers, C-18 (12):1131–1132, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671190.

Preston:1955:ACS

[Pre55] F. S. Preston. An analog computer for the solution of tangents. IRE Transactions on Electronic Computers, EC-4(3): 101–106, September 1955. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407908.

Press:1963:RLD

[Pre63] R. J. Press. R63-37 A logic design translator. IEEE Transactions on Electronic Computers, EC-12(2):157, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037837.

Preiss:1964:ACC

[Pre64a] R. J. Preiss. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-13(2):165–176, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038123.

Preparata:1964:SLR

[Pre64b] Franco P. Preparata. State-logic relations for autonomous sequential networks. IEEE Transactions on Electronic Computers, EC-13(5):542–548, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038247.

Preparata:1965:RSC

[Pre65] Franco P. Preparata. On the realizability of special classes of autonomous sequential networks. IEEE Transactions on Electronic Computers, EC-14(6): 791–797, December 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038597.

Preiss:1966:ACA

[Pre66a] R. J. Preiss. Annual cumulative author index. IEEE Transactions on Electronic Computers, EC-15(6):997–999, December 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038952.

Preparata:1966:CTB

[Pre66b] Franco P. Preparata. Convolutional transformations of binary sequences: Boolean functions and their resynchronizing properties. IEEE Transactions on Electronic Computers, EC-15(6): 898–908, December 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038933.

Preparata:1968:CTR

[Pre68] F. P. Preparata. Convolutional transformation and recovery of binary sequences. IEEE Transactions on Computers, C-17(7):649–655, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687421.

Price:1965:CML

[Pri65] J. E. Price. Counting with majority-logic networks. IEEE Transactions on Electronic Computers, EC-14(2):256–260, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038413. See comment

[Mis66].

Proctor:1964:LDT

[Pro64] Roy M. Proctor. A logic design translator experiment demonstrating relationships of language to systems and logic design. IEEE Transactions on Electronic Computers, EC-13(4):422–430, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038207.

Perles:1963:TDA

[PRS63] M. Perles, M. O. Rabin, and E. Shamir. The theory of definite automata. IEEE Transactions on Electronic Computers, EC-12(3):233–243, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037857.

Prywes:1958:DMS

[Pry58] Noah S. Prywes. Diodeless magnetic shift registers utilizing transfluxors. IRE Transactions on Electronic Computers, EC-7(4): 316–324, December 1958. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222666.

Plisch:1966:NEC

[PS66] D. C. Plisch and A. K. Scidmore. The number of equivalence classes of functions realizable by generalized tributary networks. IEEE Transactions on Electronic Computers, EC-15(2):244–245, April 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038721.

Petrick:1968:DCS

[PS68] S. R. Petrick and G. C. Sethares. On the determination of complete sets of logical functions.

IEEE Transactions on Computers, C-17(3):273, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687330.

Pleshko:1966:IPM

[PT66] P. Pleshko and L. M. Terman. An investigation of the potential of MOS transistor memories. IEEE Transactions on Electronic Computers, EC-15(4):423–427, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038814.

Paull:1959:MNS

[PU59] M. C. Paull and S. H. Unger.

Minimizing the number of states in incompletely specified sequential switching functions. IRE Transactions on Electronic Computers, EC-8(3):356–367, September 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222697.

Pugsley:1965:SFL

[Pug65] James H. Pugsley. Sequential functions and linear sequential machines. IEEE Transactions on Electronic Computers, EC-14 (3):376–382, June 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038455.

Pritsker:1959:SOS

[PVW59] A. A. B. Pritsker, R. C. Van Buskirk, and J. K. Wetherbee. Simulation to obtain a systems measure of an air duel environment. IRE Transactions on Electronic Computers, EC-8(1):55–59, March 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222762.

Paull:1967:NSM

[PW67] M. C. Paull and G. Waldbaum. A note on state minimization of asynchronous sequential functions. IEEE Transactions on Electronic Computers, EC-16(1): 94–97, February 1967. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038993.

Quatse:1967:PAG

[QK67] Jesse T. Quatse and Roy A. Keir. A parallel accumulator for a generalpurpose computer. IEEE Transactions on Electronic Computers, EC-16(2):165–171, April 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039024.

Quastler:1957:CBC

[Qua57] H. Quastler. The complexity of biological computers. IRE Transactions on Electronic Computers, EC-6(3):192–194, September 1957. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222020.

Quaife:1969:G

[Qua69] H. J. Quaife. On (d, k, μ) graphs.

IEEE Transactions on Computers, C-18(3):270–272, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671235.

Rosenheim:1959:IRU

[RA59] D. E. Rosenheim and R. B. Ash.

Increasing reliability by the use of redundant machines. IRE Transactions on Electronic Computers, EC-8(2):125–130, June 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219513.

Rabinovici:1963:AN

[Rab63] B. Rabinovici. AFIPS news. IEEE Transactions on Electronic Computers, EC-12(2):176, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037827.

Raffel:1964:N

[Raf64] J. I. Raffel. Notices. IEEE Transactions on Electronic Computers, EC-13(3):341–342, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038176.

Ralston:1964:N

[Ral64] A. Ralston. Notices. IEEE Transactions on Electronic Computers, EC-13(4):502–503, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038225.

Ramamoorthy:1965:CCG

[Ram65] C. V. Ramamoorthy. Connectivity considerations of graphs representing discrete sequential systems. IEEE Transactions on Electronic Computers, EC-14(5): 724–727, October 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038561.

Randolph:1963:PC

[Ran63] John F. Randolph. A polynomial counter. IEEE Transactions on Electronic Computers, EC-12(3):320, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037875.

Randery:1964:SPA

[Ran64] Vijay K. Randery. Study of a parametron on an analog computer. IEEE Transactions on Electronic Computers, EC-13(5): 612–614, October 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038259.

Rao:1958:NTI

[Rao58] P. Venkata Rao. A novel type of isograph (algebraic equation solver). IRE Transactions on Electronic Computers, EC-7 (2):97–103, June 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222517.

Rao:1968:ECL

[Rao68] T. R. N. Rao. Error-checking logic for arithmetic-type operations of a processor. IEEE Transactions on Computers, C-17(9):845–849, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687471.

Raphael:1966:RPD

[Rap66] B. Raphael. R66-52 A programmer’s description of LLLLLL, Bell Telephone Laboratories Low Level Linked List Language. IEEE Transactions on Electronic Computers, EC-15(4):681–682, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038859.

Raphael:1967:C

[Rap67] B. Raphael. Contributors. IEEE Transactions on Electronic Computers, EC-16(6):873–876, December 1967. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4039210.

Ratz:1967:ACF

[Rat67] Alfred G. Ratz. Analog computation of Fourier series and integrals. IEEE Transactions on Electronic Computers, EC-16(4):515, August 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039120.

Rawdin:1959:TMA

[Raw59] Eugene Rawdin. Time multiplexing as applied to analog computation. IRE Transactions on Electronic Computers, EC-8 (1):42–47, March 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222760.

Roth:1967:PAC

[RBS67] J. Paul Roth, Willard G. Bouricius, and Peter R. Schneider. Programmed algorithms to compute tests to detect and distinguish between failures in logic circuits. IEEE Transactions on Electronic Computers, EC-16(5):567–580, October 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039146.

Rajchman:1957:CSM

[RC57] J. A. Rajchman and H. D. Crane. Current steering in magnetic circuits. IRE Transactions on Electronic Computers, EC-6 (1):21–30, March 1957. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221557.

Raffel:1964:PAM

[RC64] J. I. Raffel and T. S. Crowther. A proposal for an associative memory using magnetic films. IEEE Transactions on Electronic Computers, EC-13(5):611, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038258.

Roy:1966:NTR

[RC66] K. K. Roy and A. K. Choudhury. A note on testing and realization of threshold functions. IEEE Transactions on Electronic Computers, EC-15(2):242–244, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038720.

Read:1963:DLS

[Rea63] Alvin A. Read. A dynamic large signal model for a single-domain thin magnetic film inductor. IEEE Transactions on Electronic Computers, EC-12(5):517–521, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037966.

Reed:1965:SRS

[Ree65] I. S. Reed. Some remarks on state reduction of asynchronous circuits by the Paull–Unger method. IEEE Transactions on Electronic Computers, EC-14(2):262–265, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038416.

Reitwiesner:1960:DCP

[Rei60] George W. Reitwiesner. The determination of carry propagation length for binary addition. IRE Transactions on Electronic Computers, EC-9(1):35–38, March 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221602.

Renton:1964:AI

[Ren64] C. A. Renton. Author index. IEEE Transactions on Electronic Computers, EC-13(1):80, February 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038088.

Renschler:1968:VCD

[Ren68] E. L. Renschler. A variable counter design technique. IEEE Transactions on Computers, C-17(7):694–696, July 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687428. See comment

[Tae69, Jon69a].

Renschler:1969:ARa

[Ren69a] E. L. Renschler. Author’s reply. IEEE Transactions on Computers, C-18(3):281, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671241.

Renschler:1969:ARb

[Ren69b] E. L. Renschler. Author’s reply. IEEE Transactions on Computers, C-18(3):282–284, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671243.

Rubin:1965:ART

[RG65] A. I. Rubin and G. F. Graber.

Acoustic ray tracing on the general-purpose electronic-analog computer. IEEE Transactions on Electronic Computers, EC-14 (3):443–455, June 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038464.

Rosen:1966:PRE

[RH66] C. A. Rosen and D. J. Hall. A pattern recognition experiment with near-optimum results. IEEE Transactions on Electronic Computers, EC-15(4):666–667, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038845. See comments

[FF67, Kno67a].

Ratz:1967:MPN

[RH67a] Herbert C. Ratz and J. V. Hildebrand. Modified pseudorandom number generators. IEEE Transactions on Electronic Computers, EC-16(6):854–856, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039202.

Robinson:1967:GRT

[RH67b] Stanley U. Robinson and Robert W.

House. Gimpel’s reduction technique extended to the covering problem with costs. IEEE Transactions on Electronic Computers, EC-16(4):509–514, August 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039118.

Rosen:1967:AR

[RH67c] C. A. Rosen and D. J. Hall. Authors’ reply. IEEE Transactions on Electronic Computers, EC-16(2):231, April 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039040.

Ribeiro:1964:CPD

[Rib64] Sergio T. Ribeiro. Comments on pulsed-data hybrid computers. IEEE Transactions on Electronic Computers, EC-13(5): 640–642, October 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038272.

Ribeiro:1967:RPM

[Rib67] Sergio T. Ribeiro. Randompulse machines. IEEE Transactions on Electronic Computers, EC-16(3):261–276, June 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039064.

Richards:1967:DPD

[Ric67] R. K. Richards. “Datum Processing,” “Digital datum transmission,” and “Automaton theory”. IEEE Transactions on Electronic Computers, EC-16(2):231, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039041.

Rideout:1966:RDA

[Rid66] V. C. Rideout. R66-27 delay approximations for correlation measurements using analog computers. IEEE Transactions on Electronic Computers, EC-15(2):271, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038745.

Riseman:1967:RAU

[Ris67] Edward M. Riseman. A realization algorithm using three-input majority elements. IEEE Transactions on Electronic Computers, EC-16(4):456–462, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039110.

Riordan:1965:UAT

[RM65] R. H. S. Riordan and R. R. A. Morton. The use of analog techniques in binary arithmetic units. IEEE Transactions on Electronic Computers, EC-14(1):29–35, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038346.

Robertson:1955:TCM

[Rob55a] J. E. Robertson. Two’s complement multiplication in binary parallel digital computers. IRE Transactions on Electronic Computers, EC-4(3):118–119, September 1955. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407911.

Robinson:1955:EAC

[Rob55b] A. S. Robinson. An electronic analog computing technique for the solution of trigonometric problems. IRE Transactions on Electronic Computers, EC-4(3): 95–101, September 1955. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407907.

Robertson:1956:OBA

[Rob56] J. E. Robertson. Odd binary asynchronous counters. IRE Transactions on Electronic Computers, EC-5(1):12–15, March 1956. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219785.

Robertson:1958:NCD

[Rob58] James E. Robertson. A new class of digital division methods. IRE Transactions on Electronic Computers, EC-7(3):218– 222, September 1958. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222579.

Robinson:1966:ACC

[Rob66] J. P. Robinson. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-15(1):133–144, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038690.

Roberts:1967:CDG

[Rob67] Lawrence G. Roberts. Conic display generator using multiplying digital-analog converters. IEEE Transactions on Electronic Computers, EC-16(3):369–370, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039081.

Rogers:1966:RDH

[Rog66] A. E. Rogers. R66-39 A demonstration hybrid computer for realtime flight simulation. IEEE Transactions on Electronic Computers, EC-15(3):394, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038796.

Rossing:1958:SCP

[ROK58] T. D. Rossing, W. M. Overn, and V. J. Korkowski. The switching characteristics of 4-79 permalloy cores with different anneals. IRE Transactions on Electronic Computers, EC-7(3): 228–231, September 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222581.

Rosenfeld:1958:MCP

[Ros58] Jack L. Rosenfeld. Magnetic core pulse-switching circuits for standard packages. IRE Transactions on Electronic Computers, EC-7(3):223–228, September 1958. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222580.

Rosenthal:1961:CMA

[Ros61] C. W. Rosenthal. Computing machine aids to a development project. IRE Transactions on Electronic Computers, EC-10(3): 400–406, September 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219228.

Rosenfeld:1962:RBF

[Ros62] A. Rosenfeld. Review: Design Factors in the Development of an Optical Character Recognition Machine, by J. J. Leimer. IRE Transactions on Electronic Computers, EC-11(423–423), June 1962. CODEN IRELAO. ISSN 0367-9950.

Rosenfeld:1964:RBP

[Ros64] A. Rosenfeld. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-13(3):316–322, June 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038171.

Rose:1965:LPF

[Ros65a] Gordon A. Rose. “Light-Pen” facilities for direct view storage tubes — an economical solution for multiple man-machine communication. IEEE Transactions on Electronic Computers, EC-14(4):637–639, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038520.

Rosin:1965:SPC

[Ros65b] Robert F. Rosin. A special purpose computer for solution of partial differential equations and other iterative algorithms. IEEE Transactions on Electronic Computers, EC-14(3):488–490, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038474.

Rosenkrantz:1966:SSI

[Ros66a] Daniel J. Rosenkrantz. Synchronizing sequences for incompletely specified flow tables. IEEE Transactions on Electronic Computers, EC-15(1):104–105, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038673.

Rosin:1966:RDR

[Ros66b] R. F. Rosin. R66-26 A destructivereadout associative memory. IEEE Transactions on Electronic Computers, EC-15(2):270, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038744.

Rose:1967:IMG

[Ros67a] Gordon A. Rose. “Intergraphic,” A microprogrammed graphicalinterface computer. IEEE Transactions on Electronic Computers, EC-16(6):773–784, December 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039194.

Rosene:1967:MAM

[Ros67b] A. Frederick Rosene. Memory allocation for multiprocessors. IEEE Transactions on Electronic Computers, EC-16(5):659–665, October 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039157.

Rosko:1968:CHC

[Ros68] J. S. Rosko. Comments on “Hybrid Computer Solution of Optimal Control Problems by the Maximum Principle”. IEEE Transactions on Computers, C-17(9):899, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687482. See [MTI67].

Rothstein:1959:RBN

[Rot59] Jerome Rothstein. Residues of binary numbers modulo three. IRE Transactions on Electronic Computers, EC-8(2):229, June 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219529.

Rothbart:1961:BMD

[Rot61] Arthur Rothbart. Bibliography on magnetostrictive delay lines. IRE Transactions on Electronic Computers, EC-10(2):285, June 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219200.

Rouche:1958:SPB

[Rou58] N. Rouche. Some properties of Boolean equations. IRE Transactions on Electronic Computers, EC-7(4):291–298, December 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222662.

Roy:1961:MMA

[Roy61] Rob Roy. Multichannel minimumamplitude comparator. IRE Transactions on Electronic Computers, EC-10(3):533–534, September 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219249.

Roy:1967:SSS

[Roy67] P. K. Sinha Roy. Synthesis of symmetric switching functions using threshold logic elements. IEEE Transactions on Electronic Computers, EC-16(3):359–364, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039078. See comments

[Roy68a].

Roy:1968:FCS

[Roy68a] P. K. Sinha Roy. Further comments on “Synthesis of Symmetric Switching Functions Using Threshold Logic Networks”. IEEE Transactions on Computers, C-17(9):899–900, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687483. See [Roy67].

Roy:1968:SRD

[Roy68b] P. K. Sinha Roy. A slide rule device for checking 2-summability. IEEE Transactions on Computers, C-17(3):279–283, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687332.

Rozier:1962:DBC

[Roz62] Charles P. Rozier. Decimalto-binary conversion using octal radix arithmetic. IRE Transactions on Electronic Computers, EC-11(5):708–709, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219436.

Rabinovici:1962:TFF

[RR62a] B. Rabinovici and C. A. Renton. A transistor flip-flop full binary adder. IRE Transactions on Electronic Computers, EC-11(6): 789–791, December 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219465.

Renton:1962:TDF

[RR62b] C. A. Renton and B. Rabinovici. Tunnel-diode full binary adder. IRE Transactions on Electronic Computers, EC-11 (2):213–217, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219354.

Renton:1966:LAI

[RR66] C. A. Renton and B. M. Rabinovici.

Light amplifiers and inverters for digital application employing negative conductance devices. IEEE Transactions on Electronic Computers, EC-15(4):664–665, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038844.

Roder:1967:MPM

[RR67] John Roder and A. Frederick Rosene. Memory protection in multiprocessing systems. IEEE Transactions on Electronic Computers, EC-16(3):320–326, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039070.

Rubin:1962:CRT

[Rub62a] A. I. Rubin. Continuous regression techniques using analog computers. IRE Transactions on Electronic Computers, EC-11(5):691–699, October 1962.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219432.

Rubin:1962:RBL

[Rub62b] A. I. Rubin. Review: Two-Level Correlation on an Analog Computer, by C. L. Becker and J. V. Wait. IRE Transactions on Electronic Computers, EC-11(3):421, June 1962. CODEN IRELAO. ISSN 0367-9950.

Rubin:1965:RHC

[Rub65] A. I. Rubin. R65-42 A hybrid computer for adaptive nonlinear process identification. IEEE Transactions on Electronic Computers, EC-14(3):501, June 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038488.

Rubin:1967:RAA

[Rub67] A. I. Rubin. R67-27 an automatic acoustic ray tracing computer. IEEE Transactions on Electronic Computers, EC-16(2): 240–241, April 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039053.

Rubin:1968:RST

[Rub68] A. I. Rubin. R68-39 simulation of the transfer function of a crustacean muscle bundle. IEEE Transactions on Computers, C-17 (8):819–820, August 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687463.

Rudeanu:1965:TDL

[Rud65] Sergiu Rudeanu. On Tohma’s decompositions of logical functions. IEEE Transactions on Electronic Computers, EC-14(6): 929–931, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038616.

Russo:1965:SET

[Rus65] Roy L. Russo. Synthesis of error-tolerant counters using minimum distance three state assignments. IEEE Transactions on Electronic Computers, EC-14 (3):359–366, June 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038453.

Russell:1968:VIP

[Rus68] J. K. Russell. A visual image processor. IEEE Transactions on Computers, C-17(7):635–639, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687419.

Rux:1969:GDL

[Rux69] P. T. Rux. A glass delay line content-addressed memory system. IEEE Transactions on Computers, C-18(6):512–520, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671296.

Russo:1968:PPC

[RV68] F. A. Russo and R. J. Valek. Process performance computer for adaptive control systems. IEEE Transactions on Computers, C-17 (11):1027–1037, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687260.

Rochester:1953:CD

[RW53] Nathaniel Rochester and Willis H. Ware. Computer definitions. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(4):2, December 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407686.

Rossing:1967:CCC

[RW67] Jakob F. Rossing and Joachim E. Walther. Calculation of crosscoupled noise in digital systems. IEEE Transactions on Electronic Computers, EC-16(1): 14–17, February 1967. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038979.

Reiffen:1966:NBS

[RY66] B. Reiffen and H. L. Yudkin. On nonlinear binary sequential circuits and their inverses. IEEE Transactions on Electronic Computers, EC-15(4):586–596, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038832.

Santos:1964:BVB

[SA64] J. Santos and H. Arango. Base 3 vs base 2 synchronous arithmetic units. IEEE Transactions on Electronic Computers, EC-13(5):608–609, October 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038256.

Sackman:1965:AI

[Sac65] H. Sackman. Annual index 1965. IEEE Transactions on Electronic Computers, EC-14(6): 1019–1030, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038639.

Sagalovitch:1963:PN

[Sag63] Yu. L. Sagalovitch. PGEC news.

IEEE Transactions on Electronic Computers, EC-12(3):352, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037887.

Sage:1966:RDS

[Sag66] A. P. Sage. R66-61 digital simulation of large scale systems. IEEE Transactions on Electronic Computers, EC-15(5):836–837, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038906.

Salter:1960:HST

[Sal60] Forrest Salter. High-speed transistorized adder for a digital computer. IRE Transactions on Electronic Computers, EC-9(4): 461–464, December 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219885.

Saltman:1961:RCT

[Sal61] Roy G. Saltman. Reducing computing time for synchronous binary division. IRE Transactions on Electronic Computers, EC-10 (2):169–174, June 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219186.

Salton:1963:RSS

[Sal63] G. Salton. R63-21 the Stenowriter — A system for the lexical processing of stenotypy. IEEE Transactions on Electronic Computers, EC-12(1):38, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037790.

Salter:1964:TME

[Sal64] Forrest Salter. A ternary memory element using a tunnel diode. IEEE Transactions on Electronic Computers, EC-13(2):155–156, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038117.

Salton:1965:RNL

[Sal65] G. Salton. R65-13 natural language and the computer. IEEE Transactions on Electronic Computers, EC-14(1):93–94, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038373.

Santos:1966:TST

[SAL66] J. Santos, H. Arango, and F. Lorenzo.

Threshold synthesis of ternary digital systems. IEEE Transactions on Electronic Computers, EC-15(1):105–107, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038674.

Saltzer:1968:RDE

[Sal68] J. H. Saltzer. R68-41 development of executive routines, both hardware and software. IEEE Transactions on Computers, C-17(9):923–924, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687487.

Sammet:1965:RAL

[Sam65] J. E. Sammet. R65-23 aspects of language design for combinatorial computing. IEEE Transactions on Electronic Computers, EC-14(2):277–278, April 1965.

CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038431.

Sammon:1969:NMD

[Sam69a] J. W. Sammon, Jr. A nonlinear mapping for data structure analysis. IEEE Transactions on Computers, C-18(5):401–409, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671271.

Samuel:1969:SAF

[Sam69b] I. N. Samuel. Synthesis and analysis: A flexible technique for processing command language. IEEE Transactions on Computers, C-18(11):1053–1061, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671173.

Sander:1964:RCE

[San64] Wendell B. Sander. “Ratchet” circuits for electronic analog storage. IEEE Transactions on Electronic Computers, EC-13(4): 474–475, August 1964. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038217.

Sandoval:1966:HTS

[San66] Victor H. Sandoval. On the hyperplanes of the taylor-steinbuch linear separator. IEEE Transactions on Electronic Computers, EC-15(6):935, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038942.

Santos:1965:TSE

[SAP65] J. Santos, H. Arango, and M. Pascual. A ternary storage element using a conventional ferrite core. IEEE Transactions on Electronic Computers, EC-14(2):248, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038408.

Sasseen:1957:EAC

[Sas57] J. H. Sasseen. An electronic analog cross correlator for dip logs. IRE Transactions on Electronic Computers, EC-6(3):182– 187, September 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222017.

Sasaki:1967:ASR

[Sas67] Akio Sasaki. Addition and subtraction in the residue number system. IEEE Transactions on Electronic Computers, EC-16 (2):157–164, April 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=4039023;   http://
ieeexplore.ieee.org/xpl/RecentIssue.
jsp?punumber=4037753;   http:
//ieeexplore.ieee.org/xpl/tocresult.
jsp?isnumber=4039015.

Sasaki:1968:BIA

[Sas68] A. Sasaki. The basis for implementation of additive operations in the residue number system. IEEE Transactions on Computers, C-17(11):1066–1073, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=1687264;   http://
                                        ieeexplore.ieee.org/xpl/tocresult.
                                        jsp?isnumber=35582.

Sasaki:1969:RTO

[Sas69] F. T. Sasaki. R69-6 trajectory optimization by a direct descent process. IEEE Transactions on Computers, C-18(3):303–304, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671246.

Saucier:1967:EAS

[Sau67] Gabriele Saucier. Encoding of asynchronous sequential networks. IEEE Transactions on Electronic Computers, EC-16(3):365–369, June 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039080.

Savitt:1959:HSA

[Sav59] Donald Savitt. A high-speed analog to digital converter. IRE Transactions on Electronic Computers, EC-8(1):31–35, March 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222758.

Single:1961:CMP

[SB61] C. H. Single and J. A. Brussolo. Copper-mandrel potentiometer dynamic error and compensation. IRE Transactions on Electronic Computers, EC-10(3): 516–523, September 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219241.

Sammet:1964:IF

[SB64] J. E. Sammet and E. R. Bond. Introduction to FORMAC. IEEE Transactions on Electronic Computers, EC-13(4):386–394, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038202.

Slotnick:1963:CHJ

[SBM63] D. L. Slotnick, W. C. Borck, and R. C. McReynolds. Comments on H. J. Heijn’s review of “The SOLOMON Computer”. IEEE Transactions on Electronic Computers, EC-12(4):410, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037922. See [Hei63].

Schlaeppi:1960:SCM

[SC60] H. P. Schlaeppi and I. P. V.

Carter. Submicrosecond core memories using multiple coincidence. IRE Transactions on Electronic Computers, EC-9(2):

192–198, June 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219818.

Shani:1967:OTS

[SC67] Alexander Shani and Doron J. Cohen. Optimum time sharing of a digital communication channel. IEEE Transactions on Electronic Computers, EC-16(4):397–402, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039103.

Shanks:1968:UDC

[SC68] J. L. Shanks and T. W. Cairns. Use of a digital convolution device to perform recursive filtering and the Cooley–Tukey algorithm. IEEE Transactions on Computers, C-17(10):943–949, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687239.

Scarbrough:1953:ADC

[Sca53] A. D. Scarbrough. An analogto-digital converter. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(3):5–7, September 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407732.

Schmid:1958:TFQ

[Sch58] Hermann Schmid. A transistorized four-quadrant time-division multiplier with an accuracy of 0.1 per cent. IRE Transactions on Electronic Computers, EC-7 (1):41–47, March 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222094.

Schmid:1962:LSF

[Sch62] Hermann Schmid. Linear-segment function generator. IRE Transactions on Electronic Computers, EC-11(6):780–788, December 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219463.

Schmid:1963:OHC

[Sch63a] Hermann Schmid. An operational hybrid computing system provides analog-type computation with digital elements. IEEE Transactions on Electronic Computers, EC-12(6):715–732, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038022.

Schubert:1963:RMM

[Sch63b] Ernest J. Schubert. The relative merits of minimization techniques for switching circuits. IEEE Transactions on Electronic Computers, EC-12(3):321–322, June 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037877.

Schlaeppi:1964:FLD

[Sch64a] H. P. Schlaeppi. A formal language for describing machine logic, timing, and sequencing (LOTIS). IEEE Transactions on Electronic Computers, EC-13(4):439–448, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038209.

Schmid:1964:SI

[Sch64b] H. Schmid. Subject index. IEEE Transactions on Electronic Computers, EC-13(2):177–178, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038124.

Schorr:1964:CAD

[Sch64c] Herbert Schorr. Computer-aided digital system design and analysis using a register transfer language. IEEE Transactions on Electronic Computers, EC-13(6): 730–737, December 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038305.

Scheff:1965:UDT

[Sch65a] B. Scheff. Using a decision-table structure as the input language format for programming automatic test equipment systems. IEEE Transactions on Electronic Computers, EC-14(2):248–250, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038409.

Schwartz:1965:RCC

[Sch65b] E. S. Schwartz. R65-57 CPSS — A common programming support system. IEEE Transactions on Electronic Computers, EC-14(4):671, August 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038538.

Schmookler:1968:HSB

[Sch68] M. S. Schmookler. High speed binary to decimal conversion. IEEE Transactions on Computers, C-17(5):506–508, May 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687381.

Schisler:1969:RMR

[Sch69a] C. W. Schisler. R69-15 microwave radar and electronic environment simulator. IEEE Transactions on Computers, C-18(6):573, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671313.

Schmookler:1969:MSP

[Sch69b] M. S. Schmookler. On mod-2 sums of products. IEEE Transactions on Computers, C-18(10):957, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671146.

Schrage:1969:AOQ

[Sch69c] L. Schrage. Analysis and optimization of a queueing model of a real-time computer control system. IEEE Transactions on Computers, C-18(11):997–1003, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671162.

Scidmore:1964:RBP

[Sci64] A. K. Scidmore. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-13 (5):645–649, October 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038275.

Scidmore:1966:RML

[Sci66] A. K. Scidmore. R66-79 on measures of logic performance: Logic quantum, factor and figure of merit. IEEE Transactions on Electronic Computers, EC-15(6):957–958, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038957.

Scott:1957:URI

[Sco57] Norman R. Scott. On the use of redundant integrators in analog computers. IRE Transactions on Electronic Computers, EC-6(4):287–288, December 1957.

CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222042.

Stone:1957:VFD

[SD57] R. S. Stone and R. A. Dandl. A variable function delay for analog computers. IRE Transactions on Electronic Computers, EC-6(3):187–189, September 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222018.

Sakai:1963:ASR

[SD63] T. Sakai and S. Doshita. The automatic speech recognition system for conversational sound. IEEE Transactions on Electronic Computers, EC-12(6):835–846, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038031.

Schneider:1968:ASM

[SD68] P. R. Schneider and D. L. Dietmeyer. An algorithm for synthesis of multiple-output combinational logic. IEEE Transactions on Computers, C-17(2):117–128, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687301. See comments

[DM68, Has69].

Su:1969:CRT

[SD69] Yueh-Hsung Su and D. L. Dietmeyer. Computer reduction of two-level, multiple-output switching circuits. IEEE Transactions on Computers, C-18(1):58–63, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671118.

Sebestyen:1966:ANP

[SE66] G. Sebestyen and J. Edie. An algorithm for non-parametric pattern recognition. IEEE Transactions on Electronic Computers, EC-15(6):908–915, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038934.

Sebestyen:1957:DTP

[Seb57] George Sebestyen. A design technique for pedestal-free switching circuits. IRE Transactions on Electronic Computers, EC-6(3): 162–166, September 1957. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222013.

Sebestyen:1963:IN

[Seb63] G. S. Sebestyen. IFIP news. IEEE Transactions on Electronic Computers, EC-12(1):56, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037779.

Sedore:1968:MEU

[Sed68] S. R. Sedore. More efficient use of the F matrix in practical circuit analysis programs.

IEEE Transactions on Computers, C-17(5):503–506, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687380.

Seitzer:1964:ADC

[Sei64] D. Seitzer. Amplifier and driver circuits for thin film memories with 15 nanoseconds read cycle time. IEEE Transactions on Electronic Computers, EC-13(6): 722–729, December 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038304.

Seitzer:1967:EWD

[Sei67] Dieter Seitzer. An experimental word decode and drive system for a magnetic film memory with 20-ns READ-cycle time. IEEE Transactions on Electronic Computers, EC-16(2):172–179, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039025.

Semon:1963:RSD

[Sem63] W. Semon. R63-18 sign detection in nonredundant residue systems. IEEE Transactions on Electronic Computers, EC-12(1):37, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037787.

Serlin:1968:RIH

[Ser68a] O. Serlin. R68-32 the IADIC: A hybrid computing element. IEEE Transactions on Computers, C-17(7):718, July 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687437.

Serlin:1968:RPN

[Ser68b] O. Serlin. R68-33 PHENO: A new concept of hybrid computing elements. IEEE Transactions on Computers, C-17(7):718–719, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687438.

Seshagiri:1963:ACT

[Ses63] N. Seshagiri. Analog-computer technique for impulsive atmospheric radio-noise analysis. IEEE Transactions on Electronic Computers, EC-12(4):413–414, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037926.

Seshu:1965:RBP

[Ses65a] S. Seshu. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-14(1):89–94, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038363.

Seshu:1965:IDP

[Ses65b] Sundaram Seshu. On an improved diagnosis program. IEEE Transactions on Electronic Computers, EC-14(1):76–79, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038358.

Sethares:1966:LSH

[Set66a] G. C. Sethares. On the limit sets of the 2 × 1 homogeneous nets. IEEE Transactions on Electronic Computers, EC-15(5):810–812, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038894.

Sethares:1966:CSB

[Set66b] George Sethares. Closed sets of Boolean functions and the reliability problem for polyfunctional nets. IEEE Transactions on Electronic Computers, EC-15(1): 115–117, February 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038679.

Sethares:1967:CCB

[Set67] G. C. Sethares. On the complete convergence of bordered nets. IEEE Transactions on Electronic Computers, EC-16(1):92–93, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038992.

Seshu:1962:DAS

[SF62] S. Seshu and D. N. Freeman. The diagnosis of asynchronous sequential switching systems. IRE Transactions on Electronic Computers, EC-11(4):459–465, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219384.

Shemer:1969:SAP

[SG69] J. E. Shemer and S. C. Gupta. A simplified analysis of processor “Look-Ahead” and simultaneous operation of a multi-module main memory. IEEE Transactions on Computers, C-18(1):64–71, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671119.

Stearns:1961:SAP

[SH61] R. E. Stearns and J. Hartmanis. On the state assignment problem for sequential machines II. IRE Transactions on Electronic Computers, EC-10(4):593–603, December 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219265.

Sheng:1966:TRT

[SH66a] C. L. Sheng and H. R. Hwa. Testing and realization of threshold functions by successive higher ordering of incremental weights. IEEE Transactions on Electronic Computers, EC-15(2):212–220, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038715.

Sih:1966:CCM

[SH66b] K. Y. Sih and M. Y. Hsiao.

Cyclic codes in multiple channel parallel systems. IEEE Transactions on Electronic Computers, EC-15(6):927–930, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038938.

Smith:1966:CAM

[SH66c] D. O. Smith and K. J. Harte.

Content-addressed memory using magneto- or electro-optical interrogation. IEEE Transactions on Electronic Computers, EC-15 (1):123, February 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038685. See comment

[SH67b].

Sheng:1967:TRT

[SH67a] C. L. Sheng and H. R. Hwa. Testing and realization of threshold functions with don’t cares. IEEE Transactions on Electronic Computers, EC-16(6):868–870, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039207.

Smith:1967:CCA

[SH67b] D. O. Smith and K. J. Harte.

Comment on “Content-Addressed Memory Using Magneto-or Electro-Optical Interrogation”. IEEE Transactions on Electronic Computers, EC-16(3):372, June 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039084. See [SH66c].

Smith:1968:ATL

[SH68] John R. Smith and Cyrus O. Harbourt. An adaptive threshold logic gate using capacitive analog weights. IEEE Transactions on Computers, C-17(1):78–81, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008875.

Shaffer:1963:RTS

[Sha63] S. S. Shaffer. R63-105 time sharing on the Ferranti-Packard FP6000 computer system. IEEE Transactions on Electronic Computers, EC-12(6):935, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038056.

Shaffer:1965:ACC

[Sha65] S. S. Shaffer. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-14(3):503–516, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038479.

Shamir:1966:ICB

[Sha66a] E. Shamir. Implementation of communications-based EDP information systems. IEEE Transactions on Electronic Computers, EC-15(2):271–272, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038732.

Sharma:1966:IAS

[Sha66b] R. L. Sharma. Information about the Society. IEEE Transactions on Electronic Computers, EC-15 (2):297–299, April 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038740.

Shanks:1969:CFW

[Sha69] J. L. Shanks. Computation of the fast Walsh–Fourier transform. IEEE Transactions on Computers, C-18(5):457–459, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671278.

Sellers:1968:AEB

[SHB68] F. F. Sellers, Jr., M. Y. Hsiao, and L. W. Bearnson. Analyzing errors with the Boolean difference. IEEE Transactions on Computers, C-17(7):676–683, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687424. See correction

[SHB69].

Sellers:1969:CAE

[SHB69] F. F. Sellers, Jr., M. Y. Hsiao, and L. W. Bearnson. Correction to “Analyzing Errors with the Boolean Difference”. IEEE Transactions on Computers, C-18(4):381, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671262. See [SHB68].

Shew:1962:HDM

[She62] Lester F. Shew. High-density magnetic head design for noncontact recording. IRE Transactions on Electronic Computers, EC-11(6):764–773, December 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219461.

Sheingold:1963:CCS

[She63a] Daniel H. Sheingold. Constant current source for analog computer use. IEEE Transactions on Electronic Computers, EC-12(3):324, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037879.

Shevel:1963:ACC

[She63b] W. L. Shevel. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(3):337–349, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037884.

Shew:1963:DTS

[She63c] Lester F. Shew. Discrete tracks for saturation magnetic recording. IEEE Transactions on Electronic Computers, EC-12(4): 383–387, August 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037911.

Sheehan:1964:ACS

[She64a] Christopher J. Sheehan. An analog computer simulation of the restricted three-body problem by automatic scale-changing techniques. IEEE Transactions on Electronic Computers, EC-13 (1):41–50, February 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038075.

Sheng:1964:MTR

[She64b] C. L. Sheng. A method for testing and realization of threshold functions. IEEE Transactions on Electronic Computers, EC-13(3):232–239, June 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038147.

Sheng:1965:CST

[She65a] C. L. Sheng. Compound synthesis of threshold-logic network for the realization of general Boolean functions. IEEE Transactions on Electronic Computers, EC-14(6): 798–814, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038599.

Sheng:1965:CGI

[She65b] C. L. Sheng. Computer Group of The Institute of Electrical and Electronics Engineers, Inc. IEEE Transactions on Electronic Computers, EC-14(1):2–6, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038342.

Sheng:1965:C

[She65c] C. L. Sheng. Contributors. IEEE Transactions on Electronic Computers, EC-14(3):493–496, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038477.

Sheng:1965:DTS

[She65d] C. L. Sheng. Detection of totally symmetric Boolean functions. IEEE Transactions on Electronic Computers, EC-14(6): 924–926, December 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038614. See comment

[CD66a].

Sheng:1965:GIR

[She65e] C. L. Sheng. A graphical interpretation of realization of symmetric Boolean functions with threshold logic elements. IEEE Transactions on Electronic Computers, EC-14(1):8–18, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038344.

Sheng:1966:ACS

[She66] C. L. Sheng. Annual cumulative subject index. IEEE Transactions on Electronic Computers, EC-15(6):976–996, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038951.

Shively:1968:DPG

[Shi68] R. R. Shively. A digital processor to generate spectra in real time. IEEE Transactions on Computers, C-17(5):485–491, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687378.

Shook:1961:DSM

[Sho61] C. G. Shook. A digital static magnetic wire storage with nondestructive read-out. IRE Transactions on Electronic Computers, EC-10(1):56–62, March 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219152.

Short:1962:DCO

[Sho62] Robert A. Short. The design of complementary-output networks. IRE Transactions on Electronic Computers, EC-11(6): 743–753, December 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219458.

Short:1964:RDB

[Sho64] R. A. Short. R64-51 determining the best ordering of variables in cascade switching circuits. IEEE Transactions on Electronic Computers, EC-13(4):485–486, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038231.

Short:1966:ACC

[Sho66] R. A. Short. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-15(4):683–695, August 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038851.

Shumate:1962:RBC

[Shu62] M. S. Shumate. Review: Use of a Combined Analog-Digital System for Re-Entry Vehicle Flight Simulation, by A. N. Wilson. IRE Transactions on Electronic Computers, EC-11(3):420, June 1962. CODEN IRE-LAO. ISSN 0367-9950.

Sauter:1959:MLC

[SI59] W. Sauter and P. J. Isaacs. Microwave logic circuits using diodes. IRE Transactions on Electronic Computers, EC-8(3): 302–307, September 1959. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222687.

Sierra:1965:LPM

[Sie65] Hugh M. Sierra. Linear, passive, matched filter for digital magnetic recording. IEEE Transactions on Electronic Computers, EC-14(2):204–209, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038400.

Silberberg:1968:RHA

[Sil68a] M. Y. Silberberg. R68-34 hybrid Apollo docking simulation. IEEE Transactions on Computers, C-17(7):719, July 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687439.

Silverberg:1968:NOO

[Sil68b] M. Silverberg. Near-optimal ordering of electronic circuit equations. IEEE Transactions on Computers, C-17(12):1173–1174, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687287.

Singleton:1962:LSC

[Sin62] Richard C. Singleton. Load-sharing core switches based on block designs. IRE Transactions on Electronic Computers, EC-11(3):346– 351, June 1962. CODEN IRELAO. ISSN 0367-9950.

Singleton:1966:GSB

[Sin66] Richard C. Singleton. Generalized snake-in-the-box codes. IEEE Transactions on Electronic Computers, EC-15(4):596–602, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038833.

Singh:1969:ASC

[Sin69] S. Singh. Asynchronous sequential circuits with feedback. IEEE Transactions on Computers, C-18(5):440–450, May 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671276.

Seddon:1968:SVD

[SJ68] J. P. Seddon and R. A. Johnson. The simulation of variable delay. IEEE Transactions on Computers, C-17(1):89–94, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008879.

Stone:1969:SAF

[SJ69] H. S. Stone and C. L. Jackson. Structures of the affine families of switching functions. IEEE Transactions on Computers, C-18(3):251–257, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671231.

Sklansky:1963:TMC

[SK63] J. Sklansky and K. R. Kaplan. Transients in Markov chains. IEEE Transactions on Electronic Computers, EC-12(6):

921–922, December 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038045.

Stone:1965:CFS

[SK65] H. S. Stone and A. J. Korenjak.

Canonical form and synthesis of cellular cascades. IEEE Transactions on Electronic Computers, EC-14(6):852–862, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038604.

Sklansky:1960:CSA

[Skl60a] J. Sklansky. Conditional-sum addition logic. IRE Transactions on Electronic Computers, EC-9 (2):226–231, June 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219822.

Sklansky:1960:EST

[Skl60b] J. Sklansky. An evaluation of several two-summand binary adders. IRE Transactions on Electronic Computers, EC-9(2): 213–226, June 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219821.

Sklansky:1963:GST

[Skl63a] J. Sklansky. General synthesis of tributary switching networks. IEEE Transactions on Electronic Computers, EC-12(5):464–469, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037959.

Sklansky:1963:USA

[Skl63b] J. Sklansky. Ultimate-speed adders. IEEE Transactions on Electronic Computers, EC-12(2): 142–148, April 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037819.

Skramstad:1966:RHC

[Skr66] H. K. Skramstad. R66-81 A hybrid-code differential analyzer. IEEE Transactions on Electronic Computers, EC-15(6): 958–959, December 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038959.

Sklansky:1965:CTN

[SKS65] J. Sklansky, A. J. Korenjak, and H. S. Stone. Canonical tributary networks. IEEE Transactions on Electronic Computers, EC-14(6):961–963, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038629.

Spiegel:1965:NMT

[SL65] P. Spiegel and R. L. Luce. A nanosecond monolithic TTL gate. IEEE Transactions on Electronic Computers, EC-14(6): 944–946, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038622.

Slobodin:1962:SAD

[Slo62] R. J. Slobodin. On statistical analysis of a DTL gate. IRE Transactions on Electronic Computers, EC-11(4):577, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219406.

Slotnick:1963:CSI

[Slo63] D. L. Slotnick. The computer system issue. IEEE Transactions on Electronic Computers, EC-12(6):607–608, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038010.

Schnurmann:1969:SAC

[SM69] H. D. Schnurmann and K. Maling. A statistical approach to the computation of delays in logic circuits. IEEE Transactions on Computers, C-18(4):320–328, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671252.

Small:1968:PEW

[Sma68] A. W. Small. Partitions and edge-weighted pair-graphs. IEEE Transactions on Computers, C-17 (11):1089, November 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687267.

Smith:1961:PMD

[Smi61] Donald O. Smith. Proposal for magnetic domain-wall storage and logic. IRE Transactions on Electronic Computers, EC-10(4):708–711, December 1961. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219277.

Smith:1965:CAV

[Smi65a] Fred W. Smith. Coding analog variables for threshold logic discrimination. IEEE Transactions on Electronic Computers, EC-14(6):941–943, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038621.

Smith:1965:MTV

[Smi65b] R. A. Smith. Minimal threevariable NOR and NAND logic circuits. IEEE Transactions on Electronic Computers, EC-14(1): 79–81, February 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038359.

Smith:1966:BNT

[Smi66] David R. Smith. Bounds on the number of threshold functions. IEEE Transactions on Electronic Computers, EC-15(3): 368–369, June 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038772.

Smith:1968:PMC

[Smi68a] David R. Smith. A partitioning method for combinational synthesis. IEEE Transactions on Computers, C-17(1):72–75, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008873.

Smith:1968:PCD

[Smi68b] F. W. Smith. Pattern classifier design by linear programming. IEEE Transactions on Computers, C-17(4):367–372, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687347.

Smith:1968:RES

[Smi68c] J. L. Smith. R68-20 effects of scheduling on file memory applications. IEEE Transactions on Computers, C-17(5):520–521, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687385.

Smith:1969:DMP

[Smi69] F. W. Smith. Design of multicategory pattern classifiers with two-category classifier design procedures. IEEE Transactions on Computers, C-18(6):548–551, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671301.

Smith:1969:QSI

[SNS69] M. G. Smith, W. A. Notz, and E. Schischa. The questions of systems implementation with largescale integration. IEEE Transactions on Computers, C-18(8): 690–694, August 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671343.

Stuanciulescu:1968:MMF

[SO68] Florin Stuanciulescu and Mihai Francisc Anton Oprescu. A mathematical model of finite random sequential automata. IEEE Transactions on Computers, C-17 (1):27–31, January 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008865.

Sobol:1962:TAT

[Sob62] Harold Sobol. Time average thermal properties of a computer utilizing thin-film superconducting elements. IRE Transactions on Electronic Computers, EC-11 (2):200–212, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219353.

Songster:1963:NBN

[Son63] G. F. Songster. Negativebase number-representation systems. IEEE Transactions on Electronic Computers, EC-12(3): 274–277, June 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037863.

Songster:1964:RBP

[Son64] G. Songster. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-13(2):160–164, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038122.

Sydnor:1956:AMS

[SOS56] R. L. Sydnor, T. R. O’Meara, and J. Strathman. Analog multipliers and squarers using a multigrid modulator. IRE Transactions on Electronic Computers, EC-5(2):82–85, June 1956. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219804.

Smith:1962:NAR

[SP62] W. R. Smith and A. V. Pohm.

A new approach to resistortransistor-tunnel-diode nanosecond logic. IRE Transactions on Electronic Computers, EC-11(5): 658–664, October 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219428.

Steinbuch:1963:LMT

[SP63] K. Steinbuch and U. A. W. Piske. Learning matrices and their applications. IEEE Transactions on Electronic Computers, EC-12(6):846–862, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038032.

Sethares:1969:GCM

[SP69] G. C. Sethares and J. N. Pierce. On the generation of a class of multithreshold functions. IEEE Transactions on Computers, C-18(6):557–559, June 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671303.

Spandorfer:1966:AI

[Spa66] L. M. Spandorfer. Author index. IEEE Transactions on Electronic Computers, EC-15(3):410, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038789.

Specker:1965:CAL

[Spe65] W. H. Specker. A class of algorithms for ln x, exp x, sin x, cos x, tan^{−1} x, and cot^{−1} x. IEEE Transactions on Electronic Computers, EC-14(1):85–86, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038361.

Specht:1967:GPD

[Spe67] Donald F. Specht. Generation of polynomial discriminant functions for pattern recognition. IEEE Transactions on Electronic Computers, EC-16(3):308–319, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039069.

Squire:1962:CFA

[Squ62] Jon S. Squire. An 11-cryotron full adder. IRE Transactions on Electronic Computers, EC-11 (5):710–711, October 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219439.

Srinivasan:1964:RIR

[Sri64a] C. V. Srinivasan. R64-78 interwoven redundant logic. IEEE Transactions on Electronic Computers, EC-13(6):760–761, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038329.

Srinivasan:1964:RBP

[Sri64b] C. V. Srinivasan. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-13(6):757–762, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038320.

Schwartz:1959:ETB

[SS59] S. J. Schwartz and J. S. Sallo. Electrodeposited twistor and bit wire components. IRE Transactions on Electronic Computers, EC-8(4):465–469, December 1959. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5222059.

Simmons:1961:SCP

[SS61] P. L. Simmons and R. F. Simmons. The simulation of cognitive processes: An annotated bibliography. IRE Transactions on Electronic Computers, EC-10(3): 462–483, September 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219236.

Simmons:1962:SCP

[SS62] P. L. Simmons and R. F. Simmons. The simulation of cognitive processes, II: An annotated bibliography. IRE Transactions on Electronic Computers, EC-11(4):535–552, August 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219394.

Shemer:1966:SAP

[SS66] J. E. Shemer and G. A. Shippey. Statistical analysis of paged and segmented computer systems. IEEE Transactions on Electronic Computers, EC-15(6):855–863, December 1966. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038928.

Suryanarayanan:1968:ACA

[SS68] K. L. Suryanarayanan and A. C.

Soudack. Analog computer automatic parameter optimization of nonlinear control systems with specific inputs. IEEE Transactions on Computers, C-17(8):782–788, August 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687451.

Shemer:1969:DBS

[SSG69] J. E. Shemer, Someshwar, and C. Gupta. On the design of Bayesian storage allocation algorithms for paging and segmentation. IEEE Transactions on Computers, C-18(7):644–651, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671326.

Schauer:1960:SAM

[SSPR60] R. F. Schauer, R. M. Stewart, A. V. Pohm, and A. A. Read. Some applications of magnetic film parametrons as logical devices. IRE Transactions on Electronic Computers, EC-9(3): 315–320, September 1960. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219851.

Stanciulescu:1965:SLA

[Sta65] Florin S. Stanciulescu. Sequential logic and its application to the synthesis of finite automata. IEEE Transactions on Electronic Computers, EC-14(6): 786–791, December 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038596.

Stearns:1960:MDP

[Ste60] Sam D. Stearns. A method for the design of pattern recognition logic. IRE Transactions on Electronic Computers, EC-9 (1):48–53, March 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221604.

Steck:1962:SMB

[Ste62] G. P. Steck. Stochastic model for the Browning–Bledsoe pattern recognition scheme. IRE Transactions on Electronic Computers, EC-11(2):274–282, April 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219360.

Stearns:1963:RSG

[Ste63a] R. E. Stearns. R63-90 sequence generators, graphs, and formal languages. IEEE Transactions on Electronic Computers, EC-12(5):581, October 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037996.

Stein:1963:GPR

[Ste63b] Irving Stein. Generalized pulse recording. IEEE Transactions on Electronic Computers, EC-12(2):77–92, April 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037803.

Stein:1963:ADP

[Ste63c] Marvin L. Stein. Automatic digital programming of analog computers. IEEE Transactions on Electronic Computers, EC-12 (2):100–111, April 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037805.

Stearns:1965:IAS

[Ste65] R. E. Stearns. Information about the Society. IEEE Transactions on Electronic Computers, EC-14 (2):307–313, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038426.

Stearman:1966:STM

[Ste66] G. H. Stearman. Is switching theory mathematics or engineering? IEEE Transactions on Electronic Computers, EC-15(1):124, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038686.

Steinmeyer:1969:LFG

[Ste69] D. Steinmeyer. Logarithm function generated by parallel resistors. IEEE Transactions on Computers, C-18(4):379–381, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671261.

Stone:1963:RAM

[Sto63] P. J. Stone. R63-20 adaptive mechanisms in digital “Concept” processing. IEEE Transactions on Electronic Computers, EC-12(1):37–38, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037789.

Stone:1964:RCS

[Sto64] P. J. Stone. R64-62 computer simulation of human interaction in small groups. IEEE Transactions on Electronic Computers, EC-13(5):648, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038284.

Stover:1965:IAS

[Sto65] D. Stover. Information about the Society. IEEE Transactions on Electronic Computers, EC-14(1):118–119, February 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038368.

Stone:1966:RAM

[Sto66a] H. S. Stone. R66-50 an algorithm for the machine calculation of complex Fourier series. IEEE Transactions on Electronic Computers, EC-15(4):680–681, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038857.

Stone:1966:RPS

[Sto66b] H. S. Stone. R66-51 programming semantics for multiprogrammed computations. IEEE Transactions on Electronic Computers, EC-15(4):681, August 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038858.

Stone:1967:AI

[Sto67] H. S. Stone. Author index.

IEEE Transactions on Electronic Computers, EC-16(5):709, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039173.

Strachey:1962:BRP

[Str62] Christopher Strachey. Book reviews: Planning a Computer System: Project Stretch. Edited by Werner Buchholz. 322 pp.(London: McGraw-Hill). The Computer Journal, 5(2):152–153, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). See [Buc62].

Strauss:1966:RAC

[Str66] J. C. Strauss. R66-68 an analysis of continuous parameter identification methods. IEEE Transactions on Electronic Computers, EC-15(5):840, October 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038912.

Strauss:1968:RTC

[Str68a] J. C. Strauss. R68-43 two continuous system modelling programs. IEEE Transactions on Computers, C-17(9):924, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687489.

Strauss:1968:RRA

[Str68b] J. C. Strauss. R68-6 regression analysis and parameter identification. IEEE Transactions on Computers, C-17(2):192, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687316.

Sturman:1968:AOI

[Stu68a] Joel N. Sturman. Asynchronous operation of an iteratively structured general-purpose digital computer. IEEE Transactions on Computers, C-17(1):10–17, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008863.

Sturman:1968:ISG

[Stu68b] Joel N. Sturman. An iteratively structured general-purpose digital computer. IEEE Transactions on Computers, C-17(1):2–9, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008862.

Stucki:1969:GGT

[Stu69] P. Stucki. Generation of grey tones by computer for simulation of visual information systems. IEEE Transactions on Computers, C-18(7):642–643, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671324.

Suran:1963:PN

[Sur63] J. J. Suran. PTGEC news. IEEE Transactions on Electronic Computers, EC-12(4):440, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037932.

Sureshchander:1968:CRF

[Sur68] Sureshchander. Comments on “RST Flip-Flop Input Equations”. IEEE Transactions on Computers, C-17(7):701–702, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687432. See [GD67].

Susskind:1958:DIP

[Sus58] Alfred K. Susskind. Digital information processing for machinetool control. IRE Transactions on Electronic Computers, EC-7 (2):136–140, June 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222525.

Sussenguth:1961:AAD

[Sus61] E. H. Sussenguth. An algorithm for automatic design of logical cryogenic circuits. IRE Transactions on Electronic Computers, EC-10(4):623–630, December 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219268.

Sutherland:1968:RSI

[Sut68] W. R. Sutherland. R68-53 A system for interactive graphical programming. IEEE Transactions on Computers, C-17(12): 1181–1182, December 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687293.

Sutherland:1969:MSA

[Sut69] I. E. Sutherland. A method for solving arbitrary-wall mazes by computer. IEEE Transactions on Computers, C-18(12): 1092–1097, December 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671185.

Svoboda:1963:ASB

[Svo63a] Antonin Svoboda. An algorithm for solving Boolean equations. IEEE Transactions on Electronic Computers, EC-12(5):557–559, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037975.

Svoboda:1963:SLS

[Svo63b] Antonin Svoboda. Synthesis of logical systems of given activity. IEEE Transactions on Electronic Computers, EC-12(6): 904–910, December 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038038.

Svoboda:1967:OI

[Svo67] A. Svoboda. Ordering of implicants. IEEE Transactions on Electronic Computers, EC-16(1): 100–105, February 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038996.

Svoboda:1969:DAS

[Svo69] A. Svoboda. Decimal adder with signed digit arithmetic. IEEE Transactions on Computers, C-18 (3):212–215, March 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671226.

Stone:1961:HOP

[SW61] R. B. Stone and G. M. White. High-order probability generators. IRE Transactions on Electronic Computers, EC-10(1):92– 93, March 1961. CODEN IRE-LAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219158.

Schorr:1963:SDS

[SW63] H. Schorr and N. E. Wiseman. System design of a small, fast digital computer. IEEE Transactions on Electronic Computers, EC-12(6):698–706, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038020.

Steinbuch:1965:CCT

[SW65] K. Steinbuch and B. Widrow. A critical comparison of two kinds of adaptive classification networks. IEEE Transactions on Electronic Computers, EC-14(5): 737–740, October 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038566.

Swift:1963:ACS

[Swi63] G. W. Swift. Analog computer simulation of a curved hysteresis loop. IEEE Transactions on Electronic Computers, EC-12(3):324, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037880.

Su:1968:USR

[SY68] C. C. Su and S. S. Yau. Unitary shift-register realizations of sequential machines. IEEE Transactions on Computers, C-17(4):312–324, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687342.

Szabo:1962:SDN

[Sza62] Nicholas Szabó. Sign detection in nonredundant residue systems. IRE Transactions on Electronic Computers, EC-11(4): 494–500, August 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219388.

Szentirmai:1967:RBP

[Sze67] G. Szentirmai. R67-42 bandpass quadrature filters. IEEE Transactions on Electronic Computers, EC-16(3):380, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039095.

Tabory:1967:AI

[Tab67] R. Tabory. Author index. IEEE Transactions on Electronic Computers, EC-16(3):396, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039089.

Taebel:1969:CVC

[Tae69] T. A. Taebel. Comment on “A Variable Counter Design Technique”. IEEE Transactions on Computers, C-18(1):80, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671126. See [Ren68,

Dav69b, Jon69a].

Tainiter:1963:RBP

[Tai63] M. Tainiter. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-12(6): 931–936, December 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038047.

Talkin:1964:SAP

[Tal64] Albert I. Talkin. A systematic approach to programming an analog computer to generate a large class of trajectories. IEEE Transactions on Electronic Computers, EC-13(5):594–597, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038253.

Tammaru:1967:RAC

[Tam67] E. Tammaru. R67-59 the analogue computer simulation of an alloy junction transistor wide band equivalent circuit. IEEE Transactions on Electronic Computers, EC-16(5):714–715, October 1967.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039179.

Tanaka:1965:A

[Tan65] Richard I. Tanaka. In appreciation.

IEEE Transactions on Electronic Computers, EC-14(5):682, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038552.

Tarjan:1965:CGN

[Tar65] R. Tarjan. Computer group news.

IEEE Transactions on Electronic Computers, EC-14(5):762, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038574.

Taube:1962:RBR

[Tau62] M. Taube. Review: Information Retrieval: State of the Art, by D. R. Swanson. IRE Transactions on Electronic Computers, EC-11(3): 426, June 1962. CODEN IRELAO. ISSN 0367-9950.

Tavernini:1965:ASB

[Tav65] Lucio Tavernini. The automatic sequencing of block-operators in the digital simulation of analog and analog-hybrid computers. IEEE Transactions on Electronic Computers, EC-14(6): 963–965, December 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038630.

Taylor:1962:MIN

[Tay62] G. W. Taylor. A method of increasing the number of orders in a digital computer. IRE Transactions on Electronic Computers, EC-11(3): 416–418, June 1962. CODEN IRE-LAO. ISSN 0367-9950.

Taylor:1965:UPS

[Tay65] George W. Taylor. Utilization of the t∗ partial switching properties of ferroelectrics in memory devices. IEEE Transactions on Electronic Computers, EC-14(6):881–886, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038608.

Taylo:1969:NSN

[Tay69] J. T. Taylo. Note on “Solution of Nonlinear Equations”. IEEE Transactions on Computers, C-18(12):1138–1140, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671194. See [Kuo68,

Ogb69, Vic69b, Tho69].

Thornley:1964:EBR

[TBS64] R. F. M. Thornley, A. V. Brown, and A. J. Speth. Electron beam recording of digital information. IEEE Transactions on Electronic Computers, EC-13(1): 36–40, February 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038074.

Teichmann:1955:CLC

[Tei55] T. Teichmann. Closed-loop control systems containing a digital computer. IRE Transactions on Electronic Computers, EC-4(3):106–117, September 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407909.

Teoste:1962:DRR

[Teo62] Rein Teoste. Design of a repairable redundant computer. IRE Transactions on Electronic Computers, EC-11(5):643–649, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219425.

Terasaki:1962:ACG

[Ter62] Richard M. Terasaki. Analog computation of Green’s function for integrating two-point boundary value problems. IRE Transactions on Electronic Computers, EC-11(1):57–63, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219317.

Townsend:1964:CFM

[TF64] C. J. Townsend and P. E. Fox.

Cylindrical film memory device characteristics. IEEE Transactions on Electronic Computers, EC-13(3):261–268, June 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038152.

Thiney:1963:RFT

[Thi63] A. Thiney. Rise and fall times of transistors in switching operation regardless of the driving source impedance. IEEE Transactions on Electronic Computers, EC-12(1):23, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037767.

Thiney:1964:RFT

[Thi64] A. Thiney. Rise and fall times of transistors in switching operation regardless of the driving source impedance. IEEE Transactions on Electronic Computers, EC-13(5):616–620, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038261.

Thomson:1955:TDC

[Tho55] W. E. Thomson. Time-delay circuits. IRE Transactions on Electronic Computers, EC-4(2):74, June 1955. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5407896.

Thomae:1960:NTA

[Tho60] M. A. Thomae. A new technique for analog integration and differentiation. IRE Transactions on Electronic Computers, EC-9(4):507–509, December 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219894.

Thorpe:1962:CCS

[Tho62] R. W. Thorpe. Constant current source for analog computer use. IRE Transactions on Electronic Computers, EC-11(6):792– 793, December 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219467.

Thomas:1967:AAC

[Tho67] Lee C. Thomas. An application of the analog computer to electronic circuit simulation. IEEE Transactions on Electronic Computers, EC-16(4):515–516, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039121.

Thomas:1968:TTD

[Tho68a] C. H. Thomas. Transport timedelay simulation for transmission line representation. IEEE Transactions on Computers, C-17 (3):205–214, March 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687322.

Thomas:1968:RCS

[Tho68b] L. C. Thomas. R68-9 A computer simulation of electrical loss and loading effect in magnetic recording. IEEE Transactions on Computers, C-17(3):298–300, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687336.

Thomson:1969:CSN

[Tho69] W. E. Thomson. Comment on “Solution of Nonlinear Equations”. IEEE Transactions on Computers, C-18(12):1138, December 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671193. See [Kuo68,

Ogb69, Vic69b, Tay69].

Tillman:1960:FNR

[Til60] Robert M. Tillman. Fluxlok — A nondestructive, random-access electrically alterable: High-speed memory technique using standard ferrite memory cores. IRE Transactions on Electronic Computers, EC-9(3):323–328, September 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219853.

Tison:1967:GCT

[Tis67] Pierre Tison. Generalization of consensus theory and application to the minimization of Boolean functions. IEEE Transactions on Electronic Computers, EC-16 (4):446–456, August 1967. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039109.

Tan:1969:SSD

[TMF69] Chung-Jen Tan, P. R. Menon, and A. D. Friedman. Structural simplification and decomposition of asynchronous sequential circuits. IEEE Transactions on Computers, C-18(9):830–838, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671373.

Tommerdahl:1964:MMP

[TNM64] J. B. Tommerdahl, A. C. Nelson, and K. K. Mazuy. Mathematical models for predicting pulse characteristics in digital logic systems. IEEE Transactions on Electronic Computers, EC-13(6): 705–710, December 1964. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038302.

Toda:1962:NTS

[Tod62] Iwao Toda. On the number of types of self-dual logical functions. IRE Transactions on Electronic Computers, EC-11 (2):282–284, April 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219361.

Todd:1963:ABN

[Tod63] Carl D. Todd. An annotated bibliography on NOR and NAND logic. IEEE Transactions on Electronic Computers, EC-12(5): 462–464, October 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037958.

Toda:1965:TSL

[Tod65] Iwao Toda. The tree set of a linear machine. IEEE Transactions on Electronic Computers, EC-14(6):954–957, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038626.

Tohma:1964:DLF

[Toh64] Y. Tohma. Decompositions of logical functions using majority decision elements. IEEE Transactions on Electronic Computers, EC-13(6):698–705, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038301.

Tompkins:1956:UDB

[Tom56] Howard E. Tompkins. Unit-distance binary-decimal codes for two-track commutation. IRE Transactions on Electronic Computers, EC-5 (3):139, September 1956. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219934.

Tomovich:1957:PMA

[Tom57] Rajko Tomovich. Picard’s method and analog computation. IRE Transactions on Electronic Computers, EC-6(1):34, March 1957. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221559.

Tomovich:1958:NAE

[Tom58] Rajko Tomovich. New applications of an electronic function generator. IRE Transactions on Electronic Computers, EC-7(1):48–51, March 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222095.

Topham:1967:ACC

[Top67] W. S. Topham. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-16(6):881–893, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039212.

Torng:1966:ARL

[Tor66] H. C. Torng. An approach for the realization of linearly-separable switching functions. IEEE Transactions on Electronic Computers, EC-15(1):14–20, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038662.

Torng:1968:AFS

[Tor68] H. C. Torng. An algorithm for finding secondary assignments of synchronous sequential circuits. IEEE Transactions on Computers, C-17(5):461–469, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687375.

Torng:1969:RSS

[Tor69] H. C. Torng. R69-16 satisfaction of sensitivity requirements by threshold logic elements. IEEE Transactions on Computers, C-18(7):668, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671332.

Tou:1960:DPI

[Tou60] Julus T. Tou. Digital programming for the inversion of Z transforms. IRE Transactions on Electronic Computers, EC-9 (3):373, September 1960. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219864.

Tomovic:1960:SIE

[TP60] R. Tomovic and N. Parezanovic. Solving integral equations on a repetitive differential analyzer. IRE Transactions on Electronic Computers, EC-9(4):503–506, December 1960. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219893.

Tomovic:1965:SDS

[TPM65] R. Tomovic, N. S. Parezanovic, and M. J. Merritt. Sensitivity of dynamic systems to parameters which increase the order of mathematical models. IEEE Transactions on Electronic Computers, EC-14(6):890–897, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038610.

Turner:1958:RRT

[TR58] L. R. Turner and J. H. Rawlings.

Realization of randomly timed computer input and output by means of an interrupt feature. IRE Transactions on Electronic Computers, EC-7(2):141–149, June 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222526.

Tracey:1966:ISA

[Tra66] James H. Tracey. Internal state assignments for asynchronous sequential machines. IEEE Transactions on Electronic Computers, EC-15(4):551–560, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038828.

Tracey:1967:RBP

[Tra67] J. H. Tracey. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-16(2):234–242, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039044.

Truitt:1962:ADR

[Tru62] T. D. Truitt. An analog-digital real-time computer. IRE Transactions on Electronic Computers, EC-11(1):46–52, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219315.

Truitt:1963:RAI

[Tru63] T. D. Truitt. R63-36 an analog input and output system for a real-time process control computer system. IEEE Transactions on Electronic Computers, EC-12(2):157, April 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037836.

Tsui:1961:FIM

[Tsu61] Frank F. Tsui. A flexible and inexpensive method of monitoring program execution in a digital computer. IRE Transactions on Electronic Computers, EC-10 (2):253–259, June 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219196.

Tsui:1962:IPS

[Tsu62] Frank F. Tsui. Improving the performance of the sense-amplifier circuit through pre-amplification strobing and noise-matched clipping. IRE Transactions on Electronic Computers, EC-11(5): 677–683, October 1962. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219430.

Tsybakov:1963:N

[Tsy63] B. S. Tsybakov. Notices. IEEE Transactions on Electronic Computers, EC-12(3):353, June 1963.

CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037888.

Tung:1968:DAS

[Tun68] Chin Tung. A division algorithm for signed-digit arithmetic. IEEE Transactions on Computers, C-17(9):887–889, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687477.

Uhr:1961:PMC

[Uhr61] Leonard Uhr. A possibly misleading conclusion as to the inferiority of one method for pattern recognition to a second method to which it is guaranteed to be superior. IRE Transactions on Electronic Computers, EC-10 (1):96–97, March 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219163.

Uhr:1962:RBR

[Uhr62] L. Uhr. Review: Statistical Recognition Functions and the Design of Pattern Recognizers, by T. Marill and D. M. Green. IRE Transactions on Electronic Computers, EC-11(3): 423, June 1962. CODEN IRELAO. ISSN 0367-9950.

Ulbrich:1963:SAT

[Ulb63] Egbert Ulbrich. Struktur und Arbeitsweise der Telefunken-Digitalrechenanlage TR 4. IEEE Transactions on Electronic Computers, EC-12(6):

613–618, December 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038012.

Ullman:1969:REE

[Ull69a] J. D. Ullman. R69-19 on the elimination of endmarkers. IEEE Transactions on Computers, C-18(7):669, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671335.

Ullmann:1969:ETM

[Ull69b] J. R. Ullmann. Experiments with the n-tuple method of pattern recognition. IEEE Transactions on Computers, C-18(12): 1135–1137, December 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671192.

Ulzurrun:1963:TDT

[Ulz63] Eduardo T. Ulzurrun. Tunneldiode threshold discriminator tolerance analysis. IEEE Transactions on Electronic Computers, EC-12(3):296–299, June 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037866.

Ulzurrun:1966:SET

[Ulz66] Eduardo Tomas Ulzurrun. Stress effects on thin-film parametrons. IEEE Transactions on Electronic Computers, EC-15(4):647–657, August 1966. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038839.

Urbano:1956:TMD

[UM56] R. H. Urbano and R. K. Mueller.

A topological method for the determination of the minimal forms of a Boolean function. IRE Transactions on Electronic Computers, EC-5(3):126–132, September 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219932.

Uncapher:1959:PMS

[Unc59] K. W. Uncapher. 1958 PGEC membership survey report. IRE Transactions on Electronic Computers, EC-8(1):60–67, March 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222763.

Uncapher:1961:PMR

[Unc61] Keith W. Uncapher. 1960 PGEC membership report. IRE Transactions on Electronic Computers, EC-10(1):81–91, March 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219157.

Uncapher:1964:MNC

[Unc64] Keith W. Uncapher. Message from the new Chairman. IEEE Transactions on Electronic Computers, EC-13(3):184, June 1964.

CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038142.

Unger:1963:ACC

[Ung63a] S. H. Unger. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-12(1):39–51, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037774.

Unger:1963:AN

[Ung63b] S. H. Unger. AFIPS news. IEEE Transactions on Electronic Computers, EC-12(1):55–56, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037778.

Unger:1963:C

[Ung63c] S. H. Unger. Contributors. IEEE Transactions on Electronic Computers, EC-12(1):28–29, February 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037772.

Unger:1965:FTS

[Ung65] Stephen H. Unger. Flow table simplification-some useful aids. IEEE Transactions on Electronic Computers, EC-14(3):472–475, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038469.

Unger:1968:RAD

[Ung68] S. H. Unger. A row assignment for delay-free realizations of flow tables without essential hazards. IEEE Transactions on Computers, C-17(2):146–151, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687305.

Urbano:1963:BMS

[Urb63] R. H. Urbano. Boolean matrices and the stability of neural nets. IEEE Transactions on Electronic Computers, EC-12(2):61–66, April 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037801.

Urbano:1964:CUR

[Urb64] R. H. Urbano. On the convergence and ultimate reliability of iterated neural nets. IEEE Transactions on Electronic Computers, EC-13(3):204–225, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038145.

Urbano:1965:MCA

[Urb65a] Rocco H. Urbano. Matrix criteria for arbitrary reliability in iterated neural nets. IEEE Transactions on Electronic Computers, EC-14(4):627–629, August 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038517.

Urbano:1965:SNR

[Urb65b] Rocco H. Urbano. Some new results on the convergence, oscillation, and reliability of polyfunctional nets. IEEE Transactions on Electronic Computers, EC-14(6):769–781, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038594.

Urbano:1966:RBP

[Urb66] R. H. Urbano. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-15(5):833–841, October 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038900.

Urbano:1968:SFP

[Urb68] R. H. Urbano. Structure and function in polyfunctional nets. IEEE Transactions on Computers, C-17(2):152–173, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687306.

Uzunoglu:1962:CUT

[Uzu62] Vasil Uzunoglu. Circuits using tunnel diode flip-flops and PNPN diodes. IRE Transactions on Electronic Computers, EC-11(6):793–796, December 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219468.

Vairavan:1969:LBM

[Vai69] K. Vairavan. On the lower bound to the memory of finite state machines. IEEE Transactions on Computers, C-18(9):856–861, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671375.

Valentinuzzi:1967:TVP

[Val67] M. Eugenio Valentinuzzi. Threevalued propositional calculus of Lukasiewicz. and three-position double switches. IEEE Transactions on Electronic Computers, EC-16(1):39–44, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038981.

VanHorne:1957:AMS

[Van57] Thomas B. Van Horne. An analog method for the solution of probability of hit and related statistical problems. IRE Transactions on Electronic Computers, EC-6(3):170–175, September 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222015.

VanHeerden:1959:ABT

[Van59] P. J. Van Heerden. Analysis of binary time series in periodic functions. IRE Transactions on Electronic Computers, EC-8(2):228–229, June 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219528.

VanHeerden:1960:PBT

[Van60a] P. J. Van Heerden. Periodic binary time series and their relation to Boolean functions. IRE Transactions on Electronic Computers, EC-9(4):510, December 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219897.

Vandling:1960:SMO

[Van60b] G. C. Vandling. The simplification of multiple-output switching networks composed of unilateral devices. IRE Transactions on Electronic Computers, EC-9(4):477–486, December 1960. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219889.

VanderVelde:1966:RHC

[Van66] W. E. Vander Velde. R66-85 hybrid computer Monte Carlo techniques. IEEE Transactions on Electronic Computers, EC-15(6): 960–961, December 1966. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038963.

Vidal:1965:FAT

[VB65] J. J. Vidal and S. M. Bawin. Frequency analysis of truncation errors in RC networks. IEEE Transactions on Electronic Computers, EC-14(2):229–233, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038403.

VanDeGoor:1969:CUD

[VB69] A. Van De Goor and C. G. Bell.

A control unit for a DEC PDP-8 computer and a Burroughs disk. IEEE Transactions on Computers, C-18(11):1044–1048, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671171.

VanDeGoor:1969:DBT

[VBW69] A. Van De Goor, C. G. Bell, and D. A. Witcraft. Design and behavior of TSS/8: a PDP-8 based time-sharing system. IEEE Transactions on Computers, C-18 (11):1038–1043, November 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671170.

Verster:1964:MIA

[Ver64] T. C. Verster. A method to increase the accuracy of fast-serial-parallel analog-to-digital converters. IEEE Transactions on Electronic Computers, EC-13(4):471–473, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038216.

Vichnevetsky:1967:RTF

[Vic67a] R. Vichnevetsky. R67-31 transfer functions and the matric computor.

IEEE Transactions on Electronic Computers, EC-16(2):242, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039057.

Vichnevetsky:1967:EAC

[Vic67b] Robert Vichnevetsky. Error analysis in the computer simulation of dynamic systems: Variational aspects of the problem. IEEE Transactions on Electronic Computers, EC-16(4):403–411, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039104.

Vichnevetsky:1969:UFA

[Vic69a] R. Vichnevetsky. Use of functional approximation methods in the computer solution of initial value partial differential equation problems. IEEE Transactions on Computers, C-18(6):499–512, June 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671295.

Vickers:1969:CSN

[Vic69b] V. E. Vickers. Comment on “Solution of Nonlinear Equations”. IEEE Transactions on Computers, C-18(3):277, March 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671237. See [Kuo68,

Ogb69, Tho69, Tay69].

Vidal:1967:RAP

[Vid67] J. Vidal. R67-60 ADAC-A programmed direct analog computer. IEEE Transactions on Electronic Computers, EC-16(5):715, October 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039180.

Varshavsky:1965:NCT

[VO65] V. Varshavsky and B. Ovsievich.

Networks composed of ternary majority elements. IEEE Transactions on Electronic Computers, EC-14(5):730–733, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038564.

Volder:1959:CTC

[Vol59] Jack E. Volder. The CORDIC trigonometric computing technique. IRE Transactions on Electronic Computers, EC-8(3): 330–334, September 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222693.

Vivatson:1962:CRTa

[VS62a] A. L. Vivatson and W. C. Slemmer.

Comments on “Real-Time Analog-Digital Computation”. IRE Transactions on Electronic Computers, EC-11(5):707, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219434. See [Con62,

VS62b].

Vivatson:1962:CRTb

[VS62b] A. L. Vivatson and W. C. Slemmer. Comments on “Real-Time Analog-Digital Computation”. IRE Transactions on Electronic Computers, EC-11(5):707–708, October 1962. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219435. See [Con62,

VS62a].

Waaben:1967:HSP

[Waa67] Sigurd Waaben. High-speed platedwire memory system. IEEE Transactions on Electronic Computers, EC-16(3):335–343, June 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039073.

Waaben:1968:HSI

[Waa68] S. Waaben. High-speed, interlaced WRITE and READ-only operation of a plated-wire memory system. IEEE Transactions on Computers, C-17(11):1062–1065, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687263.

Wadel:1957:NBN

[Wad57] Louis B. Wadel. Negative base number systems. IRE Transactions on Electronic Computers, EC-6(2):123, June 1957. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5221586.

Wadel:1961:CCN

[Wad61] Louis B. Wadel. Conversion from conventional to negative-base number representation. IRE Transactions on Electronic Computers, EC-10(4):779, December 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219290.

Wadhwa:1964:STO

[Wad64] L. K. Wadhwa. Simulation of third order systems by a single operational amplifier. IEEE Transactions on Electronic Computers, EC-13(2):128–147, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038110.

Wadhwa:1968:SNT

[Wad68] L. K. Wadhwa. Simulation of nonquiescent third-order systems by a single operational amplifier. IEEE Transactions on Computers, C-17(10):986–987, October 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687244.

Wagner:1966:CMT

[Wag66] Eric G. Wagner. On connecting modules together uniformly to form a modular computer. IEEE Transactions on Electronic Computers, EC-15(6):864–873, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038929.

Wagner:1967:RTO

[Wag67] W. E. Wagner. R67-39 trajectory optimization using fast-time repetitive computation. IEEE Transactions on Electronic Computers, EC-16(3):379, June 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039092.

Wait:1964:AI

[Wai64a] J. V. Wait. Author index. IEEE Transactions on Electronic Computers, EC-13(5):668, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038278.

Waite:1964:PCS

[Wai64b] William M. Waite. The production of completion signals by asynchronous, iterative networks. IEEE Transactions on Electronic Computers, EC-13(2):83–86, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038103.

Waite:1965:IAS

[Wai65] W. M. Waite. Information about the Society. IEEE Transactions on Electronic Computers, EC-14(4):677–681, August 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038531.

Wait:1966:RES

[Wai66] J. V. Wait. R66-12 experimental study of a new method of time delay for analog computers. IEEE Transactions on Electronic Computers, EC-15(1):130–131, February 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038698.

Wait:1967:SSM

[Wai67] J. V. Wait. State-space methods for designing digital simulations of continuous fixed linear systems. IEEE Transactions on Electronic Computers, EC-16(3):351–355, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039075. See correction

[Wai68a].

Wait:1968:CSS

[Wai68a] J. V. Wait. Correction to “State-Space Methods for Designing Digital Simulations of Continuous Fixed Linear Systems”. IEEE Transactions on Computers, C-17(4):385, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687352. See [Wai67].

Waite:1968:RPD

[Wai68b] W. M. Waite. R68-24 A proposal for definitions in ALGOL. IEEE Transactions on Computers, C-17(5):522–523, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687389.

Wakshman:1969:RBA

[Wak69] A. Wakshman. R69-10 bounded action machines: Toward an abstract theory of computer structure. IEEE Transactions on Computers, C-18(5):477, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671286.

Walters:1953:HRL

[Wal53] Louis G. Walters. Hidden regenerative loops in electronic analog computers. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(2):1–4, June 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Wallace:1964:SFM

[Wal64] C. S. Wallace. A suggestion for a fast multiplier. IEEE Transactions on Electronic Computers, EC-13(1):14–17, February 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038071.

Wallace:1967:AI

[Wal67] V. L. Wallace. Author index. IEEE Transactions on Electronic Computers, EC-16(2):260, April 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039047.

Wanlass:1955:TCD

[Wan55] C. L. Wanlass. Transistor circuitry for digital computers. IRE Transactions on Electronic Computers, EC-4(1):11–16, March 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407947.

Wang:1965:INS

[Wan65] F. L. Wang. An isomorphic notation of switching circuits. IEEE Transactions on Electronic Computers, EC-14(6):952–954, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038625.

Wang:1966:AGB

[Wan66] Michael C. Wang. An algorithm for Gray-to-binary conversion. IEEE Transactions on Electronic Computers, EC-15(4):659–660, August 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038841.

Wang:1967:GLF

[Wan67] K. C. Wang. Graphs of linear forms on modules over Boolean rings. IEEE Transactions on Electronic Computers, EC-16(5): 670–671, October 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039159.

Wang:1969:SLS

[Wan69] Kung-Chi Wang. Synthesis of linear sequential machines with unspecified outputs. IEEE Transactions on Computers, C-18(2): 145–153, February 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671210.

Warfield:1958:SCT

[War58a] J. N. Warfield. Switching circuits as topological models in discrete probability theory. IRE Transactions on Electronic Computers, EC-7(3):251–252, September 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222586.

Warfield:1958:NRS

[War58b] John N. Warfield. A note on the reduction of switching functions. IRE Transactions on Electronic Computers, EC-7 (2):180–181, June 1958. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222531.

Ware:1960:SCT

[War60] Willis H. Ware. Soviet computer technology — 1959. IRE Transactions on Electronic Computers, EC-9(1):72–120, March 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221607.

Ward:1965:CGN

[War65] D. M. Ward. Computer group news. IEEE Transactions on Electronic Computers, EC-14(1): 110–117, February 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038367.

Warfield:1966:RBP

[War66] J. N. Warfield. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-15 (1):127–131, February 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038688.

Wassel:1967:MRP

[Was67] Gustav N. Wassel. Multiple reflections on pulse signal transmission lines-model for computer solution. IEEE Transactions on Electronic Computers, EC-16 (2):193–202, April 1967. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039028.

Wassel:1968:MRR

[Was68] G. N. Wassel. Multiple reflections from RC loading of pulsesignal transmission lines. IEEE Transactions on Computers, C-17 (8):729–737, August 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687446.

Waters:1959:AMC

[Wat59] William E. Waters. An analog method for constructing a curve. IRE Transactions on Electronic Computers, EC-8(1):68, March 1959. CODEN IRELAO. ISSN 0367-9950. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5222764.

Wattenburg:1966:DAC

[Wat66] W. H. Wattenburg. Design automation for computer software. IEEE Transactions on Electronic Computers, EC-15(3):378–380, June 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038777.

Watts:1967:RRP

[Wat67] D. G. Watts. R67-14 randomprocess simulation and measurements. IEEE Transactions on Electronic Computers, EC-16 (1):112, February 1967. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039010.

Williams:1962:TAG

[WB62] D. S. Williams and J. T. Beckett. Teaching aid for “Games that Teach the Fundamentals of Computer Operation”. IRE Transactions on Electronic Computers, EC-11(3):415, June 1962. CODEN IRE-LAO. ISSN 0367-9950.

Weiner:1968:DSF

[WD68] P. Weiner and T. F. Dwyer. Discussion of some flaws in the classical theory of two-level minimization of multiple-output switching networks. IEEE Transactions on Computers, C-17(2):184–186, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687310.

Weiner:1969:MMT

[WD69] P. Weiner and T. A. Dolotta.

Mixed memory type realizations of sequential machines. IEEE Transactions on Computers, C-18 (3):272–277, March 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671236.

Weeg:1958:FSF

[Wee58] G. P. Weeg. Folding of symmetric functions. IRE Transactions on Electronic Computers, EC-7(4):325, December 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222668.

Weeg:1960:UWC

[Wee60] G. P. Weeg. Uniqueness of weighted code representations. IRE Transactions on Electronic Computers, EC-9(4):487–489, December 1960. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219890.

Weeks:1963:CSE

[Wee63] W. T. Weeks. Computer simulation of the electrical properties of memory arrays. IEEE Transactions on Electronic Computers, EC-12(6):874–887, December 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038035.

Weeg:1966:C

[Wee66] G. P. Weeg. Contributors. IEEE Transactions on Electronic Computers, EC-15(1):125–126, February 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038687.

Wee:1968:GIA

[Wee68] W. G. Wee. Generalized inverse approach to adaptive multiclass pattern classification. IEEE Transactions on Computers, C-17 (12):1157–1164, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687285.

Weeks:1969:MAF

[Wee69] W. T. Weeks. Mathematical analysis of ferrite core memory arrays. IEEE Transactions on Computers, C-18(5):409–416, May 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671272.

Weiss:1952:ACD

[Wei52] E. Weiss. Applications of CRC-105 decimal DDA. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-1(??): 19–24, December 1952. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (electronic).

Weibel:1957:EAM

[Wei57] Erich S. Weibel. An electronic analog multiplier using carriers. IRE Transactions on Electronic Computers, EC-6(1):30–34, March 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221558.

Weinstein:1963:HSC

[Wei63a] H. Weinstein. A high-speed “Compare” circuit. IEEE Transactions on Electronic Computers, EC-12(4):410–411, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037923.

Weinstein:1963:POS

[Wei63b] Hillel Weinstein. Proposals for ordered sequential detection of simultaneous multiple responses. IEEE Transactions on Electronic Computers, EC-12(5):564–567, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037978.

Weiss:1969:CPC

[Wei69a] C. D. Weiss. The characterization and properties of cascade realizable switching functions. IEEE Transactions on Computers, C-18(7):624–633, July 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671322.

Weiss:1969:OSA

[Wei69b] C. D. Weiss. Optimal synthesis of arbitrary switching functions with regular arrays of 2input 1-output switching elements. IEEE Transactions on Computers, C-18(9):839–856, September 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671374.

Wells:1964:ALD

[Wel64] Mark B. Wells. Aspects of language design for combinatorial computing. IEEE Transactions on Electronic Computers, EC-13(4):431–438, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038208.

Wells:1968:RSM

[Wel68] C. Wells. R68-52 the structure of the “The”-multiprogramming system. IEEE Transactions on Computers, C-17(12):1181, December 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687292.

Weller:1969:HSC

[Wel69] C. W. Weller. A high-speed carry circuit for binary adders. IEEE Transactions on Computers, C-18(8):728–732, August 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671348.

Wee:1968:APM

[WF68] W. G. Wee and K. S. Fu. An adaptive procedure for multiclass pattern classification. IEEE Transactions on Computers, C-17(2): 178–182, February 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687308.

Wendland:1962:IRF

[WH62] P. H. Wendland and P. M. Hanlet. An infinite-resolution function generator. IRE Transactions on Electronic Computers, EC-11(1):26–30, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219312.

Wedel:1956:ADA

[WHB56] J. J. Wedel, A. Huntington, and M. B. Bain. Automatic dataaccumulation system for wind tunnels. IRE Transactions on Electronic Computers, EC-5(1):7–11, March 1956. CODEN IRELAO. ISSN 0367-9950. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219784.

Whigham:1965:FAC

[Whi65] Robert H. Whigham. A fast analog comparison for hybrid computation. IEEE Transactions on Electronic Computers, EC-14(5): 717–719, October 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038558.

White:1966:ACT

[Whi66] M. E. White. An analog computer technique for solving a class of nonlinear ordinary differential equations. IEEE Transactions on Electronic Computers, EC-15(2):157–163, April 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038712.

White:1967:EDC

[Whi67] R. C. White. Experiments with digital computer simulations of pseudorandom noise generators. IEEE Transactions on Electronic Computers, EC-16(3):355–357, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039076.

Whipple:1969:CIC

[Whi69a] W. L. Whipple. Calculation of integrated circuit yields. IEEE Transactions on Computers, C-18(3):268, March 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671233.

Whipple:1969:CHR

[Whi69b] W. L. Whipple. Comments on “Higher-Radix Division Using Estimates of the Divisor and Partial Remainders”. IEEE Transactions on Computers, C-18(2):183, February 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671216. See [Atk68].

White:1969:DAE

[Whi69c] S. A. White. Digital adaptiveelement building blocks for MOS large-scale integration. IEEE Transactions on Computers, C-18 (8):699–706, August 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671345.

Wholey:1963:RAG

[Who63] J. S. Wholey. R63-19 an approach to general patteme recognition. IEEE Transactions on Electronic Computers, EC-12(1):37, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037788.

Widrow:1954:RFN

[Wid54] Bernard Widrow. A radiofrequency nondestructive readout for magnetic-core memories. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-3(4):12–15, December 1954. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

                                        tronic). URL http://isl-www.
                                        stanford.edu/~widrow/papers/
                                        j1954aradio.pdf.

Widrow:1964:SI

[Wid64] B. Widrow. Subject index. IEEE Transactions on Electronic Computers, EC-13(4):498–500, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038222.

Widrow:1967:SI

[Wid67] B. Widrow. Subject index. IEEE Transactions on Electronic Computers, EC-16(5):706–709, October 1967. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4039172.

Wier:1955:HSP

[Wie55] Joseph M. Wier. A high-speed permanent storage device. IRE Transactions on Electronic Computers, EC-4(1):16–20, March 1955. CODEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5407948.

Wierwille:1965:ESN

[Wie65a] Walter W. Wierwille. Experimental study of a new method of time delay for analog computers. IEEE Transactions on Electronic Computers, EC-14(4):617–623, August 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&

arnumber=4038514. See comments

[Bri66].

Wierwille:1965:TMC

[Wie65b] Walter W. Wierwille. A theory and method for correlation analysis of nonstationary signals. IEEE Transactions on Electronic Computers, EC-14(6):909–919, December 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038612.

Wiederhold:1968:RBC

[Wie68] G. Wiederhold. R68-19 bulk core in a 360/67 Time Sharing System. IEEE Transactions on Computers, C-17(5):520, May 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687384.

Wigington:1963:MOG

[Wig63] Ronald L. Wigington. A machine organization for a general purpose list processor. IEEE Transactions on Electronic Computers, EC-12(6):707–714, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038021.

Wijnhoven:1968:SHS

[Wij68] R. M. G. Wijnhoven. A simple high-speed magnetic access switch matrix. IEEE Transactions on Computers, C-17(6):542–550, June 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687398.

Wilkes:1961:SRC

[Wil61] M. V. Wilkes. Self-repairing computers. IRE Transactions on Electronic Computers, EC-10 (1):93–94, March 1961. CO-DEN IRELAO. ISSN 0367-

                                        9950. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5219159.

Wilkinson:1963:MGF

[Wil63] R. H. Wilkinson. A method of generating functions of several variables using analog diode logic. IEEE Transactions on Electronic Computers, EC-12(2):112–129, April 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037806. See comments

[Nat64].

Wilkinson:1964:NRI

[Wil64] R. H. Wilkinson. On the number of regions in input domain of a free distributive lattice. IEEE Transactions on Electronic Computers, EC-13(6):753–754, December 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038317.

Wilkes:1965:ACC

[Wil65a] M. V. Wilkes. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-14(2):281–293, April 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038422.

Wilkes:1965:SMD

[Wil65b] M. V. Wilkes. Slave memories and dynamic storage allocation. IEEE Transactions on Electronic Computers, EC-14(2):270–271, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038419.

Winett:1962:ASF

[Win62] Joel M. Winett. An alpha-state finite automaton for multiplication by alpha. IRE Transactions on Electronic Computers, EC-11(3): 412–413, June 1962. CODEN IRE-LAO. ISSN 0367-9950.

Winder:1963:BTG

[Win63] Robert O. Winder. Bounds on threshold gate realizability. IEEE Transactions on Electronic Computers, EC-12(5):561–564, October 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037977.

Winograd:1964:C

[Win64] S. Winograd. Contributors. IEEE Transactions on Electronic Computers, EC-13(6):755–756, December 1964. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038319.

Winder:1965:ESA

[Win65a] R. O. Winder. Enumeration of seven-argument threshold functions. IEEE Transactions on Electronic Computers, EC-14(3): 315–325, June 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038448. See correction

[Win67a].

Winder:1965:PTF

[Win65b] R. O. Winder. Properties of threshold functions. IEEE Transactions on Electronic Computers, EC-14(2):252–254, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038411.

Winder:1965:SI

[Win65c] R. O. Winder. Subject index. IEEE Transactions on Electronic Computers, EC-14(2):294–298, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038424.

Winograd:1965:C

[Win65d] S. Winograd. Contributors. IEEE Transactions on Electronic Computers, EC-14(1):87–88, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038362.

Winder:1966:CGN

[Win66a] R. O. Winder. Computer group news. IEEE Transactions on Electronic Computers, EC-15(2):295, April 1966. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038738.

Winder:1966:C

[Win66b] R. O. Winder. Contributors.

IEEE Transactions on Electronic Computers, EC-15(6):950–952, December 1966. CODEN IEECA8. ISSN 0367-7508. URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038948.

Winder:1967:CES

[Win67a] R. O. Winder. Correction to “Enumeration of Seven-Argument Threshold Functions”. IEEE Transactions on Electronic Computers, EC-16(2):231, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039042. See [Win65a].

Wingrove:1967:ACS

[Win67b] R. C. Wingrove. Annual cumulative subject index. IEEE Transactions on Electronic Computers, EC-16(6):893–910, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039213.

Winder:1968:STT

[Win68a] Robert O. Winder. Symmetry types in threshold logic. IEEE Transactions on Computers, C-17 (1):75–78, January 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008874.

Winograd:1968:NAI

[Win68b] S. Winograd. A new algorithm for inner product. IEEE Transactions on Computers, C-17(7):693–694, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687427.

Winder:1969:TGA

[Win69] R. O. Winder. Threshold gate approximations based on Chow parameters. IEEE Transactions on Computers, C-18(4):372–375, April 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671258.

Wirth:1965:RIF

[Wir65] N. Wirth. R65-24 introduction to FORMAC. IEEE Transactions on Electronic Computers, EC-14(2):278, April 1965. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038432.

Witsenhausen:1963:RAB

[Wit63] H. Witsenhausen. R63-68 APACHE — A breakthrough in analog computing. IEEE Transactions on Electronic Computers, EC-12(4): 419–420, August 1963. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037938.

Witsenhausen:1967:RNN

[Wit67] H. S. Witsenhausen. R67-43 nonobservable or noncontrollable analog schemes for the solution of a special class of differential equations. IEEE Transactions on Electronic Computers, EC-16(3):380, June 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039096.

Wierwille:1968:LCA

[WK68] W. W. Wierwille and J. R.

Knight. Off-line correlation analysis of nonstationary signals. IEEE Transactions on Computers, C-17(6):525–536, June 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687396.

Wilson:1961:ARB

[WL61] J. B. Wilson and R. S. Ledley. An algorithm for rapid binary division. IRE Transactions on Electronic Computers, EC-10(4):662–670, December 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219271.

Wallace:1967:RGT

[WL67] C. S. Wallace and O. Longe. Reading gapless tape. IEEE Transactions on Electronic Computers, EC-16(4):517–518, August 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039124.

Wilson:1953:MIU

[WM53] Louis D. Wilson and Saul Meyer. The Model II Unityper. Transactions of the I.R.E. Professional Group on Electronic Computers, EC-2(4):19–27, December 1953. CODEN ???? ISSN 2168-1740 (print), 2168-1759 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407690.

Wallmark:1959:IDU

[WM59] J. T. Wallmark and S. M. Marcus. Integrated devices using direct-coupled unipolar transistor logic. IRE Transactions on Electronic Computers, EC-8 (2):98–107, June 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219509.

Wolff:1962:CMA

[Wol62] C. H. Wolff. Coding for multiple asymmetric errors in one channel of a multichannel system. IRE Transactions on Electronic Computers, EC-11(5):655–658, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219427.

Wolinsky:1969:UIC

[Wol69] A. Wolinsky. Unified interval classification and unified 3classification for associative memories. IEEE Transactions on Computers, C-18(10):899–911, October 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671138.

Womack:1965:SPA

[Wom65] Charles P. Womack. Schmoo plot analysis of coincident-current memory systems. IEEE Transactions on Electronic Computers, EC-14(1):36–44, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038347.

Wong:1955:HDW

[Won55] S. Y. Wong. High density Williams storage. IRE Transactions on Electronic Computers, EC-4(4): 156–158, December 1955. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219486.

Wood:1963:RTT

[Woo63a] P. E. Wood. R63-107 ternary threshold logic. IEEE Transactions on Electronic Computers, EC-12(6):936, December 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038058.

Wood:1963:NTD

[Woo63b] Paul E. Wood. A note on threshold device error analysis. IEEE Transactions on Electronic Computers, EC-12(4):403–405, August 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037917.

Woo:1964:PIH

[Woo64a] P. W. Woo. A proposal for input of hand-drawn information to a digital system. IEEE Transactions on Electronic Computers, EC-13(5):609–611, October 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038257.

Wood:1964:HPS

[Woo64b] Paul E. Wood. Hazards in pulse sequential circuits. IEEE Transactions on Electronic Computers, EC-13(2):151–153, April 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038113.

Woodger:1964:A

[Woo64c] Michael Woodger. Algol. IEEE Transactions on Electronic Computers, EC-13(4):377–381, August 1964. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038200.

Wood:1965:DDA

[Woo65] Paul E. Wood. Digital differential analyzers with arbitrary stored interconnections. IEEE Transactions on Electronic Computers, EC-14(6):936–941, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038620.

Williams:1962:SCO

[WP62] R. W. Williams and G. M. Parker.

Synchro circuit with output proportional to input angle. IRE Transactions on Electronic Computers, EC-11(1):82, February 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219323.

Wald:1964:SIW

[WP64] B. Wald and F. A. Polknghorn.

Simple instructions which enhance the error control capabilities of a programmed communications processor. IEEE Transactions on Electronic Computers, EC-13(6):747–748, December 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038311.

Wallmark:1963:RUF

[WR63] J. T. Wallmark and A. G. Revesz.

Redundancy in unipolar field-effect transistor circuits. IEEE Transactions on Electronic Computers, EC-12(1):23–25, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4037768.

Wray:1962:WCD

[Wra62] W. J. Wray, Jr. Worst case design of variable-threshold TRL circuits. IRE Transactions on Electronic Computers, EC-11(3):382– 389, June 1962. CODEN IRELAO. ISSN 0367-9950.

Weinberger:1956:OMA

[WS56] A. Weinberger and J. L. Smith. A one-microsecond adder using onemegacycle circuitry. IRE Transactions on Electronic Computers, EC-5(2):65–73, June 1956. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219801.

Watts:1967:RCD

[WS67a] D. G. Watts and H. S. Stone. R67-15 A contribution to the development of ALGOL. IEEE Transactions on Electronic Computers, EC-16(1):112–113, February 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039011.

Weiner:1967:NDS

[WS67b] Peter Weiner and Edward J. Smith. On the number of distinct state assignments for synchronous sequential machines. IEEE Transactions on Electronic Computers, EC-16(2):220–221, April 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039033.

Weiner:1967:ORD

[WS67c] Peter Weiner and Edward J. Smith. Optimization of reduced dependencies for synchronous sequential machines. IEEE Transactions on Electronic Computers, EC-16(6):835–847, December 1967. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=4039200.

Wadel:1955:SEA

[WW55] L. B. Wadel and A. W. Wortham.

A survey of electronic analog computer installations. IRE Transactions on Electronic Computers, EC-4(2):52–55, June 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5407891.

Wyner:1964:SI

[Wyn64] A. D. Wyner. Subject index. IEEE Transactions on Electronic Computers, EC-13(3):337–339, June 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038173.

Yamada:1962:DLL

[Yam62a] Hisao Yamada. Disjunctively linear logic nets. IRE Transactions on Electronic Computers, EC-11(5):623–639, October 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219423.

Yamada:1962:RTC

[Yam62b] Hisao Yamada. Real-time computation and recursive functions not real-time computable. IRE Transactions on Electronic Computers, EC-11(6):753–760, December 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219459.

Yao:1963:AST

[Yao63] F. C. Yao. Analysis of signal transmission in ultra high speed transistorized digital computers. IEEE Transactions on Electronic Computers, EC-12(4):372–382, August 1963. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037910.

Yao:1965:INI

[Yao65] Frank C. Yao. Interconnection and noise immunity of circuitry in digital computers. IEEE Transactions on Electronic Computers, EC-14(6):875–880, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038607.

Yao:1966:PSS

[Yao66] Frank C. Yao. Propagation of sense signals in large-scale magnetic thin film memories. IEEE Transactions on Electronic Computers, EC-15(4):468–474, August 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038820.

Yau:1965:ACS

[Yau65] S. S. Yau. Autonomous clocks in sequential machines. IEEE Transactions on Electronic Computers, EC-14(3):467–472, June 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038468.

Yau:1966:RBP

[Yau66] S. S. Yau. Reviews of books and papers in the computer field. IEEE Transactions on Electronic Computers, EC-15(3):390–394, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038786.

Yourke:1961:EDL

[YBS61] H. S. Yourke, S. A. Butler, and W. G. Strohm. Esaki diode NOT– OR logic circuits. IRE Transactions on Electronic Computers, EC-10(2):183–190, June 1961. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219188.

Yen:1968:MMC

[Yen68a] Y. T. Yen. A mathematical model characterizing four-phase MOS circuits for logic simulation. IEEE Transactions on Computers, C-17(9):822–826, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687469.

Yen:1968:STP

[Yen68b] Yao Tung Yen. Some theoretical properties of multithreshold realizable functions. IEEE Transactions on Computers, C-17 (11):1081–1088, November 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687266.

Yen:1969:CAT

[Yen69] Yao Tung Yen. Computer-aided test generation for four-phase MOS LSI circuits. IEEE Transactions on Computers, C-18(10): 890–894, October 1969. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671136.

Yajima:1965:LBN

[YI65] S. Yajima and T. Ibaraki. A lower bound of the number of threshold functions. IEEE Transactions on Electronic Computers, EC-14(6):926–929, December 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038615.

Yajima:1968:RAL

[YI68a] S. Yajima and T. Ibaraki. Realization of arbitrary logic functions by completely monotonic functions and its applications to threshold logic. IEEE Transactions on Computers, C-17(4):338–351, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1687345.

Yajima:1968:TCM

[YI68b] S. Yajima and T. Ibaraki. A theory of completely monotonic functions and its applications to threshold logic. IEEE Transactions on Computers, C-17(3):214–229, March 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687323.

Yajima:1968:ALN

[YIK68] S. Yajima, T. Ibaraki, and I. Kawano. On autonomous logic nets of threshold elements. IEEE Transactions on Computers, C-17(4):385–391, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687353.

Yau:1968:RCS

[YO68] S. S. Yau and D. L. Ostapko. Realization of a class of switching functions by threshold-logic networks. IEEE Transactions on Computers, C-17(4):391–399, April 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687354.

Yoeli:1961:CDS

[Yoe61] M. Yoeli. The cascade decomposition of sequential machines. IRE Transactions on Electronic Computers, EC-10(4): 587–592, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219264.

Yoeli:1963:CPD

[Yoe63a] Michael Yoeli. Cascade-parallel decompositions of sequential machines. IEEE Transactions on Electronic Computers, EC-12(3): 322–324, June 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037878.

Yoeli:1963:CNB

[Yoe63b] Michael Yoeli. Counting with nonlinear binary feedback shift registers. IEEE Transactions on Electronic Computers, EC-12(4): 357–361, August 1963. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4037907.

Yoeli:1965:GTA

[Yoe65] Michael Yoeli. A group-theoretical approach to two-rail cascades. IEEE Transactions on Electronic Computers, EC-14(6):815–822, December 1965. CODEN IEECA8. ISSN 0367-7508. URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4038600.

Yoeli:1968:TCC

[Yoe68] Michael Yoeli. Ternary cellular cascades. IEEE Transactions on Computers, C-17(1):66–67, January 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008871.

Younker:1957:TDM

[You57] E. Leroy Younker. A transistordriven magnetic-core memory. IRE Transactions on Electronic Computers, EC-6(1):14–20, March 1957. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5221556.

Young:1964:N

[You64] F. H. Young. Notices. IEEE Transactions on Electronic Computers, EC-13(2):181–182, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038127.

Yovits:1959:OSM

[Yov59] Marshall C. Yovits. ONR Symposium on Microwave Techniques for Computing Systems. IRE Transactions on Electronic Computers, EC-8(3):262, September 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222680.

Yoeli:1965:LDT

[YR65a] M. Yoeli and G. Rosenfeld. Logical design of ternary switching circuits. IEEE Transactions on Electronic Computers, EC-14(1):19–29, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038345.

Yoeli:1965:NRS

[YR65b] M. Yoeli and G. Rosenfeld. A new reader service — publication of informational retrieval catalog cards. IEEE Transactions on Electronic Computers, EC-14(1):7–8, February 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038343.

Yau:1968:DPC

[YS68] S. S. Yau and J. M. Schumpert. Design of pattern classifiers with the updating property using stochastic approximation techniques. IEEE Transactions on Computers, C-17(9):861–872, September 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1687473.

Yau:1966:LSM

[YW66] S. S. Yau and K. C. Wang. Linearity of sequential machines. IEEE Transactions on Electronic Computers, EC-15(3):337–354, June 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038770.

Yang:1966:CCA

[YY66a] C. C. Yang and S. S. Yau. A cutpoint cellular associative memory. IEEE Transactions on Electronic Computers, EC-15(4): 522–528, August 1966. CO-DEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038825.

Yau:1966:NAT

[YY66b] S. S. Yau and C. C. Yang. A nonbulk addition technique for associative processors. IEEE Transactions on Electronic Computers, EC-15(6):938–941, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038944.

Yau:1966:PRU

[YY66c] S. S. Yau and C. C. Yang. Pattern recognition by using an associative memory. IEEE Transactions on Electronic Computers, EC-15(6):944–947, December 1966. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038946.

Zaitzeff:1959:RVU

[ZA59] E. M. Zaitzeff and M. M. Astrahan. Russian visit to U.S. computers. IRE Transactions on Electronic Computers, EC-8(4): 489–497, December 1959. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5222064.

Zackon:1961:VTD

[Zac61] David L. Zackon. Variable time delay by Padé approximation. IRE Transactions on Electronic Computers, EC-10(4): 783, December 1961. CO-DEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5219294.

Zahle:1966:CSS

[Zah66] Torben U. Zahle. On coding the states of sequential machines with the use of partition pairs. IEEE Transactions on Electronic Computers, EC-15(2):249–253, April 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038723.

Zuffo:1966:NTD

[ZP66] Joao A. Zuffo and Walter Del Picchia. New transistorized dcac amplitude converter. IEEE Transactions on Electronic Computers, EC-15(2):262, April 1966. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038728.

Zucker:1965:LEM

[Zuc65] M. S. Zucker. LOCS: An EDP machine logic and control simulator. IEEE Transactions on Electronic Computers, EC-14 (3):403–416, June 1965. CO-DEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038459.